A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1995–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.10

Title word cross-reference

#5824 [736].

(1 + 1) [653, 858, 1861]. (1 + 2) [1574, 1007]. (2 + 1) [78, 598, 695, 719, 1535, 2110, 1611, 1499, 1094, 953, 1834, 597, 1621]. (2, 0) [546]. (3/2) [214]. (a, b, c) [859]. (α, β) [662]. (D, X) = Id [1402]. (D = 1, 2, 3,) [2106]. (I, q) [164]. (N(≤ 2) + 1) [234]. (p^{2} + m^{2})^{1/2} − α/r [1192]. (ψ_{i}, ψ_{i})_{D}^{2} [1593]. (su(1, 1)) [1689]. (T^{∗}B)_{q}, q [912]. ∗ [323]. 1 [17]. 1 + 1 [1927, 2253, 1400]. 1 + 2 [528, 514, 1432]. 1/2 [1111, 1405, 581, 391, 1070]. 1/D [1692]. 1/r^{2} [21]. 2 [450, 410, 320, 915, 976, 889, 420, 1290, 746]. 2 + 1 [2253, 1542, 69, 1911, 2185, 1674, 307, 445, 1202, 1798, 1242]. {2, 2} [479]. 2N [1109]. 2 × 2 [2052]. 3 [188, 430, 1619, 1393, 1394, 428, 54]. 3 + 1 [1229, 2168, 2185]. 3 + 2 [1786]. 4 [439, 430, 379]. 5 [625]. 6 [1619, 1844, 81, 1099]. 90 [950]. 9j [1812, 2258]. ^{∗} [275, 778, 1763, 505, 282]. +(p, q) [1764]. _{h}(A_{∞}) [1744]. _{q} [1979]. A_{n}^{(1)} [1199]. A_{∞} [136]. A_{L} [2073]. A_{n} [1590]. A_{n}^{(1)} [1707]. A_{q,p}(sl_{n}) [1648]. α_{c} = 2/π [1192]. B_{0} [2145]. BF [430, 740]. R^{3} [145]. C [2057]. C(n + 1) ⊃ B(0, n) [774]. C^{∗} [1844]. C_{2}^{(1)} [2131]. U_{h,w}(H(4)) [1336]. cam [950]. CM(n) [1773]. CP [22]. CP^{1} [598, 72]. CP^{2} [1090]. CP^{n} [911, 226]. CP^{N−1} [1364]. D [1571, 2164, 556, 2106, 573, 625, 67, 104, 1824, 28, 2199, 1375, 1902, 1742]. d = 2 [1588]. d = 2, 3 [665]. D = 3 [1961]. D = 4 [696]. D > 0 [1269]. δ′ [332]. Δ = −1XXZ [385]. δ = 2 [313]. δ′ [2109]. DYh(glN)k [1509]. DYh(slN)k [1509]. E [2222]. e(4) [770]. η [100]. G [1510]. g > 1 [679]. G_{2} [995, 326]. γ [2090]. }↕_{\} [135]. gl(1|∞) [1921]. GL(2|1) [501]. GL(4, C) [973]. GL(8, R) [973]. gl(∞) [1247]. gl(m|n) ↓ osp(m|n) [2183]. gl(n.m) [1544]. GL_{q}(2) [497]. GL_{q}(n) [1141]. gl\sqrt{q}(n) [1060]. Gr(1|1) [2017]. Gr_{2}(C^{m+2}) [1404]. h [1586, 2216, 2017]. H^{2} [2160]. H_{n}(q) [370]. h̄ → 0 [57]. I [164]. I^{+} [1978]. IGL_{q}(n) [490]. ISO_{q}(2, 1) [532]. ISO_{q}(3) [532]. J [2200, 1619, 1844, 81, 1099]. K [125, 748, 1424, 2209]. κ [696, 475]. L [45, 1472, 379]. l^{P} [884]. L_{1} [921]. m [343]. N [2126, 2127, 366, 3, 736, 613, 1820, 484, 1464, 1608, 1930, 1260, 2241, 1740, 1990, 1530, 564, 1070, 1239, 1546, 589, 1627, 2068, 276, 735, 583, 2215, 1060, 1920, 1008, 257, 471, 1446, 1255, 663]. N = 1 [1001]. N = 2 [1939, 590, 1009, 1618, 1302, 1343, 1588, 1961, 1936, 2093, 2237, 2025, 1203]. N = 2SU(2) [893, 1294, 980]. N = 4 [590, 1009]. n^{−1} [1337]. O(3, 3) [291]. O(n) [1877, 573, 1255]. O(N) ⊂ O(N − 1) [1407]. O^{∗} [275]. O_{h} [655]. osp(1/2) [328]. osp(2|2) [913]. osp(D/d) [1735]. OSP_{q}(1|2n) [2023]. p [45, 1550, 2250, 2091, 2270, 1016, 454]. p + q [43]. ϕ^{4} [91, 745]. P T [2136, 1960]. Q [1682, 1184, 739, 1059, 557, 1060, 1139, 896, 1399, 604, 804, 191, 668, 560, 367, 279, 224, 770, 562, 617, 616, 1954, 369, 422, 530, 159, 1252, 923, 139, 52, 53, 345, 126, 2134, 1981, 164, 1350, 1791, 84, 342, 1790]. q_{t} = f(q, q_{x}, q_{xx}, q_{xxx}) [697]. q ∈ R^{+} [2021]. q ∈ S^{1} [2021]. q → 1 [804]. R [351, 2020, 1699, 2104, 721, 793, 1576, 1408, 1755, 1547, 1845]. R = 0 [685]. R^{1,1} [766]. R^{2} [1085]. R^{2}n [1952]. R^{3} [1633]. r^{D} [1269, 1747]. R_{4} [1220]. C^{∞}[SL(2, C)] [732]. ρ_{1} = ρ_{2} ≠ ρ_{3} [999]. ρ_{1} = ρ_{2} ≠ ρ_{3} [765]. W [731]. Z [728]. S [821, 1338, 1778, 379, 1115, 2112, 1992, 648]. S2 [1391]. S = 1/2 [506]. S^{2} [2160]. S^{n} [386, 2128]. S_{4} [1844]. S_{n} [370, 1737]. U_{⟨}(∫⊓(N)) [1005]. W [486]. sdiff(T^{2})_{loc} [808]. σ [2169]. sl(1|2) [380, 913]. sl(2) [1471, 1408, 1618]. SL(2, R) [975]. SL(2, R) [1487, 605, 239]. sl(m/n) [55]. sl(n)_{1} ⊗ sl(n)_{1}/sl(n)_{2} [561]. sl(n, C) [1957, 2264]. SL_{h}(2) [2212]. so [2157]. SO(1, 2) [500]. SO(1, 4) [1508]. SO(3) [1761, 603, 1403]. SO(4) × U(1) [1220]. SO(N) [1980, 1774, 1359, 1347, 1565]. so_{q}(3) [1954]. Sp(1, 2) [241]. Sp(4) ⊃ SU(2) × U(1) [694]. sp(4, R) [50]. Sp(6, R) [1693, 1101]. Sp(n, R) [1773]. spl(m/n) [286]. spl(N, 1) [1679]. SU(1, 1) [1138, 136, 1156, 707, 1689]. SU(2) [1647, 1741, 1967, 1936, 1138, 337, 609, 1216, 798, 596, 1360, 343, 958]. su(2) ⊕ ⋯ su(2) [62]. SU(2) × SU(2) [163]. SU(3) [1766, 2056, 2262, 924, 1253, 1101, 1216, 162]. SU(3) ⊃ U(2) [1728]. SU(4) [2093]. SU(4) ⊃ Sp(4) ⊃ SU(2) × U(1) [2077]. SU(N) [2232, 1980, 1727, 1728, 1296, 1327, 1987, 2027, 1303, 370, 1166, 80, 1675]. SU(n)/S(U(1) × U(n − 1)) [538]. SU(r + 1) [2237]. SU3 [775]. SU_{3} [161].

SU_{π} ⊗ SU_{ν}(3) [1454]. SU_{q}(2) [155, 2021]. SU_{q}(N) [369, 370]. t [2200]. T^{∗}M [1865]. T^{2} [1840]. T^{†} [1720]. T ⊃ C3 [1398]. T ⊃ D_{2} ⊃ C_{2} [1719]. U [1779]. U(1, 2) [241]. U(1, 3) ⊗_{s} H(1, 3) [1065]. U(2) [9, 230]. U(2, n − 2) [1727]. U(5) ⊃ SO(5) ⊃ SO(3) [340]. U(m|n) [220, 701, 1286, 700]. U(n) [1036]. U(N) ⊃ SO(N) ⊃ SO(N_{a}) ⊕ SO(N_{b}) [1365]. U(N) ⊃ U(N_{a}) ⊕ U(N_{b}) ⊃ SO(N_{a}) ⊕ SO(N_{b}) [1365]. U(ν + 1) ⊃ SO(ν + 1) ⊃ SO(ν) [672]. U(ν + 1) ⊃ U(ν) ⊃ SO(ν) [672]. U_{h}(sl(2)) [2212]. U_{q} [1689]. u_{q}(3) [2213, 872]. U_{q}(e^{N}) [321]. U_{q}(g^{l_{n+1}}) [2024]. U_{q}(∫↕_{∈}) [777]. U_{q}(gl(m/n)) [960]. U_{q}(gl(m|n)) [1182]. U_{q}(igl(n)) [490]. U_{q}(mathfrakb_{+}) [2055]. U_{q}(n) [49]. U_{q}(osp(1|2)) [81, 1099]. U_{q}(osp(2|2n)) [51]. U′_{q}(s^{l_{n+1}}) [2024]. U_{q}(sl(3)) [1096, 1175]. U_{q}(su(2)) [645]. U_{q}[gl(2/2)] [422]. U_{q}[gl(m|n)] [47]. U_{q}[gl(N|N)] [2223]. U_{q}osp(1|2) [1139]. U_{q}u(M) [1881]. U_{q}u(m, n) [1881]. ua [1393, 1394]. Uq(e(3)) [1368]. Uq(e(3, 1)) [1368]. Uq(e(N)) [1368]. Uq(iso(N)) [1368]. W [1588, 1345, 2114]. W (∞) [1518]. W_{KP}^{(n)} [923]. W_{1+∞} [1738]. W_{1+∞}(gl_{s}) [648]. W_{∞} [1300, 254, 2062, 2061]. wa [1393, 1394]. XXZ [1859, 512]. Z [2080]. Z_{2} [381, 964, 1776, 1174]. Z_{3} [1032, 529]. Z_{n} [1668, 235]. ζ [2075].

-adic [45, 454]. -algebra [505, 1518]. -algebras [778, 1763, 282, 663, 275, 1588]. -analog [912]. -body [589, 583, 1446, 3, 736, 613, 1820, 1740, 1990]. -braids [2253]. -branes [1550, 2250, 2091, 2270, 1016]. -cocycles [188, 420, 1290]. -coherent [1350]. -component [648, 1546]. -conformal [1510]. -cosymplectic [1424]. -Coulomb [739, 668]. -coupled [340]. -covariance [1300]. -covariant [1060, 1856, 2230]. -d [450, 889, 625]. -deformation [2017, 604, 560, 139]. -deformed [1060, 896, 1399, 770, 1954, 369, 696, 2134]. -derivations [275, 662]. -difference [224]. -dimensional [276, 78, 653, 735, 2215, 858, 1574, 695, 1920, 1008, 1393, 1394, 1902, 1094, 257, 953, 1834, 597, 1621, 1007, 1432, 2164, 556, 366, 2106, 234, 67, 104, 2199]. -dynamics [1489]. -entropies [617, 616, 345]. -epsilon [159]. -exponential [52, 53]. -exponentials [126]. -extended [1608]. -face [2131]. -Fock [1252]. -fold [564, 1239]. -forms [17, 320]. -Fourier [804]. -function [2075]. -gamma [52, 53]. -gauge [486, 2145]. -geometries [54]. -graded [1776, 1032, 529, 964, 164, 1174]. -harmonic [1790]. -Hermite [367]. -instantons [226]. -invariance [22]. -invariant [1737, 1980, 1255]. -isotropic [1059]. -Jacobi [770]. -level [2068]. -like [291]. -manifolds [428]. -matrices [721, 1408, 1755, 1845]. -matrix [793, 1576, 2104, 1547, 821, 1338, 351, 2020, 2112, 125, 1699]. -model [2169]. -models [91]. -oscillator [1184, 557, 191, 1682]. -parametric [1255]. -particle [471]. -phase-coherent [1791]. -phase-difference [1350]. -piece [1109]. -pseudodifferential [923]. -sheeted [343]. -soliton [484]. -space [2110, 475]. -spaces [921]. -sphere [1627, 2109]. -state [915]. -structures [323]. -superdimension [530]. -superplane [1586, 2216]. -supertrace [562].

-symbols [1812, 2258]. -symmetric [2136, 1960, 1840]. -symmetries [648]. -symmetry [1345]. -Trinomial [1981]. -type [1375]. -ultraspherical [279]. -wave [1992]. -Zetlin [1096, 1175].

1 [837]. 12 [736]. 13 [1775]. 19 [837]. 1PI [2244].

2 [1279]. 2D [1666, 1307].

34 [776, 166, 85]. 3435-3445 [1335]. 35 [568, 567, 374, 814, 776, 229, 230, 231, 376]. 36 [2271, 1477, 735, 1721, 1885, 1887, 1888, 1886, 1105, 375, 700, 2061, 815, 330, 1147, 634]. 37 [927, 2063, 969, 1009, 926, 928, 2064, 1218, 1256, 1592, 1106, 2273]. 38 [1335, 2272, 1448, 1521, 1729, 1776, 1294, 1959]. 39 [1958, 2263, 2262, 2261, 2101, 1731, 2225, 1730, 2266, 2104]. 3D [2033].

40 [2224, 2176, 2027, 2267, 2103, 2270, 2268, 2102, 2062, 2177, 2264, 2265]. 42 [2269]. 43 [736].

805g [837].

93m [776]. 94e [166, 85]. 95d [568, 567]. 95e [230, 231]. 95f [374]. 95g [776, 229]. 95h [814]. 95i [376]. 95k [330]. 96c [634]. 96g [735, 1147]. 96j [1105]. 96k [928, 815]. 97d [969, 375]. 97e [1009]. 97i [1106]. 97j [1256].

abelian [128, 1013, 92, 96, 936, 1119, 168, 1559, 2098, 1563, 1626, 2072, 2218, 1382, 1383, 1743]. abelianization [614]. Abelianness [1691]. Ablowitz [2119, 107]. above [1748]. Absence [2142, 332]. absolute [881]. absolutely [332]. absorption [1490]. absorptive [90]. Abstract [688, 699]. ac [1476]. acceleration [397]. Accepted [1183, 1219]. Access [1107, 1148]. according [411]. accumulating [60, 284]. Accuracy [1054, 1692]. acoustic [297]. across [1937]. action [1932, 2185, 796, 1652, 2019, 1262, 1026, 1155, 2198, 1255, 1356, 1505]. action-angle [1026, 1356]. actions [1651, 2024, 979]. activity [853]. adapted [1640, 655, 1789]. Addendum [969, 1721, 166, 634]. addition [1103]. Adelic [470]. Adiabatic [2187, 1056, 346, 347, 1339, 1453, 2034, 905]. adiabatically [2226]. adic [45, 454, 1457]. adjoint [574]. adjointness [1796, 1144]. Adler [926, 1315, 1274, 1598, 356]. admissibility [914]. admit [154]. admitting [1976, 467, 1207]. Advection [1606, 951, 994]. Advection-diffusion [994]. AF [1781]. affiliated [1880]. Affine [248, 1095, 153, 375, 2200, 2214, 960, 1618, 993, 1222, 1874, 420, 476, 1685, 2208, 1002, 609, 415, 1687, 968, 2223, 473]. again [1650]. against [76]. Aharonov [1375, 1898, 26, 391, 983, 242]. air [1748]. AKNS [947, 1543, 647, 2121]. Aktosun [1256]. Alfaro [700]. Algebra [1236, 1336, 850, 56, 482, 1469, 427, 1678, 1949, 958, 77, 728, 1552, 998, 464, 1508, 1979, 44, 2154, 557, 1299, 501, 2099, 2157, 1097, 118, 1648, 1176, 321, 915, 1441, 1682, 468, 420, 1954, 504, 505, 783, 1211, 2021, 2091, 2270, 1685, 370, 50, 1698, 1859, 609, 1848, 254, 2062, 2061, 922, 130, 1967, 923, 509, 1555, 1744, 660, 326, 2057, 2134, 490, 2035, 162, 1030, 632, 812, 1103, 1518, 164, 895, 84, 997, 534, 1442]. algebra-subalgebra [812]. Algebraic [1152, 1335, 959, 1720, 1719, 1398, 89, 730, 1553, 1737, 352, 1683, 2158, 1972, 402, 1495, 1774, 1422, 1407, 373, 248, 128, 364, 522, 1477, 911, 858, 403, 1560, 432, 235, 1694, 367, 486, 1845, 1211, 2233, 1544, 1686, 147, 924, 1293, 1216, 565, 1475, 2211, 1836, 164, 1297, 655]. Algebraically [2207, 1871]. algebras [275, 1523, 778, 1763, 418, 873, 1616, 1124, 1617, 1451, 1095, 2217, 1208, 875, 688, 800, 2098, 731, 559, 190, 502, 1680, 1618, 769, 1015, 1329, 1780, 1588, 1510, 1781, 2254, 1845, 796, 1482, 96, 209, 1377, 1920, 599, 280, 684, 2132, 1686, 1002, 282, 1116, 1514, 1590, 202, 664, 82, 964, 1776, 1725, 940, 941, 1348, 1687, 1849, 1180, 810, 1334, 1445, 541, 1368, 163, 1204, 662, 925, 966, 663, 329, 1146, 1863, 773]. algorithm [207, 147, 1216, 2156, 1127]. aliasing [1403]. allowed [333]. Almost [2044, 1305]. Almost-complex [2044]. almost-product [2044]. alone [1992]. Alternative [1955, 1536]. amended [1814]. Amir [567]. among [424]. Ampère [2078]. Amplitude [1658, 2012, 355, 1020, 1158]. amplitudes [90, 2145]. analog [728, 912, 894]. analogous [24]. analogs [1224]. analyses [1070]. analysis [970, 1926, 56, 2043, 1661, 2163, 1928, 1660, 1632, 2184, 752, 2140, 307, 1443, 104, 2005, 686, 1705, 204, 615, 1237, 1386, 1516, 2059, 1623, 813, 1641, 2060, 295]. Analytic [1593, 1665, 1666, 558, 1066, 2096, 172, 830, 1530, 1003, 2038]. Analytic-bilinear [1665, 1666]. Analytical [1889, 2224, 39, 1099, 631, 1914, 779, 1195, 2005]. analyticity [1520]. and/or [3, 736]. Andrews [2071]. anew [1177]. angle [30, 1645, 1026, 1356]. Angular [524, 396, 132, 243, 2088, 298, 1971, 1882, 1741]. angular-momentum [1741]. anharmonic [1258, 1297, 1142, 2140, 641, 1218, 1486, 2266]. Anisotropic [1976, 249, 1907, 39, 1205, 206, 207, 1466, 316, 1752]. annihilation [134, 1587, 1455]. Announcement [1183, 1219, 1107, 1148, 1257, 1295, 1816]. Anomalies [2096, 486, 1683, 1751, 997]. Anomalous [1470, 789, 1064, 581]. anomalously [464, 1048]. anomaly [366, 1438, 1223, 1652, 1963, 546, 997]. anomaly-free [997]. Anosov [1656]. ansatz [1859, 1024, 2200]. ansätze [839]. Anti [1589, 2045, 892, 261, 2217]. anti- [2045]. anti-BRST [2217]. anti-de [892]. Anti-isospectral [1589]. anti-self-dual [261]. antibracket [838]. anticommutation [635]. anticommuting [182]. anticontinuous [2066]. antidot [833]. antiferromagnet [1664]. antiferromagnetic [714, 2155]. antifield [1878]. Antisymmetric [702, 1414]. antiunitary [2192]. any [1232, 22, 217]. anyon [1307]. anyons [1934]. apparent [1028]. Application [1695, 2261, 1884, 1573, 268, 544, 2022, 203, 293, 98, 391, 829, 1216, 746, 1829, 1532, 149, 1845, 1850, 2058]. Applications [1763, 1322, 673, 1948, 822, 1766, 2262, 1768, 916, 1635, 587, 1277, 1179, 1307, 491]. applied [1126, 1080]. approach [1926, 1108, 1479, 128, 522, 778, 737, 2067, 233, 612, 116, 117, 1665, 1666, 870, 828, 1746, 558, 1595, 288, 574, 172, 1049, 2245, 486, 1352, 830, 1739, 2073, 297, 1323, 1761, 2103, 1943, 2139, 1841, 1427, 937, 2159, 1784, 238, 743, 2247, 1381, 1529, 1383, 1495, 1195, 1119, 98, 1728, 391, 255, 256, 845, 1486, 2266, 326, 15, 376, 2035, 1102, 813, 149, 723, 1310, 394, 212]. approaches [2075]. approximants [1692, 1904]. approximate [1233]. approximation [750, 1054, 1172, 451, 1595, 236, 334, 1073, 1018, 1604, 641, 1218, 393]. arbitrarily [2000]. arbitrary [593, 16, 814, 714, 1595, 1189, 1222, 639, 640, 1664, 2151, 727, 1269]. arbitrary-range [714, 1664]. area [1802]. arising [1690, 1363, 1778, 915, 2131, 2189, 588, 1884, 813]. Arithmetic [1904, 1488]. Arnold [1746, 1603]. array [833]. arrays [631, 832]. arrow [187]. Ashtekar [1322, 2092]. aspect [555]. aspects [1523, 31, 1855, 1223, 445]. associated [1258, 1625, 32, 2047, 209, 919, 2158, 1250, 1275, 1156, 1717, 1069, 649, 2041]. associativity [889]. Astronomy [493, 930, 1369, 1814]. asymmetric [1548]. asymmetries [1899]. asymmetry [508]. Asymptotic [34, 1287, 504, 908, 64, 1840, 827, 428, 1194, 1071, 483, 1691, 350, 1353, 1619, 718, 1800, 298, 1482, 1098, 2078, 646, 1439]. Asymptotically [1945, 624, 1758, 1872, 1879, 1804, 2225, 581, 624]. Asymptotics [2065, 390, 2146, 882, 1436, 2263, 1561, 706, 141]. atom [103, 1854, 172, 1385]. atomic [1363]. atoms [2068]. attractor [1913, 2123, 1133]. Attractors [1085, 1975]. attributes [1835]. Augmented [1412]. Author [496, 933, 1372]. auto [1317, 472]. auto-Bácklund [1317, 472]. automata [121, 950]. Automorphisms [1408, 1678, 1841, 939, 1849]. average [1427]. averages [1243]. averaging [1531]. Axially [1319, 798, 750, 1573, 1160, 1094]. axion [362, 2137]. axisymmetric [1135]. azimuthal [371].

B [935, 1289]. Bach [1241, 272]. background [726, 725, 2075, 316, 1397, 76, 642, 764, 1423, 2130]. backgrounds [479]. Bäcklund [2006, 1951, 1670, 2221, 1726, 1671, 1239, 472, 723, 1546, 1317]. Backscattering [2117]. backward [1214]. backward-forward [1214]. Bailey [2071]. ball [556, 2235]. band [576, 2009, 1086]. bands [602]. bang [644]. Barbero [2092]. Barenblatt [1470]. Bargmann [333, 338, 1517, 1924, 1925]. Bargmann- [338]. barrier [2033]. barriers [886]. Barut [1221]. based [1412, 136, 754, 2019, 2053, 1630, 2060]. bases [1640, 655, 959, 1957, 2264]. basic [1861, 1516, 2259]. basis [129, 1585, 1096, 1175, 333, 1210, 1511]. Batalin [2179, 838, 1525, 1932, 742, 1155]. Baxter [425]. Bayesian [1070]. be [1304, 615]. Becchi [1732, 840, 841]. before [1575]. behavior [2004, 908, 1840, 217, 2079, 861, 646, 1278]. behaviour [891]. Bel [1500]. Belavin [1325, 1668]. Belavkin [198]. Bell [24]. Beltrami [17, 2162]. Bender [285]. Benjamin [1277, 1799, 406]. Berezin [1961]. Berry [453, 1013, 538, 201]. Bertrand [1700]. Besieris [1147]. Bessel [631]. beta [1694]. Bethe [2200, 1859, 1024]. between [838, 2082, 1542, 913, 1618, 1047, 505, 1365, 630, 2075, 2257, 672, 1788, 925]. beyond [824, 1068]. Bhabha [2077]. bi [1572, 1274, 945, 1725, 1356, 2161, 2124, 348, 647]. bi-crossproduct [1725]. bi-Hamiltonian [1572, 1274, 1356, 2161, 348, 647]. bi-isotropic [945]. bialgebra [418]. bialgebras [58, 1684, 1264]. Bianchi [2084, 2085, 1841, 1093, 526, 1285]. bianisotropic [1351]. Bicrossproduct [696]. Biedenharn [1405]. Bifurcation [1278, 1647]. Bifurcations [1799]. big [644, 367]. bilateral [8, 374]. bilinear [1665, 1666, 502, 2103, 1943]. billiard [1128, 1390, 1747, 833, 294]. billiards [821, 105, 1386, 1825, 861, 1613]. bilocal [1738]. bimodules [801]. Binary [1671, 759, 969, 2000, 1431, 1357, 2249]. bind [2138]. binomial [1190]. biosuspensions [1466]. Birkhoff [1616, 670]. Birman [1929]. BKP [1912, 356]. Black [1661, 1839, 481, 212, 2164, 593, 357, 594, 39, 1205, 928, 552, 956, 1802, 910, 1169, 1675, 650, 596, 1805]. Blaszak [1670, 2221]. Bleuler [1786]. Bloch [1524, 1520, 1997]. block [1543, 2119]. blocks [264, 499]. blood [1466]. BMS [1848]. bodies [1971, 1909, 462]. body [589, 3, 736, 613, 583, 1702, 2106, 1820, 39, 1205, 782, 135, 1668, 21, 1729, 23, 1064, 1740, 1990, 74, 1380, 203, 1491, 2000, 1084, 1446, 1312, 1730, 628, 848, 2156, 407]. Bogoliubov [783]. Bogoyavlensky [758, 1204]. Bohm [1898, 1375, 26, 391, 983, 242]. Bokhari [567]. Bolkmann [1493]. Boltzmann [678, 831, 786, 1541, 898]. Bombieri [1646]. Bondi [1917]. Bonnet [221]. Borcherds [2020]. Borel [1182, 1650]. Born [1817, 1018]. Bose [1927, 585, 2191, 1900]. Boson [96, 487, 848, 655, 1509, 691, 516, 1345, 1253]. Boson-fermion [487, 848]. bosonic [131, 499, 1189, 1694]. bosonization [1648, 1191]. bosons [733]. Bott [1182]. Bound [20, 972, 820, 1642, 331, 132, 571, 1259, 1929, 2272, 97, 1029, 143]. boundaries [2105]. Boundary [468, 1214, 1031, 1041, 68, 460, 1436, 2263, 1793, 350, 1425, 1201, 1448, 820, 876, 1401, 1153, 1668, 2206, 1713, 948, 1512, 199, 2174, 1027, 246, 1306, 1620, 2038, 1078, 228, 512, 1235]. Bounded [1927, 2150, 146, 1991]. Bounds [1260, 687, 884, 90, 657, 1803, 338]. Boussinesq [2009, 469]. brachistochrone [1352]. brachistochrones [1767]. bracket [838, 1315, 1210, 746, 1307, 1174]. brackets [1399, 1074, 1365, 1935, 1100, 541, 672]. braid [223, 202, 664]. Braided [922, 2215, 1894, 1539, 1329, 2100, 159, 323, 1252]. braids [2253]. Branching [1773, 2183, 812, 925]. brane [1873]. branes [1550, 2250, 2091, 2270, 1016]. Brans [409]. Breakdown [1437]. breaking [450, 1340, 603, 1580, 1050, 1489]. Bremmer [711]. bridge [1847]. Brillouin [2261, 1595, 1112, 334, 1695, 294]. broadening [1152, 1335]. broken [667, 193, 1049]. Brownian [377, 2004, 1847, 1540, 1466, 716, 1231]. Brownian-like [1466]. BRST [1452, 2217, 1110, 1224, 1211, 2139, 743, 843, 239]. Buckingham [1306]. Building [264, 757]. bulk [1715]. bulk-viscous [1715]. bundle [1311, 2063, 322, 1683, 580, 2064]. bundles [1188, 916, 227, 1217, 533]. Burgers [2009, 158, 517]. Buslaev [1905].

C [1256]. C. [2159]. Cabled [740]. Calabi [531, 1406]. calculating [522, 1216]. calculation [128, 364, 1477, 21, 606]. calculations [1585, 1466]. calculi [802, 877, 2220, 2055, 773]. Calculus [1764, 1631, 1586, 2216, 119, 281, 1156, 1630, 963, 924, 1253, 1293]. Calogero [1451, 132, 571, 1411, 2240, 1669, 1544, 1598, 1426, 684, 338, 259, 1833, 2204, 110]. Calogero-type [338]. Camassa [1868]. Can [1304, 26, 2138, 720, 615]. cannot [635, 481]. Canonical [232, 2271, 777, 103, 667, 1614, 2046, 1735, 974, 2088, 1118, 1690, 872, 2213, 1852, 185, 635, 1151, 1795, 1173, 1464, 1999, 505, 67, 540, 1848, 1157, 459, 2058, 1603, 847, 1583, 1584, 1534, 1523]. Canonically [1065]. CAR [1482]. Carlo [439]. carried [1101]. carrier [688]. carrying [501]. cascades [1636]. case [872, 840, 841, 1189, 1471, 1112, 833, 1780, 2090, 1824, 1922, 76, 1000]. cases [1298]. Casimir [2105, 380, 738, 1279, 1109, 1890, 1271, 913, 656, 1247, 1331, 1332, 2235, 122, 1724, 987]. Casimirs [793]. Casorati [305]. cat [461]. Catalan [385, 424]. categories [1618, 1844, 1289]. category [432]. Cauchy [2043, 1757, 2118, 2167, 1880, 1082, 2010]. causal [1473, 1715, 686, 2252, 2265, 2035, 1268]. causality [890, 1147, 398]. cause [1793]. caustics [315]. cavity [1271]. Cayley [1390, 1747]. CCR [1482, 783, 1442]. CCR-algebra [1442, 783]. cellular [121, 950]. censorship [357]. Center [2020, 245]. Central [2071, 59, 559, 803, 1992, 808, 2199]. certain [1985, 102, 1275, 282, 519, 1716, 2026]. Cesàro [489]. chain [380, 737, 1572, 1719, 1398, 1892, 1618, 385, 1570, 2255, 1859, 658, 2077, 1391]. chains [497, 2202, 1985, 385, 506, 1698, 1348, 110]. Chalker [1959, 1051]. Chamseddine [1262]. change [264, 864, 244]. changes [863]. channel [1562]. Chaos [1759, 1022, 1797, 579, 2122, 1050, 15, 376, 2050, 1613]. Chaotic [1826, 857, 1793, 1049, 879, 1855]. Character [694]. Characteristic [1883, 442, 360, 359, 1885, 1887, 1886, 1247, 1428, 82]. characterization [1362, 2107, 1683, 1422]. characterized [1992]. characterizes [370]. characters [1138, 369, 1409]. Charge [23, 1819, 2071]. Charged [382, 1562, 833, 504, 607, 74, 1914, 708, 2130]. charges [1272, 385, 2002]. chemical [104]. chemistry [601]. Chern [168, 2072, 1382, 1383, 221, 1819, 2180, 1647, 503, 536, 1343, 369, 1697, 788, 747]. chiral [1088, 915, 1223, 1652, 719, 1535, 292, 1980, 1120, 162, 163]. chirality [2133]. chosen [217]. chromo [1489]. chronological [1713]. circle [1031, 1552, 1754]. circuits [1810]. circular [1386, 1825]. Civita [2090, 266]. cKP [1317]. Clarkson [158]. class [151, 450, 1232, 907, 2118, 682, 1314, 34, 1573, 303, 206, 102, 384, 614, 2149, 908, 1653, 1381, 631, 1871, 1620, 2016, 1102, 1476, 987]. classes [2084, 2085, 1208, 1255, 663, 646]. Classical [380, 58, 890, 209, 1971, 1484, 582, 2113, 344, 2147, 482, 1185, 583, 44, 896, 1464, 7, 229, 1189, 1112, 2009, 2120, 715, 1999, 298, 1729, 1550, 1083, 2220, 1055, 1395, 1117, 1344, 82, 2055, 1903, 981, 1492, 716, 83, 1312, 1730, 2002, 183, 1196, 673, 1077, 812, 925, 54, 2153]. Classification [493, 930, 1369, 1814, 2126, 1167, 47, 1915, 918, 1684, 51, 1433, 2055, 219, 162, 163, 425, 2063, 1678, 1951, 411, 1554, 62, 1359, 1165, 580, 2064, 1726, 2001, 697, 1591]. Clebsch [912, 155, 1720, 49, 924, 2146, 1216, 483, 1439]. Clifford [873, 502, 1680, 2057, 987]. Clock [432, 1549]. clones [2032]. cloning [2032]. Closed [2053, 1583, 146, 1698, 216, 1859]. Closure [1691]. cluster [333]. clusters [1789]. cmKP [1317]. coaction [800]. coadjoint [1785]. coarsely [1569]. cocycle [516]. cocycles [188, 420, 1290, 1725]. Coddington [1959, 1051]. coding [1378]. coefficient [253]. coefficients [2065, 363, 556, 912, 155, 1720, 1543, 48, 1036, 49, 81, 1099, 1728, 924, 2146, 1216, 1774, 1103, 483, 1439]. Coexistent [1784]. coherence [1056]. Coherent [1057, 781, 537, 236, 1266, 232, 2271, 131, 452, 1986, 1190, 888, 1221, 1112, 2107, 1455, 578, 337, 137, 819, 1156, 744, 1862, 1969, 953, 707, 160, 660, 161, 2059, 1624, 343, 1350, 1791, 2214]. coherent-state [337]. Cohn [571]. Cohomology [916, 1688, 17, 840, 841, 2217, 1364, 1347, 1565, 895]. coincidences [564]. Coisotropic [395]. Coleman [943]. collapse [39, 1205, 213, 2049, 270, 2194]. Collapsing [594]. Collective [2229, 966]. Colliding [1760, 152]. collineations [11, 568, 567, 724, 1167, 411]. Collins [86, 85]. collision [2005, 2000, 898]. collisions [1363, 610]. Colombeau [1030]. color [1616, 958, 1544]. Colored [1428, 1334, 1445, 425]. combinatorial [1310]. combined [1530]. Comment [701, 2063, 567, 2176, 2101, 1105, 700, 2267, 1729, 1008, 2268, 2269, 837, 2102, 2062, 2061, 1982, 1147, 2177, 1730, 1256, 1106, 2273]. Comments [927, 735, 926, 457, 86]. commutation [1690, 185, 777, 1523, 1140, 703, 1741, 1534]. commutative [803, 237]. Commutator [606, 320]. commutators [970]. Commuting [2131, 1927]. comoving [339, 815]. Compact [1006, 2217, 688, 800, 79, 920, 1724, 1396, 1516, 1102, 957, 1361, 967, 1391]. compactifications [531]. compared [124]. comparison [97]. compatible [468]. Complementary [1727, 1728]. Complete [574, 1997, 1296, 840, 333, 2101, 1762, 2102, 615, 942]. completely [2039, 682, 683, 1314, 258]. Completeness [341, 1949, 2087, 48, 1071]. completion [794]. Complex [1734, 1601, 2044, 1173, 1950, 1574, 2176, 1097, 1001, 2227, 2228, 1082, 917, 1846, 1604, 1628, 1870, 2177, 1037, 673, 2040, 473]. complexes [1193]. complexification [1702]. complexified [1072, 1418]. component [1992, 1655, 2248, 648, 1546]. components [371]. composed [1520, 1997]. composite [207, 1155, 463]. compressible [1892]. Computation [657, 176, 1262]. computational [2150, 677]. concentrated [1869]. concentric [1113]. concept [1916]. concerning [855, 760]. condensate [747]. condensation [854, 2191, 1900]. condensed [1041, 1044, 1054, 1056, 1039, 1045, 1049, 1047, 1040, 1043, 1055, 1052, 1048, 1042, 1053, 1050, 1046, 1051]. Condition [1340, 1259, 2272, 1793, 1747, 997]. Conditional [709]. conditionally [254, 2062, 2061]. Conditions [1390, 761, 1366, 1396, 1031, 1041, 795, 350, 1201, 1448, 1358, 876, 1401, 1153, 1668, 157, 842, 1955, 1512, 1811, 509, 1027, 2016, 1503, 228, 2162]. conductance [817]. conducting [401]. conduction [886, 1575, 1, 837, 2]. conductivity [1228, 145]. conductors [817, 828, 1048]. cone [1088, 382, 1697, 272]. Cones [2252, 2265, 1978]. configuration [381, 739, 1644, 1188, 510, 1021]. configurations [365, 714, 213]. confinement [2142, 627]. confocal [2082]. Conformal [2079, 597, 566, 555, 1678, 1279, 499, 366, 2014, 1738, 1510, 1915, 94, 2048, 209, 922, 1178, 1179, 325, 327, 1387, 481]. conformally [1467, 2168, 1803, 1805]. conformally-decomposed [2168]. congruence [46]. Congruences [2058]. conharmonic [1577]. conical [907, 962, 809]. conjecture [1374, 731, 1213, 2051, 1646]. connected [185, 1515]. Connection [2092, 2014, 1650, 540, 626, 472]. connection-formula [1650]. Connections [2029, 1548, 444, 1326, 801, 802, 1659, 461, 916, 419, 94, 414, 533]. Connes [1012, 435, 1262]. consequence [502]. consequences [1560]. Conservation [2100, 652, 1159, 111]. conservations [2197]. Conserved [1770, 850, 2137, 7, 229, 1189, 385, 94, 1938]. considered [1177]. consistent [1483, 845, 1285]. constant [265, 2004, 2126, 2127, 2039, 765, 999, 1819, 1353, 1160, 1649, 607, 639, 581, 1657]. constants [1203, 1079, 304, 1833]. Constrained [1025, 1564, 1114, 248, 2226, 1864, 1200, 209, 1609, 210, 952, 1276, 904, 2011, 955, 2249]. constraint [1872, 1263, 1064, 991, 491, 1127]. constraints [151, 1012, 1574, 1614, 1908, 614, 1330, 1270, 1354, 1824, 2248, 164, 1432, 1127]. construct [403]. constructing [490]. Construction [383, 2184, 942, 1877, 911, 1855, 2071, 1222, 290, 1274, 2246, 351, 1290, 1331, 1332, 2128, 1704, 1861, 423, 1333, 753, 1182, 1709]. constructions [960, 1356, 846]. Constructive [757, 1857, 1701]. contact [1388, 1415, 66]. containing [1035]. content [2077]. context [2194]. continuation [874, 558]. continued [768, 1228, 2049]. Continuous [1706, 1626, 772, 387, 1483, 1233, 1733, 779, 1272, 2070, 79, 332, 367, 279, 770, 251, 715, 834, 638, 25, 1069, 1386, 2198]. continuous-time [2198]. continuously [198]. Continuum [1172, 1959, 1051]. contraction [118, 707, 127]. Contractions [1920, 329, 58, 418, 1736, 327, 966]. contribution [63]. Contributors [1622]. control [944]. convection [1081]. convector [415]. conventional [642]. convergence [1692, 489]. converging [911]. conversion [1321]. convexity [1784]. convolution [378]. coordinate [2101, 1963, 198, 1762, 2102]. coordinates [105, 1573, 1467, 1696, 96, 695, 518, 784, 1243, 339, 815, 2238, 1106]. core [1889, 2224]. coregular [1516]. Coriolis [1446]. Corrections [1690]. correlated [1663]. Correlation [666, 30, 1464, 505, 24, 161]. correlations [824, 851, 1570, 1047]. correspondence [487, 981, 2238]. coset [98]. Cosmic [409, 1320, 271, 313]. Cosmological [1240, 2250, 1137, 553, 265, 2126, 2127, 38, 217, 1715, 686, 316]. cosmologies [2084, 2085, 2163, 1280, 358, 908, 1718]. cosmology [1759, 412, 1550, 1841, 113]. cosymplectic [1424]. cotangent [1311]. Coulomb [2224, 1220, 971, 103, 1889, 739, 1735, 1561, 2227, 2228, 604, 668, 64, 1965, 640, 1742, 1898, 1967, 1530, 1491, 1902, 242]. Coulombic [1692]. counterparts [161]. counterterms [702]. counting [819, 314]. Couple [1543]. Coupled [1316, 331, 2068, 1081, 1562, 820, 503, 1129, 1497, 291, 1165, 891, 647, 1380, 2121, 1918, 1743, 893, 1294, 980, 1936, 340, 1672, 2162]. Coupling [1649, 1103, 1298, 1201, 1448, 21, 1896, 379, 508, 711]. covariance [1300, 763]. Covariant [128, 364, 1477, 618, 2143, 1676, 1309, 1060, 1587, 1856, 2230, 1588, 48, 1155, 1380, 372, 447]. covector [415]. covering [343]. coverings [46]. Coxeter [1289]. creation [134, 1587]. creation-annihilation [134]. Cremona [1545]. Criteria [854]. criterion [1705]. critical [585, 1427, 1192, 2123]. criticality [1754]. cross [2151]. crossed [2031]. crossing [267, 1428]. crossings [2187]. crossover [1047]. crossproduct [1725]. Crum [2022]. crystals [830]. cubic [2251]. Cummings [451, 196]. curl [1474]. Current [494, 931, 1370, 1311, 2137, 559, 819, 162, 163]. currents [1, 837, 2, 1748]. Curvature [1548, 724, 2016, 1167, 2013, 2039, 840, 841, 767, 993, 1803, 607, 1972, 623, 1217, 1996]. Curvature-free [1548]. curvatures [765, 999, 961]. curve [1325, 107]. curved [1606, 410, 92, 751, 782, 2138, 1988, 1549, 726, 725, 2141, 2075, 1910]. curves [105, 235, 1694, 1611, 1836]. curvilinear [1573, 784, 1106]. cusp [1955]. Cuspidal [46]. cusps [687]. Cusson [839]. cutoff [1896]. cutoffs [1937]. cuts [272]. cycle [748]. cyclically [1954]. cylinder [2235, 2151, 240]. cylinders [1582]. cylindrical [1011, 2087, 1829]. cylindrically [762].

D [1105, 450, 439, 889, 625, 746]. d’Alembert [491]. Damazure [1471]. Damped [278, 1022, 1913, 2123, 1133]. Darboux [2006, 1950, 403, 1911, 1330, 1769, 354, 1430, 1671, 1239, 1431, 1357]. Darboux/Bäcklund [1239]. Darmois [868]. data [874, 558, 1883, 246]. Davey [1554, 308, 2083, 1085]. Debye [634, 1227, 247]. Decay [851, 1464, 576, 638, 1281]. decoherence [1308, 1745, 341, 985, 388]. decomposed [2168]. Decomposition [1722, 1647, 1316, 1687, 1849]. decompositions [2047, 1229, 920, 458]. Deconstructing [1860]. decreasing [1450]. deep [718]. defects [536, 1392, 215, 216, 962]. defined [1569, 1130]. definite [1518]. definite-parity [1518]. definition [437, 245]. deflection [477]. deformable [586, 587]. Deformation [287, 1601, 911, 729, 2017, 2052, 604, 560, 139, 632, 84, 342]. Deformations [601, 1878, 1786, 58, 418, 1679, 507, 259, 2204, 328, 2040]. deformed [839, 958, 1060, 896, 1399, 770, 1954, 369, 608, 696, 201, 698, 2134]. Deforming [1587]. Degeneracy [2110, 34, 1349]. Degenerate [2161, 1522, 682, 1673, 1746, 947, 1892, 890, 1685, 114, 41, 1709, 753]. degenerated [1572]. degree [856, 139, 2054]. degrees [444, 1832, 1188]. delays [1049]. delta [886, 391]. delta-barriers [886]. delta-function [391]. delta-wells [886]. deltas [1948]. Demazure [1409]. denominator [2213]. densities [1496]. Density [842, 24, 1052, 1593, 1338, 601, 1819, 1647, 67, 1194, 76]. dependence [2140, 456, 898]. dependent [2004, 855, 856, 1527, 2030, 1908, 1142, 1637, 544, 543, 108, 1381, 1602, 2034, 397, 941, 1833, 1537, 1531, 394, 940]. Derivation [652, 2194, 2069, 1425, 1389, 1652, 97, 1554]. derivations [275, 541, 662]. Derivative [1651, 2185, 308, 503, 743, 1006]. derivatives [30, 153, 375, 1456]. derived [615, 1005]. Deriving [1463, 748, 1949]. describing [718, 198]. description [2226, 1737, 1389, 1261, 208]. design [816]. designer [173, 1721]. Desingularization [2050]. Details [1262]. Determinant [806, 1753, 305, 82]. Determinants [876, 123, 330, 619, 1327, 14, 87, 702, 2219, 979, 227, 1217, 1239]. Determination [145, 1902, 1297, 1367]. Develop [837]. Development [1321]. developments [1179]. Deviation [2127, 1916, 1504]. deviations [344]. devices [203]. devil’s [1664]. devil’s-staircase [1664]. DeWitt [760, 1655]. Dia [1014]. Dia- [1014]. diagonal [1053]. diagonalization [170, 670]. diagram [1664]. Diagrammatic [822]. Diagrams [1492, 2182, 1687, 846]. diatomic [1018]. Dicke [400, 409]. Dielectric [515, 1271, 2151]. Diffeomorphism [17, 1410, 444, 1449, 1614, 1874, 1868, 1242, 1310]. diffeomorphism-invariant [1449]. diffeomorphisms [102]. Difference [110, 741, 485, 224, 1402, 2131, 2103, 1943, 2005, 1003, 1350, 1701, 2205]. different [1576]. differentiability [1757, 2167]. differentiable [2252, 2265, 85]. Differential [1766, 2262, 1586, 802, 1097, 877, 1936, 1974, 534, 652, 1558, 1551, 2172, 318, 1552, 1436, 2263, 2216, 1089, 1110, 1951, 277, 1607, 1940, 1229, 1553, 1315, 621, 2254, 120, 2036, 2103, 1943, 280, 2220, 963, 1726, 2055, 2025, 1293, 101, 813, 1146, 1863, 1104, 1813, 773]. differential-difference [2103, 1943]. differential-geometric [2036]. diffraction [1660, 489]. Diffuse [712]. Diffusion [1539, 1105, 465, 1606, 451, 917, 549, 951, 994, 1634, 402, 2059, 813, 517]. diffusions [1643, 550, 339, 815]. Diffusive [2148]. diffusivity [1681]. Dikii [1315]. dilatation [44]. dilation [1626]. dilaton [2137, 362, 1094]. dilute [2073, 1143]. Dimension [2123, 55, 1490, 1232, 971, 1470, 1045, 1416, 891, 2220, 1846, 1193, 402, 943, 141, 1913, 774, 1133]. dimensional [883, 1642, 1149, 555, 777, 1523, 151, 1311, 427, 37, 2068, 59, 680, 129, 276, 1571, 2164, 170, 556, 78, 653, 855, 1889, 2224, 1437, 765, 999, 366, 2045, 735, 3, 736, 2106, 874, 34, 2215, 1573, 858, 1574, 1228, 38, 501, 1425, 1561, 1341, 863, 234, 1607, 714, 1668, 1660, 404, 576, 172, 136, 1135, 1129, 47, 890, 1712, 195, 1915, 40, 1322, 297, 945, 695, 2021, 1377, 1920, 1684, 1292, 1858, 1964, 996, 2155, 1008, 1513, 761, 1136, 67, 1664, 1393, 1394, 1501, 422, 104, 200, 1898, 1708, 1980, 1283, 1723, 1499, 1345, 706, 1787, 1902, 391, 1348]. dimensional [1094, 788, 1444, 478, 257, 1237, 953, 2160, 2222, 1825, 260, 340, 1132, 1834, 219, 143, 1256, 429, 1255, 1914, 565, 1475, 2211, 597, 1851, 262, 1621, 2113, 1996, 27, 2041, 1534, 2199, 127, 2051, 709, 1007, 407, 512, 1432, 2003]. dimensionality [6, 166]. Dimensionally [2049]. dimensions [449, 1296, 1927, 43, 1854, 1831, 430, 1542, 69, 1060, 16, 814, 559, 1894, 573, 1911, 1464, 1987, 2027, 691, 914, 22, 1062, 528, 514, 598, 2185, 1260, 1191, 157, 1415, 545, 936, 1963, 2049, 719, 1535, 919, 1714, 640, 990, 292, 1611, 1742, 1122, 1197, 1202, 1798, 371, 1530, 1206, 1502, 511, 436, 394]. dimer [365]. dipolar [1909]. Dirac [2104, 377, 862, 1373, 1227, 88, 1011, 1673, 644, 876, 1401, 192, 2074, 2108, 64, 2110, 937, 1513, 50, 200, 1380, 338, 1967, 1404, 1195, 1333, 706, 581, 98, 2251, 1225, 316, 790, 142, 733, 1993, 2196, 2210, 1397, 2079, 1699, 1505, 1121, 242, 849, 1460, 554]. Dirac-like [849]. Direct [1669, 458, 558, 158, 255, 256, 2010, 149]. Directly [612]. Dirichlet [1796, 853, 1512, 1144]. disc [2113]. disclination [1808]. Disconnected [1215]. discontinuous [205, 710, 20, 1855]. Discrete [277, 1919, 1453, 638, 1771, 749, 1625, 1337, 1733, 2118, 681, 1493, 1985, 91, 107, 603, 1695, 2261, 692, 1998, 834, 305, 1989, 860, 1972, 389, 707, 1357, 1104]. discretization [1672]. discretizations [1806, 1238, 758]. discretized [436]. discussion [807]. disk [1756, 1401, 228]. disorder [829, 149, 551, 646]. disordered [1041, 824, 828, 1048]. dispersion [681, 1605]. dispersionless [262]. Dispersive [1023, 1820, 945, 1850, 953]. displaced [183]. Displacement [940, 941, 2067]. Displacement-operator [940, 941, 2067]. dissipation [108, 551]. Dissipative [1999, 408, 1273, 173, 1721, 405, 311, 1281, 178]. distances [1927]. Distortion [1430]. distortions [173, 1721]. Distributed [1250]. distribution [1777, 2255, 1114, 1143, 909, 1902, 342]. distributional [868, 1519]. distributions [1629, 1190, 544, 543, 807, 1532, 1476]. disturbed [677]. Divergences [1313, 60, 284, 2115, 2267]. Dixon [628]. domain [185, 409, 1351, 297, 1392]. domains [1450, 1991]. dominated [595]. Donaldson [1224]. Dorfman [1356]. dots [746]. double [2099, 91, 1135, 1685, 181, 27]. double-well [91, 27]. doublet [636, 1594]. doubling [1223]. Drach [1203]. dressing [1199, 1707, 1353, 993]. drift [1098]. Drinfeld [1667, 960, 1472]. dromion [1132, 1834]. droplets [586, 587]. DS [1353]. DS-II [1353]. dual [665, 58, 1249, 69, 741, 751, 769, 1343, 476, 1377, 1918, 1180, 623, 261, 1431]. Duality [1488, 1881, 599, 2055, 967, 2094]. Dually [607]. Duals [83, 1445]. Dubrovin [889]. due [883, 1088, 1, 837, 2, 687, 1434, 1256]. Duistermaat [535]. duration [2000]. dust [1756, 267, 1393, 1394, 1501]. dyadic [1636]. Dynamic [817, 149]. Dynamical [438, 1903, 1348, 1706, 1233, 1831, 1580, 2120, 1998, 1845, 1839, 1130, 146, 1083, 721, 902, 860, 217, 1429, 792, 1545, 879, 1408, 551, 646, 1755, 212, 1646]. dynamically [1765, 440]. Dynamics [1074, 397, 1231, 957, 850, 1478, 778, 1538, 2116, 737, 2066, 233, 467, 828, 277, 107, 1680, 1569, 442, 1885, 405, 548, 2229, 1465, 1564, 834, 1597, 516, 1442, 1264, 178, 1344, 1862, 355, 716, 1709, 183, 2079, 72, 1489, 2162, 407]. dynamo [1027]. Dyon [1891]. Dyson [1340, 988]. early [2163]. earth [1748]. easy [1680]. eccentric [1751]. edge [789]. Edmonds [1619]. Effect [1124, 285, 1644, 133, 1375, 1209, 26, 1345, 845, 1126, 1828]. Effective [1155, 128, 1228, 1651, 2185, 979, 2169, 463, 1516]. effectively [148]. Effects [243, 1579, 1088, 1817, 669, 298, 138, 1784]. Efimov [787]. Eichler [661]. eigenfunction [1287]. eigenfunctions [1297, 518, 1997]. eigenrays [4, 1775]. eigenstates [636, 1594, 1153, 1267]. Eigenvalue [222, 1143, 1991, 34, 656, 2187, 1512]. eigenvalues [1031, 156, 1260, 1376, 1331, 1332, 225, 1799]. eight [1164]. eikonal [1883, 491, 1879, 1843]. Einstein [1888, 1592, 862, 2044, 593, 1872, 2165, 2166, 1438, 2137, 38, 1240, 2047, 864, 361, 2168, 362, 796, 1092, 1393, 1394, 1501, 218, 2191, 1675, 1900, 1716, 1717, 1094, 1207, 142, 526, 274, 2210, 317, 596, 1360, 261, 1505, 799, 764, 480, 1324, 1805]. elastic [90, 2036, 990, 217, 462, 791]. electric [1108, 474, 1819, 1013, 2031, 819, 1810, 511]. electrical [789]. electrodynamic [1462, 1730]. electrodynamics [2147, 186, 2180, 2185, 1481, 298, 1482, 2231, 1729, 1312, 2002, 1053]. Electromagnetic [1748, 2068, 1321, 2150, 93, 63, 351, 945, 807, 1076, 1602, 1194, 1536, 677, 2151, 1830, 1161, 869, 764, 708, 1387]. Electron [819, 1817, 1778, 379, 1787, 746, 1914]. Electronic [1257, 1295, 1042, 1585]. electrons [886, 1055, 1052]. electrostatic [1658]. electrostatics [207]. electrovacuum [1758, 313]. electroweak [9, 230]. elegant [589]. elementary [26]. elements [1782, 694, 1036, 1965, 80, 122, 1583]. Elliott [1405]. ellipsoid [1128, 1747]. ellipsoidal [1390]. Elliptic [1311, 1717, 2204, 1648, 2240, 2131, 506, 1770, 1922, 2175]. elongated [1801]. embeddability [737]. embedded [886]. embedding [774]. embeddings [521]. Emch [927]. emergent [1690]. empirical [558]. employing [20]. enclosed [1930]. encountered [368, 373]. end [1757, 1753]. endomorphisms [599]. energies [1297, 1884]. Energy [824, 714, 1358, 1930, 27, 144, 1559, 364, 1477, 2148, 738, 1279, 624, 88, 2013, 1298, 1109, 1890, 1271, 1287, 1793, 90, 91, 1562, 384, 1629, 1570, 1857, 1597, 2037, 67, 2235, 312, 909, 910, 962, 1917, 456, 706, 1385, 991, 72, 1937, 393, 149, 987, 1828, 988]. energy-momentum [144]. Enlarged [1442]. Enneper [2039]. ensemble [850, 1795, 1953]. ensembles [1528, 67]. entries [827]. Entropic [2199]. entropies [1561, 617, 616, 345, 1267, 301]. Entropy [836, 1517, 301, 1538, 440, 1809, 67, 2201, 792, 1348, 1367, 1254, 15, 376, 2199, 1198, 1600]. entropy-minimum [1367]. envelope [303]. enveloping [280, 810, 1334, 1445, 1368, 490, 534, 773]. epsilon [159]. equation [619, 710, 1642, 2065, 2178, 167, 2147, 1108, 186, 1939, 68, 1325, 1362, 1227, 1297, 43, 88, 678, 513, 106, 1507, 1889, 2224, 2165, 2166, 33, 250, 874, 2082, 852, 741, 1470, 1793, 2007, 1340, 1951, 2176, 2119, 1313, 1527, 484, 1607, 1353, 1911, 2030, 404, 718, 1879, 1843, 1328, 889, 2008, 754, 1129, 1942, 1082, 1838, 252, 386, 1953, 695, 158, 2074, 2108, 948, 937, 305, 806, 306, 891, 1513, 198, 1868, 721, 308, 784, 1612, 254, 1354, 1708, 2062, 2061, 903, 1798, 2083, 786, 1177, 1909, 338, 1967, 1277, 371, 1811, 1973, 1495, 1867, 1541, 354]. equation [355, 1131, 1157, 1655, 2034, 109, 1726, 148, 1883, 1901, 255, 256, 309, 1799, 1237, 1870, 2177, 1850, 1306, 953, 2076, 406, 479, 2079, 1591, 425, 1121, 1431, 242, 1133, 1975, 1357, 1700, 2130, 1106, 723, 554, 491, 394, 709, 310, 988, 1086, 1432]. equations [883, 331, 652, 1159, 1558, 1162, 1232, 1524, 862, 1983, 2171, 2172, 59, 222, 2069, 318, 276, 2118, 1734, 2243, 302, 1011, 1872, 570, 735, 467, 1072, 2070, 1567, 1542, 69, 451, 1574, 1089, 303, 1125, 1951, 1658, 1240, 644, 1425, 2047, 1892, 1821, 1607, 1940, 864, 782, 1659, 70, 278, 102, 993, 485, 224, 1554, 192, 2168, 136, 1681, 1135, 2009, 1315, 13, 63, 621, 1803, 120, 1248, 1274, 2149, 468, 253, 1090, 145, 918, 1164, 1241, 1377, 196, 1064, 1008, 2022, 1275, 2100, 990, 1092, 1393, 1394, 1501, 1916, 600, 1611, 2078, 200, 1633]. equations [1366, 1634, 519, 1380, 952, 1276, 591, 1972, 903, 1202, 218, 1703, 272, 1169, 390, 2054, 1726, 1499, 1195, 745, 148, 1716, 1936, 2093, 2237, 1236, 255, 256, 2081, 584, 257, 316, 1850, 1207, 790, 2242, 1446, 274, 1003, 1256, 1913, 2123, 317, 1360, 1568, 2161, 73, 1699, 2104, 2059, 165, 799, 764, 1592, 517, 1672, 1126, 1463, 2040, 480, 1079, 649, 2041, 1324, 727, 1755, 2249, 1007, 1813, 1792, 2239, 1278, 2003]. equatorial [218]. equilibria [1749]. Equilibrium [1816, 1056, 1373]. equipartition [1287]. Equivalence [2084, 2085, 1618, 1291, 112, 1985, 1942, 1043, 868, 642, 592]. equivalent [1917, 1499]. equivalents [1948]. Ergodic [830, 939, 831]. Ernst [1711]. Errata [776]. Erratum [2271, 568, 1477, 1958, 2263, 2224, 2262, 736, 1335, 2272, 374, 814, 284, 1009, 1448, 87, 229, 928, 2027, 1521, 1885, 1887, 1888, 1886, 2261, 375, 2103, 2270, 1775, 230, 231, 664, 1731, 492, 815, 2225, 2064, 1776, 330, 1294, 2264, 2265, 1218, 2266, 376, 2104, 1592, 929, 1959]. escape [857]. Essential [581, 921, 1144, 1796]. Estabrook [404]. estimate [1254, 1828]. Estimates [2195, 88, 440, 899, 1082, 1416, 1850]. estimating [1692]. estimation [836, 2107]. Euclidean [378, 276, 321, 412, 919, 1858, 996, 200, 271, 1242, 811, 2153]. Euler [1162, 1806, 1437, 1567, 1803, 2001]. Evaluation [2182, 1519, 1884]. evanescent [1147, 398]. evaporation [854]. evaporation-condensation [854]. Even [2255, 32, 1664]. event [1696]. Evolution [265, 569, 1232, 1325, 2118, 1734, 187, 1339, 718, 1248, 2149, 2187, 1275, 1611, 254, 2062, 1972, 2061, 1202, 641, 1218, 1486, 2266, 1120, 1755]. Evolutional [1941]. evolutionary [1366, 1909]. evolutions [618]. Exact [2147, 2176, 1493, 1341, 21, 2074, 2108, 196, 1989, 1990, 80, 519, 525, 316, 1870, 2177, 679, 2210, 799, 1592, 242, 849, 2130, 988, 1235, 276, 303, 119, 1328, 1248, 158, 2158, 113, 97, 1381, 1157, 255, 256, 2002, 1285, 1591, 848, 1596]. Exactly [2109, 906, 498, 319, 1527, 756, 1111, 1729, 1589, 2236, 1530, 423, 1312, 1730, 610]. Exactness [334, 537, 888, 1112]. example [645, 1603, 675]. Examples [695, 2189, 1656, 763, 941]. exchange [1597, 506, 1116]. Excitations [1045, 1787, 1876]. excitonic [1193]. Existence [1819, 2206, 2189, 1675, 1515, 1259, 2272, 2150, 1767, 120, 252, 1281, 1366, 677, 1540]. exit [1718]. Exotic [953, 1205]. expanding [1467, 2207]. Expansion [1733, 18, 300, 189, 1651, 1035, 2185, 606, 64, 147, 1867, 2034, 76, 1337]. expansions [1341, 1692, 2022, 80, 428, 1602, 1485, 1269]. expectation [1436, 2263, 1520]. expected [1070]. explain [26]. Explicit [900, 2240, 1129, 1331, 1332, 2140]. exponent [2123]. Exponential [1597, 1280, 1281, 2053, 52, 53, 1006, 646]. exponentially [1653]. exponentials [768, 1722, 126]. exponents [1233]. expressible [553]. expressions [959, 1720, 30, 1774, 1974, 1407]. Extended [2135, 1135, 2132, 1901, 1878, 446, 1608, 1389, 209, 414, 609]. Extension [855, 1221, 1700, 2017, 2099, 158, 744, 2111, 808, 294]. extensions [2200, 57, 559, 1176, 1738, 308, 1498, 1966, 162, 163, 164]. extensive [2156]. exterior [119, 2101, 1762, 2102]. external [482, 1596, 2108, 761, 755, 1052, 2191]. Extremal [1512]. extremes [714]. extremum [2058]. Extrinsic [116].

F [700]. face [2131, 1598]. factor [692, 97]. Factorization [883, 1173, 1256, 2159, 1115, 2112, 1700]. factors [1407, 775]. Faddeev [1905, 326]. Failure [174]. Falling [461]. false [720]. Families [1755, 2124, 1769, 1703, 2059, 812]. family [38, 1248, 95, 246, 1180, 1656, 471, 1386, 226]. Farey [1904]. Fast [1468]. faster [2115, 2267]. Fateev [1024]. Fay [1973]. features [85]. Fermat [478]. Fermi [104]. Fermion [1223, 979, 511, 884, 2183, 487, 1657, 848, 988, 1693]. Fermionic [503, 1252, 300]. fermions [1449, 1010, 2154, 286]. Ferretti [22]. ferroelectric [1427]. ferromagnet [344, 16, 814]. ferromagnetic [1985]. few [746, 1307]. few-anyon [1307]. few-electron [746]. Feynman [2069, 934, 1645, 548, 1142, 1481, 1963, 1653, 238, 2190, 1822, 1704, 1519, 1000]. fiber [916]. fibrations [1717]. Fictitious [828]. field [1258, 2135, 344, 1756, 1479, 427, 1601, 2004, 431, 2068, 1279, 2028, 382, 2164, 170, 1734, 1410, 131, 499, 1691, 2137, 474, 1013, 1280, 972, 1400, 213, 1659, 7, 229, 1189, 1229, 358, 93, 2245, 174, 889, 1135, 1596, 1738, 13, 63, 992, 1931, 916, 726, 725, 94, 2073, 298, 1896, 516, 908, 2233, 938, 1017, 1305, 307, 74, 1091, 1264, 1394, 199, 755, 2221, 1474, 1966, 1382, 1992, 1731, 1419, 1602, 1019, 140, 2033, 1918, 1536, 1968, 2145, 581, 832, 584, 677, 1385, 1207, 1446, 1994, 447, 1285, 2122, 991, 1657, 429, 511, 1050, 443]. field [1532, 1422, 1030, 1476, 1533, 642, 247, 634, 764, 1121, 1126, 2080, 1423, 2130, 1324, 727, 407, 295]. fields [569, 750, 378, 167, 377, 1732, 1166, 1927, 1227, 1450, 1853, 870, 1673, 1072, 2046, 751, 1301, 1659, 702, 1014, 1694, 1988, 1651, 2018, 1738, 244, 1511, 290, 2184, 1946, 2031, 298, 2108, 176, 545, 2207, 2189, 1393, 1501, 1, 837, 2, 1155, 785, 208, 1055, 416, 1168, 1052, 1748, 1169, 1980, 978, 1021, 1743, 1147, 1910, 553, 2196, 797, 753, 392, 398, 674, 708, 228, 242, 1752, 2153, 296, 212]. filaments [2001]. Files [1257, 1295]. filled [1285]. filters [1403]. Finite [1056, 2009, 422, 1723, 1593, 1777, 1150, 365, 2154, 576, 485, 1129, 47, 949, 658, 787, 1066, 1705, 1384, 2222, 143, 2041, 1086, 110, 2205, 1330]. Finite-band [2009, 1086]. finite-difference [485]. Finite-dimensional [422, 1723, 1129, 47, 2041]. Finite-rank [1330]. finitely [1113]. Finsler [533]. First [1003, 2179, 856, 1644, 2018, 1465, 614, 902, 859, 2000, 1230]. first-order [1644, 1465]. Fisher [673]. five [154, 1571, 1712, 1206, 1502, 340, 219, 2051]. five-dimensional [1571, 1712, 340, 219, 2051]. fixation [2129]. Fixed [1434, 1562, 1464, 384]. fixing [1478]. flag [1005]. Flat [1320, 94, 521, 624, 1571, 1758, 1984, 1467, 214, 1879, 1843, 1803, 625, 695, 586, 1804, 2225, 2038, 760, 1899]. flat-space [1984]. Flavor [1993, 1489]. flight [1080]. Floquet [1928]. flow [1506, 676, 1160, 176, 1868, 995, 1491, 1654, 247, 634, 1547]. flows [899, 1999, 1305, 210, 1656, 36, 955, 2249]. Fluctuating [1283]. fluctuation [551]. fluctuation-dissipation [551]. Fluctuations [776, 5, 12, 464, 1427, 104]. Fluid [1577, 1976, 1571, 474, 39, 1205, 718, 412, 625, 1092, 1281, 1093, 1582, 1284, 1835, 1285]. Fluids [71, 311, 2089, 1575, 269, 1871, 954]. flux [1937]. flux-across-surfaces [1937]. Fock [1558, 1054, 333, 280, 1116, 65, 1252, 963]. Fock-space [280]. focus [1147, 398]. Focusing [306]. Fokker [34, 1495, 1867]. fold [564, 1239]. folding [1569]. folia [1442]. Foliation [1765]. force [2004, 1946, 208, 1474, 1462]. force-free [208, 1474]. forced [252]. forces [1491]. forcing [355, 1131]. Form [1089, 1732, 1567, 2030, 1414, 670, 847, 480, 472]. form-preserving [2030]. Formal [1634, 2247, 793]. formalism [1878, 1573, 2007, 688, 1412, 134, 7, 229, 1189, 235, 1694, 2183, 1932, 2218, 742, 1155, 353, 2111, 471, 274, 1120, 2244, 1996]. formalisms [800]. formation [594]. forms [17, 1249, 1796, 320, 875, 502, 1229, 2168, 2053, 1380, 1215, 372, 1750, 2058, 1533, 2026]. formula [1245, 285, 2180, 1208, 1650, 805, 656, 1670, 530, 591, 1867, 1628, 924, 1812, 2258, 1905, 698, 1772, 1583, 1753, 2205, 356]. formulas [1286, 2228, 135, 1619, 1930, 2019, 600, 1099, 1772, 1006, 55, 774, 1103]. Formulation [2084, 348, 1507, 435, 112, 1186, 754, 577, 1609, 1655, 1716, 733, 1699, 2104, 1077, 2003]. formulations [623]. forward [1214]. Foundations [2208, 238, 572]. Four [863, 2226, 40, 1322, 936, 1714, 1846, 200, 2221, 565, 1475, 2211, 1996, 436, 2153, 407]. Four-dimensional [863, 40, 1322, 200, 565, 1475, 2211, 1996, 407]. four-field [2221]. four-particle [2226]. four-vector [2153]. Fourier [2215, 804, 371, 124, 126]. fractal [1637, 1635, 658, 149]. fraction [1228]. fractional [1234, 1894, 1828, 572]. fractionally [1914]. fractions [8, 374]. Fradkin [1525, 2179, 787]. Fradkina [2179]. frame [1614, 2046, 1955, 1841, 1804, 2225]. Frames [2170, 1254, 1625, 914, 1549, 1379]. framework [1176, 2184, 352, 1120]. Free [1509, 1570, 2073, 674, 1548, 1318, 1010, 1185, 1271, 206, 585, 1780, 992, 290, 1442, 104, 282, 208, 1474, 2193, 339, 815, 1284, 541, 1226, 997, 1600, 1826]. freedom [444, 856, 1832, 1188, 139]. Frequency [1351, 2060, 1631, 1629, 1073, 1630, 1867, 1623, 1944]. Freud [301]. Friedmann [595]. Frobenius [1844, 2161]. fronts [75, 1581]. Fuchs [2237, 2025]. Fuchsian [2054]. Fuchsian-type [2054]. fulfill [635]. Full [1715]. function [1478, 1184, 2213, 777, 1777, 103, 20, 1246, 30, 1494, 189, 21, 703, 1442, 2139, 64, 990, 2235, 353, 371, 1630, 1867, 2075, 1020, 123, 330, 1346, 1902, 2256, 391, 2194, 2195, 1835, 662, 262]. Functional [1852, 14, 87, 177, 2219, 180, 169, 170, 171, 702, 172, 1383, 1829, 1006, 149]. functional-analytic [172]. functional-integral [172]. functionals [779, 1680, 1265, 1308, 1745, 341, 985, 388]. functions [750, 377, 821, 57, 1551, 1170, 1171, 1707, 1337, 1733, 601, 779, 1585, 1760, 19, 250, 1438, 45, 1209, 500, 134, 1288, 2227, 2228, 135, 823, 1464, 993, 1440, 62, 796, 606, 1635, 1403, 1114, 454, 2219, 1002, 271, 52, 53, 1704, 1973, 1654, 631, 1605, 294, 340, 2056, 746, 1922, 1923, 2244, 2197, 161, 698, 36, 1937, 1161, 1103, 1689, 241, 2060, 1638, 1520, 1997]. Fundamental [318, 321, 370, 50, 1725]. Further [140, 1179, 2038]. fusion [1409]. future [1713]. Fuzzy [434].

G [927, 1105]. gain [1794, 1958]. gains [1070]. Galilean [2214, 78, 972, 234]. Galilean-invariant [972]. Galilei [2214, 528]. Gambier [1951, 1726]. gamma [1694, 50, 52, 53, 1225]. Gamow [1522, 1557, 187, 1338, 779]. gaps [1632, 1376, 1512, 706]. gas [1220, 2243, 467, 585, 516, 104, 831, 65, 1787]. gateway [438]. Gauge [1478, 185, 151, 1821, 928, 552, 956, 1932, 1263, 210, 1865, 1503, 392, 2129, 1863, 2114, 2135, 1852, 449, 169, 128, 1732, 170, 593, 795, 1452, 1200, 667, 1807, 1579, 1647, 603, 7, 229, 1189, 1414, 1343, 486, 1856, 2230, 1942, 1874, 1652, 1683, 2189, 9, 230, 1697, 238, 1155, 743, 1698, 239, 2145, 1021, 293, 1743, 893, 1119, 788, 733, 1566, 597, 642, 592, 929, 611, 472, 895, 1310, 512, 1420]. Gauge-invariant [1865]. gauges [974]. Gauging [389, 1387]. Gauss [221, 2180, 342]. Gaussian [1643, 1663, 1250, 2033, 1520]. Gel′fand [1096, 1175]. Gel′fand- [1096, 1175]. Gel’fand [1315, 2119, 1356]. Gel’fand-Dorfman [1356]. General [2171, 656, 505, 1094, 1756, 1325, 1450, 1873, 852, 1089, 1614, 2046, 1647, 1694, 2088, 1802, 1875, 2149, 752, 867, 2139, 1359, 379, 1155, 75, 1380, 1581, 844, 416, 312, 1917, 868, 1282, 1027, 1284, 478, 1884, 2050, 1624, 2060, 987, 491, 512, 1468, 1867]. Generalization [571, 1929, 1441, 943, 711, 1198, 1032, 1149, 1227, 760, 529, 2192, 2112]. Generalizations [22, 1411, 2097, 1607, 1688, 592, 646]. Generalized [1948, 1524, 1571, 2119, 286, 1455, 66, 254, 2062, 1590, 2061, 2034, 1605, 1922, 1923, 126, 1487, 850, 1478, 2065, 378, 221, 1559, 1056, 1616, 1958, 250, 542, 2070, 1208, 452, 1110, 1412, 134, 1747, 1154, 1521, 1809, 1977, 1782, 1769, 1074, 387, 137, 1770, 1741, 1798, 1630, 792, 1348, 1969, 1367, 763, 1435, 1603, 2011, 533, 328, 329, 1278, 1174, 1665]. generally [447]. generate [720]. generated [2253, 77, 875, 1931]. generates [1333]. Generating [10, 231, 1265, 149, 912, 2157, 1680, 1689]. generation [2125, 1928, 936, 1610]. generator [694]. generators [850, 1472, 455]. generic [2021, 422]. Gentle [585]. genus [60, 284, 679]. Geodesic [2087, 1491, 627, 1756, 865, 1946, 4, 1775, 1868, 1916, 1504]. Geodesics [1615, 1504, 2127, 1473, 2257]. geodesy [1620]. Geometric [1108, 1479, 2063, 578, 658, 580, 2064, 2035, 1469, 1339, 1866, 112, 1985, 1023, 194, 95, 2036, 1420, 1884, 2113, 1342]. Geometrical [555, 937, 843, 944, 2126, 1022, 1645, 1045, 288, 902]. Geometries [1016, 671, 180, 54]. Geometro [1301]. Geometro-stochastically [1301]. Geometry [1525, 2023, 520, 426, 1134, 1249, 1766, 2262, 1388, 433, 1919, 767, 1986, 22, 1141, 1352, 153, 375, 2254, 725, 157, 1322, 237, 280, 197, 743, 922, 1252, 1966, 659, 978, 1861, 315, 2257, 1068, 1568, 566, 101, 1146, 717]. Ghost [2006]. Giacomini [855]. Gibbs [851, 1781, 2155]. Ginzburg [2176, 2177, 1081, 1082, 1991, 1870, 1126]. Girardello [1221]. given [856]. Gleason [1384]. Global [1088, 302, 252, 516, 1281, 68, 1878, 676, 1401, 1442, 1359, 531, 1913, 2123, 2129, 1133, 315]. GNS [282]. Gödel [2086, 1502]. Gödel-type [2086]. golden [1754]. Good [457]. Gordan [912, 155, 1720, 49, 924, 2146, 1216, 483, 1439]. Gordon [302, 1975, 68, 276, 1072, 404, 695, 354, 892, 1133]. governed [451]. Gowdy [1677]. GR [1887, 360, 1979]. grade [51]. Graded [2254, 327, 966, 1032, 1033, 998, 501, 2183, 529, 964, 1776, 329, 164, 1146, 1863, 1174]. graded-fermion [2183]. gradient [1681, 553]. Grammians [1276]. grand [67]. graph [203]. Graphical [1359]. Graphics [1257, 1295]. graphs [1326]. Grassmann [729, 300, 2154, 136, 537, 1191, 130, 712]. Grassmannian [2113]. gravisolitons [866]. Gravitating [1873, 1168]. gravitation [1321, 1397]. gravitation-electromagnetic [1321]. Gravitational [2126, 2127, 213, 910, 962, 869, 2012, 2125, 1756, 2043, 870, 435, 759, 969, 2014, 776, 5, 2049, 2207, 1136, 74, 1091, 1916, 416, 909, 492, 273, 417, 313, 314, 315, 687, 1029, 1434, 1504, 2094]. Gravity [42, 438, 151, 906, 522, 431, 1578, 409, 439, 1673, 2086, 38, 432, 446, 1579, 1874, 1801, 1674, 1322, 1323, 1761, 476, 437, 996, 1714, 2208, 761, 477, 445, 415, 1395, 1242, 1243, 1918, 2169, 447, 443, 596, 597, 1583, 1584, 436]. great [1031]. Greechie [846]. Green [1551, 750, 1184, 1170, 839, 500, 21, 606, 64, 2219, 1251, 1965, 271, 371, 1704, 1867, 2244, 2197, 1161, 241]. Green-functions [750]. Green’s [391]. Green’s-function [391]. Groenewold [1300, 1952]. Ground [2231, 1055, 21, 1990, 1024, 1828]. ground-state [21]. Group [1939, 335, 336, 1674, 1443, 510, 1021, 572, 651, 927, 2200, 2214, 777, 1311, 2063, 1877, 1150, 318, 678, 78, 653, 822, 155, 1452, 729, 44, 741, 655, 959, 1720, 1719, 1398, 223, 688, 800, 500, 2047, 234, 118, 926, 560, 1587, 136, 1135, 1874, 528, 726, 420, 2019, 387, 49, 579, 1868, 2159, 80, 9, 230, 200, 1065, 732, 531, 1215, 580, 2064, 372, 615, 1727, 1728, 99, 1462, 679, 2056, 326, 898, 161, 1487, 490, 2077, 1556, 880, 327, 674, 813, 1046, 895, 127]. Group-theoretical [1674, 726, 579]. Groupoid [1323]. groups [482, 1469, 1853, 593, 1831, 1719, 1398, 766, 1807, 79, 1626, 1329, 1881, 1768, 1417, 369, 1290, 920, 1393, 1394, 1501, 2220, 1846, 2053, 488, 324, 2144, 1724, 1178, 1179, 810, 1293, 1516, 126, 1004, 699, 1851, 1102, 2259, 241, 967]. growing [1653]. GUE [1047]. Gupta [1786]. Gustavson [670]. Gyldén [1797]. gyrostat [1160].

H [567]. hair [650, 481]. hairy [1675]. half [1631, 874, 2008, 805, 1289, 677]. half-Hartley [805]. half-plane [1631]. half-space [874, 677]. Hall [1959, 487, 1002, 1345, 1828, 1051]. Hallén [299]. Hamilton [1362, 466, 1463]. Hamiltonian [735, 1521, 32, 2136, 348, 1572, 2202, 1734, 106, 682, 349, 683, 780, 1314, 636, 1594, 1574, 2007, 947, 1832, 1154, 889, 1274, 1769, 1896, 1609, 2110, 1026, 1008, 1610, 1251, 2234, 1192, 647, 507, 1381, 456, 581, 257, 1349, 260, 1537, 1356, 2161, 847, 1361, 2011, 2124, 1424]. Hamiltonians [380, 129, 856, 1746, 1820, 1649, 172, 2158, 201, 706, 1069, 247, 634]. handles [60, 284]. handling [2115, 2267]. Hankel [1585]. Hard [1613, 29, 1889, 2224]. Harmonic [1034, 204, 232, 2271, 501, 1561, 766, 1807, 2150, 2101, 830, 1782, 291, 1761, 2173, 2186, 2241, 1610, 684, 1742, 1540, 1762, 2102, 2190, 892, 677, 340, 746, 1267, 301, 1790]. harmonics [1928, 1955, 1307, 2199]. Harper [1895]. Harper-like [1895]. Hartley [805]. Hartree [1054]. having [366, 940, 941]. Headings [494, 931, 1370, 1815]. Heat [871, 363, 556, 1035, 364, 522, 1477, 1038]. Hecke [1845, 1685, 370]. Heckman [535]. Heckmann [535]. Heisenberg [1521, 1336, 778, 737, 16, 814, 1985, 1892, 118, 1154, 196, 1859, 1065, 1555, 2134, 1391, 749, 512]. Heisenberg/Schrödinger [1521, 1154]. Helicity [111]. Helmholtz [371]. hereditary [591]. Hermite [134, 367, 1267]. Hermitian [2136]. Herrera [2147]. heterostructures [1662, 818]. Heun [1171]. Hidden [1451, 1961, 512, 1457, 50, 259]. hierarchical [1464, 147]. Hierarchies [1275, 73, 248, 1025, 1665, 1666, 1234, 2203, 900, 210, 353, 1904, 623, 1974, 722, 1317, 262, 2011, 2248]. Hierarchy [310, 2006, 32, 713, 348, 1665, 1200, 590, 1009, 2157, 107, 620, 251, 209, 1609, 469, 35, 308, 1379, 647, 1912, 2121, 1316, 904, 2237, 1236, 1709, 258, 1671, 261, 356, 648]. Higgs [1220, 1559, 1819, 92, 434, 10, 231, 843, 1994, 747, 1505]. High [88, 1944, 300, 90, 1073, 990, 147, 706, 393, 84, 997]. High- [88]. high-energy [90, 706]. High-frequency [1944]. high-order [84, 997]. high-temperature [147]. Higher [427, 1010, 2217, 1328, 1136, 1708, 1132, 905, 1551, 37, 1338, 1854, 1928, 38, 1607, 1894, 718, 503, 914, 157, 1963, 761, 308, 1366, 943, 1237, 325, 262, 304]. Higher-dimensional [427, 1136, 1132, 37, 38, 1607, 761, 1237, 262]. Higher-order [2217, 905, 1551, 1338, 718, 308, 304]. Highest [1744, 1921, 1552]. highly [405]. Hilbert [927, 1505, 651, 167, 2095, 56, 777, 233, 636, 1594, 766, 926, 321, 1135, 336, 25, 2192, 421, 879, 880, 1078, 882]. Hilbert-space [321]. Hirota [2103, 1129, 1943, 1798]. histories [1412, 1739, 387, 1483, 845, 2198, 388]. histories-based [1412]. history [387, 1745, 894]. Hodograph [467]. Hofstadter [835]. Hofstadter-type [835]. hole [1661, 2164, 594, 39, 1205, 1839, 1169]. holes [593, 357, 928, 552, 956, 1802, 910, 1675, 650, 596, 481, 1805, 212]. Holm [1868]. hologram [441]. holonomic [717]. holonomies [907, 763]. Holonomy [440, 324]. Holstein [1056, 2066, 1597]. HOMFLY [630]. Homogeneity [463]. Homogeneous [1280, 1479, 1232, 1765, 654, 334, 179, 423, 1444, 1101, 2122, 957, 1361].

Homothetic [1087, 408]. Hopf [958, 1979, 2099, 1097, 321, 696, 1686, 732, 1725, 83, 2175, 534]. Hopfield [2081]. horizon [1757, 1028]. horizons [2167]. horizontal [95]. Hove [1952]. Howe [1881]. Hubbard [1054, 1298, 100, 1789]. huge [722]. Hurwitz [61]. Hydrodynamic [2004]. hydrodynamical [1541]. hydrogen [1479, 1337, 1733, 103, 1854, 1152, 1335, 1644, 172, 1385]. hydrogen-like [1479, 1337, 1733]. Hyperbolic [2091, 2270, 1923, 366, 2096, 874, 2245, 1281, 148, 1849, 1628, 2160, 1309, 1613, 1038]. hyperboloid [1292, 1964, 1491]. hyperboloids [396]. Hypercomplex [1261]. Hyperfinite [1534]. Hyperfinite-dimensional [1534]. hyperfunction [1691]. Hypergeometric [1363, 1061, 2259]. hyperkahler [1322]. hypermultiplet [1936]. hyperspherical [1955, 1307, 2199]. hypersurface [1410]. hypersurfaces [1364]. Hypersymmetry [1032].

I-matrices [1445]. ibid [776]. IBM [837]. Ichikawa [258]. icosahedral [959]. ideal [831, 1835]. ideals [619]. identical [1930, 1070]. identification [1047, 1076]. Identifying [809]. Identities [40, 155, 852, 561, 368, 771, 1247, 82, 565, 1475, 2211, 1981]. identity [1405, 1973, 1000]. II [2213, 1538, 17, 2127, 117, 1666, 2166, 583, 841, 1594, 1951, 1353, 1189, 956, 503, 1112, 771, 63, 336, 157, 1416, 606, 806, 1964, 1036, 1018, 994, 1394, 422, 1383, 53, 1179, 941, 1093, 1728, 256, 1709, 315, 1923, 1181, 1475, 263]. III [305, 1501, 1029, 1531, 2211]. ill [874]. ill-posed [874]. image [314]. images [1029]. Imaginary [1526]. Imaging [547]. imbedding [752, 1038]. impedance [789]. Implementation [1855]. implementations [1417]. implications [890]. Imploding [797]. imprimitive [2019]. imprimitivity [2170]. Improved [90, 2075]. improvement [294]. improvements [1634]. Improving [1692]. impulsive [776, 5, 12, 1916, 1504]. impurities [1349]. impurity [2031]. incidence [951, 994]. incident [983]. Inclusion [733]. incompleteness [865]. incompressible [954]. Incorporation [1064, 1324]. increasing [1600]. indecomposable [2222, 458]. independence [1987, 2027, 1015]. independent [1803, 505, 1359, 827, 1501, 551]. independent-oscillator [505]. indeterminacy [1384]. Index [496, 495, 494, 933, 932, 931, 1372, 1371, 1370, 1815, 970, 453, 2257, 226, 1701]. indicators [1844]. Induced [521, 197, 160, 2086, 2018]. inequalities [1416, 24]. inequality [1962, 1144, 1925]. inequivalent [2229]. Infeld [1817]. Infinite [1494, 404, 478, 1384, 1349, 344, 777, 2116, 467, 1796, 60, 284, 1425, 136, 1135, 1073, 830, 516, 783, 1377, 225, 1348, 2151, 1851, 2113]. Infinite-dimensional [404, 478, 777, 136, 1135, 1377, 1851]. infinite-genus [60, 284]. Infinitely [1354, 2187, 306]. infinitesimal [1431]. infinity [1353, 1800, 1938]. inflationary [1433, 1718]. Information [1622, 1794, 1958, 1389, 1070, 1561, 1367, 1158, 1267, 673, 301]. Information-theoretical [1389]. informational [48]. infrared [1416]. infravacua [1783, 1599]. ingoing [269]. Inhibition [857]. Inhomogeneous [762, 737, 265, 874, 16, 814, 1461, 358, 908, 1686, 2220, 810, 2076, 490]. Initial [1277, 569, 68, 250, 1214, 526, 406]. initial-boundary [68, 1214]. initialization [1468]. injection [1793]. Inlaying [1020]. inner [1678, 239]. Inönü [329]. Inönü-Wigner [329]. Inquiry [1183, 1219]. inspired [2171, 354]. instabilities [720]. Instability [593, 2018, 1931, 2122, 263]. instanton [2093]. Instantons [986, 1220, 1855, 1322, 226]. instruments [1456]. integer [21, 325, 1051, 1959]. Integrability [1339, 621, 1075, 1355, 829, 106, 33, 16, 814, 1940, 718, 1942, 1669, 1026, 307, 1122, 1545, 1624, 1813]. Integrable [1806, 1702, 500, 107, 1140, 412, 1497, 1426, 1238, 259, 758, 1672, 1025, 1524, 1665, 1666, 1410, 2039, 682, 683, 780, 1314, 1831, 69, 1411, 2181, 858, 1208, 1985, 1128, 190, 1668, 1608, 993, 1129, 2203, 468, 2103, 1943, 719, 901, 210, 254, 1708, 2062, 591, 2061, 903, 1202, 2083, 1019, 1499, 255, 256, 470, 258, 1576, 471, 1834, 1003, 1196, 262, 241, 1837, 649, 2041]. Integral [710, 1507, 30, 135, 2173, 1169, 2005, 148, 2244, 220, 701, 186, 170, 1984, 131, 781, 2119, 60, 284, 1425, 2228, 974, 172, 1930, 537, 1221, 236, 703, 251, 1142, 700, 917, 1481, 1063, 1933, 2218, 337, 137, 784, 1697, 1898, 2190, 1822, 1810, 178, 1529, 1383, 140, 1704, 2000, 463, 183, 1077, 954, 1230, 1106]. Integral-difference [2005]. integrals [168, 167, 1171, 1245, 1072, 300, 1363, 1645, 535, 193, 2018, 715, 175, 1989, 1653, 902, 859, 1956, 179, 744, 661, 1309, 1519, 2199, 296]. Integrating [1288]. Integration [862, 852, 811, 169, 822, 155, 171, 2182, 177, 742, 281, 180, 182, 429, 698, 633, 2273]. integrators [2037]. integro [2172]. integro-partial [2172]. integrodifferential [222, 1355]. interacting [884, 612, 300, 992, 1839, 1896, 516, 1055, 1903, 313]. interacting-free [992]. Interaction [1559, 2125, 851, 1373, 971, 1889, 2224, 350, 1375, 714, 269, 2155, 1664, 519, 1884, 2041]. interactions [344, 2116, 1150, 1729, 1113, 2109, 1380, 456, 1312, 1730, 1834, 183, 2152, 848]. Interchannel [885]. interest [1476]. interior [2101, 1762, 2102, 416, 1811]. Intermediate [1528, 309]. internal [243, 1188, 1301, 718, 148]. interpolating [548]. interpolation [884, 934]. interpretation [2127, 1894, 1454, 367, 2036, 454, 1092, 843, 75, 658, 2085]. intersecting [1550, 2250, 2091, 2270]. intersection [2026]. intersections [693]. Intertwining [1685, 1460]. Intrinsic [1362, 117]. Introduction [1039, 1623]. Invariance [2259, 151, 318, 1089, 22, 1049, 2072, 1420, 1077, 1161, 512]. Invariant [1440, 1315, 200, 792, 1724, 2059, 1038, 2200, 105, 444, 1939, 1449, 678, 1410, 856, 998, 972, 690, 928, 552, 956, 1553, 1737, 62, 752, 918, 370, 1865, 1980, 456, 1120, 1516, 161, 1255, 1079, 1805, 1310]. invariants [150, 1949, 855, 1950, 1681, 1596, 656, 1247, 1224, 1359, 920, 1697, 428, 1123, 1516, 565, 1475, 2211, 436]. Inverse [205, 1842, 249, 2076, 897, 710, 1108, 1479, 460, 1337, 1733, 2118, 1319, 1562, 288, 384, 297, 945, 948, 937, 199, 489, 1367, 1254, 1710, 2117, 1772, 1447, 1524]. inversely [3, 736]. inversion [805, 1857, 1962, 1755]. inverted [854]. inviscid [1125]. involution [1116]. Involutions [509]. Involutive [1849, 2082]. involving [1998, 631]. ionization [759, 969]. irrational [692, 94]. Irreducible [800, 1555, 1877, 959, 688, 1856, 528, 1723, 1744, 1921, 2057, 1407]. Irregular [1096]. irreps [47, 51, 530]. Irreversible [2120, 1163]. Ising [714, 1664, 147, 679]. Isogroups [276]. Isolated [1825, 1827]. isometries [154, 153, 375]. isomonodromic [2204]. isomonodromy [1524]. isomorphic [940, 941]. Isomorphism [365]. Isomorphisms [838]. isoscalar [775]. Isospectral [1941, 1589, 1421]. isotopic [578]. Isotropic [157, 1059, 385, 945, 1291]. Israel [868]. issue [1039, 1623]. iterated [734]. iteration [1706]. iterative [1855]. itinerant [1055, 1052]. Itzykson [701, 700, 220, 917]. Iwasawa [1687, 1849].

J [927, 2271, 568, 1477, 2063, 567, 1958, 2263, 2224, 2262, 735, 736, 1335, 2272, 969, 374, 2176, 814, 1009, 1448, 926, 776, 229, 928, 2027, 1721, 1521, 1885, 1887, 1888, 1886, 2261, 2101, 1105, 375, 700, 2267, 1729, 2103, 2270, 1775, 2268, 2269, 230, 231, 166, 837, 2102, 2062, 2061, 1731, 815, 2225, 2064, 1776, 330, 1294, 1147, 2177, 2264, 2265, 1730, 1218, 2266, 1256, 376, 2104, 634, 1592, 85, 1106, 2273, 1959]. Jacobi [1362, 1760, 770, 1953, 1251, 520, 2050, 2152, 1755, 1342]. Jacobi-matrix [1251]. Jaulent [1086]. Jaynes [451, 196]. jet [1851]. jets [794]. Jimbo [2256, 1132]. joint [1070]. jointed [1971]. Jordan [1338]. Jordanian [2132]. Josephson [832]. Jost [1115]. Journal [1107, 1148]. junctions [832]. Justification [1346]. JWKB [1650].

Kac [2270, 934, 548, 2091, 609, 1687, 1849]. Kadometsev [2082]. Kadomtsev [1132, 2006, 308, 1276, 1708, 903, 1131, 904, 1236, 1974]. Kähler [871, 2196]. Kählerian [2260]. Kaluza [265, 89, 1875, 1137, 477, 762, 1503, 1970, 219]. Kantowski [211, 113]. Karman [990]. Kasner [762]. Kasner-type [762]. Kauffman [630]. Kaup [2119, 2121, 1726]. Kaya [776]. Kazakov [181]. Kazhdan [731]. KdV [619, 1234, 590, 1009, 2157, 1274, 1497, 721, 1498, 2010, 1671, 1672, 2011]. Kepler [146, 539, 615, 1266, 1824, 1230]. Keplerian [759, 969]. kernel [871, 364, 363, 522, 1477, 556, 1035, 489, 1924]. kernels [934, 548]. Kerr [1548, 1134, 870, 956, 416, 1435, 2130]. Kerr-interior [416]. key [2203]. Kibble [675]. Kicked [233]. Killing [624, 1227, 524, 1393, 1394, 1501, 2128, 1207, 479]. kinematic [1976, 1027]. kinematical [1072, 1841]. kinematics [1741]. kinetic [1162, 1235]. kink [1358, 745]. Kinks [865]. Kirchhoff [2036, 2001]. Kirkwood [1306]. Klaus [1256]. Klein [1072, 68, 265, 276, 302, 89, 1875, 695, 1137, 477, 892, 762, 1503, 1970, 219, 1975, 987]. Klein-type [219]. Kleinfolds [1406]. Kleinian [629]. Knizhnik [1524]. Knot [446]. Knots [54, 2253, 693, 177, 630]. Knotted [70]. Kochen [846]. Kodaira [1601]. Kolmogorov [1746]. Konno [258]. Kontsevich [1697]. Korteweg [1939, 348, 250, 620, 1607, 1129, 253, 1165, 469, 35, 1612, 1177, 1316, 1973, 1974, 36, 73]. Kostant [1274, 1598, 2251]. Koszul [1110]. Kowalewski [946]. KP [248, 1025, 32, 1665, 1666, 1200, 251, 209, 1609, 900, 210, 952, 1354, 353, 1709, 722, 2011, 1357, 2248, 648, 1546]. KP-hierarchies [722]. Kramers [1595, 1112, 334, 1695, 2261, 294, 1591]. Krein [2022, 1555, 967]. Kronecker [1948, 1305]. Kruskal [158]. Kupershmidt [1726]. Kuramoto [106, 1081]. KZ [741].

L [927, 926, 700]. label [125]. laboratory [518]. LaChapelle [1106]. Ladder [1421, 840, 841]. ladders [332]. Ladik [107]. Lagrange [533]. Lagrangian [2135, 1601, 1659, 1856, 2230, 98, 716, 2242, 274, 304, 717]. Lagrangians [2014, 395, 1137, 1136, 397, 142, 1127]. Laguerre [1561]. Lamb [606]. Lamé [386, 1922, 1923]. Lämmerzahl [166]. Lanczos [1548, 1134, 1711, 28, 2156]. Landau [2176, 2177, 602, 1081, 1082, 201, 1991, 1870, 1349, 1126, 1638]. Landau-type [1081]. Langevin [1119]. language [844]. Laplace [1842, 556, 1319, 1035, 82, 712]. Laplacian [1031, 1933]. Laplacians [123, 330]. lapse [582]. Large [2012, 344, 2136, 1869, 1464, 1416, 827, 1143, 1278]. Large-order [2136]. laser [400, 1538]. laterally [820]. Lattice [669, 1403, 938, 1604, 1830, 851, 1172, 1666, 781, 576, 174, 1223, 1652, 1670, 2173, 949, 2189, 1366, 2221, 995, 793, 293, 1837, 649, 1790]. lattices [365, 121, 679, 758, 1204]. Lavrov [1155]. law [1273, 1388, 1098]. laws [652, 1159, 2100]. Lax [32, 1573, 858, 757, 2246, 901, 1354, 788, 462, 471, 260, 2161, 946, 1837, 472, 310, 1547]. layer [2038]. layered [344]. Lazutkin [105]. leading [924]. leads [821]. Lee [849]. Left [508, 801, 118]. Left-right [508]. Legendre [1978, 698]. length [1056, 1584]. lengths [121, 950]. lens [1091, 123, 330]. lenses [687, 1434]. lensing [314, 315, 1029]. Level [2223, 824, 2068, 1010, 828, 1648, 1222, 1528, 1428, 2024, 673, 1638]. Level-one [2223]. levels [91, 1930, 335, 27]. Levi [2090, 266]. Levi-Civita [2090, 266]. Levinson [891, 1827]. Levitan [2119]. Lévy [1080]. Lewis [1615]. Lie [2266, 150, 248, 482, 58, 418, 1616, 1617, 1552, 464, 1508, 858, 2217, 1208, 875, 2098, 501, 286, 190, 1176, 769, 1587, 485, 657, 153, 375, 420, 1920, 1684, 2132, 1686, 1846, 1698, 2053, 122, 1590, 82, 1724, 615, 1921, 1180, 2251, 1486, 1516, 326, 1688, 722, 399, 632, 1851, 812, 925, 966, 55, 774, 1103, 329, 164, 1146, 1863, 895, 1813, 534]. Lie-algebraic [1686]. Lie-series [2266, 1486]. Lifetimes [2031]. Light [477, 1091, 1697, 272, 610]. light-cone [1697]. lightcone [2182]. like [1479, 450, 1337, 1733, 412, 917, 291, 1380, 239, 1466, 456, 1360, 1239, 1944, 1079, 1895, 849]. Limit [1660, 2200, 57, 2066, 1172, 1337, 1246, 60, 284, 803, 289, 804, 992, 1896, 549, 1098, 1117, 83, 2156, 1051, 1959, 1439]. limiting [2253, 833, 2090]. Limits [1206, 1162, 665, 1125, 2120, 1484, 1541, 2055]. Lindstedt [2245]. line [2105, 222, 583, 756, 2008, 1418, 2244, 2197, 227, 247, 634]. Linear [1151, 998, 1326, 414, 779, 1542, 2046, 802, 935, 1328, 1329, 757, 2100, 658, 1474, 1630, 1021, 898]. Linearization [2165, 2166]. linearizations [2010]. Linearized [361, 1888, 151, 2043, 1612, 1277, 799, 1592, 988]. linearizing [1607]. linearly [1501]. lines [1152, 1335, 1644, 1946]. link [562]. Linking [443, 168]. Liouville [2095, 276, 251, 1942, 2022, 1973, 745, 1069].

Liouville-type [276]. Liouvillian [1654]. liquids [1496]. Littlewood [487, 1002, 1727]. Lobachevsky [704]. Local [1779, 101, 444, 702, 876, 1696, 1977, 2111, 1710, 1145, 162, 163, 1533, 512]. Localization [170, 887, 602, 312, 982, 48, 637, 1933, 1783, 705, 1520]. Localized [1082, 790, 1232, 1, 837, 2, 1052, 1048, 1638]. located [2151]. log [1144, 1925]. log-Sobolev [1144, 1925]. Logarithmic [1267, 1082, 181]. logic [1412, 1384]. logics [846]. Long [1187, 298, 641, 1218, 2155, 35, 178, 255, 256, 953, 1905, 1053]. Long-range [1187, 298, 2155, 1905]. Long-time [641, 1218]. long-wave [255, 256]. longitudinally [1663]. Loop [907, 128, 738, 766, 1987, 2027, 1651, 2159, 609, 2075, 1144, 1657, 1447, 1519]. loops [738, 740, 446, 1935, 541]. Lorentz [560, 760, 2234, 65, 732, 372, 399, 699, 2079, 566]. Lorentzian [1886, 999, 102, 359, 114, 2051]. Lorenz [1278]. Lotka [150, 735, 1950, 1769, 1008, 859, 257]. Lott [435]. Low [72, 364, 1477, 88, 170, 853, 91, 1684, 104, 1055, 1217]. low-dimensional [170, 1684]. Low-energy [72, 364, 1477, 88]. low-temperature [1055]. lump [72]. lumps [1891]. Lusztig [731]. Lyapunov [1233].

M [567, 735, 776, 1147, 1256]. Machian [2079]. Macías [166]. Macroscopic [1662, 713, 1243]. Magnetic [2069, 167, 1258, 1108, 1479, 382, 131, 1014, 1931, 2031, 413, 1376, 66, 1064, 199, 208, 1474, 1731, 1419, 581, 1657, 1004, 861, 1613, 247, 634, 1121, 2152, 1126, 1046, 242]. magnetism [829]. magnetofluid [206]. Magnetohydrodynamic [2038]. magnitudes [1233]. magnon [506]. Magri [1356]. Mahan [1105]. Majorana [527]. Mandula [943]. Manev [1907, 676]. Manifestly [372, 1079, 791]. manifold [871, 1764, 1765, 102, 1191, 40, 1861, 622]. manifolds [629, 1627, 154, 1033, 116, 117, 1410, 2044, 1436, 2263, 765, 999, 1438, 2096, 440, 277, 1595, 1553, 119, 136, 537, 538, 514, 1274, 66, 1114, 742, 563, 428, 179, 659, 1515, 1899, 1283, 1565, 2252, 2265, 753, 1005, 2161, 263, 1391, 436, 1104, 2051, 1342, 1424]. Manin [1671, 2011]. Manuscript [1183, 1219]. Many [1084, 1702, 2106, 135, 1668, 21, 1113, 2187, 306, 1354, 628, 2156]. Many-body [1084, 1702, 2106, 135, 1668, 21, 628, 2156]. map [1855, 461, 1266, 1204, 649]. mapped [96]. mapping [96, 2053]. mappings [1403]. maps [1706, 766, 1807, 1754, 1034, 1587, 2186, 2241, 1484, 734]. Marangoni [1081]. Marciniak [1670, 2221]. Markov [815, 339, 1348, 246]. Markovian [548, 618]. Marsden [1033]. martingale [1558]. mas [2269]. Mass [13, 63, 1527, 936, 434, 2268, 1897, 10, 231, 245, 687, 898, 2077, 2051]. Massive [523, 976, 1852, 151, 2043, 410, 667, 1154, 1521, 1988, 2206, 1856, 2230, 518, 893, 1294, 980, 1936, 2130]. Massless [978, 1279, 612, 244, 1786, 785, 2033, 727]. Master [754, 1932, 1972, 1833]. materials [463]. Math [927, 2271, 568, 1477, 2063, 567, 1958, 2263, 2224, 2262, 735, 736, 1335, 2272, 969, 374, 2176, 814, 1009, 1448, 926, 776, 229, 928, 2027, 1721, 1521, 1885, 1887, 1888, 1886, 2261, 2101, 1105, 375, 700, 2267, 1729, 2103, 2270, 1775, 2268, 2269, 230, 231, 166, 2102, 2062, 2061, 1731, 815, 2225, 2064, 1776, 330, 1294, 1147, 2177, 2264, 2265, 1730, 1218, 2266, 1256, 376, 2104, 634, 1592, 85, 1106, 2273, 1959]. Mathematical [1107, 1148, 1652, 1778, 2174]. mathematics [1906]. matrices [883, 1949, 912, 2052, 1845, 827, 1143, 1115, 721, 50, 1052, 2112, 99, 1225, 1910, 1445, 1256, 1408, 826, 1755]. Matrix [2097, 1782, 1609, 325, 1690, 821, 1025, 363, 1338, 1552, 729, 2181, 403, 1045, 1049, 1441, 2254, 1845, 796, 694, 351, 1528, 2020, 1036, 563, 1251, 1965, 1115, 308, 80, 600, 755, 743, 2053, 793, 2112, 181, 1576, 125, 2057, 2058, 1699, 2104, 1583, 1547, 825, 1251]. matter [521, 1041, 1044, 2135, 1054, 1056, 1039, 870, 2086, 1045, 1189, 503, 1049, 1931, 1047, 1839, 1040, 1043, 1055, 416, 217, 1052, 1282, 1048, 893, 1294, 980, 1042, 1053, 1050, 1046, 1051, 2046]. Maximal [1681, 637, 872, 1773, 925]. Maximally [1664]. Maximum [1809, 2243, 1367, 1254]. Maxwell [1159, 1524, 1227, 2069, 1011, 1673, 1072, 1567, 644, 751, 102, 362, 145, 831, 218, 1094, 790, 1568, 764, 1752, 1792, 2003]. Maxwell-dilaton [1094]. Maxwell-dilaton-axion [362]. McCoy [368]. mean [344, 2039, 1754, 516, 832, 2133]. mean-field [344]. meander [1327]. means [102, 707]. Measure [1892, 698]. Measure-valued [1892]. measurement [638]. measurements [666, 1070]. measures [1558, 575, 1637, 1968, 1195, 2133, 1367, 846]. Mechanical [1270, 187, 896, 1453, 1810, 140, 716]. mechanics [1057, 1690, 1926, 29, 839, 1960, 1151, 1060, 1412, 1821, 289, 1779, 605, 975, 754, 1908, 715, 401, 1999, 290, 175, 1564, 1739, 1857, 1962, 1063, 2109, 1971, 2141, 586, 587, 67, 2143, 540, 550, 1065, 26, 2236, 139, 1117, 1344, 1654, 1347, 1565, 466, 2194, 845, 641, 1218, 1486, 2266, 1537, 1459, 1906, 582, 2035, 36, 708, 2113]. media [643, 750, 1351, 249, 1461, 1663, 945, 217, 2117, 1161, 1080, 791]. Medina [700]. medium [1271, 1660, 897, 1944]. Mee [1256]. Meissner [1126]. Melnikov [905]. memory [950, 2015]. mesoatom [607]. Mesoscopic [816, 822, 817, 823, 141, 829]. metal [886]. metallic [1, 837, 2]. metals [824]. Method [2115, 2267, 970, 144, 364, 1477, 1877, 1279, 680, 822, 1950, 1929, 1319, 403, 1321, 1045, 1353, 1562, 2150, 1695, 2261, 577, 158, 720, 902, 2053, 2236, 868, 299, 1466, 2034, 2112, 677, 622, 1306, 490, 1829, 1078, 1547]. methods [710, 400, 1163, 287, 1519]. Metric [1899, 1548, 2014, 1874, 1674, 419, 2074, 415, 42, 1772, 1145, 520]. metric-affine [1874, 415]. metric-connection [2014]. metric-torus [1674]. Metrics [801, 1319, 2029, 359, 1886, 760, 890, 4, 1775, 2201, 114, 41, 1283, 1444, 753, 2050]. MIC [1266]. MIC-Kepler [1266]. Microcanonical [850]. Microscopic [1538, 1569, 1454]. middle [1793]. midi [906]. midi-superspaces [906]. Migdal [181]. Mikhailov [723]. Milburn [451]. Mills [2137, 2027, 843, 2210, 1505, 1149, 1296, 381, 1166, 1506, 593, 570, 69, 741, 1240, 2047, 575, 1987, 1651, 1322, 237, 545, 1377, 1675, 893, 1294, 980, 1936, 2093, 2237, 2025, 240, 1446, 274, 798, 317, 596, 1360, 165, 1431, 1996, 1423]. Mills-dilaton [2137]. minimal [637, 1017, 508]. minimally [608]. minimally-deformed [608]. minimum [1629, 1367]. minisuperspaces [1566, 2129]. Minkowski [1246, 320, 875, 877, 2048, 1225, 1268, 327, 2153]. Miodek [1086]. mirror [1930]. missing [125]. Miura [1542, 1317]. Miwa [2256, 1132]. Mixed [19, 228]. mixing [1644]. mixtures [1823]. MKdV [948]. mode [783, 1876, 1350]. model [400, 2200, 1593, 1220, 1311, 1054, 1538, 2066, 265, 2068, 635, 499, 1298, 2039, 1272, 912, 989, 1012, 435, 451, 1186, 644, 92, 1201, 1448, 714, 1081, 573, 1693, 1153, 1668, 1464, 1580, 1454, 279, 22, 915, 757, 1062, 1669, 385, 1931, 598, 95, 2131, 1597, 505, 2231, 719, 1535, 23, 437, 196, 1427, 518, 2155, 1457, 563, 684, 903, 787, 1066, 1193, 1491, 147, 1900, 2193, 1346, 181, 2081, 1444, 1068, 2169, 1024, 679, 142, 100, 1753, 610, 848, 849, 675, 748, 512, 1051, 1959, 1790]. modeled [1150, 892]. Models [835, 1690, 344, 1025, 450, 665, 1056, 31, 1707, 713, 1451, 2165, 2166, 781, 300, 350, 896, 1493, 1400, 1662, 91, 206, 1471, 691, 796, 2073, 351, 1895, 352, 2109, 2232, 1544, 1598, 1426, 1714, 1136, 1770, 595, 1122, 1708, 2247, 415, 217, 1715, 2236, 1019, 686, 1433, 1093, 1743, 762, 1238, 553, 1255, 966, 296, 295]. modern [1503]. modes [542, 1147, 398]. Modified [250, 484, 1768, 455, 1325, 1562, 1535, 406, 1672]. Modular [1853, 2072, 62]. modulations [857]. modules [1552, 1471, 25, 2024, 2192]. Moduli [693, 60, 284, 575]. Moduli-space [693]. modulo [712]. Moerbeke [356]. molecular [601, 1862, 1306]. Molien [1440]. Moment [1416, 1163, 1809, 577, 581, 2243]. momenta [637, 1017]. moments [1245, 1644, 1494, 2255, 1064]. Momentum [1995, 144, 132, 243, 557, 461, 524, 298, 1971, 2037, 1741, 396, 991, 2197]. monodromy [1204]. monoidal [1844]. monopole [2069, 570, 89, 192, 1121, 242]. Monopoles [2186, 192, 1303, 66]. Monotone [2201]. monotony [1705]. Monte [439]. Moody [2270, 2091, 609, 1687, 1849]. Morosi [1356]. Morse [478, 314, 2257, 1487]. Moser [1598, 1411, 1746, 1669, 2159, 259, 1833, 2204, 110]. Motion [989, 1611, 2147, 408, 405, 74, 1540, 1491, 627, 1203, 1833, 1463, 1079, 304]. motions [1087, 377, 1756, 1410, 2048, 76]. move [2115, 2267]. Moving [1379, 513]. Moyal [1983, 2070, 1399, 1300, 194, 1210]. MR [568, 567, 735, 969, 374, 814, 1009, 776, 229, 928, 1105, 375, 230, 231, 166, 815, 330, 1147, 1256, 376, 634, 85, 1106, 736, 837]. Multi [888, 2248, 682, 2202]. multi-component [2248]. Multi-Hamiltonian [682, 2202]. Multi-periodic [888]. Multibody [1036]. multichannel [149]. multicollision [1541]. Multicomponent [1595, 412, 296, 295, 1666]. Multidimensional [2181, 1550, 294, 288, 412, 1433, 2169, 897]. Multifold [146, 539]. multifractal [1635]. Multilinear [1469]. multiloop [45, 2182]. multimode [400]. multinomial [1190]. multiparticle [1069, 2035]. Multiperiodic [1112]. multiphoton [196]. Multiplane [314, 315, 1029]. Multiple [747, 1079, 970, 1135, 469, 1658]. Multiple-time [1079, 469]. multiplet [893]. multiplets [1294, 980]. multiplicative [1438]. Multiplicity [920, 2206, 1727, 1728]. multipole [1091, 1602]. multipoles [544]. multiscale [547]. Multivariable [300, 1685]. Multivector [1659, 1764]. Muskhelishvili [186].

Nahm [1983, 2070]. Nambu [1074]. nanoelectronic [203]. nanostructures [816]. Natural [984, 165, 1210, 1511, 744]. nature [685]. Navier [1162, 1125, 1633]. near [1192, 1628]. nearby [861]. Negative [1190, 934, 607, 310]. Nelson [2194]. Nester [795]. net [392]. nets [1533]. network [1947]. networks [1578, 277]. Neumann [2095, 386, 599, 2159, 388, 1547]. Neumann-type [1547]. neutral [1765, 581]. neutrinos [2108]. Neveu [881]. Newell [2119, 2121]. Newman [2207, 2130]. Newton [1562, 1316, 1703, 2242]. Newton-Sabatier [1562]. Newtonian [76, 1079, 475]. Newtonian-like [1079]. NGT [523]. Nijenhuis [901, 1100, 1174]. Nikodym [1456]. nilpotent [2098]. no [927, 568, 567, 735, 969, 374, 814, 1009, 926, 776, 229, 928, 1105, 375, 700, 1092, 230, 231, 166, 815, 330, 1294, 1147, 650, 1218, 1256, 376, 634, 85, 1106]. no-scalar-hair [650]. nodal [331, 1287]. Noether [111]. noise [378, 817, 278, 1880, 2184, 1098, 1428, 645]. Non [934, 1013, 617, 936, 763, 185, 128, 2136, 1338, 636, 1594, 1542, 803, 206, 702, 574, 1563, 757, 1663, 948, 237, 1290, 2218, 141, 1207, 717]. Non-abelian [1013, 936, 128, 1563, 2218]. non-commutative [803, 237]. Non-covariance [763]. non-Gaussian [1663]. non-Hermitian [2136]. non-Hilbert [636, 1594]. non-holonomic [717]. non-linear [1542, 757]. non-local [702]. Non-negative [934]. non-null [1207]. non-reducible [1338]. non-self-adjoint [574]. non-simply [185]. Non-standard [617, 1290]. non-static [206]. non-trivial [141]. non-vanishing [948]. nonabelian [293]. Nonalgebraic [1545]. nonanalytic [681]. Nonassociative [2141]. nonautonomous [1357]. nonbilinear [61]. Noncanonical [2208]. nonclassical [1061, 120]. Noncommutative [433, 867, 281, 1068, 1146, 1012, 767, 1141, 1210, 1511, 95, 1323, 280, 197, 743, 2201, 1529, 732, 659, 963, 126]. noncommutativity [1413]. noncompact [136, 174, 1112]. noncomplete [841]. Nonconducting [1161]. nonconservative [1493, 1463]. nonconstant [1258]. nondynamical [1576]. Nonequilibrium [1816, 2116, 551, 1538, 754, 1465]. Nonexistence [68, 1819, 252, 846]. Nonextensive [1900, 1794]. nonfactorizable [2074]. Nonholonomic [1908, 1270, 1127]. nonhomogeneous [1014]. nonidentical [1188]. noninformative [1070]. noninteger [1928]. nonintegrability [621]. nonintegrable [519, 2247]. Noninvertible [1001]. nonisolated [1825]. Nonisospectral [2203]. Nonlinear [681, 1874, 2019, 1083, 1285, 1181, 965, 1196, 76, 1432, 1232, 31, 68, 2171, 958, 33, 2007, 303, 2014, 1821, 484, 1023, 1607, 1940, 573, 1328, 2008, 1248, 252, 1670, 306, 1611, 2078, 108, 588, 254, 2062, 2061, 1381, 1202, 1168, 1429, 355, 1499, 309, 878, 1913, 813, 533, 1755, 491, 212]. Nonlocal [1159, 1204, 1827, 1852, 652, 94, 584, 1532]. nonminimal [871]. nonminimally [1918]. nonorientable [938]. Nonperturbative [1186, 2034, 443]. nonquadratic [851]. nonreciprocal [1078]. Nonrelativistic [2234, 1178, 1179, 583, 1481, 2231, 2109, 307, 1226, 1519]. nonrenormalizable [296, 295]. nonresonant [641, 1218, 1486, 2266]. nonsimply [1515]. nonsimultaneous [685]. Nonsingular [1585]. nonspherically [195]. Nonstandard [2190, 488, 233, 2052, 1703, 1195]. nonstationary [643, 897]. Nonsymmetric [492, 273, 2043, 1736, 502, 417]. Nontrivial [296, 350, 1201, 1448, 719, 1683, 1824]. nontwisting [2126, 2127]. Nontypical [422]. nonultralocal [352]. Nonunitary [1967]. nonzero [132, 1464, 1017]. Nordstrom [2209, 1360, 1030]. Nordstrom-like [1360]. norm [1367]. normal [1736, 502, 1935, 670, 951]. normal-ordered [1935]. Normalization [1171, 1414]. Note [410, 211, 1234, 1467, 2257, 226]. Nöther [229, 7, 1189]. Novel [1149, 828]. nuclear [1693]. Null [569, 2045, 1241, 1978, 872, 1318, 870, 361, 1888, 269, 1473, 1207, 764, 1843]. null-tetrad [870]. Number [67, 1642, 331, 1337, 1733, 467, 571, 1929, 2138, 1570, 1740, 687, 1029, 143, 1488]. number-theoretical [1570]. numbers [168, 1855, 1261, 424, 1278]. Numerical [178, 2005, 1930]. Nyström [207].

Oblique [994]. oblivious [1304]. observable [1741]. observables [574, 1015, 48, 684, 1784, 2143, 1956, 509, 1826]. observation [1304]. observed [198]. observer [1229]. obstacle [1606, 2150]. Obstructions [629, 188]. obtained [378, 155]. obtaining [591]. occurrence [2016]. occurring [1152, 1335]. Octonionic [973, 385]. octonions [2097]. odd [2049, 157]. Off [1053]. Off-diagonal [1053]. Oikawa [647]. One [1394, 2075, 706, 2093, 883, 1642, 1490, 449, 128, 2068, 971, 129, 855, 856, 1889, 2224, 3, 736, 1561, 1341, 1045, 714, 1160, 1648, 1668, 1660, 576, 172, 1154, 1521, 1651, 1528, 1417, 2155, 891, 1664, 1193, 139, 1283, 402, 2112, 1496, 141, 1069, 1545, 143, 1256, 1255, 1914, 642, 1519, 2223, 709, 512, 1980]. one- [1496, 1980]. one-degree-of-freedom [856]. One-dimensional [706, 883, 1642, 2068, 129, 855, 1889, 2224, 3, 736, 1561, 1341, 714, 1668, 1660, 576, 172, 2155, 1664, 1283, 143, 1256, 1255, 1914, 709, 512]. One-instanton [2093]. One-loop [2075, 128, 1519]. one-parameter [1528, 1417]. one-particle [1069]. ongoing [644]. Online [1107, 1148]. Ono [1277, 1799, 406]. Onset [1450]. Open [1297, 1153, 1668, 216]. Operator [970, 56, 235, 1694, 1956, 1485, 2134, 1699, 2104, 295, 619, 2095, 363, 2067, 32, 1707, 1617, 556, 383, 1299, 91, 1941, 1401, 1680, 1154, 1521, 1222, 1455, 1035, 1224, 1512, 1036, 199, 1991, 1404, 1194, 940, 941, 2251, 1225, 2195, 1308, 898, 2198, 1006, 1583, 1584, 1505, 1350]. Operator-theoretical [56]. operators [569, 871, 872, 2213, 1044, 380, 1551, 1763, 1338, 1552, 1436, 2263, 998, 1691, 34, 2181, 1796, 2217, 557, 688, 800, 134, 913, 876, 1111, 1587, 1563, 1441, 1588, 1035, 1402, 2246, 2131, 1472, 1376, 2140, 1685, 1331, 1332, 1018, 2189, 1770, 1965, 901, 921, 1630, 2005, 423, 82, 1724, 1265, 1485, 1458, 981, 1236, 1144, 421, 1069, 161, 1181, 965, 722, 1421, 227, 1217, 547, 1102, 393, 1161, 1460, 1701, 2223, 2205]. Oppenheimer [2049, 1018]. optical [144, 1466, 1568]. optics [1023].

Optimal [2032]. optimization [297]. optimized [884]. optimum [657]. orbi [1406]. orbi-Kleinfolds [1406]. Orbit [1516, 812, 925, 1386]. Orbit-orbit [812, 925]. orbital [829]. orbitals [1520]. orbits [70, 1785, 1628, 1825]. order [2179, 1551, 2136, 1338, 1436, 2263, 1644, 2217, 718, 1554, 1465, 253, 308, 138, 1823, 2054, 1003, 219, 1053, 905, 84, 997, 310, 304]. Ordered [1518, 1736, 1722, 1935]. ordering [970, 502]. ordinary [621, 917, 130, 2025, 1566, 1813]. origin [248, 825]. original [837]. Orrick [368]. Orthogonal [949, 418, 1877, 1171, 222, 498, 1831, 134, 756, 1685, 1589, 1083, 1846, 126, 301, 1689, 1518]. Orthogonality [1286]. orthogonally [2087]. Orthonormal [1638, 1841, 1403]. Oscillation [1635]. oscillations [346, 347]. oscillator [1184, 232, 2271, 1258, 1058, 1297, 1059, 557, 1299, 501, 1561, 2047, 191, 1737, 1682, 1782, 505, 640, 977, 1742, 2190, 1530, 892, 1824, 340, 746, 1267, 101, 301, 1997, 1790]. oscillators [1060, 1341, 1142, 291, 2140, 608, 1540, 641, 1218, 1486, 2266, 1421, 1269]. Oscillatory [167, 168, 882]. osp [1856, 2230, 1405]. other [1016, 2010]. Outer [1028, 105, 1727, 1728]. outgoing [269]. Outline [447]. Overall [628]. Overcoming [173, 1721]. Overcomplete [129]. Overview [448].

p [1457]. p-adic [1457]. packet [576, 834]. PACS [494, 931, 1370, 1815]. Painlevé [1165, 307, 1237, 1951, 1953, 351, 305, 806, 1708, 1498, 109, 2054, 1726, 2059]. pair [886, 3, 736, 757, 788, 511, 310]. pairing [100]. pairs [801, 46, 1402, 1354, 260, 812, 1837]. pairwise [884]. palindromic [1044]. paper [927, 926]. para [2110]. para-supersymmetry [2110]. parabolic [1681, 1214, 1773]. parabosons [328]. parafermions [527]. parallel [1515, 149]. paramagnetism [1014]. parameter [729, 2216, 2052, 1528, 1417, 2207, 218, 1120]. parameters [935, 2227, 2228]. Parametric [116, 117, 2017, 857, 258, 1255]. parametrization [17, 99]. parametrized [974, 335, 336, 726, 725]. paraoscillator [287]. parity [32, 1518]. Partial [8, 374, 138, 275, 1558, 2172, 318, 1951, 1607, 1940, 621, 120, 505, 1726, 1488]. Partially [2106, 255, 256]. Particle [74, 2116, 2226, 680, 1185, 131, 2045, 1796, 1186, 21, 1154, 1521, 1930, 833, 1062, 1564, 1263, 563, 198, 1934, 10, 231, 1197, 2193, 339, 815, 581, 1496, 1069, 259, 471, 1226, 1397, 2077, 2079, 1423, 1600, 1826, 1995]. Particles [1731, 1419, 2004, 612, 382, 1188, 1562, 535, 2138, 2206, 504, 1896, 245, 397, 982, 716, 1970, 1231, 708, 1079]. particular [1610]. partition [1478, 103, 1494, 62]. partitioning [883, 1256]. partitions [2028]. parts [1526, 90]. Passage [25]. past [951, 994]. Pasting [336]. Path [131, 535, 193, 137, 784, 1898, 1529, 179, 1195, 182, 429, 1106, 1984, 2180, 781, 974, 537, 1221, 236, 703, 175, 1063, 1933, 337, 1722, 2190, 1822, 178, 744, 140, 1704, 1309, 183, 1077, 954, 633, 2273, 296]. path-integral [781]. path-ordered [1722]. pattern [385]. Pauli [333, 1901]. Pauli-allowed [333]. pendulum [1022, 1989]. pendulum-integrals [1989]. Penner [181]. Penrose [1838]. Perennials [726]. perfect [474, 412, 1092, 1093, 1582, 1871, 1285]. perhaps [1314]. Period [121, 950]. Periodic [1386, 1314, 1022, 1855, 759, 969, 303, 2176, 1034, 70, 1649, 1660, 576, 1930, 888, 506, 2188, 1305, 995, 1194, 793, 706, 1870, 2177, 1825, 143, 2060, 256]. periodical [1747]. permanent [1948]. permittivity [145]. perspective [171, 125]. Perturbation [144, 1513, 1117, 18, 2136, 1297, 1842, 1298, 1893, 289, 1697, 1492, 1269]. Perturbations [1718, 764, 2084, 2085, 2137, 1128, 928, 552, 956, 585, 2149, 734, 869, 2094]. perturbative [2172, 2140]. perturbed [1649, 1376, 1098, 1177]. Peter [293]. Petviashvili [2006, 2082, 308, 1276, 1708, 903, 1131, 904, 1236, 1132, 1974]. Phase [1480, 2067, 134, 2228, 1160, 355, 2042, 750, 57, 2063, 18, 887, 1013, 1125, 853, 535, 1930, 1049, 1465, 538, 1801, 401, 453, 578, 1664, 2143, 1956, 1705, 1420, 1982, 580, 2064, 1430, 2129, 1350, 1791, 1701]. Phase-amplitude [355]. phase-difference [1701]. Phase-integral [2228, 1930]. Phase-space [1480, 2067, 134, 18, 1982]. phases [1339, 201, 1055]. phenomena [816, 813]. Phillips [462]. phonon [1914]. photon [887, 1385]. photons [786, 1118]. Phys [927, 2271, 568, 1477, 2063, 567, 1958, 2263, 2224, 2262, 735, 736, 1335, 2272, 969, 374, 2176, 1009, 1448, 926, 776, 229, 928, 2027, 1721, 1521, 1885, 1887, 1888, 1886, 2261, 2101, 1105, 375, 700, 2267, 1729, 2103, 2270, 1775, 2268, 2269, 230, 231, 166, 2102, 2062, 2061, 1731, 2225, 2064, 1776, 330, 1294, 1147, 2177, 2264, 2265, 1730, 1218, 2266, 1256, 2104, 634, 1592, 1106, 2273, 1959]. Phys. [814, 815, 376, 85]. Physical [1968, 1458, 1124, 2007, 509, 1265, 2145, 1179, 1620]. Physics [1816, 1041, 1044, 1054, 1056, 1039, 1363, 86, 1045, 1049, 1696, 1047, 268, 1040, 1043, 1055, 2174, 1052, 1048, 1042, 1053, 1050, 1046, 85, 1051, 493, 930, 1107, 1148, 1369, 1814]. Picard [2237, 2025]. picture [1154, 1521, 691, 579]. piece [1109]. piecewise [1890]. Pin [629]. Planar [564]. Planck [34, 639, 1495, 1867]. plane [1631, 1351, 152, 1760, 613, 989, 1702, 411, 1660, 1639, 1289, 543, 1213, 1882, 1602, 978, 2160, 1829, 1556, 1613]. plane-wave [543, 1829]. planes [704]. Plank [735]. plasma [1313, 1605, 515]. plate [2038]. plates [625]. Platonic [1303]. plus [1530]. Poincaré [130, 1616, 653, 1508, 1567, 44, 696, 147, 632, 674, 164, 1079, 127]. Poincaré-invariant [1079]. Point [2105, 655, 1749, 1279, 971, 513, 1720, 1719, 1398, 1464, 549, 2075, 687, 898, 1079, 748]. Point-group [655]. point-splitting [1279, 2075]. points [1757, 1287, 1434]. Poiseuille [176]. Poisson [2078, 482, 58, 838, 1572, 1437, 715, 1047, 1074, 1935, 1634, 25, 1703, 1811, 1349, 541, 399, 632, 1851, 1204, 2042, 1174]. Pol [1793]. polar [193]. polarization [1067, 1420]. polarized [1644, 1840]. poles [1338, 1049, 1709]. Polyakov [1325, 1213]. polycrystal [1228]. polyharmonic [2195]. Polyhedral [1104]. Polylogarithmic [104]. polymer [2255]. polymers [1272]. Polynomial [1175, 328, 998, 446, 561, 1953, 1260, 919, 859, 1709, 1516, 1974]. polynomials [222, 498, 1298, 134, 1561, 756, 367, 279, 770, 1440, 562, 487, 1685, 1250, 949, 1589, 1083, 630, 1101, 1267, 301, 1689, 1518, 373]. Ponzano [437]. Poschl [1487]. posed [874, 2168]. position [2004, 557, 1527, 1187]. position-dependent [2004, 1527]. positions [637, 1017]. Positive [2051, 624, 19, 1780, 2058]. Positivity [1500, 1629]. possessing [218]. Possibly [1673]. Postscript [184]. potential [128, 971, 1134, 857, 1259, 2272, 2181, 1574, 1280, 1494, 1711, 1128, 1647, 1414, 1930, 1415, 1990, 684, 104, 2191, 1967, 1194, 181, 260, 1905, 1267, 1884, 1231, 28, 301, 242, 407]. potentials [883, 2178, 1044, 1548, 1523, 1108, 1227, 855, 3, 736, 1561, 1321, 1188, 189, 935, 756, 384, 524, 195, 386, 1260, 1962, 1075, 2188, 1018, 1653, 1742, 1530, 2193, 1536, 391, 1203, 1516, 143, 1256, 1487, 1937, 247, 634, 1827, 27, 2199]. Potts [915]. power [768, 2253, 657]. Preliminary [697]. preon [1150]. Prepotential [893]. Prepotentials [1294, 980, 2093]. presence [864, 637, 2108, 1397, 1126, 242, 791]. Present [1815]. preserving [2030, 1473, 2037, 1243, 472]. Price [97]. Principal [1617, 1337, 1733, 765, 999]. principle [1538, 2044, 20, 112, 1461, 2268, 1897, 2269, 2221, 466, 478, 2058, 2152, 1463]. principles [1429]. Probabilistic [1816, 176]. probabilities [246, 1037, 1788, 85]. probability [1190, 2201, 1884]. Probing [906]. problem [710, 2213, 589, 128, 68, 460, 934, 2118, 249, 1889, 2224, 739, 613, 874, 1411, 558, 1907, 1735, 676, 1797, 1562, 1941, 604, 1569, 668, 384, 1650, 1809, 1135, 1352, 1082, 1977, 1952, 386, 577, 297, 1729, 506, 1428, 2159, 199, 588, 416, 1076, 1277, 299, 2112, 1902, 1027, 1727, 1728, 391, 1367, 1824, 2151, 2076, 1254, 1446, 1312, 1730, 2010, 406, 897, 2117, 592, 1230, 149, 1755, 1235]. problems [1041, 1044, 1054, 1056, 59, 3, 736, 583, 1702, 2106, 1321, 1045, 135, 1778, 1049, 2245, 2203, 1047, 156, 468, 752, 1211, 1040, 1043, 1214, 1589, 1740, 1990, 587, 721, 755, 455, 1865, 1991, 1055, 2174, 1052, 1992, 1048, 1306, 1620, 1042, 125, 1307, 1053, 1050, 1080, 1078, 1046, 882, 1051]. procedure [1468, 784, 2075, 1106]. process [1304]. processes [854, 1880, 25, 551, 646]. produced [1904]. Product [2218, 2044, 1438, 1595, 1482, 920, 239, 2034, 1485, 1006]. production [511]. products [1625, 911, 1399, 1736, 2022, 1722, 1518]. projection [405, 153, 375, 2112, 981]. projection-tensor [153, 375]. projections [77]. Projective [651, 927, 169, 653, 926, 1244, 528, 1418, 880]. Prolongation [723]. prolongations [404]. Proof [1374, 1213, 593, 1152, 1335, 133, 730, 561, 1669, 2258, 642, 1646]. Propagation [1663, 1091, 1850, 1872, 244, 549, 1078]. propagator [377, 103]. propagators [18, 976, 965]. Proper [703, 2268, 1897, 2269]. Properties [1153, 345, 1500, 2126, 666, 1061, 833, 866, 830, 669, 1845, 1895, 505, 2233, 827, 939, 307, 831, 217, 1405, 1968, 818, 1904, 1042, 463, 1791, 373, 1635]. property [1890, 1498]. proposal [457]. propositions [842]. protein [1569].

Proving [1949]. pruning [1209]. Pseudo [1831, 412, 1931, 722]. pseudo-Euclidean [412]. Pseudo-orthogonal [1831]. pseudo-relativistic [1931]. pseudo-superdifferential [722]. Pseudoclassical [1062]. pseudodifferential [1441, 923, 1236]. pseudomeasures [1130]. pseudounitary [600]. pulses [1022]. pumping [1524]. pure [2032, 1122]. Purely [413]. purification [2029]. purported [2111]. Putzer [2053].

Qadir [567]. QCD [1191, 197, 844]. QED [1596]. QFT [1289]. quadratic [1950, 3, 736, 1746, 761, 1137, 1203, 1204, 407]. quadratures [1250]. Qualitative [2163, 686]. quantal [1930]. Quantifying [1876]. quantitative [2133]. quantities [7, 229, 1189, 94, 1938]. Quantization [444, 1449, 93, 1988, 539, 609, 1067, 326, 1566, 1423, 773, 2179, 1852, 1524, 58, 911, 667, 1110, 1300, 1856, 2230, 1961, 194, 335, 336, 726, 1674, 577, 1323, 352, 1114, 2208, 1785, 743, 239, 339, 815, 1118, 964, 1776, 457, 510, 1021, 98, 2238, 183, 1196, 642, 2113, 929, 611, 944, 1342]. quantizations [2229]. Quantized [1329, 386, 1305, 1373, 1301, 1896, 2020, 1212, 280, 939, 122, 1368, 675]. Quantum [497, 1558, 1296, 426, 2226, 431, 873, 2028, 2164, 654, 44, 118, 190, 224, 1696, 385, 1427, 1378, 639, 579, 704, 1418, 645, 445, 1395, 786, 1966, 1344, 1157, 2145, 981, 1656, 1531, 459, 15, 376, 1556, 708, 610, 241, 970, 850, 1057, 1478, 438, 2200, 344, 777, 1523, 1162, 380, 906, 186, 522, 427, 418, 1480, 2067, 1578, 839, 836, 1960, 1337, 1733, 1984, 187, 601, 1853, 1151, 728, 1795, 1818, 822, 439, 1691, 666, 817, 780, 2106, 816, 2215, 1339, 2180, 1411, 741, 1060, 972, 688, 800, 1299, 432, 1412, 1341, 446, 863, 2029]. quantum [1821, 79, 1097, 1453, 135, 820, 2138, 960, 1648, 289, 575, 1779, 172, 1987, 2027, 21, 605, 975, 1049, 833, 1801, 877, 992, 1999, 656, 562, 830, 890, 1881, 1768, 290, 175, 2184, 1564, 669, 1739, 2185, 1481, 1845, 1857, 1962, 1224, 842, 1063, 1482, 2231, 618, 1456, 1761, 516, 369, 2109, 506, 2021, 387, 1550, 1971, 437, 1290, 2233, 1364, 1114, 2140, 1304, 1989, 1017, 2141, 454, 2208, 1305, 49, 198, 2234, 1741, 422, 2220, 138, 2143, 1002, 2023, 595, 831, 1484, 540, 1540, 530, 1859, 507, 1065, 25, 1514, 1604, 65, 26, 1810, 159, 488, 178, 2236]. quantum [1382, 732, 139, 140, 1117, 2111, 1861, 81, 1099, 963, 1654, 1982, 1862, 1968, 1345, 670, 1903, 1347, 1565, 1084, 1744, 1348, 716, 99, 671, 810, 1225, 141, 1334, 1445, 1293, 294, 447, 746, 845, 1308, 1745, 1003, 641, 1218, 1486, 2266, 429, 161, 126, 1459, 1906, 698, 894, 392, 490, 1050, 673, 443, 1532, 2035, 811, 1005, 1851, 36, 846, 662, 1583, 1584, 1476, 1533, 674, 1489, 967, 436, 388, 1828, 1310, 373, 127, 1182, 968, 2223, 2175, 1051, 1959, 1600, 1826, 1041, 1044, 1054, 1056, 1045, 1049, 1047, 1040, 1043, 1055, 1052, 1048, 1042, 1053, 1050]. Quantum [1046, 1051]. quantum-chromo [1489]. Quantum-classical [981]. quantum-group [741]. quartic [1258, 1058, 1297, 1136, 977, 260]. Quasi [498, 756, 1111, 1427, 2242, 2175, 778, 418, 1318, 319, 1855, 1780, 1589, 1530, 423, 1386, 2124]. Quasi-average [1427]. quasi-bi-Hamiltonian [2124]. quasi-circular [1386]. Quasi-exactly [498, 756, 1111, 319, 1589, 1530, 423]. quasi-free [1780]. Quasi-Hopf [2175]. Quasi-Lagrangian [2242]. quasi-orthogonal [418]. quasi-spherical [1318]. quasicanonical [1735]. quasicrystal [1640]. quasicrystals [1042]. quasiexactly [1740, 2188]. Quasifinite [1552]. quasifree [1788]. quasilinear [690, 1681]. Quasilinearization [2236]. quasinormal [1876]. quasiorthogonal [373]. Quasispin [2183]. quasitriangular [958, 83]. Quasitriangularity [810]. Quaternionic [1004, 651, 927, 1057, 1244, 926, 673, 880]. Quaternions [689]. question [37]. Quommutator [1679]. quotient [654, 659].

R [776, 700, 1147]. R. [85]. Rabi [847]. Racah [81, 1099, 1405, 660, 1774]. Radial [2048, 640, 59, 347, 474, 64, 390, 1884, 1700]. radiating [408, 74]. Radiation [1461, 595, 2115, 1758, 759, 969, 2267, 504, 544, 543, 1729, 1462, 1312, 1730, 373]. Radiation-dominated [595]. radiative [1389, 1704]. Radon [1456]. Radul [1671, 2011]. Rajeev [22]. Ramond [1458, 881]. Random [1636, 2189, 2015, 734, 378, 2148, 1272, 1494, 1049, 405, 1663, 2149, 1528, 827, 1143, 549, 1098, 645, 2117, 1829, 708, 1080, 551, 826, 825]. random-matrix [825]. range [2116, 1187, 714, 298, 2155, 1664, 1905, 1053, 1937]. rank [1217, 1330]. rapid [355, 1131]. Rarita [479]. rate [1549, 551]. Rational [1738, 1709, 806, 1075, 684, 2161]. Ray [234, 123, 330]. Rayleigh [460, 1071]. reaction [1729, 1462, 1312, 1730, 2059]. reaction-diffusion [2059]. Real [1918, 222, 1768, 2021, 1846, 1418, 1822, 1067, 2057]. Reality [473, 433, 157, 509]. realization [1175, 1693, 1848]. Realizations [1689, 619, 1095, 1735, 1874]. Realizing [188]. rearrangement [610]. reciprocal [550, 791]. Reconstruction [1298]. Recoupling [1728]. Recovery [1158, 246]. Rectangular [1893]. recursion [2246]. recursions [878]. Recursive [246, 1877]. Recursively [608]. Reduced [605, 606, 200, 98, 1728, 1431, 2248]. reducible [1338, 2230]. Reduction [1245, 1864, 1033, 1667, 788, 1871, 1407, 675, 2042]. reductions [2006, 69, 1912]. Reductive [2047, 925]. reference [1549]. reflecting [1753]. reflection [680, 318]. reflectionless [2205]. Reformulation [844, 1496]. Regge [1090, 437]. regime [585]. regions [594]. Regular [317, 800, 118, 1396]. Regularity [417, 1047]. regularization [1593, 1984, 1246, 573, 395, 1066, 705]. regularizations [2139]. regularized [1481, 255, 256]. Reissner [2209, 1360, 1030]. rejoinder [1244]. related [1627, 155, 1979, 1574, 1845, 280, 587, 455, 225, 2054, 1565, 471, 1974, 126, 1851, 241]. relating [1329]. Relation [2082, 630, 2075, 424, 1936, 1206, 1534]. Relations [913, 1690, 1286, 185, 777, 1523, 635, 681, 1140, 434, 1741, 202, 664, 551]. relative [2067, 2201, 1749]. relative-state [2067]. Relativistic [1557, 1578, 971, 2243, 1185, 975, 311, 1751, 475, 232, 2271, 1756, 1889, 2224, 1109, 467, 852, 1299, 1560, 1186, 644, 1400, 1778, 1779, 1154, 1521, 691, 1352, 1767, 1931, 290, 1803, 2184, 268, 1263, 1064, 1016, 1965, 1281, 245, 1898, 1065, 1823, 339, 815, 892, 397, 982, 942, 1462, 1947, 606]. relativity [1089, 1614, 2046, 689, 2088, 1875, 1712, 867, 75, 1581, 844, 416, 312, 1917, 868, 1282, 1284, 478, 2050, 1468]. relaxation [1575, 1040, 149, 646]. relaxing [1944]. relics [50]. remark [122]. Remarks [808, 421, 929, 611, 446, 419, 1633]. Removal [456]. removing [1896]. Renormalization [1496, 2075, 1346, 984]. Renormalized [289, 1063, 1415]. renormgroup [1443]. reparametrization [1077]. repellers [1793]. replica [2081]. Reply [701, 700, 85]. Representation [377, 1523, 1405, 340, 1037, 56, 2063, 1480, 2067, 1707, 636, 1594, 781, 655, 1228, 2097, 501, 1187, 1509, 405, 715, 2246, 65, 1704, 963, 1982, 580, 2064, 1725, 258, 1368, 325, 2057, 101, 946, 1547]. Representation-theoretic [1523]. Representations [78, 1508, 1954, 1211, 1101, 2056, 1005, 283, 651, 927, 1244, 2212, 1286, 185, 777, 1469, 1877, 873, 1617, 653, 1139, 223, 973, 1096, 234, 559, 79, 1140, 118, 926, 135, 190, 1618, 321, 1135, 548, 703, 528, 1682, 543, 2173, 2021, 1212, 280, 370, 422, 50, 2023, 282, 1116, 1065, 1967, 324, 1555, 1723, 372, 1744, 1921, 1957, 2264, 707, 2222, 1773, 1308, 2244, 463, 2134, 1004, 1407, 880, 1103, 1534, 1104, 968, 2223]. reprint [837]. reproducing [1924]. reptation [587]. repulsive [1491]. Res [837]. Reshetikhin [428, 326]. Resolution [1778, 125, 1727, 1728]. resolutions [421]. resolvent [456]. resolvent-like [456]. resonance [1022, 1049]. resonances [821, 1522, 885, 1526]. resonant [2125, 818, 641, 1218, 1486, 2266]. respect [1678, 1229]. respecting [2029]. Response [927, 86, 2269, 2102, 2062, 2177, 1730]. Rest [1804, 2225]. restorer [173, 1721]. restricted [1547]. Restricting [1564]. restriction [1773]. result [760]. results [855, 2206, 1670, 97, 1055, 631, 1027, 240, 515, 2038, 848, 849]. retarded [1729, 1312, 1730]. reversal [1049, 1050]. reverse [1925]. reversible [1545]. review [170, 1055]. revisited [1650, 1112, 379, 622, 459, 566]. revolution [156]. Rham [1347, 1565]. Ribbon [1002]. Riccati [1297, 600]. Ricci [568, 567, 11, 765, 999, 411, 214, 760, 1899, 114, 41]. Ricci-flat [760, 1899]. ridges [990]. Riemann [521, 1949, 575, 1135, 251, 639, 123, 330, 565, 1475, 2211, 1078, 882]. Riemannian [1410, 765, 1919, 1141, 1352, 760, 40, 587, 2201, 1515, 1899, 41, 2257, 1502, 825]. Rigged [879]. right [801, 118, 508]. rigid [624, 1844, 1971, 2000, 1346, 407]. rigidity [881]. rigidly [1756, 2226]. Rigorous [899, 1829, 1984, 499, 1916, 1822, 1055, 1027, 240, 515]. Rindler [1118]. rings [1364, 1869]. Robertson [1759, 2165, 2166, 554, 727]. Robin [1620]. Robinson [1500]. Rodrigues [698]. Rodriguez [698]. rods [2036]. role [1414, 1841, 1367, 983]. root [1139, 1563, 1035, 2158, 2133]. roots [1096]. Rössler [70]. Rotating [1393, 1394, 1501, 1756, 1013, 776, 1575, 12, 910, 1282, 1169]. rotation [1469, 1855, 243, 628, 1278, 1741]. Rotation-observable [1741]. rotational [1862, 1161]. rotationally [542]. rotor [1160, 1454]. Rouet [1732, 840, 841]. rough [1425]. rowed [1407]. Rudin [1044]. Ruijsenaars [1311, 1770, 1238]. rule [1152, 1335, 121, 1405]. rules [452, 2183, 1963, 457, 1727, 1773, 812, 925, 1409].

S [927, 1279]. S-2 [1279]. S. [926]. Sabatier [1562]. Sachs [211, 113]. Salpeter [2224, 1507, 1889]. same [1576]. sandwich [1977]. Sasa [1942]. Satake [1687]. satisfying [1402, 1955]. Sato [313, 2205]. Satsuma [1129, 1942]. Scalar [358, 2105, 1279, 2164, 2013, 1280, 1186, 726, 725, 1896, 545, 908, 292, 1168, 1918, 1910, 797, 340, 650, 242, 2130, 481, 1805, 212]. scalar-valued [340]. scalars [40]. Scale [2089, 1541]. scaled [464]. scales [899, 2245, 1047]. scaling [1890, 692, 1705, 1936]. scatterers [1, 837, 2]. Scattering [1524, 189, 935, 195, 1113, 891, 198, 755, 141, 205, 883, 1490, 2065, 1522, 1108, 1479, 2148, 2118, 249, 542, 1319, 45, 90, 30, 1425, 1187, 1562, 288, 820, 2150, 384, 1049, 1073, 2203, 1637, 752, 945, 948, 719, 23, 937, 1934, 807, 455, 203, 402, 1495, 1020, 2145, 942, 99, 2151, 462, 983, 1256, 1829, 1071, 610, 791]. scheme [1598, 493, 930, 1369, 1814]. schemes [1172]. Schild [1548, 870, 416, 1435]. Schlesinger [109, 1546]. Schmidt [2095, 421, 1145]. Schneider [1311, 1238]. Schouten [1100, 1174]. Schrödinger [1044, 2104, 883, 1642, 2065, 2178, 167, 331, 1108, 363, 18, 934, 33, 2181, 2007, 1527, 91, 1821, 484, 1941, 1111, 2030, 1154, 1521, 691, 1328, 548, 2008, 252, 1260, 1376, 306, 1512, 891, 2189, 199, 784, 108, 1380, 1823, 423, 1194, 355, 1157, 2034, 1499, 1195, 1901, 309, 1306, 2076, 1256, 227, 1217, 1699, 393, 1975, 1700, 1106, 394, 709, 1007, 1792, 2239]. Schrödinger-like [1380]. Schrödinger-type [883, 1256]. Schur [1844]. Schwarz [881]. Schwarzschild [2101, 2102, 928, 552, 1945, 685, 1762, 2209, 869]. Schweinler [2058]. Schwinger [377, 1929, 1340, 479, 988]. screening [1313, 515]. sea [1373]. search [1950, 148]. Second [1610, 1823, 2054, 151, 856, 1436, 2263, 1273, 1388, 1554, 219]. second-degree [2054]. Second-order [1823, 2054, 2263]. section [366, 2151]. sections [1396]. sector [2043, 435, 1674, 1882, 1918, 1458]. sectors [1010, 14, 87, 575, 1783]. secular [1298, 1612]. seed [1319]. Segal [1517, 1924, 1925]. segment [1031, 738]. Segrè [1206]. Segur [2119]. Seiberg [192, 745]. Self [751, 1343, 1168, 2094, 665, 1976, 1640, 1249, 69, 741, 1796, 574, 769, 1931, 476, 1377, 1918, 1144, 1180, 1285, 261, 1431, 988]. self-adjointness [1796, 1144]. self-consistent [1285]. Self-dual [751, 1343, 665, 1249, 69, 741, 769, 476, 1377, 1918, 1180, 1431]. Self-duality [2094]. self-energy [988]. self-generated [1931]. Self-gravitating [1168]. self-similar [1640]. self-similarity [1976]. semi [344, 1672, 2257]. semi-discretization [1672]. semi-infinite [344]. semi-Riemannian [2257]. Semiclassical [452, 156, 227, 1217, 393, 1926, 57, 1341, 289, 236, 706, 1492, 2238]. semiconductor [886, 816, 1662, 1235]. semiconductors [713, 1541]. Semidirect [1482, 1625, 1065]. semigroup [187]. semigroups [1903]. semilinear [2123]. Semilocal [536]. semispin [2144]. Separability [1629, 260, 1572, 2202, 807, 1750, 479, 2124]. separable [386, 1214, 1792, 2239]. separate [1070]. separating [695]. Separation [727, 955, 2249, 1007, 1362, 1920, 2247, 394, 1547]. Serban [1374]. series [768, 1298, 853, 289, 2245, 657, 707, 1486, 2266, 2259, 711]. set [2253, 1949, 942, 1906]. sets [574, 1955, 1664, 1638, 1997]. Setting [1659]. several [1990, 2002]. Shaarawi [1147]. Shabat [592, 723]. Shadow [1413]. Shannon [1198]. shape [156]. Shapiro [1044]. sharing [1576]. Shear [1318, 206]. Shear-free [1318, 206]. shearing [4, 1775]. sheeted [343]. shells [267, 216, 868]. shift [606, 2195, 160]. shifts [1049]. Shiota [356]. shocks [1803]. short [2000, 1937]. Sigma [1743, 31, 2039, 573, 757]. sign [1681]. sign-invariants [1681]. signals [1632]. signature [864, 2026]. Signs [1233]. similar [1640]. Similarity [625, 1871, 1460, 1976, 1571]. Simons [1819, 747, 168, 2180, 503, 536, 1343, 2072, 1697, 1382, 1383, 788]. Simple [690, 1839, 428, 203, 1724, 1496, 1516, 490, 925]. simplest [1499, 748]. Simplicial [1801, 436, 439]. simply [185]. simulating [1654]. simulations [439]. Simultaneous [2162]. sine [404, 354, 1133]. sine-Gordon [404, 354, 1133]. Singer [123, 330]. single [602, 287, 21, 866, 2032]. single-particle [21]. Singular [620, 834, 1523, 1288, 189, 935, 395, 1731, 1419, 622, 1069, 1127, 1018]. Singularities [1882, 826, 789, 1843, 1803, 867, 1581, 809, 2016, 1158, 547]. Singularity [1284, 37, 1209, 685, 526]. Singularity-free [1284]. singularly [1098]. Sitter [2045, 1246, 976, 1786, 909, 892]. Sivashinsky [106, 1081]. Six [218, 1153, 1197, 425, 1753]. Six-parameter [218]. six-vertex [1153, 1753]. six-vertex-type [425]. size [1705]. skew [2043]. Sklyanin [915, 82]. Skyrme [450, 665, 2232, 1122]. Skyrme-like [450]. skyrmion [322, 1891]. skyrmions [2241]. slab [1105, 465]. Slater [1992]. slicing [2209]. slow [1468]. slowly [1013, 1575]. small [857, 30, 899, 195, 1212]. small-angle [30]. smoke [1869]. smooth [302, 1437, 1527, 753, 1718]. Snapback [1793]. Snyder [2049]. Sobolev [1144, 1925]. Sokolov [1667]. Solitary [1787, 469, 519, 2191, 255]. solitary-wave [469, 2191]. solitinos [900]. Soliton [1942, 719, 961, 592, 2080, 144, 1707, 1574, 484, 2121, 1237, 1944, 2040, 649, 2249, 1432]. soliton-like [1944]. solitons [1199, 593, 1201, 1448, 1343, 2229, 306, 255, 1914, 473]. soluble [906, 1527, 1530]. Solution [2118, 795, 3, 736, 350, 506, 1121, 589, 2147, 1296, 2171, 302, 513, 1889, 2224, 655, 1470, 1340, 1892, 484, 1569, 384, 2101, 252, 379, 469, 113, 1916, 784, 1634, 1859, 1762, 2102, 1381, 218, 299, 1157, 1367, 679, 2002, 869, 242, 1106, 491, 1086, 1235]. Solutions [43, 1400, 136, 1838, 1090, 1535, 1621, 619, 2006, 450, 1232, 1939, 68, 318, 276, 1571, 2126, 2127, 678, 1437, 152, 2045, 1314, 1719, 1398, 303, 38, 1125, 2176, 1493, 39, 1313, 92, 620, 690, 1353, 1911, 102, 1328, 1596, 2240, 1129, 2009, 1942, 1712, 1803, 598, 1248, 625, 158, 2074, 2108, 900, 719, 2232, 1468, 2250, 196, 305, 806, 919, 2022, 1610, 761, 1137, 1092, 1393, 1394, 1501, 2078, 1612, 952, 1276, 525, 1530, 1973, 745, 904, 313, 1094, 255, 256, 1709, 316, 1870, 2177, 258, 790, 526, 2210, 2010, 1285, 798, 317, 1360, 1591]. solutions [425, 747, 799, 764, 1592, 592, 2130, 988, 1432, 2049]. solvability [1977, 2158].

Solvable [583, 780, 498, 319, 349, 613, 989, 1702, 2106, 756, 1111, 1729, 2109, 1026, 1589, 1740, 2188, 455, 2236, 1193, 423, 1180, 1312, 1730, 610]. solve [175]. Solving [1200, 1811, 878, 2172, 2034, 1078]. Some [57, 69, 368, 771, 419, 1670, 1633, 1027, 240, 2222, 276, 724, 152, 853, 1560, 446, 559, 1778, 334, 1090, 428, 97, 75, 1565, 1849, 1774, 515, 28, 85, 373, 1873]. sound [2125]. source [513, 544, 543, 903, 1076, 2197]. sourceless [2130]. sources [2115, 348, 2267, 1378, 1316, 799, 1592]. Space [215, 216, 1243, 1577, 521, 651, 927, 1087, 872, 2095, 1558, 56, 57, 777, 381, 1480, 2067, 37, 1318, 18, 1927, 624, 410, 1976, 907, 276, 1571, 1758, 2164, 1984, 724, 1167, 2013, 998, 887, 912, 155, 739, 2106, 874, 2215, 2086, 1573, 2180, 2007, 1246, 320, 875, 1125, 688, 730, 134, 60, 284, 863, 1711, 751, 782, 1808, 1358, 411, 1580, 928, 552, 956, 2087, 333, 575, 321, 214, 1879, 1843, 691, 975, 93, 1988, 976, 1549, 1596, 1945, 1767, 877, 890, 1803, 693, 625, 413, 1915, 420, 1615, 2090, 2048, 695, 1761, 1473, 1963, 1840]. space [607, 2110, 280, 919, 1858, 414, 2141, 1513, 761, 1956, 6, 166, 200, 1786, 271, 396, 65, 1395, 1252, 909, 1529, 130, 963, 1982, 626, 1804, 2225, 1195, 510, 892, 1021, 1717, 1396, 627, 1206, 762, 677, 1207, 1502, 1430, 699, 1924, 1925, 811, 1677, 880, 327, 28, 987, 1752, 2042, 1324, 554, 727, 394, 266, 475, 2003]. space- [1927]. spaces [167, 233, 654, 366, 2045, 1188, 926, 1539, 334, 336, 268, 1442, 1635, 1075, 1275, 2100, 961, 921, 2201, 159, 323, 239, 423, 1555, 123, 330, 98, 1144, 1225, 1910, 1516, 126, 879, 520, 533, 241, 2153]. Spacetime [1392, 2016, 1268]. spacetimes [11, 568, 1548, 567, 2126, 2127, 2209, 1435]. span [1793]. spanned [1393, 1394]. spanning [365]. spannors [204]. Spatial [1, 837, 2, 1048, 366, 1614, 1396, 1938, 141, 2199]. Spatiality [275]. spatially [2122]. Special [892, 1039, 644, 689, 2207, 1115, 1871, 2057, 1623, 663, 1103, 1000]. Special-relativistic [892]. specializations [1923]. species [1658, 1998]. specified [1031, 1333]. Specker [846]. Specker-type [846]. spectra [1733, 13, 63, 834, 124, 1904, 826]. Spectral [1201, 1448, 833, 1047, 1040, 1661, 828, 1820, 484, 1895, 1989, 2005, 1771, 309, 2195, 373]. Spectrum [1404, 2095, 1058, 346, 347, 2157, 332, 543, 921, 25, 581, 1385, 1069, 709]. speed [205, 710]. Spencer [1601]. sphere [1031, 1627, 1095, 1288, 1210, 1511, 772, 1113, 2109, 2021, 1920, 919, 1749, 2160, 72, 811, 393, 343]. spheres [29, 408, 14, 87, 702, 1113, 1284]. spherical [1318, 1271, 776, 2089, 5, 1474, 1444, 2238, 294]. Spherically [154, 1205, 2232, 1087, 11, 568, 1166, 1976, 567, 43, 724, 1167, 523, 16, 814, 39, 1358, 269, 2250, 270]. spiked [1782]. Spin [1835, 2077, 449, 31, 778, 737, 1578, 410, 1764, 1984, 1985, 1579, 714, 1301, 1111, 214, 1154, 1521, 976, 1695, 2261, 1570, 1261, 506, 578, 1698, 1859, 1899, 1499, 581, 391, 1238, 1947, 325, 1676, 1120, 1255, 1070, 71, 1621, 727, 1995, 1817]. spin- [214, 976, 391, 1070]. Spin-mass [2077]. spin-models [1255]. spin-one [449]. spin-torsion [1579]. spinless [1507, 1889, 2224, 339, 815]. Spinning [1197, 131, 535, 1961]. Spinor [533, 569, 1948, 873, 1511, 113, 271, 1333, 1285, 753]. spinors [200, 2128, 1515, 479]. Split [1875]. Splitting [1416, 1279, 1351, 91, 145, 2075, 986]. Spontaneous [603]. spontaneously [667, 193]. Square [1641, 365, 1563, 121, 631, 2133, 1825, 1624]. square-integrability [1624]. squared [518, 1973]. squeezed [940, 941]. Squeezing [783, 1417, 1791]. stability [128, 2165, 2166, 1597, 761, 787, 2080, 1749]. stabilization [558]. Stable [1581, 263, 1232, 1872, 1854]. Stäckel [1075, 1836]. staircase [1664]. Standard [616, 1533, 1012, 435, 1097, 617, 1290, 531, 1215, 1068, 748]. star [911, 1399, 1595, 270, 51, 525]. Stark [1335, 1526, 1152, 332]. starting [2135]. state [1480, 2067, 20, 1766, 2262, 781, 852, 21, 1221, 915, 1990, 337, 744, 1982, 1024, 160, 161, 1828]. states [1057, 1642, 232, 2271, 331, 2214, 1056, 779, 1853, 131, 19, 1818, 571, 1259, 1929, 2272, 452, 972, 1186, 446, 2029, 820, 1778, 1986, 333, 1737, 1061, 1190, 537, 888, 1112, 2107, 236, 1455, 1780, 1781, 1210, 2031, 694, 1261, 1482, 2231, 578, 2155, 2032, 137, 282, 819, 1156, 1055, 1048, 1194, 1862, 940, 941, 1266, 942, 1969, 671, 707, 1226, 143, 1788, 1789, 1624, 343, 1827, 1350, 1791]. Static [1712, 1917, 1582, 11, 568, 567, 2164, 724, 1167, 474, 206, 411, 93, 1284, 1094]. Stationary [721, 1282, 1093, 1435, 106, 2157, 535, 362, 1135, 1767, 1761, 1091, 4, 1775, 588, 1805, 2003]. Statistical [1816, 29, 1125, 401, 586, 587, 454, 1690, 1124, 1015, 754, 715, 67, 1604, 1254]. Statistics [823, 1049, 31, 828, 617, 616, 1528, 831, 819]. Status [1183, 1219, 1242]. stellar [346, 347]. Stephani [115]. steplike [2178]. Stewartson [1554, 308, 2083, 1085]. Stewartson-type [2083]. Stochastic [550, 1704, 1861, 1558, 1245, 1880, 1465, 992, 1457, 1466, 339, 815, 466, 2194, 1829, 517, 944]. Stochastically [1780, 1301]. Stockum [627]. Stokes [1162, 1125, 2218, 1633]. Stora [1732, 840, 841]. straight [271]. straightforward [784, 1106]. stratified [750, 1351, 1850]. Stress [991, 990, 2075, 1485, 584]. Stress-energy-momentum [991]. stress-tensor [2075]. stretched [646]. Strict [1785]. String [1847, 2171, 738, 1109, 1890, 1302, 1222, 2089, 2139, 271, 2144, 1346, 313, 661]. stringlike [1712]. strings [409, 989, 1320, 1016, 1698]. strip [1073, 951, 994]. Strong [1561, 186, 1298]. strongly [1892, 1473]. Structure [563, 292, 2006, 482, 381, 1166, 958, 1585, 106, 1643, 319, 1691, 1979, 1866, 947, 2014, 1209, 1097, 1647, 1808, 1001, 769, 889, 890, 693, 867, 1769, 1544, 806, 1426, 2141, 1686, 1847, 97, 1972, 732, 531, 1048, 402, 1727, 99, 1293, 1676, 261, 2161, 1556, 880, 1268, 388, 895, 2042]. Structured [268]. structures [400, 629, 1538, 1601, 32, 1572, 682, 735, 858, 1553, 1302, 486, 1780, 1875, 1768, 1609, 1008, 1610, 647, 323, 1703, 257, 953, 2252, 2265, 1132, 1851, 1361, 2059, 2011, 533, 1174, 534]. Stückelberg [675]. Stückelberg-Kibble [675]. Study [1541, 1024, 1834, 1914]. Sturm [2022, 1973]. Sturmian [1965]. Sturmians [1269]. sub [1352]. sub-Riemannian [1352]. subalgebra [836, 96, 812]. Subalgebras [911, 1678, 2098, 1846, 925]. subduction [1036]. subdynamics [1540]. subgroup [1773]. subgroups [603, 46, 200]. Subject [495, 932, 1371]. subjected [1022, 1127]. Submanifolds [659, 1978, 1291, 2257]. Submitting [1257, 1295]. subspaces [1980, 456]. substitutions [1355]. successive [680]. sufficient [1259, 2272]. sum [1152, 1335, 452, 1992, 225, 1405, 1812, 2258, 458]. summability [489]. Summary [493, 494, 930, 931, 1369, 1370, 1814]. Summation [768, 1692, 1650]. Summational [1123]. summations [8, 374]. sums [2173, 631]. Super [2013, 1939, 32, 1451, 590, 1009, 1588, 1684, 722, 1671, 2011, 84, 283, 1174]. super-bialgebras [1684]. Super-energy [2013]. super-KdV [2011]. super-KP [32]. super-Lax [32]. super-Poisson [1174]. superalgebra [555, 1735, 960, 422, 81, 1099, 1921, 722, 55, 774, 2223]. superalgebras [380, 913, 286, 656, 562, 2020, 696, 122, 530, 2251, 83, 1688, 968, 2175, 881]. superanalog [1001]. supercharacter [1286]. supercoherent [1226]. supercomplex [1213]. superconductive [1053]. superconductivity [1450]. superconformal [1618, 546]. superconformally [1588]. Superconnections [1994]. superdense [525]. superdifferential [722]. superdimension [530]. superelliptic [1608]. superfield [292]. supergroup [729, 696, 2023, 1182]. supergroups [2175]. Superintegrability [919, 1292, 1858, 1964, 1833]. Superintegrable [1203, 2160]. superization [488]. supermanifold [197]. supermanifolds [2260]. supermatrices [917, 529]. supermomentum [2088]. superparticles [1961]. superplane [1586, 2216]. Superposition [474, 600, 2019, 1670, 2221]. superselection [1599, 1242]. supersolitons [900]. superspace [1732, 1996]. superspaces [906, 1735, 590, 1009, 157, 825]. superstring [1265, 2145]. Supersymmetric [2006, 1667, 839, 33, 781, 896, 501, 1343, 900, 453, 353, 1936, 2093, 2237, 2025, 1347, 1565, 36, 101]. supersymmetries [1901]. Supersymmetry [1226, 1032, 56, 2226, 59, 1560, 1894, 1961, 2110, 130, 1860, 182, 1422, 1488, 572]. supertrace [562]. Supertraces [684]. support [481]. supported [1113]. Surface [737, 1425, 575, 1564, 1883, 1829, 596]. Surface-embeddability [737]. surfaces [569, 2148, 2039, 45, 1208, 442, 360, 361, 359, 1885, 1843, 1887, 1888, 1886, 156, 1241, 437, 938, 586, 587, 639, 961, 1028, 1611, 1309, 1937]. Sutherland [1451, 1544, 1598]. Sutherland-type [1598]. Suzuki [970]. symbol [1156]. Symbolic [1631, 1630]. symbols [1619, 1844, 923, 81, 1099, 1812, 2258]. Symes [1274, 1598]. symmetric [750, 1087, 11, 568, 154, 1166, 2136, 1960, 1976, 567, 886, 43, 724, 1167, 1949, 542, 1573, 1319, 523, 730, 16, 814, 39, 1205, 1358, 411, 235, 769, 1930, 1135, 195, 1954, 269, 2232, 1840, 2250, 1275, 270, 1094, 2081, 762, 219, 798, 241, 825]. symmetrical [243, 1160]. Symmetries [1150, 1758, 1832, 362, 1998, 49, 860, 2174, 99, 1745, 1750, 985, 262, 304, 1577, 150, 652, 1159, 1878, 1864, 467, 1452, 485, 224, 1328, 2089, 1932, 120, 1224, 291, 1377, 1443, 1366, 591, 658, 1972, 389, 1899, 1021, 1236, 1771, 259, 1833, 325, 399, 165, 1324, 648]. Symmetrizable [968]. symmetrized [655]. Symmetry [1912, 894, 1591, 1789, 1489, 497, 2006, 450, 1199, 1011, 655, 1340, 603, 1580, 1779, 193, 1844, 1800, 468, 1915, 95, 1761, 1303, 476, 1291, 1092, 1393, 1394, 1501, 50, 843, 455, 1859, 785, 254, 1354, 2062, 2061, 218, 147, 1345, 940, 941, 2133, 564, 1993, 1050, 261, 100, 2114, 709, 1813]. symmetry-adapted [655]. symmetry-breaking [1340]. Symplectic [1643, 2037, 1714, 794, 2196, 100, 1033, 1595, 1176, 119, 514, 725, 96, 1114, 75, 1067, 966, 675, 649]. Symplectic-energy-momentum [2037]. Symplecton [2212]. synchronizable [1549]. System [1183, 1219, 150, 1524, 460, 2039, 1437, 349, 683, 780, 2082, 1273, 89, 759, 969, 1543, 1128, 21, 362, 2240, 2101, 1330, 1839, 2103, 1943, 379, 539, 1026, 859, 245, 1898, 1762, 2102, 1530, 1804, 2225, 1266, 258, 259, 1833, 526, 1545, 260, 1132, 628, 1914, 813, 1463, 27, 2041, 1805, 212]. systems [1041, 331, 2179, 2170, 851, 1523, 1479, 1777, 778, 2116, 2226, 884, 498, 1233, 681, 319, 1873, 822, 682, 1766, 2262, 1950, 1314, 1831, 1864, 840, 841, 69, 858, 1574, 1796, 523, 947, 2119, 1832, 500, 1034, 690, 1747, 1453, 206, 287, 1692, 190, 823, 1778, 1608, 974, 1553, 333, 1389, 214, 1049, 405, 412, 1129, 2120, 715, 1998, 621, 1497, 335, 336, 1130, 1769, 516, 146, 2187, 2158, 1165, 901, 640, 1083, 1270, 902, 108, 860, 507, 25, 1381, 1202, 178, 1630, 1429, 792, 1499, 940, 941, 1876, 2025, 1903, 1084, 470, 141, 1576, 259]. systems [471, 1203, 2160, 2242, 829, 1834, 2210, 1003, 1531, 1196, 1566, 673, 1070, 905, 879, 2204, 1621, 1836, 241, 149, 646, 1085, 1391, 1837, 2156, 2124, 304, 717, 1424, 110]. syzygies [1949].

T [1256]. tableaux [1002]. Takhtajan [326]. Tamburino [2207]. Taming [1046]. tangent [1452]. Tau [993, 1639, 353, 36, 1707, 1973, 2256, 262]. Tau-function [353]. Tau-functions [993]. Tau-wavelets [1639]. Tauber [1467]. Taylor [1646]. technique [750, 18, 2172, 2235]. Techniques [1816, 169, 170, 1351, 178, 1884]. teleparallel [1917]. Teller [1487]. temperature [1593, 300, 2154, 1604, 1055, 787, 1066, 2191, 147, 898]. temperatures [854, 104]. temporal [1412]. ten [292]. Tensor [1289, 1948, 872, 2213, 144, 1500, 1949, 2013, 688, 800, 702, 153, 375, 920, 159, 2075, 584, 733, 1993, 1446, 1397, 565, 1475, 2211, 1102]. Tensoring [1212]. tensors [2088, 114, 41, 1724, 160, 660, 219, 991]. terms [377, 1760, 446, 1738, 437, 807, 272, 512]. test [779, 1930, 1442]. test-function [1442]. testing [2032]. Tests [33]. tetrad [870, 1673, 1333, 142]. tetrahedral [1719, 1398]. tetrahedron [1164]. their [1948, 2115, 1167, 559, 411, 1843, 1061, 486, 2267, 268, 544, 1028, 282, 1380, 1499, 1206, 1771, 1709, 161, 1688, 965, 699, 1791, 2249, 968, 825]. them [1851]. theorem [221, 331, 1616, 624, 1525, 1746, 730, 803, 535, 1736, 2150, 1802, 1769, 2218, 891, 1378, 807, 1810, 2192, 293, 943, 2257, 677, 650, 894, 226, 1532, 1937, 1624, 1827, 1700, 111, 1701, 1198]. theorems [1506, 1873, 1453, 1675, 417, 1384, 2153]. theoretic [1523]. theoretical [56, 1389, 2107, 1570, 726, 1674, 579, 572]. theories [2135, 169, 444, 381, 1449, 1124, 170, 1410, 430, 740, 1659, 7, 229, 1189, 1302, 503, 1112, 174, 311, 1856, 2230, 1875, 916, 2139, 2233, 454, 1137, 1136, 1281, 1155, 1698, 239, 1019, 1675, 1294, 980, 2025, 240, 1745, 2122, 219, 991, 1422, 1699, 2104, 929, 611, 1863, 895, 1310]. Theory [2202, 1465, 521, 2105, 168, 400, 1852, 449, 1149, 185, 2115, 1296, 426, 128, 427, 1538, 2043, 1601, 431, 409, 2136, 2028, 1842, 1734, 1185, 499, 1691, 2086, 2137, 1819, 741, 667, 972, 1820, 1579, 1647, 603, 191, 1987, 2027, 361, 1888, 1343, 2245, 889, 1738, 1874, 1767, 992, 2072, 94, 2267, 1652, 1113, 237, 545, 387, 1040, 1263, 1064, 938, 177, 1017, 742, 1513, 761, 1305, 307, 198, 1934, 1599, 1697, 238, 588, 743, 843, 1379, 1418, 807, 1076, 1382, 1383, 2191, 1823, 1630, 2144, 140, 1117, 1495, 492, 273, 417, 793, 1968, 2145, 892, 293]. theory [893, 1119, 1936, 2093, 2237, 1496, 1492, 832, 1094, 788, 942, 99, 478, 314, 462, 141, 733, 1993, 1386, 447, 160, 125, 1368, 1503, 1970, 628, 661, 1120, 898, 429, 1906, 894, 1050, 798, 443, 1532, 905, 1488, 1145, 597, 747, 1830, 1476, 1533, 610, 2080, 1996, 1387, 791, 2114, 825, 1051, 1959, 1235]. there [1854, 1802]. Thermal [1575, 1264, 1853, 617, 616, 2233, 1040, 148]. Thermalization [1818]. Thermodynamic [1657, 2200, 1373, 852, 2244, 2156]. thermodynamical [1389, 1092]. Thermodynamics [115, 1661, 1163, 1273, 1866, 1388]. thermoelasticity [1909]. thermostatistics [1794, 1958]. these [241]. thin [1977, 868, 462]. thin-sandwich [1977]. third [1554, 253]. Thirring [635, 1746, 518]. Thomas [133]. Three [765, 999, 945, 1444, 2003, 449, 1149, 151, 1373, 59, 680, 1437, 1766, 2262, 2045, 874, 34, 1573, 350, 1201, 1448, 1464, 1778, 333, 691, 2245, 195, 1915, 297, 545, 506, 23, 1858, 1714, 1501, 1898, 371, 114, 41, 1496, 391, 259, 673, 1407, 27, 2041, 1432]. three-body [23]. three-cluster [333]. Three-dimensional [765, 999, 945, 1444, 2003, 1149, 151, 59, 680, 1437, 2045, 874, 34, 1573, 195, 1915, 297, 1858, 1501, 1898, 391, 27]. three-electron [1778]. three-level [673]. three-magnon [506]. three-metrics [114, 41]. three-particle [1496, 259]. three-rowed [1407]. three-wave [1373, 1201, 1448, 2041, 1432]. threshold [885, 891]. thresholds [789, 1649]. Tikekar [525]. tilings [1043]. Time [2150, 1803, 1637, 544, 543, 2140, 2078, 108, 1602, 1630, 1531, 183, 1050, 37, 1631, 1233, 187, 855, 856, 2007, 320, 875, 1453, 782, 1808, 1358, 928, 552, 956, 2030, 576, 1629, 691, 975, 1988, 1049, 2245, 703, 1596, 1908, 1142, 297, 1615, 2090, 2048, 695, 1483, 2110, 414, 2141, 469, 2268, 1897, 2269, 1722, 6, 166, 215, 216, 271, 1822, 396, 1381, 1395, 178, 1243, 626, 2034, 941, 627, 762, 677, 1833, 641, 1218, 1537, 2198, 2129, 327, 1623, 1079, 1324, 554, 727, 394, 266, 335, 940]. Time-asymptotic [2078]. Time-dependent [1637, 544, 543, 108, 1602, 1531, 855, 856, 2030, 1908, 1142, 1381, 2034, 941, 1833, 1537, 394, 940]. Time-frequency [1630, 1631, 1623]. Time-independent [1803]. time-ordered [1722]. Time-reversal [1050, 1049]. time-varying [1630]. Timelike [1756, 1800, 1927]. Times [1396, 1577, 521, 1087, 1318, 624, 1976, 907, 1758, 2164, 724, 1167, 2013, 2086, 863, 1711, 751, 411, 2087, 214, 1879, 93, 1549, 1945, 1767, 890, 1803, 413, 1915, 1761, 1473, 1840, 1529, 1804, 2225, 1717, 1206, 1207, 1502, 699, 1677, 28, 987, 1752, 475, 2003]. tip [1882]. Toda [589, 1199, 1707, 1666, 2070, 1034, 136, 412, 949, 995, 793, 471, 1487, 473, 110]. Toda-like [412]. Toda-related [471]. Tomimatsu [313]. tomography [789, 712]. tools [1516]. top [1806, 243, 946]. Topological [31, 357, 430, 575, 866, 237, 545, 1043, 1019, 987, 427, 1763, 431, 170, 570, 1188, 1647, 1808, 889, 1223, 936, 453, 938, 843, 1382, 42, 1119, 2252, 2265, 429, 443, 1422, 1830]. topologically [555, 667]. Topology [244, 322, 264, 357, 863, 1459]. toral [939]. tori [91, 2245, 201, 1067]. Toroidal [2024, 1170, 1617]. Torsion [415, 2096, 1579, 1392, 996, 2092, 42, 123, 330]. torus [2072, 1674, 630, 1484]. total [1971, 456]. toy [796]. TQFTs [458]. trace [1628, 1905, 1772]. traces [1604]. tracial [1308]. tractability [633, 2273]. trajectories [1747, 2206, 1857, 146, 861]. trajectory [1628]. transcendents [351]. Transfer [1910, 1045, 23, 1630, 1704, 99, 149]. transform [1842, 2215, 484, 1201, 1448, 804, 805, 948, 2022, 1474, 354, 489, 309, 1517, 1447, 1524]. Transformation [1325, 102, 191, 1365, 746, 1307, 672, 383, 1542, 1951, 61, 1607, 1911, 420, 1670, 2221, 1381, 1726, 1430, 1603, 472, 723]. Transformations [1473, 1707, 1151, 1173, 1200, 1614, 1821, 993, 2030, 657, 1330, 1955, 783, 1263, 1589, 210, 1157, 109, 2054, 126, 1317, 1671, 459, 399, 1239, 1431, 2259, 1357, 1546]. transformed [1634]. transforms [1625, 1626, 772, 1710, 749]. Transient [643, 874]. Transition [1754, 1788, 1680, 804, 246, 1161, 1051, 1959]. transitions [1465, 401, 1705]. Transitive [514, 2087]. translational [624]. translations [1046]. Transmission [886]. Transport [2149, 818, 1041, 821, 680, 822, 921, 588, 1466, 148, 1042]. trapped [1028]. traveling [2078, 1748]. traveling-wave [2078, 1748]. travelling [1248]. treatment [1138, 499, 89, 1778, 1742, 1810]. tree [385]. trees [1636, 365]. trial [20]. triangular [1725]. triangulated [440]. triangulations [438, 1090]. triaxial [1454]. trigonometric [368, 771, 2158, 1923]. trigonometric-type [2158]. trimerized [1914]. Trinomial [1981]. triple [1812, 2258]. triples [1771]. triplets [1786]. trivial [1865, 141]. truncation [657]. Tsallis [1198]. Tuning [217]. Tunneling [2033, 1695, 818]. tunnelling [2261]. Turaev [428]. Turbiner [1530]. turbulence [954]. turbulent [899]. turning [549]. twist [1746, 71]. Twisted [1041, 1336, 555, 738, 44, 1153, 1723]. Twisting [1290, 1725, 2207, 626]. twistor [612, 157, 354]. twistor-inspired [354]. twistors [157]. twists [2132]. Two [1732, 729, 2216, 2017, 2052, 501, 782, 691, 2233, 1934, 1393, 1380, 1703, 203, 2193, 1655, 1345, 555, 1523, 1311, 1296, 738, 1249, 1949, 958, 77, 1831, 1542, 1228, 320, 875, 1493, 1543, 2119, 1832, 1186, 1425, 1561, 559, 1454, 1987, 2027, 1998, 562, 890, 1191, 1415, 1729, 545, 2021, 1263, 1064, 919, 1292, 1964, 996, 1714, 2022, 1513, 1122, 1980, 1491, 1787, 1496, 391, 2160, 1312, 1730, 1825, 260, 429, 511, 673, 2059, 393, 1518, 848, 748, 1350, 127, 394]. two- [1496, 391]. two-block [1543]. Two-body [782, 1380, 203, 1729, 1064, 1491, 1312, 1730, 848]. two-color [958]. Two-component [1655]. Two-dimensional [501, 1345, 555, 1523, 1311, 1228, 1425, 1561, 890, 2021, 1292, 1964, 996, 1513, 1980, 1787, 2160, 1825, 260, 429, 127]. Two-form [1732]. two-forms [1249, 875]. two-level [673]. two-mode [1350]. Two-parameter [2216, 2052]. Two-parametric [2017]. Two-particle [1934, 1186, 1263]. two-point [748]. two-sphere [393]. two-variable [562, 1518]. two-velocity [1493]. Type [562, 883, 186, 276, 1571, 2126, 2127, 2086, 350, 1375, 1390, 1081, 1987, 2027, 561, 1554, 656, 1946, 625, 253, 1289, 1598, 2158, 835, 1115, 721, 530, 1590, 2083, 338, 2054, 1093, 762, 1502, 526, 1545, 1905, 1285, 219, 326, 1256, 479, 425, 73, 846, 28, 1547]. Type-I [562, 1285]. types [2084, 2085, 1206]. Tyutin [2179, 1732, 1155]. Tzitzéica [1951].

U [1744]. ua [1501]. Uehling [1162]. Ueno [2256]. Uhlenbeck [1162]. ultra [1140]. ultra-commutation [1140]. ultralocal [296, 295]. ultrarelativistic [1030]. ultraspherical [279]. ultraviolet [1896]. unbounded [1763, 574]. Uncertain [1130]. uncertainties [637, 1017]. Uncertainty [2268, 1897, 2269, 1629]. underspread [1630]. Unification [142, 1821]. Unified [1742, 1138]. Uniform [1073, 738, 1890, 1596, 549, 76]. uniformization [1506]. uniformly [513]. unipotent [1002]. Uniqueness [2155, 596, 2153, 851, 1532]. Unitary [223, 79, 1417, 2021, 1381, 2040, 822, 1508, 47, 528, 51, 1065, 2192, 372]. unity [1139, 1096, 1995]. Universal [1202, 2020, 1262, 808, 1334, 1445, 1368, 1906, 490, 997, 534, 773]. Universality [1713, 646, 824, 174, 1255]. universe [2163, 1467, 2206, 1285]. universes [211, 115, 957, 1361]. Unobstructedness [1406]. unperturbed [76]. unsolvable [1211]. unstable [620, 263]. Unti [2207]. Upper [1337, 1029]. Uq [1405]. Urrutia [700, 700]. use [1810]. Used [1815]. uses [140]. using [1984, 857, 145, 981, 489, 1687, 1463].

V [2084, 2085]. Vacuum [1436, 2263, 1420, 409, 2126, 2127, 1691, 1711, 213, 1135, 1840, 4, 1775, 787, 272]. value [68, 1436, 2263, 468, 948, 1214, 1192, 2174, 1277, 246, 1306, 1620, 406, 1078, 1235]. valued [1892, 796, 454, 340]. values [569, 250, 1353, 2227, 2228, 1989, 1520]. Vandermonde [1239]. Vandermonde-like [1239]. vanishing [2013, 948, 2207, 1971]. Variable [253, 1362, 2227, 2228, 562, 1457, 2247, 1518, 394]. variables [907, 1301, 193, 1920, 1026, 148, 182, 1356, 955, 2249, 1007, 1547]. variants [1825]. variation [1, 837, 2]. Variational [2014, 405, 577, 720, 1429, 2044, 20, 133, 2037, 1306, 1145, 1463]. variationally [1865]. varieties [1002, 1677]. varying [1630, 511, 935]. Vassiliev [1697]. VCS [340]. Vector [1188, 228, 1852, 1523, 1464, 1414, 1988, 2018, 1511, 545, 553, 160, 1422, 533, 2153, 1624]. vector-gauge [1852]. vectors [1522, 1557, 1480, 2067, 624, 187, 1338, 2180, 1982]. velocity [750, 2243, 1493, 176, 547]. verification [2236]. Verma [2264, 1957]. version [780, 1746, 61, 1646]. versus [1852, 172, 1741, 148, 1825, 329]. Vertex [1707, 502, 1617, 1153, 1471, 1563, 1222, 1265, 1020, 1458, 425, 1753, 2223]. vertices [1947]. Veselov [2159]. VI [1953, 109, 2054]. via [168, 1524, 103, 131, 1702, 1096, 1229, 2182, 360, 361, 1887, 1888, 1650, 1654, 111, 2249]. vibration [1793]. view [1045, 898]. viewpoint [86, 1941, 1830, 85]. Vilkovisky [838, 1525, 1932, 742]. violation [2234]. Virasoro [84, 997]. viscous [1715]. Vlasov [1567, 2078, 526]. Vogt [86]. Volterra [150, 735, 1950, 1769, 1008, 859, 257]. Volume [496, 495, 933, 932, 1372, 1371, 544, 543, 1243, 1750, 1583]. volume-preserving [1243]. Vortex [450, 542, 92, 745, 2162]. vortices [1559, 542, 2142, 1749]. vorticity [1125, 1869]. Vries [1939, 348, 250, 620, 1607, 1129, 253, 1165, 469, 35, 1612, 1177, 1316, 1973, 1974, 36, 73].

W [1147]. wa [1501]. Wadati [258]. Wahlquist [404]. walk [645, 2015]. Walker [1759, 2165, 2166, 554, 727]. walks [1494]. walls [409, 1392]. Wannier [1526, 332, 1520, 1997]. water [35]. water-waves [35]. Wave [1490, 1072, 205, 710, 821, 1373, 43, 1337, 1733, 2243, 513, 1351, 152, 1793, 1201, 1448, 782, 2227, 2228, 278, 776, 135, 823, 1660, 576, 834, 1248, 752, 145, 918, 543, 5, 12, 1064, 454, 549, 1018, 469, 2078, 75, 1581, 1992, 1602, 1748, 978, 390, 1147, 255, 256, 2194, 1850, 953, 340, 1835, 983, 398, 1913, 2123, 1397, 1829, 1937, 1078, 1641, 2041, 1432, 189, 2191]. wave-fronts [1581]. wave-splitting [1351]. waveform [173, 1721]. waveguide [2138]. waveguides [820]. Wavelet [1625, 1632, 914, 124, 1636, 1626, 772, 1403, 1634, 1623]. wavelet-transformed [1634]. Wavelets [1627, 1640, 1514, 2214, 692, 1639]. Waves [1970, 643, 2012, 2125, 2115, 2126, 2127, 1760, 350, 1658, 690, 718, 1663, 2267, 35, 1916, 519, 313, 1504, 1944, 1038]. WDVV [2093]. Weak [1896, 1298, 2120]. weakly [2148]. wedge [1118, 2151]. Weierstrass [777, 2039]. weight [1552, 1744, 1921]. Weighted [1628, 1102]. weights [301]. Weil [1773, 1182]. Weinstein [1033]. Well [2168, 857, 91, 1893, 27]. Well-posed [2168]. wells [886, 1930]. Wentzel [2261, 1595, 1112, 334, 1695, 294]. Wess [22]. Weyl [1096, 1175, 1300, 1580, 1780, 1062, 1946, 419, 1963, 415, 293, 2238, 2056, 1387, 749]. Weyl-type [1946]. Wheeler [1655]. which [154, 2098]. white [378, 2184]. Whitham [2237]. whose [1314, 146]. Wick [1736, 1156, 202, 664]. Wiener [1984]. Wigner [1300, 2104, 57, 1777, 19, 48, 1654, 2192, 1902, 1084, 1728, 2238, 1812, 2258, 660, 2056, 894, 2058, 1699, 329]. wild [1046]. Wilson [740, 1652, 1935, 541]. window [820]. within [1756, 1469, 1794, 1958, 1128, 1747, 1861]. without [44, 451, 10, 231, 1052, 42, 1937]. Witt [1616]. Witten [192, 22, 428, 745]. WKB [537, 888, 225]. World [2197, 2244, 441]. World-line [2197, 2244]. wormholes [1320]. Wronski [2219]. Wronskian [1973, 904]. Wu [285]. WZNW [605]. WZW [1010, 1409].

Yajima [647]. Yang [593, 2137, 1240, 2047, 2027, 1675, 274, 2210, 317, 1505, 1149, 1296, 381, 1166, 1506, 570, 69, 741, 575, 1987, 1651, 1322, 237, 545, 1377, 843, 893, 1294, 980, 1936, 2093, 2237, 2025, 240, 1446, 798, 596, 1360, 425, 165, 1431, 1996, 1423]. Yang/Mills [596, 1360]. Yangian [283]. Yangians [1723]. Yano [1227]. Yau [531, 1406]. yield [1772]. Yukawa [1186, 1066, 1505]. Yukawian [384].

Zakharov [302, 592]. Zamolochikov [2099]. Zamolodchikov [1524, 1024]. Zeeman [285]. Zero [542, 623, 1996, 840, 841, 993, 1972, 2191, 596]. zerocurvature [993]. Zeros [2256]. Zeta [1438, 1246, 573, 777, 1209, 2139, 2235, 123, 330, 1346, 294]. Zeta-regularization [573]. Zetlin [1096, 1175]. Zhiber [723]. Ziolkowski [1147]. zone [263]. Zuber [701, 700, 220, 917]. Zuber-like [917]. Zumino [22].

References

Landauer:1957:SVC

[1] R. Landauer. Spatial variation of currents and fields due to localized scatterers in metallic conduction. IBM Journal of Research and Development, 1(3):223–231, July 1957. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Landauer:2000:SVC

[2] R. Landauer. Spatial variation of currents and fields due to localized scatterers in metallic conduction. IBM Journal of Research and Development, 44(1/2):251–??, ???? 2000. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.research.ibm.com/ journal/rd/441/landaueri.pdf. Special issue: reprints on Evolution of information technology 1957–1999. See comment [837].

Calogero:1971:SOD

[3] F. Calogero. Solution of the one-dimensional N-body problems with quadratic and/or inversely quadratic pair potentials. Journal of Mathematical Physics, 12(3):419–436, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [736].

Kota:1972:ASV

[4] J. Kota and Z. Perjes. All stationary vacuum metrics with shearing geodesic eigenrays. Journal of Mathematical Physics, 13(11):1695–1699, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1775].

Hortacsu:1993:FSG

[5] M. Hortaçsu. Fluctuations of a spherical gravitational impulsive wave. Journal of Mathematical Physics, 34(2):690–699, February 1993. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See errata [776].

Lammerzahl:1993:DST

[6] Claus Lämmerzahl and Alfredo Macías. On the dimensionality of space– time. Journal of Mathematical Physics, 34(10):4540–4553, October 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See addendum [166].

Fatibene:1994:NFC

[7] L. Fatibene, M. Ferraris, and M. Francaviglia. Nöther formalism for conserved quantities in classical gauge field theories. Journal of Mathematical Physics, 35(4):1644–1657, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [229].

Chu:1994:PFB

[8] Wen Chang Chu. Partial fractions and bilateral summations. Journal of Mathematical Physics, 35(4):2036–2042, April 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [374].

LaChapelle:1994:EGG

[9] J. LaChapelle. U(2) as the electroweak gauge group. Journal of Mathematical Physics, 35(5):2186–2198, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [230].

LaChapelle:1994:GMH

[10] J. LaChapelle. Generating mass without the Higgs particle. Journal of Mathematical Physics, 35(5):2199–2209, May 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [231].

Amir:1994:RCS

[11] M. Jamil Amir, Ashfaque H. Bokhari, and Asghar Qadir. Ricci collineations of static spherically symmetric spacetimes. Journal of Mathematical Physics, 35(6):3005–3012, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [568] and comment [567].

Hortacsu:1994:FRI

[12] M. Hortaçsu and R. Kaya. Fluctuations of a rotating impulsive wave. Journal of Mathematical Physics, 35(6):3043–3050, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See errata [776].

Good:1994:MSF

[13] R. H. Good, Jr. Mass spectra from field equations. Journal of Mathematical Physics, 35(7):3333–3339, July 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comments [457].

Dowker:1994:FDS

[14] J. S. Dowker. Functional determinants on spheres and sectors. Journal of Mathematical Physics, 35(9):4989–4999, September 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [87].

Slomczynski:1994:QCE

[15] Wojciech Słomczyński and Karol Życzkowski. Quantum chaos: an entropy approach. Journal of Mathematical Physics, 35(11):5674–5700, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [376].

Daniel:1994:IIS

[16] M. Daniel, K. Porsezian, and M. Lakshmanan. On the integrability of the inhomogeneous spherically symmetric Heisenberg ferromagnet in arbitrary dimensions. Journal of Mathematical Physics, 35(12):6498– 6510, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [814].

Bandelloni:1995:DCB

[17] Giuseppe Bandelloni and Serge Lazzarini. Diffeomorphism cohomology in Beltrami parametrization. II. The 1-forms. Journal of Mathematical Physics, 36(1):1–29, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barvinsky:1995:PST

[18] A. O. Barvinsky, T. A. Osborn, and Yu. V. Gusev. A phase-space technique for the perturbation expansion of Schrödinger propagators. Journal of Mathematical Physics, 36(1):30–61, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brocker:1995:MSP

[19] T. Bröcker and R. F. Werner. Mixed states with positive Wigner functions. Journal of Mathematical Physics, 36(1):62–75, January 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Brownstein:1995:BSV

[20] K. R. Brownstein. Bound state variational principle employing a discontinuous trial function. Journal of Mathematical Physics, 36(1):76–85, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forrester:1995:ECG

[21] P. J. Forrester. Exact calculation of the ground-state single-particle Green’s function for the 1/r^{2} quantum many-body system at integer coupling. Journal of Mathematical Physics, 36(1):86–96, January 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujii:1995:GWZ

[22] Kazuyuki Fujii. Generalizations of the Wess–Zumino–Witten model and Ferretti–Rajeev model to any dimensions: CP-invariance and geometry. Journal of Mathematical Physics, 36(1):97–114, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ito:1995:CTM

[23] Hiroshi T. Ito. Charge transfer model and three-body scattering. Journal of Mathematical Physics, 36(1):115–132, January 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kanter:1995:DCI

[24] Marek Kanter. Density correlation inequalities analogous to Bell’s. Journal of Mathematical Physics, 36(1):133–141, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:1995:PQS

[25] Yun Gang Lu. Passage from quantum systems with continuous spectrum to quantum Poisson processes on Hilbert modules. Journal of Mathematical Physics, 36(1):142–176, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Magni:1995:CEQ

[26] C. Magni and F. Valz-Gris. Can elementary quantum mechanics explain the Aharonov–Bohm effect. Journal of Mathematical Physics, 36(1):177– 186, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Witwit:1995:ELD

[27] M. R. M. Witwit. Energy levels of double-well potentials in a threedimensional system. Journal of Mathematical Physics, 36(1):187–194, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1995:LPS

[28] G. F. Torres del Castillo. Lanczos potential for some type D space–times. Journal of Mathematical Physics, 36(1):195–200, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baeyens:1995:SMH

[29] Bruno Baeyens and Henri Verschelde. Statistical mechanics of hard spheres. Journal of Mathematical Physics, 36(1):201–218, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciccariello:1995:IED

[30] S. Ciccariello. Integral expressions of the derivatives of the small-angle scattering correlation function. Journal of Mathematical Physics, 36(1): 219–246, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baez:1995:TAS

[31] John C. Baez, Michael S. Ody, and William Richter. Topological aspects of spin and statistics in nonlinear sigma models. Journal of Mathematical Physics, 36(1):247–257, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barcelos-Neto:1995:HSS

[32] J. Barcelos-Neto, Sasanka Ghosh, and Shibaji Roy. The Hamiltonian structures of the super-KP hierarchy associated with an even parity super-Lax operator. Journal of Mathematical Physics, 36(1):258–267, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brunelli:1995:TIS

[33] J. C. Brunelli and Ashok Das. Tests of integrability of the supersymmetric nonlinear Schrödinger equation. Journal of Mathematical Physics, 36 (1):268–280, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carlini:1995:AED

[34] Alessandra Carlini, Angelo Alessandro Mazzotti, Maria Cristina Recchioni, and Francesco Zirilli. Asymptotic eigenvalue degeneracy for a class of three-dimensional Fokker–Planck operators. Journal of Mathematical Physics, 36(1):281–306, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kraenkel:1995:KVH

[35] R. A. Kraenkel, M. A. Manna, and J. G. Pereira. The Korteweg–de Vries hierarchy and long water-waves. Journal of Mathematical Physics, 36 (1):307–320, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sukumar:1995:TFK

[36] C. V. Sukumar and P. Guha. Tau functions, Korteweg–de Vries flows, and supersymmetric quantum mechanics. Journal of Mathematical Physics, 36(1):321–330, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banerjee:1995:QSH

[37] A. Banerjee, D. Panigrahi, and S. Chatterjee. On the question of singularity in higher-dimensional space–time. Journal of Mathematical Physics, 36(1):331–334, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coley:1995:FCS

[38] Alan A. Coley and Des J. McManus. A family of cosmological solutions in higher-dimensional Einstein gravity. Journal of Mathematical Physics, 36 (1):335–339, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:1995:ASS

[39] A. Das, N. Tariq, and T. Biech. Analytical solutions of spherically symmetric collapse of an anisotropic fluid body into a black hole. Journal of Mathematical Physics, 36(1):340–355, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harvey:1995:ISF

[40] Alex Harvey. Identities of the scalars of the four-dimensional Riemannian manifold. Journal of Mathematical Physics, 36(1):356–361, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McManus:1995:RTM

[41] Des J. McManus. Riemannian three-metrics with degenerate Ricci tensors. Journal of Mathematical Physics, 36(1):362–369, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mullick:1995:GMT

[42] Lipika Mullick and Pratul Bandyopadhyay. Gravity without metric, torsion, and topological gravity. Journal of Mathematical Physics, 36(1): 370–382, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bietenholz:1995:SSS

[43] W. Bietenholz and J. J. Giambiagi. Solutions of the spherically symmetric wave equation in p + q dimensions. Journal of Mathematical Physics, 36(1):383–397, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chaichian:1995:QPG

[44] M. Chaichian and A. P. Dëmichev. Quantum Poincaré group without dilatation and twisted classical algebra. Journal of Mathematical Physics, 36(1):398–413, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chekhov:1995:FSA

[45] L. Chekhov. L functions in scattering on p-adic multiloop surfaces. Journal of Mathematical Physics, 36(1):414–425, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gilbert:1995:CCP

[46] George T. Gilbert and Matthew P. Richey. Cuspidal coverings for pairs of congruence subgroups. Journal of Mathematical Physics, 36(1):426–434, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gould:1995:CFD

[47] M. D. Gould and M. Scheunert. Classification of finite-dimensional unitary irreps for U_{q}[gl(m|n)]. Journal of Mathematical Physics, 36(1): 435–452, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Healy:1995:ICC

[48] D. M. Healy, Jr. and F. E. Schroeck, Jr. On informational completeness of covariant localization observables and Wigner coefficients. Journal of Mathematical Physics, 36(1):453–507, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimyk:1995:SCG

[49] A. U. Klimyk. Symmetries of Clebsch–Gordan coefficients of the quantum group U_{q}(n). Journal of Mathematical Physics, 36(1):508–523, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1995:DGM

[50] Dae-Gyu Lee. The Dirac gamma matrices as “relics” of a hidden symmetry?: As fundamental representations of the algebra sp(4, R). Journal of Mathematical Physics, 36(1):524–530, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Links:1995:CUG

[51] J. R. Links and M. D. Gould. Classification of unitary and grade star irreps for U_{q}(osp(2|2n)). Journal of Mathematical Physics, 36(1):531– 545, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McAnally:1995:EGFa

[52] D. S. McAnally. q-exponential and q-gamma functions. I. q-exponential functions. Journal of Mathematical Physics, 36(1):546–573, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McAnally:1995:EGFb

[53] D. S. McAnally. q-exponential and q-gamma functions. II. q-gamma functions. Journal of Mathematical Physics, 36(1):574–595, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Toh:1995:KCG

[54] Tze-Chuen Toh and Malcolm R. Anderson. Knots and classical 3geometries. Journal of Mathematical Physics, 36(1):596–604, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanderJeugt:1995:DFL

[55] J. Van der Jeugt. Dimension formulas for the Lie superalgebra sl(m/n). Journal of Mathematical Physics, 36(1):605–611, January 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arai:1995:OTA

[56] Asao Arai. Operator-theoretical analysis of a representation of a supersymmetry algebra in Hilbert space. Journal of Mathematical Physics, 36 (2):613–621, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arai:1995:SES

[57] Takahiro Arai. Some extensions of semiclassical limit h̄ → 0 for Wigner functions on phase space. Journal of Mathematical Physics, 36(2):622– 630, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ballesteros:1995:CDP

[58] A. Ballesteros, F. J. Herranz, M. A. del Olmo, and M. Santander. Classical deformations, Poisson–Lie contractions, and quantization of dual Lie bialgebras. Journal of Mathematical Physics, 36(2):631–640, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beckers:1995:SRE

[59] J. Beckers, N. Debergh, and L. Ntibashirakandi. On supersymmetry and radial equations for three-dimensional central problems. Journal of Mathematical Physics, 36(2):641–647, February 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davis:1995:DMS

[60] Simon Davis. Divergences in the moduli space integral and accumulating handles in the infinite-genus limit. Journal of Mathematical Physics, 36 (2):648–663, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [284].

Davtyan:1995:HTN

[61] Levon S. Davtyan, Aleksey N. Sissakian, and Valery M. Ter-Antonyan. The Hurwitz transformation: nonbilinear version. Journal of Mathematical Physics, 36(2):664–674, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gannon:1995:TCM

[62] Terry Gannon. Towards a classification of su(2) ⊕ ⋯ ⊕ su(2) modular invariant partition functions. Journal of Mathematical Physics, 36(2): 675–706, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Good:1995:MSF

[63] R. H. Good, Jr. Mass spectra from field equations. II. The electromagnetic contribution. Journal of Mathematical Physics, 36(2):707–713, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Indelicato:1995:AED

[64] P. Indelicato and Peter J. Mohr. Asymptotic expansion of the Dirac– Coulomb radial Green’s function. Journal of Mathematical Physics, 36 (2):714–724, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maassen:1995:FSR

[65] H. Maassen and A. Tip. A Fock space representation for the quantum Lorentz gas. Journal of Mathematical Physics, 36(2):725–741, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Itoh:1995:GMM

[66] Mitsuhiro Itoh. Generalized magnetic monopoles over contact manifolds. Journal of Mathematical Physics, 36(2):742–749, February 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kostin:1995:NDE

[67] M. D. Kostin. Number density, entropy density, and energy density of statistical mechanics for D-dimensional grand canonical ensembles. Journal of Mathematical Physics, 36(2):750–755, February 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bainov:1995:NGS

[68] D. D. Baĭnov and E. Minchev. Nonexistence of global solutions of the initial-boundary value problem for the nonlinear Klein–Gordon equation. Journal of Mathematical Physics, 36(2):756–762, February 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Chakravarty:1995:SRS

[69] S. Chakravarty, S. L. Kent, and E. T. Newman. Some reductions of the self-dual Yang–Mills equations to integrable systems in 2+1 dimensions. Journal of Mathematical Physics, 36(2):763–772, February 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

El-Rifai:1995:KPO

[70] E. A. El-Rifai and E. Ahmed. Knotted periodic orbits in Rössler’s equations. Journal of Mathematical Physics, 36(2):773–777, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smalley:1995:FST

[71] Larry L. Smalley and Jean P. Krisch. Fluids with spin and twist. Journal of Mathematical Physics, 36(2):778–795, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Speight:1995:LED

[72] J. M. Speight. Low-energy dynamics of a CP^{1} lump on the sphere. Journal of Mathematical Physics, 36(2):796–813, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Svinin:1995:HKV

[73] Andrei K. Svinin. Hierarchies of Korteweg–de Vries type equations. Journal of Mathematical Physics, 36(2):814–827, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Koberlein:1995:PMG

[74] Brian D. Koberlein and Ronald L. Mallett. Particle motion in the gravitational field of a charged, radiating body. Journal of Mathematical Physics, 36(2):828–836, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liern:1995:SIS

[75] V. Liern and J. Olivert. A symplectic interpretation of some wave fronts in general relativity. Journal of Mathematical Physics, 36(2):837–846, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silbergleit:1995:NMA

[76] A. S. Silbergleit. Nonlinear motions against the Newtonian uniform expansion background: the case of the unperturbed density. Journal of Mathematical Physics, 36(2):847–862, February 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borac:1995:AGT

[77] Silviu Borac. On the algebra generated by two projections. Journal of Mathematical Physics, 36(2):863–874, February 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bose:1995:RDG

[78] S. K. Bose. Representations of the (2 + 1)-dimensional Galilean group. Journal of Mathematical Physics, 36(2):875–890, February 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Drabant:1995:UCR

[79] Bernhard Drabant and Wolfgang Weich. Unitary continuous representations of compact quantum groups. Journal of Mathematical Physics, 36 (2):891–897, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kusnezov:1995:EME

[80] Dimitri Kusnezov. Exact matrix expansions for group elements of su(n). Journal of Mathematical Physics, 36(2):898–906, February 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Minnaert:1995:RCS

[81] Pierre Minnaert and Marek Mozrzymas. Racah coefficients and 6-j symbols for the quantum superalgebra U_{q}(osp(1|2)). Journal of Mathematical Physics, 36(2):907–922, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molev:1995:SDL

[82] Alexander Molev. Sklyanin determinant, Laplace operators, and characteristic identities for classical Lie algebras. Journal of Mathematical Physics, 36(2):923–943, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pittner:1995:DQH

[83] L. Pittner and P. Uray. Duals of quasitriangular Hopf superalgebras and the classical limit. Journal of Mathematical Physics, 36(2):944–966, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zha:1995:DSH

[84] Chao Zheng Zha and Wei Zhong Zhao. The q deformation of super highorder Virasoro algebra. Journal of Mathematical Physics, 36(2):967–979, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vogt:1995:RCN

[85] Andrew Vogt. On R. E. Collins’s new viewpoint in physics: “On some features of differentiable probabilities; a new viewpoint in physics”, [J. Math. Phys. 34 (1993), no. 7, 3210–3221; MR 94e:82061]. With a reply by Collins. Journal of Mathematical Physics, 36(2):980–987, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [?] and response [86].

Collins:1995:RVC

[86] R. E. Collins. Response to Vogt’s comments on Collins’s new viewpoint in physics. Journal of Mathematical Physics, 36(2):983–987, February 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [?, 85].

Dowker:1995:EFD

[87] J. S. Dowker. Erratum: “Functional determinants on spheres and sectors”. Journal of Mathematical Physics, 36(2):988, February 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [14].

Blancarte:1995:HLE

[88] Herminio Blancarte, Benoî t Grébert, and Ricardo Weder. High- and low-energy estimates for the Dirac equation. Journal of Mathematical Physics, 36(3):991–1015, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chetouani:1995:ATK

[89] L. Chetouani, L. Guechi, T. F. Hammann, and A. Lecheheb. Algebraic treatment of the Kaluza–Klein monopole system. Journal of Mathematical Physics, 36(3):1016–1028, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:1995:IHE

[90] B. K. Chung and Y. D. Kang. Improved high-energy bounds on absorptive parts of elastic scattering amplitudes. Journal of Mathematical Physics, 36(3):1029–1037, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrokhotov:1995:DWS

[91] Sergei Yu. Dobrokhotov and Vassili N. Kolokol′tsov. The double-well splitting of the low energy levels for the Schrödinger operator of discrete ϕ^{4}-models on tori. Journal of Mathematical Physics, 36(3):1038–1053, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrowolski:1995:CVS

[92] Tomasz Jacek Dobrowolski. On curved vortex solutions in the abelian Higgs model. Journal of Mathematical Physics, 36(3):1054–1062, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Furlani:1995:QEF

[93] E. P. Furlani. Quantization of the electromagnetic field on static space– times. Journal of Mathematical Physics, 36(3):1063–1079, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Halpern:1995:FCN

[94] M. B. Halpern and N. A. Obers. Flat connections and nonlocal conserved quantities in irrational conformal field theory. Journal of Mathematical Physics, 36(3):1080–1110, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hanlon:1995:NGM

[95] B. E. Hanlon and G. C. Joshi. A noncommutative geometric model with horizontal family symmetry. Journal of Mathematical Physics, 36(3): 1111–1122, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hess:1995:BMS

[96] P. O. Hess and J. C. López. Boson mapping of symplectic algebras with abelian subalgebra mapped as coordinates. Journal of Mathematical Physics, 36(3):1123–1135, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1995:PBS

[97] M. Howard Lee. Price’s bound on the structure factor: derivation and comparison with some exact results. Journal of Mathematical Physics, 36 (3):1136–1145, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ordonez:1995:DRQ

[98] C. R. Ordóñez and J. M. Pons. Dirac and reduced quantization: a Lagrangian approach and application to coset spaces. Journal of Mathematical Physics, 36(3):1146–1165, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pereyra:1995:SPG

[99] Pedro Pereyra. Symmetries, parametrization, and group structure of transfer matrices in quantum scattering theory. Journal of Mathematical Physics, 36(3):1166–1176, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:1995:SSH

[100] C. C. Sun, B. F. Li, and Z. S. Li. Symplectic symmetry of Hubbard model and η pairing. Journal of Mathematical Physics, 36(3):1177–1191, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thienel:1995:LDG

[101] H.-P. Thienel. Local differential geometry as a representation of the supersymmetric oscillator. Journal of Mathematical Physics, 36(3):1192– 1197, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fecko:1995:TSM

[102] Marián Fecko. Transformation of the solutions of the Maxwell equations on a Lorentzian manifold by means of a certain class of diffeomorphisms. Journal of Mathematical Physics, 36(3):1198–1207, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blinder:1995:CPF

[103] S. M. Blinder. Canonical partition function for the hydrogen atom via the Coulomb propagator. Journal of Mathematical Physics, 36(3):1208– 1216, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1995:PAC

[104] M. Howard Lee. Polylogarithmic analysis of chemical potential and fluctuations in a D-dimensional free Fermi gas at low temperatures. Journal of Mathematical Physics, 36(3):1217–1231, March 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amiran:1995:LCI

[105] Edoh Y. Amiran. Lazutkin coordinates and invariant curves for outer billiards. Journal of Mathematical Physics, 36(3):1232–1241, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouquet:1995:HSI

[106] S. Bouquet. Hamiltonian structure and integrability of the stationary Kuramoto–Sivashinsky equation. Journal of Mathematical Physics, 36 (3):1242–1258, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doliwa:1995:IDD

[107] Adam Doliwa and Paolo Maria Santini. Integrable dynamics of a discrete curve and the Ablowitz–Ladik hierarchy. Journal of Mathematical Physics, 36(3):1259–1273, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lange:1995:TDD

[108] H. Lange, B. Toomire, and P. F. Zweifel. Time-dependent dissipation in nonlinear Schrödinger systems. Journal of Mathematical Physics, 36(3): 1274–1283, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mugan:1995:STP

[109] U. Muğan and A. Sakka. Schlesinger transformations for Painlevé VI equation. Journal of Mathematical Physics, 36(3):1284–1298, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanDiejen:1995:DCM

[110] J. F. van Diejen. Difference Calogero–Moser systems and finite Toda chains. Journal of Mathematical Physics, 36(3):1299–1323, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yahalom:1995:HCN

[111] Asher Yahalom. Helicity conservation via the Noether theorem. Journal of Mathematical Physics, 36(3):1324–1327, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coleman:1995:GFE

[112] Robert Alan Coleman and Heinz-Jürgen Schmidt. A geometric formulation of the equivalence principle. Journal of Mathematical Physics, 36 (3):1328–1346, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krori:1995:ESS

[113] K. D. Krori, A. K. Goswami, and Anuradha Das Purkayastha. An exact spinor solution in Kantowski–Sachs cosmology. Journal of Mathematical Physics, 36(3):1347–1352, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McManus:1995:LTM

[114] Des J. McManus. Lorentzian three-metrics with degenerate Ricci tensors. Journal of Mathematical Physics, 36(3):1353–1364, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quevedo:1995:TSU

[115] Hernando Quevedo and Roberto A. Sussman. Thermodynamics of the Stephani universes. Journal of Mathematical Physics, 36(3):1365–1377, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boersma:1995:PME

[116] Stuart Boersma and Tevian Dray. Parametric manifolds. I. Extrinsic approach. Journal of Mathematical Physics, 36(3):1378–1393, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boersma:1995:PMI

[117] Stuart Boersma and Tevian Dray. Parametric manifolds. II. Intrinsic approach. Journal of Mathematical Physics, 36(3):1394–1403, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ellinas:1995:QHG

[118] Demosthenes Ellinas and Jan Sobczyk. Quantum Heisenberg group and algebra: contraction, left and right regular representations. Journal of Mathematical Physics, 36(3):1404–1412, March 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frescura:1995:ECE

[119] F. A. M. Frescura and G. Łubczonok. An exterior calculus for exact symplectic manifolds. Journal of Mathematical Physics, 36(3):1413–1425, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:1995:ENS

[120] A. M. Grundland and J. Tafel. On the existence of nonclassical symmetries of partial differential equations. Journal of Mathematical Physics, 36 (3):1426–1434, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kawahara:1995:PLC

[121] Yasuo Kawahara, Satoru Kumamoto, Yoshihiro Mizoguchi, Masaya Nohmi, Hiroshi Ohtsuka, and Takayoshi Shoudai. Period lengths of cellular automata on square lattices with rule 90. Journal of Mathematical Physics, 36(3):1435–1456, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lesniewski:1995:RCE

[122] Andrzej Lesniewski. A remark on the Casimir elements of Lie superalgebras and quantized Lie superalgebras. Journal of Mathematical Physics, 36(3):1457–1461, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nash:1995:DLR

[123] Charles Nash and Denjoe O’Connor. Determinants of Laplacians, the Ray–Singer torsion on lens spaces and the Riemann zeta function. Journal of Mathematical Physics, 36(3):1462–1505, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [330].

Perrier:1995:WSC

[124] Valérie Perrier, Thierry Philipovitch, and Claude Basdevant. Wavelet spectra compared to Fourier spectra. Journal of Mathematical Physics, 36(3):1506–1519, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1995:RML

[125] D. J. Rowe. Resolution of missing label problems; a new perspective on K-matrix theory. Journal of Mathematical Physics, 36(3):1520–1530, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schirrmacher:1995:GER

[126] Arne Schirrmacher. Generalized q-exponentials related to orthogonal quantum groups and Fourier transformations of noncommutative spaces. Journal of Mathematical Physics, 36(3):1531–1546, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zaugg:1995:QTD

[127] Philippe Zaugg. The quantum two-dimensional Poincaré group from quantum group contraction. Journal of Mathematical Physics, 36(3): 1547–1553, March 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramidi:1995:CAC

[128] I. G. Avramidi. Covariant algebraic calculation of the one-loop effective potential in non-abelian gauge theory and a new approach to stability problem. Journal of Mathematical Physics, 36(4):1557–1571, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergeron:1995:OBO

[129] H. Bergeron and A. Valance. Overcomplete basis for one-dimensional Hamiltonians. Journal of Mathematical Physics, 36(4):1572–1592, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mankoc-Borstnik:1995:PAO

[130] Norma Mankoč-Borštnik. Poincaré algebra in ordinary and Grassmann space and supersymmetry. Journal of Mathematical Physics, 36(4):1593– 1601, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boudjedaa:1995:PIS

[131] T. Boudjedaa, A. Bounames, L. Chetouani, T. F. Hammann, and Kh. Nouicer. Path integral for spinning particle in magnetic field via bosonic coherent states. Journal of Mathematical Physics, 36(4):1602–1615, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chadan:1995:CBN

[132] K. Chadan, A. Martin, and J. Stubbe. The Calogero bound for nonzero angular momentum. Journal of Mathematical Physics, 36(4):1616–1624, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coutinho:1995:VPT

[133] F. A. B. Coutinho, J. Fernando Perez, and Walter F. Wreszinski. A variational proof of the Thomas effect. Journal of Mathematical Physics, 36 (4):1625–1635, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dattoli:1995:PSF

[134] G. Dattoli and A. Torre. Phase-space formalism: generalized Hermite polynomials, orthogonal functions and creation-annihilation operators. Journal of Mathematical Physics, 36(4):1636–1644, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Etingof:1995:IFW

[135] Pavel Etingof. Integral formulas for wave functions of quantum manybody problems and representations of }↕_{\}. Journal of Mathematical Physics, 36(4):1645–1651, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujii:1995:STE

[136] Kazuyuki Fujii. Solutions of A_{∞} Toda equations based on noncompact group SU(1, 1) and infinite-dimensional Grassmann manifolds. Journal of Mathematical Physics, 36(4):1652–1665, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kochetov:1995:PIG

[137] E. A. Kochetov. Path integral over the generalized coherent states. Journal of Mathematical Physics, 36(4):1666–1672, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lahti:1995:POQ

[138] Pekka J. Lahti and Maciej J. Maczyński. Partial order of quantum effects. Journal of Mathematical Physics, 36(4):1673–1680, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maximov:1995:DQM

[139] V. Maximov and A. Odzijewicz. The q-deformation of quantum mechanics of one degree of freedom. Journal of Mathematical Physics, 36 (4):1681–1690, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McKeon:1995:FUQ

[140] D. G. C. McKeon and C. Wong. Further uses of the quantum mechanical path integral in quantum field theory. Journal of Mathematical Physics, 36(4):1691–1706, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Potz:1995:STM

[141] W. Pötz. Scattering theory for mesoscopic quantum systems with nontrivial spatial asymptotics in one dimension. Journal of Mathematical Physics, 36(4):1707–1740, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reifler:1995:UDE

[142] Frank Reifler and Randall Morris. Unification of the Dirac and Einstein Lagrangians in a tetrad model. Journal of Mathematical Physics, 36 (4):1741–1752, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

SassolideBianchi:1995:NSB

[143] M. Sassoli de Bianchi and M. Di Ventra. On the number of states bound by one-dimensional finite periodic potentials. Journal of Mathematical Physics, 36(4):1753–1764, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arthur:1995:PME

[144] K. Arthur and J. Burzlaff. Perturbation method for the energymomentum tensor of an optical soliton. Journal of Mathematical Physics, 36(4):1765–1775, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:1995:DPC

[145] Sailing He and Vaughan H. Weston. Determination of the permittivity and conductivity in R^{3} using wave splitting of Maxwell’s equations. Journal of Mathematical Physics, 36(4):1776–1789, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwai:1995:MKS

[146] Toshihiro Iwai and Noriaki Katayama. Multifold Kepler systems— dynamical systems all of whose bounded trajectories are closed. Journal of Mathematical Physics, 36(4):1790–1811, April 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meurice:1995:HTE

[147] Y. Meurice. The high-temperature expansion of the hierarchical Ising model: from Poincaré symmetry to an algebraic algorithm. Journal of Mathematical Physics, 36(4):1812–1824, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nettleton:1995:IEV

[148] R. E. Nettleton. Integral equations versus internal variables in the search for an effectively hyperbolic thermal transport equation. Journal of Mathematical Physics, 36(4):1825–1833, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vlad:1995:GFA

[149] Marcel Ovidiu Vlad and Michael C. Mackey. Generating functional approach to multichannel parallel relaxation with application to the problem of direct energy transfer in fractal systems with dynamic disorder. Journal of Mathematical Physics, 36(4):1834–1853, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Almeida:1995:LSI

[150] M. A. Almeida, M. E. Magalhães, and I. C. Moreira. Lie symmetries and invariants of the Lotka–Volterra system. Journal of Mathematical Physics, 36(4):1854–1867, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arias:1995:GIS

[151] Pío José Arias and Jorge Stephany. Gauge invariance and second class constraints in three-dimensional linearized massive gravity. Journal of Mathematical Physics, 36(4):1868–1876, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Breton:1995:SCP

[152] Nora Bretón. On some colliding plane wave solutions. Journal of Mathematical Physics, 36(4):1877–1881, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gowdy:1995:APT

[153] Robert H. Gowdy. Affine projection-tensor geometry: Lie derivatives and isometries. Journal of Mathematical Physics, 36(4):1882–1907, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [375].

Azad:1995:SSM

[154] H. Azad and M. Ziad. Spherically symmetric manifolds which admit five isometries. Journal of Mathematical Physics, 36(4):1908–1911, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cadavid:1995:CGC

[155] A. C. Cadavid and R. J. Finkelstein. Clebsch–Gordan coefficients and related identities obtained by an integration over the group space of SU_{q}(2). Journal of Mathematical Physics, 36(4):1912–1933, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurarie:1995:SES

[156] David Gurarie. Semiclassical eigenvalues and shape problems on surfaces of revolution. Journal of Mathematical Physics, 36(4):1934–1944, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harnad:1995:IGT

[157] J. Harnad and S. Shnider. Isotropic geometry and twistors in higher dimensions. II. Odd dimensions, reality conditions, and twistor superspaces. Journal of Mathematical Physics, 36(4):1945–1970, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hood:1995:NES

[158] Simon Hood. New exact solutions of Burgers’s equation—an extension to the direct method of Clarkson and Kruskal. Journal of Mathematical Physics, 36(4):1971–1990, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:1995:ETQ

[159] S. Majid. q-epsilon tensor for quantum and braided spaces. Journal of Mathematical Physics, 36(4):1991–2007, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1995:IST

[160] D. J. Rowe and J. Repka. Induced shift tensors in vector coherent state theory. Journal of Mathematical Physics, 36(4):2008–2029, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sazdjian:1995:CSO

[161] Hagop Sazdjian, Yassen S. Stanev, and Ivan T. Todorov. SU_{3} coherent state operators and invariant correlation functions and their quantum group counterparts. Journal of Mathematical Physics, 36(4):2030–2052, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stanev:1995:CLEa

[162] Yassen S. Stanev. Classification of the local extensions of the SU(3) chiral current algebra. Journal of Mathematical Physics, 36(4):2053– 2069, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stanev:1995:CLEb

[163] Yassen S. Stanev. Classification of the local extensions of the SU(2) × SU(2) chiral current algebras. Journal of Mathematical Physics, 36(4): 2070–2084, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

WillsToro:1995:GLA

[164] Luis Alberto Wills Toro. (I, q)-graded Lie algebraic extensions of the Poincaré algebra, constraints on I and q. Journal of Mathematical Physics, 36(4):2085–2112, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Torre:1995:NSY

[165] C. G. Torre. Natural symmetries of the Yang–Mills equations. Journal of Mathematical Physics, 36(4):2113–2130, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lammerzahl:1995:ADS

[166] Claus Lämmerzahl. Addendum to: “On the dimensionality of space– time”, [J. Math. Phys. 34 (1993), no. 10, 4540–4553; MR 94e:81081] by Lämmerzahl and A. Macías. Journal of Mathematical Physics, 36(4): 2131, April 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [6].

Albeverio:1995:OIH

[167] Sergio Albeverio and Zdzisław Brzeźniak. Oscillatory integrals on Hilbert spaces and Schrödinger equation with magnetic fields. Journal of Mathematical Physics, 36(5):2135–2156, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Albeverio:1995:ACS

[168] Sergio Albeverio and Jörg Schäfer. Abelian Chern–Simons theory and linking numbers via oscillatory integrals. Journal of Mathematical Physics, 36(5):2157–2169, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashtekar:1995:PTF

[169] Abhay Ashtekar and Jerzy Lewandowski. Projective techniques and functional integration for gauge theories. Journal of Mathematical Physics, 36(5):2170–2191, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blau:1995:LDR

[170] Matthias Blau and George Thompson. Localization and diagonalization: a review of functional integral techniques for low-dimensional gauge theories and topological field theories. Journal of Mathematical Physics, 36 (5):2192–2236, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cartier:1995:NPF

[171] Pierre Cartier and Cécile DeWitt-Morette. A new perspective on functional integration. Journal of Mathematical Physics, 36(5):2237–2312, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fischer:1995:FAV

[172] Werner Fischer, Hajo Leschke, and Peter Müller. The functional-analytic versus the functional-integral approach to quantum Hamiltonians: the one-dimensional hydrogen atom. Journal of Mathematical Physics, 36 (5):2313–2323, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foong:1995:ODD

[173] S. K. Foong. Overcoming dissipative distortions by a waveform restorer and designer. Journal of Mathematical Physics, 36(5):2324–2340, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See addendum [1721].

Gallivan:1995:FUN

[174] T. E. Gallivan and Arie Kapulkin. Failure of universality in noncompact lattice field theories. Journal of Mathematical Physics, 36(5):2341–2353, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grosche:1995:HSP

[175] Christian Grosche and Frank Steiner. How to solve path integrals in quantum mechanics. Journal of Mathematical Physics, 36(5):2354–2385, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hunt:1995:PCP

[176] F. Y. Hunt, J. F. Douglas, and J. Bernal. Probabilistic computation of Poiseuille flow velocity fields. Journal of Mathematical Physics, 36 (5):2386–2401, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kauffman:1995:FIT

[177] Louis H. Kauffman. Functional integration and the theory of knots. Journal of Mathematical Physics, 36(5):2402–2429, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Makri:1995:NPI

[178] Nancy Makri. Numerical path integral techniques for long time dynamics of quantum dissipative systems. Journal of Mathematical Physics, 36(5): 2430–2457, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marinov:1995:PIH

[179] M. S. Marinov. Path integrals on homogeneous manifolds. Journal of Mathematical Physics, 36(5):2458–2469, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mottola:1995:FIG

[180] Emil Mottola. Functional integration over geometries. Journal of Mathematical Physics, 36(5):2470–2511, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paniak:1995:KMM

[181] Lori Paniak and Nathan Weiss. Kazakov–Migdal model with logarithmic potential and the double Penner matrix model. Journal of Mathematical Physics, 36(5):2512–2530, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rogers:1995:PIA

[182] Alice Rogers. Path integration, anticommuting variables, and supersymmetry. Journal of Mathematical Physics, 36(5):2531–2545, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schulman:1995:TDI

[183] L. S. Schulman. Time displaced interactions: classical dynamics and path integral quantization. Journal of Mathematical Physics, 36(5):2546– 2564, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeWitt-Morette:1995:P

[184] Cecile DeWitt-Morette. Postscript. Journal of Mathematical Physics, 36 (5):2565, May 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arai:1995:GTN

[185] Asao Arai. Gauge theory on a non-simply connected domain and representations of canonical commutation relations. Journal of Mathematical Physics, 36(6):2569–2580, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atkinson:1995:IEM

[186] D. Atkinson and V. P. Gusynin. An integral equation of Muskhelishvili type: strong quantum electrodynamics. Journal of Mathematical Physics, 36(6):2581–2592, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bohm:1995:QMA

[187] A. Bohm, I. Antoniou, and P. Kielanowski. A quantum mechanical arrow of time and the semigroup time evolution of Gamow vectors. Journal of Mathematical Physics, 36(6):2593–2604, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carey:1995:RCO

[188] A. L. Carey, H. Grundling, C. A. Hurst, and E. Langmann. Realizing 3-cocycles as obstructions. Journal of Mathematical Physics, 36(6):2605– 2620, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dolinszky:1995:SSP

[189] T. Dolinszky. Scattering by singular potentials. A. Wave function expansion. Journal of Mathematical Physics, 36(6):2621–2635, June 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Etingof:1995:QIS

[190] Pavel I. Etingof. Quantum integrable systems and representations of Lie algebras. Journal of Mathematical Physics, 36(6):2636–2651, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkelstein:1995:TTO

[191] Robert Finkelstein and Eran Marcus. Transformation theory of the q-oscillator. Journal of Mathematical Physics, 36(6):2652–2672, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Freund:1995:DMS

[192] Peter G. O. Freund. Dirac monopoles and the Seiberg–Witten monopole equations. Journal of Mathematical Physics, 36(6):2673–2674, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Friedberg:1995:PIP

[193] Richard Friedberg. Path integrals in polar variables with spontaneously broken symmetry. Journal of Mathematical Physics, 36(6):2675–2690, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gracia-Bondia:1995:GQM

[194] José M. Gracia-Bondía and Joseph C. Várilly. From geometric quantization to Moyal quantization. Journal of Mathematical Physics, 36 (6):2691–2701, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grinberg:1995:SST

[195] N. I. Grinberg. Scattering on small three-dimensional, nonspherically symmetric potentials. Journal of Mathematical Physics, 36(6):2702– 2719, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwasawa:1995:ESH

[196] H. Iwasawa and K. Matsuo. Exact solutions of Heisenberg equations for multiphoton Jaynes–Cummings model. Journal of Mathematical Physics, 36(6):2720–2726, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalkkinen:1995:IQN

[197] Jussi Kalkkinen. Induced QCD from the noncommutative geometry of a supermanifold. Journal of Mathematical Physics, 36(6):2727–2740, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kolokoltsov:1995:STB

[198] Vassili N. Kolokol′tsov. Scattering theory for the Belavkin equation describing a quantum particle with continuously observed coordinate. Journal of Mathematical Physics, 36(6):2741–2760, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kurylev:1995:IBP

[199] Yaroslav Kurylev. An inverse boundary problem for the Schrödinger operator with magnetic field. Journal of Mathematical Physics, 36(6): 2761–2776, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Legare:1995:ISR

[200] M. Légaré. Invariant spinors and reduced Dirac equations under subgroups of the Euclidean group in four-dimensional Euclidean space. Journal of Mathematical Physics, 36(6):2777–2791, June 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levay:1995:BPL

[201] Péter Lévay. Berry phases for Landau Hamiltonians on deformed tori. Journal of Mathematical Physics, 36(6):2792–2802, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marcinek:1995:WAB

[202] Władysław Marcinek and Robert Rałowski. On Wick algebras with braid relations. Journal of Mathematical Physics, 36(6):2803–2812, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [664].

Melnikov:1995:TBS

[203] Yu. B. Melnikov and B. S. Pavlov. Two-body scattering on a graph and application to simple nanoelectronic devices. Journal of Mathematical Physics, 36(6):2813–2825, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moylan:1995:HAS

[204] P. Moylan. Harmonic analysis on spannors. Journal of Mathematical Physics, 36(6):2826–2879, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:1995:IWS

[205] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. Inverse wave scattering with discontinuous wave speed. Journal of Mathematical Physics, 36(6):2880–2928, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duggal:1995:CSF

[206] K. L. Duggal, G. G. Asgekar, and S. M. Aherkar. A class of shearfree, non-static models for anisotropic magnetofluid systems. Journal of Mathematical Physics, 36(6):2929–2940, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Helsing:1995:NAE

[207] Johan Helsing. A Nyström algorithm for electrostatics of an anisotropic composite. Journal of Mathematical Physics, 36(6):2941–2950, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://scitation.aip.org/content/aip/
journal/jmp/36/6/10.1063/1.531002.

MacLeod:1995:NDF

[208] Malcolm A. MacLeod. A new description of force-free magnetic fields. Journal of Mathematical Physics, 36(6):2951–2958, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:1995:CEC

[209] Wen-Jui Huang, J. C. Shaw, and H. C. Yen. Classical extended conformal algebras associated with constrained KP hierarchy. Journal of Mathematical Physics, 36(6):2959–2971, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kundu:1995:GTC

[210] Anjan Kundu, Walter Strampp, and Walter Oevel. Gauge transformations of constrained KP flows: new integrable hierarchies. Journal of Mathematical Physics, 36(6):2972–2984, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dabrowski:1995:NKS

[211] Mariusz P. Dabrowski. Note on Kantowski–Sachs universes. Journal of Mathematical Physics, 36(6):2985–2987, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deOliveira:1995:BHN

[212] H. P. de Oliveira. Black holes and nonlinear scalar fields: a dynamical system approach. Journal of Mathematical Physics, 36(6):2988–3003, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eardley:1995:GCV

[213] Douglas M. Eardley. Gravitational collapse of vacuum gravitational field configurations. Journal of Mathematical Physics, 36(6):3004–3011, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frauendiener:1995:SSR

[214] Jörg Frauendiener. On spin-(3/2) systems in Ricci flat space–times. Journal of Mathematical Physics, 36(6):3012–3022, June 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Letelier:1995:STDa

[215] Patricio S. Letelier and An Zhong Wang. Space–time defects. Journal of Mathematical Physics, 36(6):3023–3042, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Letelier:1995:STDb

[216] Patricio S. Letelier. Space–time defects: open and closed shells. Journal of Mathematical Physics, 36(6):3043–3053, June 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Magli:1995:TPM

[217] Giulio Magli. Tuning the properties of matter to any chosen dynamical behavior in cosmological models with elastic media. Journal of Mathematical Physics, 36(6):3054–3062, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manko:1995:SPS

[218] V. S. Manko, J. Martín, and E. Ruiz. Six-parameter solution of the Einstein–Maxwell equations possessing equatorial symmetry. Journal of Mathematical Physics, 36(6):3063–3073, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Santos:1995:CSO

[219] J. Santos, M. J. Rebouças, and A. F. F. Teixeira. Classification of second order symmetric tensors in five-dimensional Kaluza–Klein-type theories. Journal of Mathematical Physics, 36(6):3074–3084, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alfaro:1995:IZI

[220] Jorge Alfaro, Ricardo Medina, and Luis F. Urrutia. The Itzykson–Zuber integral for U(m|n). Journal of Mathematical Physics, 36(6):3085–3093, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [700, 701].

Alty:1995:GGB

[221] L. J. Alty. The generalized Gauss–Bonnet–Chern theorem. Journal of Mathematical Physics, 36(6):3094–3105, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:1995:EIE

[222] Carl M. Bender and Joshua Feinberg. Eigenvalue integrodifferential equations for orthogonal polynomials on the real line. Journal of Mathematical Physics, 36(6):3106–3125, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Constantinescu:1995:UBG

[223] Florin Constantinescu. Unitary braid group representations. Journal of Mathematical Physics, 36(6):3126–3133, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Floreanini:1995:QSD

[224] Roberto Floreanini and Luc Vinet. Quantum symmetries of q-difference equations. Journal of Mathematical Physics, 36(6):3134–3156, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Millane:1995:ISR

[225] R. P. Millane. An infinite sum related to WKB eigenvalues. Journal of Mathematical Physics, 36(6):3157–3160, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scott:1995:NIF

[226] S. G. Scott. A note on CP^{n}-instantons and the family’s index theorem. Journal of Mathematical Physics, 36(6):3161–3167, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simon:1995:SDS

[227] David S. Simon. Semiclassical determinants of Schrödinger operators on line bundles. Journal of Mathematical Physics, 36(6):3168–3173, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vassilevich:1995:VFD

[228] D. V. Vassilevich. Vector fields on a disk with mixed boundary conditions. Journal of Mathematical Physics, 36(6):3174–3182, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fatibene:1995:ENF

[229] L. Fatibene, M. Ferraris, and M. Francaviglia. Erratum: “Nöther formalism for conserved quantities in classical gauge field theories”, [J. Math. Phys. 35 (1994), no. 4, 1644–1657; MR 95g:58064]. Journal of Mathematical Physics, 36(6):3183–3185, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [7].

LaChapelle:1995:EEG

[230] J. LaChapelle. Erratum: “U(2) as the electroweak gauge group”, [J. Math. Phys. 35 (1994), no. 5, 2186–2198; MR 95e:81253]. Journal of Mathematical Physics, 36(6):3186, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [9].

LaChapelle:1995:EGM

[231] J. LaChapelle. Erratum: “Generating mass without the Higgs particle”, [J. Math. Phys. 35 (1994), no. 5, 2199–2209; MR 95e:81257]. Journal of Mathematical Physics, 36(6):3187, June 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [10].

Aldaya:1995:CCS

[232] V. Aldaya and J. Guerrero. Canonical coherent states for the relativistic harmonic oscillator. Journal of Mathematical Physics, 36(7):3191–3199, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2271].

Bertoglio:1995:KDH

[233] Néstor Bertoglio and Claudio Fernández. Kicked dynamics on Hilbert spaces: a nonstandard approach. Journal of Mathematical Physics, 36 (7):3200–3209, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doebner:1995:RRD

[234] H.-D. Doebner and H.-J. Mann. Ray representations of (N(≤ 2) + 1)- dimensional Galilean group. Journal of Mathematical Physics, 36(7): 3210–3215, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferrari:1995:OFS

[235] F. Ferrari, J. Sobczyk, and W. Urbanik. Operator formalism on the Z_{n} symmetric algebraic curves. Journal of Mathematical Physics, 36 (7):3216–3231, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Funahashi:1995:CSP

[236] Kunio Funahashi, Tarō Kashiwa, Seiji Sakoda, and Kazuyuki Fujii. Coherent states, path integral, and semiclassical approximation. Journal of Mathematical Physics, 36(7):3232–3253, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hwang:1995:TYM

[237] Dae Sung Hwang and Chang-Yeong Lee. Topological Yang–Mills theory from non-commutative geometry. Journal of Mathematical Physics, 36 (7):3254–3262, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Land:1995:FAF

[238] M. C. Land, N. Shnerb, and L. P. Horwitz. On Feynman’s approach to the foundations of gauge theory. Journal of Mathematical Physics, 36 (7):3263–3288, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marnelius:1995:BQG

[239] Robert Marnelius and Ulrich Quaade. BRST quantization of gauge theories like SL(2, R) on inner product spaces. Journal of Mathematical Physics, 36(7):3289–3307, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajeev:1995:SRR

[240] S. G. Rajeev and L. Rossi. Some rigorous results for Yang–Mills theories on a cylinder. Journal of Mathematical Physics, 36(7):3308–3319, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Verdiev:1995:QIS

[241] Yi. A. Verdiev. Quantum integrable systems related with symmetric spaces of the groups U(1, 2) and Sp(1, 2) and Green functions on these spaces. Journal of Mathematical Physics, 36(7):3320–3331, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Villalba:1995:ESD

[242] Víctor M. Villalba. Exact solution of the Dirac equation for a Coulomb and scalar potential in the presence of an Aharonov–Bohm and a magnetic monopole fields. Journal of Mathematical Physics, 36(7):3332–3344, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chiang:1995:EIA

[243] R. C. Chiang. Effects of an internal angular momentum on the rotation of a symmetrical top. Journal of Mathematical Physics, 36(7):3345– 3352, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gratus:1995:TCP

[244] Jonathan Gratus and Robin W. Tucker. Topology change and the propagation of massless fields. Journal of Mathematical Physics, 36(7):3353– 3376, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lehner:1995:DCM

[245] Luis R. Lehner and Osvaldo M. Moreschi. On the definition of the center of mass for a system of relativistic particles. Journal of Mathematical Physics, 36(7):3377–3394, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Patch:1995:RRF

[246] S. K. Patch. Recursive recovery of a family of Markov transition probabilities from boundary value data. Journal of Mathematical Physics, 36 (7):3395–3412, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1995:DPH

[247] G. F. Torres del Castillo. Debye potentials and Hamiltonians for magnetic field line flow. Journal of Mathematical Physics, 36(7):3413–3418, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See addendum [634].

Aratyn:1995:ALA

[248] H. Aratyn, J. F. Gomes, and A. H. Zimerman. Affine Lie algebraic origin of constrained KP hierarchies. Journal of Mathematical Physics, 36(7):3419–3442, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

BoutetdeMonvel:1995:ISP

[249] A. Boutet de Monvel and D. Shepelsky. Inverse scattering problem for anisotropic media. Journal of Mathematical Physics, 36(7):3443–3453, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bu:1995:MKV

[250] Charles Bu. Modified Korteweg–de Vries equation with generalized functions as initial values. Journal of Mathematical Physics, 36(7):3454–3460, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gawrylczyk:1995:RLI

[251] Jarosław Gawrylczyk. The Riemann–Liouville integral and continuous KP hierarchy. Journal of Mathematical Physics, 36(7):3461–3478, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:1995:GEN

[252] Bo Ling Guo and Yong Hui Wu. Global existence and nonexistence of the solution of a forced nonlinear Schrödinger equation. Journal of Mathematical Physics, 36(7):3479–3484, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurses:1995:VCT

[253] Metin Gürses and Atalay Karasu. Variable coefficient third order Korteweg–de Vries type of equations. Journal of Mathematical Physics, 36(7):3485–3491, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:1995:GSA

[254] Sen Yue Lou and Jian Ping Weng. Generalized W_{∞} symmetry algebra of the conditionally integrable nonlinear evolution equation. Journal of Mathematical Physics, 36(7):3492–3497, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2061] and response [2062].

Parker:1995:ESRa

[255] A. Parker. On exact solutions of the regularized long-wave equation: a direct approach to partially integrable equations. I. Solitary wave and solitons. Journal of Mathematical Physics, 36(7):3498–3505, July 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parker:1995:ESRb

[256] A. Parker. On exact solutions of the regularized long-wave equation: a direct approach to partially integrable equations. II. Periodic solutions. Journal of Mathematical Physics, 36(7):3506–3519, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Plank:1995:HSD

[257] Manfred Plank. Hamiltonian structures for the n-dimensional Lotka– Volterra equations. Journal of Mathematical Physics, 36(7):3520–3534, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comments [735, 1008].

Qiao:1995:CIS

[258] Zhi Jun Qiao. A completely integrable system and parametric representation of solutions of the Wadati–Konno–Ichikawa hierarchy. Journal of Mathematical Physics, 36(7):3535–3540, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:1995:ITP

[259] Manuel F. Rañada. Integrable three-particle systems, hidden symmetries and deformations of the Calogero–Moser system. Journal of Mathematical Physics, 36(7):3541–3558, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Romeiras:1995:SLP

[260] Filipe J. Romeiras. Separability and Lax pairs for the two-dimensional Hamiltonian system with a quartic potential. Journal of Mathematical Physics, 36(7):3559–3565, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strachan:1995:SSA

[261] I. A. B. Strachan. The symmetry structure of the anti-self-dual Einstein hierarchy. Journal of Mathematical Physics, 36(7):3566–3573, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takasaki:1995:STF

[262] Kanehisa Takasaki. Symmetries and tau function of higher-dimensional dispersionless integrable hierarchies. Journal of Mathematical Physics, 36(7):3574–3607, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanikawa:1995:SUM

[263] Kiyotaka Tanikawa and Yoshihiro Yamaguchi. Stable and unstable manifolds in a zone of instability. II. Journal of Mathematical Physics, 36 (7):3608–3612, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alty:1995:BBT

[264] L. J. Alty. Building blocks for topology change. Journal of Mathematical Physics, 36(7):3613–3618, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banerjee:1995:EKK

[265] A. Banerjee, D. Panigrahi, and S. Chatterjee. Evolution of Kaluza– Klein inhomogeneous model with a cosmological constant. Journal of Mathematical Physics, 36(7):3619–3624, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daSilva:1995:LCS

[266] M. F. A. da Silva, L. Herrera, F. M. Paiva, and N. O. Santos. The Levi-Civita space–time. Journal of Mathematical Physics, 36(7):3625– 3631, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frauendiener:1995:CDS

[267] J. Frauendiener and C. Klein. On crossing dust shells. Journal of Mathematical Physics, 36(7):3632–3643, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heller:1995:SST

[268] Michael Heller and Wiesław Sasin. Structured spaces and their application to relativistic physics. Journal of Mathematical Physics, 36(7):3644– 3662, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holvorcem:1995:IOI

[269] Paulo R. Holvorcem, Patricio S. Letelier, and An Zhong Wang. The interaction of outgoing and ingoing spherically symmetric null fluids. Journal of Mathematical Physics, 36(7):3663–3675, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kriele:1995:CSS

[270] Marcus Kriele. On the collapse of a spherically symmetric star. Journal of Mathematical Physics, 36(7):3676–3693, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Linet:1995:ESG

[271] B. Linet. Euclidean spinor Green’s functions in the space–time of a straight cosmic string. Journal of Mathematical Physics, 36(7):3694– 3703, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mason:1995:VBE

[272] L. J. Mason. The vacuum and Bach equations in terms of light cone cuts. Journal of Mathematical Physics, 36(7):3704–3721, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moffat:1995:NGT

[273] J. W. Moffat. Nonsymmetric gravitational theory. Journal of Mathematical Physics, 36(7):3722–3732, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [492].

Robinson:1995:LFE

[274] D. C. Robinson. A Lagrangian formalism for the Einstein–Yang–Mills equations. Journal of Mathematical Physics, 36(7):3733–3742, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoine:1995:SDP

[275] J.-P. Antoine, A. Inoue, and C. Trapani. Spatiality of ^{∗}-derivations of partial O^{∗}-algebras. Journal of Mathematical Physics, 36(7):3743–3758, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhutani:1995:IES

[276] O. P. Bhutani and Lipika Bhattacharya. Isogroups and exact solutions for some Klein–Gordon and Liouville-type equations in n-dimensional Euclidean space. Journal of Mathematical Physics, 36(7):3759–3770, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakis:1995:DDM

[277] Aristophanes Dimakis, Folkert Müller-Hoissen, and François Vanderseypen. Discrete differential manifolds and dynamics on networks. Journal of Mathematical Physics, 36(7):3771–3791, July 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Elshamy:1995:DWE

[278] Maged Elshamy. Damped wave equations with noise. Journal of Mathematical Physics, 36(7):3792–3799, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Floreanini:1995:MCU

[279] Roberto Floreanini and Luc Vinet. A model for the continuous qultraspherical polynomials. Journal of Mathematical Physics, 36(7):

3800–3813, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jurco:1995:FSR

[280] B. Jurčo and M. Schlieker. On Fock-space representations of quantized enveloping algebras related to noncommutative differential geometry. Journal of Mathematical Physics, 36(7):3814–3821, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langmann:1995:NIC

[281] Edwin Langmann. Noncommutative integration calculus. Journal of Mathematical Physics, 36(7):3822–3835, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenczewski:1995:CSF

[282] Romuald Lenczewski. On certain states on free ^{∗}-algebras and their GNS representations. Journal of Mathematical Physics, 36(7):3836–3853, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:1995:RSY

[283] R. B. Zhang. Representations of super Yangian. Journal of Mathematical Physics, 36(7):3854–3865, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davis:1995:EDM

[284] Simon Davis. Erratum: “Divergences in the moduli space integral and accumulating handles in the infinite-genus limit”. Journal of Mathematical Physics, 36(7):3866, July 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [60].

Briet:1995:BWF

[285] Philippe Briet. Bender–Wu formula for the Zeeman effect. Journal of Mathematical Physics, 36(8):3871–3882, August 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Debergh:1995:GFL

[286] N. Debergh. Generalized fermions and Lie superalgebras spl(m/n). Journal of Mathematical Physics, 36(8):3883–3888, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunne:1995:DMS

[287] Robert S. Dunne. Deformation methods for single paraoscillator systems. Journal of Mathematical Physics, 36(8):3889–3901, August 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Enss:1995:GAM

[288] Volker Enss and Ricardo Weder. The geometrical approach to multidimensional inverse scattering. Journal of Mathematical Physics, 36(8): 3902–3921, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:1995:RPS

[289] Francisco M. Fernández. Renormalized perturbation series and the semiclassical limit of quantum mechanics. Journal of Mathematical Physics, 36(8):3922–3930, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grigore:1995:CFF

[290] D. R. Grigore. On the construction of free fields in relativistic quantum mechanics. Journal of Mathematical Physics, 36(8):3931–3939, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:1995:LSC

[291] D. Han, Y. S. Kim, and Marilyn E. Noz. O(3, 3)-like symmetries of coupled harmonic oscillators. Journal of Mathematical Physics, 36(8): 3940–3954, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kwon:1995:SCS

[292] P. S. Kwon and M. Villasante. Structure of the chiral scalar superfield in ten dimensions. Journal of Mathematical Physics, 36(8):3955–3992, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nishino:1995:APW

[293] Kazuyoshi Nishino. An application of the Peter–Weyl theorem to nonabelian lattice gauge theory. Journal of Mathematical Physics, 36(8): 3993–4004, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Romeo:1995:MEW

[294] August Romeo. Multidimensional extension of a Wentzel–Kramers– Brillouin improvement for spherical quantum billiard zeta functions. Journal of Mathematical Physics, 36(8):4005–4011, August 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:1995:OAN

[295] Chengjun Zhu and John R. Klauder. Operator analysis of nonrenormalizable multicomponent ultralocal field models. Journal of Mathematical Physics, 36(8):4012–4019, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:1995:NPI

[296] Chengjun Zhu and John R. Klauder. Nontrivial path integrals for nonrenormalizable fields—multicomponent ultralocal models. Journal of Mathematical Physics, 36(8):4020–4027, August 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:1995:OAT

[297] Sailing He and Sergei I. Kabanikhin. An optimization approach to a three-dimensional acoustic inverse problem in the time domain. Journal of Mathematical Physics, 36(8):4028–4043, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herdegen:1995:LRE

[298] Andrzej Herdegen. Long-range effects in asymptotic fields and angular momentum of classical field electrodynamics. Journal of Mathematical Physics, 36(8):4044–4086, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miano:1995:NMS

[299] G. Miano, L. Verolino, and V. G. Vaccaro. A new method of solution of Hallén’s problem. Journal of Mathematical Physics, 36(8):4087–4099, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Charret:1995:MGI

[300] I. C. Charret, E. V. Corrêa Silva, S. M. de Souza, and M. T. Thomaz. Multivariable Grassmann integrals and the high temperature expansion of interacting fermionic models. Journal of Mathematical Physics, 36(8): 4100–4105, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanAssche:1995:EOP

[301] W. Van Assche, R. J. Yañez, and J. S. Dehesa. Entropy of orthogonal polynomials and Freud weights and information entropies of the harmonic oscillator potential. Journal of Mathematical Physics, 36(8): 4106–4118, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boling:1995:GSS

[302] Guo Boling and Yuan Guangwei. Global smooth solution for the Klein– Gordon–Zakharov equations. Journal of Mathematical Physics, 36(8): 4119–4124, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chow:1995:CEP

[303] Kwok W. Chow. A class of exact, periodic solutions of nonlinear envelope equations. Journal of Mathematical Physics, 36(8):4125–4137, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:1995:SCM

[304] Manuel de León and David Martín de Diego. Symmetries and constants of the motion for higher-order Lagrangian systems. Journal of Mathematical Physics, 36(8):4138–4161, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kajiwara:1995:CDS

[305] Kenji Kajiwara, Yasuhiro Ohta, and Junkichi Satsuma. Casorati determinant solutions for the discrete Painlevé III equation. Journal of Mathematical Physics, 36(8):4162–4174, August 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamvissis:1995:FNS

[306] Spyridon Kamvissis. Focusing nonlinear Schrödinger equation with infinitely many solitons. Journal of Mathematical Physics, 36(8):4175– 4180, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Knecht:1995:PAI

[307] M. Knecht, R. Pasquier, and J.-Y. Pasquier. Painlevé analysis and integrability properties of a 2 + 1 nonrelativistic field theory. Journal of Mathematical Physics, 36(8):4181–4191, August 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kundu:1995:DHO

[308] Anjan Kundu and Walter Strampp. Derivative and higher-order extensions of Davey–Stewartson equation from matrix Kadomtsev– Petviashvili hierarchy. Journal of Mathematical Physics, 36(8):4192– 4202, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pelinovsky:1995:STI

[309] Dmitry E. Pelinovsky and Roger H. J. Grimshaw. A spectral transform for the intermediate nonlinear Schrödinger equation. Journal of Mathematical Physics, 36(8):4203–4219, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:1995:HNO

[310] Ru Guang Zhou. Hierarchy of negative order equation and its Lax pair. Journal of Mathematical Physics, 36(8):4220–4225, August 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geroch:1995:RTD

[311] Robert Geroch. Relativistic theories of dissipative fluids. Journal of Mathematical Physics, 36(8):4226–4241, August 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maluf:1995:LEG

[312] José W. Maluf. Localization of energy in general relativity. Journal of Mathematical Physics, 36(8):4242–4247, August 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Papadopoulos:1995:TSE

[313] Demetrios B. Papadopoulos and Basilis C. Xanthopoulos. Tomimatsu– Sato electrovacuum cosmic string solutions for δ = 2 interacting with gravitational waves. Journal of Mathematical Physics, 36(8):4248–4262, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petters:1995:MGLa

[314] A. O. Petters. Multiplane gravitational lensing. I. Morse theory and image counting. Journal of Mathematical Physics, 36(8):4263–4275, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petters:1995:MGLb

[315] A. O. Petters. Multiplane gravitational lensing. II. Global geometry of caustics. Journal of Mathematical Physics, 36(8):4276–4295, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Portugal:1995:ESD

[316] Renato Portugal. Exact solutions of the Dirac equations in an anisotropic cosmological background. Journal of Mathematical Physics, 36(8):4296– 4300, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smoller:1995:RSE

[317] J. A. Smoller and A. G. Wasserman. Regular solutions of the Einstein– Yang–Mills equations. Journal of Mathematical Physics, 36(8):4301– 4323, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berest:1995:FSP

[318] Yuri Berest and Yuri Molchanov. Fundamental solutions for partial differential equations with reflection group invariance. Journal of Mathematical Physics, 36(8):4324–4339, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:1995:SQE

[319] Y. Brihaye, S. Giller, C. Gonera, and P. Kosiński. The structure of quasi-exactly solvable systems. Journal of Mathematical Physics, 36(8): 4340–4349, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coll:1995:CTF

[320] Bartolomé Coll and Fernando San José Martínez. On the commutator of two 2-forms in Minkowski space–time. Journal of Mathematical Physics, 36(8):4350–4362, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fiore:1995:EHA

[321] Gaetano Fiore. The Euclidean Hopf algebra U_{q}(e^{N}) and its fundamental Hilbert-space representations. Journal of Mathematical Physics, 36(8): 4363–4405, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gross:1995:TSB

[322] Christian Gross. Topology of the skyrmion bundle. Journal of Mathematical Physics, 36(8):4406–4435, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:1995:SBS

[323] S. Majid. ∗-structures on braided spaces. Journal of Mathematical Physics, 36(8):4436–4449, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McInnes:1995:HGH

[324] Brett McInnes. Holonomy groups and holonomy representations. Journal of Mathematical Physics, 36(8):4450–4460, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saidi:1995:MRH

[325] E. H. Saidi. Matrix representation of higher integer conformal spin symmetries. Journal of Mathematical Physics, 36(8):4461–4475, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sasaki:1995:QLG

[326] Norihito Sasaki. Quantization of Lie group and algebra of G_{2} type in the Faddeev–Reshetikhin–Takhtajan approach. Journal of Mathematical Physics, 36(8):4476–4488, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tolar:1995:GCC

[327] J. Tolar and P. Trávníček. Graded contractions and the conformal group of Minkowski space–time. Journal of Mathematical Physics, 36(8):4489– 4506, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanderJeugt:1995:PDG

[328] J. Van der Jeugt and R. Jagannathan. Polynomial deformations of osp(1/2) and generalized parabosons. Journal of Mathematical Physics, 36(8):4507–4518, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weimar-Woods:1995:CLA

[329] Evelyn Weimar-Woods. Contractions of Lie algebras: generalized Inönü-Wigner contractions versus graded contractions. Journal of Mathematical Physics, 36(8):4519–4548, August 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nash:1995:EDL

[330] Charles Nash and Denjoe O’Connor. Erratum: “Determinants of Laplacians, the Ray–Singer torsion on lens spaces and the Riemann zeta function”, [J. Math. Phys. 36 (1995), no. 3, 1462–1505; MR 95k:58173]. Journal of Mathematical Physics, 36(8):4549, August 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [123].

Amann:1995:NTC

[331] Herbert Amann and Pavol Quittner. A nodal theorem for coupled systems of Schrödinger equations and the number of bound states. Journal of Mathematical Physics, 36(9):4553–4560, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:1995:AAC

[332] P. Exner. The absence of the absolutely continuous spectrum for δ′ Wannier–Stark ladders. Journal of Mathematical Physics, 36(9):4561– 4570, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Filippov:1995:CBP

[333] G. F. Filippov, I. Yu. Rybkin, S. V. Korennov, and K. Katō. On the complete basis of Pauli-allowed states of three-cluster systems in the Fock–Bargmann space. Journal of Mathematical Physics, 36(9):4571– 4589, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Funahashi:1995:EWK

[334] Kunio Funahashi, Tarō Kashiwa, Seiji Sakoda, and Kazuyuki Fujii. Exactness in the Wentzel–Kramers–Brillouin approximation for some homogeneous spaces. Journal of Mathematical Physics, 36(9):4590–4611, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hajicek:1995:GQPa

[335] P. Hájíček. Group quantization of parametrized systems. I. Time levels. Journal of Mathematical Physics, 36(9):4612–4638, September 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Hajicek:1995:GQPb

[336] P. Hájíček, A. Higuchi, and J. Tolar. Group quantization of parametrized systems. II. Pasting Hilbert spaces. Journal of Mathematical Physics, 36 (9):4639–4666, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kochetov:1995:CSP

[337] E. A. Kochetov. SU(2) coherent-state path integral. Journal of Mathematical Physics, 36(9):4667–4679, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martin:1995:BCT

[338] A. Martin and J. Stubbe. Bargmann- and Calogero-type bounds for the Dirac equation. Journal of Mathematical Physics, 36(9):4680–4690, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morato:1995:MDC

[339] Laura M. Morato and Lorenza Viola. Markov diffusions in comoving coordinates and stochastic quantization of the free relativistic spinless particle. Journal of Mathematical Physics, 36(9):4691–4710, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [815].

Rowe:1995:RFD

[340] D. J. Rowe and K. T. Hecht. Representation of the five-dimensional harmonic oscillator with scalar-valued U(5) ⊃ SO(5) ⊃ SO(3)-coupled VCS wave functions. Journal of Mathematical Physics, 36(9):4711–4734, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schreckenberg:1995:CDF

[341] Stephan Schreckenberg. Completeness of decoherence functionals. Journal of Mathematical Physics, 36(9):4735–4742, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanLeeuwen:1995:DGD

[342] Hans van Leeuwen and Hans Maassen. A q deformation of the Gauss distribution. Journal of Mathematical Physics, 36(9):4743–4756, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vourdas:1995:CSS

[343] A. Vourdas. SU(2) coherent states on an m-sheeted covering of the sphere. Journal of Mathematical Physics, 36(9):4757–4773, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Angelescu:1995:LDS

[344] N. Angelescu, M. Bundaru, and J. V. Pulé. Large deviations in a semiinfinite layered ferromagnet with mean-field interactions: classical and quantum models. Journal of Mathematical Physics, 36(9):4774–4784, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Raggio:1995:PE

[345] G. A. Raggio. Properties of q-entropies. Journal of Mathematical Physics, 36(9):4785–4791, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beyer:1995:SAS

[346] Horst R. Beyer. The spectrum of adiabatic stellar oscillations. Journal of Mathematical Physics, 36(9):4792–4814, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beyer:1995:SRA

[347] Horst R. Beyer. The spectrum of radial adiabatic stellar oscillations. Journal of Mathematical Physics, 36(9):4815–4825, September 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaszak:1995:BHF

[348] Maciej Błaszak. Bi-Hamiltonian formulation for the Korteweg–de Vries hierarchy with sources. Journal of Mathematical Physics, 36(9):4826– 4831, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1995:SHS

[349] F. Calogero. A solvable Hamiltonian system. Journal of Mathematical Physics, 36(9):4832–4840, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Claude:1995:STW

[350] Christophe Claude. Solution of three waves interaction type models with nontrivial asymptotic and boundary conditions. Journal of Mathematical Physics, 36(9):4841–4862, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harnad:1995:MCE

[351] J. Harnad and M. Routhier. R-matrix construction of electromagnetic models for the Painlevé transcendents. Journal of Mathematical Physics, 36(9):4863–4881, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hlavaty:1995:AFQ

[352] Ladislav Hlavatý. Algebraic framework for quantization of nonultralocal models. Journal of Mathematical Physics, 36(9):4882–4897, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

MartinezAlonso:1995:TFF

[353] Luis Martínez Alonso and Elena Medina Reus. Tau-function formalism for supersymmetric KP hierarchies. Journal of Mathematical Physics, 36(9):4898–4913, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moorhouse:1995:TID

[354] T. Moorhouse. A twistor-inspired Darboux transform for the sine-Gordon equation. Journal of Mathematical Physics, 36(9):4914–4922, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moroz:1995:PAD

[355] Irene M. Moroz and Paul K. Newton. Phase-amplitude dynamics of the nonlinear Schrödinger equation with rapid forcing. Journal of Mathematical Physics, 36(9):4923–4939, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandeLeur:1995:ASV

[356] Johan van de Leur. The Adler–Shiota–van Moerbeke formula for the BKP hierarchy. Journal of Mathematical Physics, 36(9):4940–4951, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Browdy:1995:TCT

[357] Steven F. Browdy and Gregory J. Galloway. Topological censorship and the topology of black holes. Journal of Mathematical Physics, 36(9):

4952–4961, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feinstein:1995:SFI

[358] A. Feinstein, J. Ibañez, and P. Labraga. Scalar field inhomogeneous cosmologies. Journal of Mathematical Physics, 36(9):4962–4974, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:1995:LMC

[359] Simonetta Frittelli, Carlos Kozameh, and Ezra T. Newman. Lorentzian metrics from characteristic surfaces. Journal of Mathematical Physics, 36(9):4975–4983, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1886].

Frittelli:1995:GCS

[360] Simonetta Frittelli, Carlos Kozameh, and Ezra T. Newman. GR via characteristic surfaces. Journal of Mathematical Physics, 36(9):4984– 5004, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1887].

Frittelli:1995:LET

[361] Simonetta Frittelli, Carlos N. Kozameh, and Ezra T. Newman. Linearized Einstein theory via null surfaces. Journal of Mathematical Physics, 36(9):5005–5022, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1888].

Galtsov:1995:SSE

[362] Dmitri Gal′tsov, Alberto García, and Oleg Kechkin. Symmetries of the stationary Einstein–Maxwell-dilaton-axion system. Journal of Mathematical Physics, 36(9):5023–5041, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramidi:1995:HKC

[363] I. G. Avramidi and R. Schimming. Heat kernel coefficients for the matrix Schrödinger operator. Journal of Mathematical Physics, 36(9):5042– 5054, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramidi:1995:CAM

[364] I. G. Avramidi. Covariant algebraic method for calculation of the lowenergy heat kernel. Journal of Mathematical Physics, 36(9):5055–5070, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1477].

Brankov:1995:IDC

[365] J. G. Brankov. Isomorphism of dimer configurations and spanning trees on finite square lattices. Journal of Mathematical Physics, 36(9):5071– 5083, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsenko:1995:CAD

[366] A. A. Bytsenko, E. Elizalde, and S. D. Odintsov. The conformal anomaly in N-dimensional spaces having a hyperbolic spatial section. Journal of Mathematical Physics, 36(9):5084–5090, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Floreanini:1995:AIC

[367] Roberto Floreanini, Jean LeTourneux, and Luc Vinet. An algebraic interpretation of the continuous big q-Hermite polynomials. Journal of Mathematical Physics, 36(9):5091–5097, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gervois:1995:STI

[368] A. Gervois and M. L. Mehta. Some trigonometric identities encountered by McCoy and Orrick. Journal of Mathematical Physics, 36(9):5098– 5109, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hou:1995:DCC

[369] Bo Yu Hou, Bo Yuan Hou, and Zhong Qi Ma. q-deformed Chern characters for quantum groups SU_{q}(N). Journal of Mathematical Physics, 36(9):5110–5138, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katriel:1995:FIH

[370] J. Katriel, B. Abdesselam, and A. Chakrabarti. The fundamental invariant of the Hecke algebra H_{n}(q) characterizes the representations of H_{n}(q), S_{n}, SU_{q}(N), and SU(N). Journal of Mathematical Physics, 36 (9):5139–5158, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matviyenko:1995:AFC

[371] Gregory Matviyenko. On the azimuthal Fourier components of the Green’s function for the Helmholtz equation in three dimensions. Journal of Mathematical Physics, 36(9):5159–5169, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mukunda:1995:MCF

[372] N. Mukunda and R. Simon. Manifestly covariant forms for Lorentz group unitary representations. Journal of Mathematical Physics, 36(9):5170– 5178, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zarzo:1995:ASP

[373] A. Zarzo, R. J. Yañez, A. Ronveaux, and J. S. Dehesa. Algebraic and spectral properties of some quasiorthogonal polynomials encountered in quantum radiation. Journal of Mathematical Physics, 36(9):5179–5197, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chu:1995:EPF

[374] Wen Chang Chu. Erratum: “Partial fractions and bilateral summations”, [J. Math. Phys. 35 (1994), no. 4, 2036–2042; MR 95f:33004]. Journal of Mathematical Physics, 36(9):5198–5199, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [8].

Gowdy:1995:EAP

[375] Robert H. Gowdy. Erratum: “Affine projection-tensor geometry: Lie derivatives and isometries”, [J. Math. Phys. 36 (1995), no. 4, 1882–1907; MR 97d:53029]. Journal of Mathematical Physics, 36(9):5200, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [153].

Slomczynski:1995:EQC

[376] Wojciech Słomczyński and Karol Życzkowski. Erratum: “Quantum chaos: an entropy approach”, [J. Math. Phys. 35 (1994), no. 11, 5674– 5700; MR 95i:81062]. Journal of Mathematical Physics, 36(9):5201, September 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [15].

Albeverio:1995:RPS

[377] S. Albeverio and Zheng Dong Wang. Representation of the propagator and Schwinger functions of Dirac fields in terms of Brownian motions. Journal of Mathematical Physics, 36(10):5207–5216, October 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Albeverio:1995:ERF

[378] Sergio Albeverio and Jiang Lun Wu. Euclidean random fields obtained by convolution from generalized white noise. Journal of Mathematical Physics, 36(10):5217–5245, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

IommiAmunategui:1995:CRG

[379] G. Iommi Amunátegui. The L-S coupling revisited: a general solution for the 4 electron system. Journal of Mathematical Physics, 36(10):5246– 5261, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arnaudon:1995:CQS

[380] Daniel Arnaudon, Chryssomalis Chryssomalakos, and Luc Frappat. Classical and quantum sl(1|2) superalgebras, Casimir operators and quantum chain Hamiltonians. Journal of Mathematical Physics, 36(10):5262–5283, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Axenides:1995:SCS

[381] Minos Axenides, Andrei Johansen, and Jesper Møller. A Z_{2} structure in the configuration space of Yang–Mills theories. Journal of Mathematical Physics, 36(10):5284–5296, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

BezerradeMello:1995:CPM

[382] E. R. Bezerra de Mello and V. B. Bezerra. Charged particles in a magnetic field on a cone. Journal of Mathematical Physics, 36(10):5297–5304, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boumenir:1995:CTO

[383] Amin Boumenir. Construction of a transformation operator. Journal of Mathematical Physics, 36(10):5305–5309, October 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fioravanti:1995:SIS

[384] R. Fioravanti and G. A. Viano. On the solution of the inverse scattering problem, at fixed energy, for the class of Yukawian potentials. Journal of Mathematical Physics, 36(10):5310–5339, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grabowski:1995:QCC

[385] M. P. Grabowski and P. Mathieu. Quantum chains with a Catalan tree pattern of conserved charges: the Δ = −1XXZ model and the isotropic octonionic chain. Journal of Mathematical Physics, 36(10):5340–5354, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurarie:1995:QNP

[386] David Gurarie. Quantized Neumann problem, separable potentials on S^{n} and the Lamé equation. Journal of Mathematical Physics, 36(10): 5355–5391, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isham:1995:CHH

[387] C. J. Isham and N. Linden. Continuous histories and the history group in generalized quantum theory. Journal of Mathematical Physics, 36(10): 5392–5408, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wright:1995:SDF

[388] J. D. Maitland Wright. The structure of decoherence functionals for von Neumann quantum histories. Journal of Mathematical Physics, 36(10): 5409–5413, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McInnes:1995:GDS

[389] Brett McInnes. Gauging discrete symmetries. Journal of Mathematical Physics, 36(10):5414–5430, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morehead:1995:ARW

[390] James J. Morehead. Asymptotics of radial wave equations. Journal of Mathematical Physics, 36(10):5431–5452, October 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:1995:GFA

[391] D. K. Park. Green’s-function approach to two- and three-dimensional delta-function potentials and application to the spin-1/2 Aharonov– Bohm problem. Journal of Mathematical Physics, 36(10):5453–5464, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Selesnick:1995:GFQ

[392] S. A. Selesnick. Gauge fields on the quantum net. Journal of Mathematical Physics, 36(10):5465–5479, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thomas:1995:SAS

[393] Lawrence E. Thomas and Stephen R. Wassell. Semiclassical approximation for Schrödinger operators on a two-sphere at high energy. Journal of Mathematical Physics, 36(10):5480–5505, October 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhdanov:1995:NAV

[394] R. Z. Zhdanov, I. V. Revenko, and W. I. Fushchych. On the new approach to variable separation in the time-dependent Schrödinger equation with two space dimensions. Journal of Mathematical Physics, 36(10):5506– 5521, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ibort:1995:CRS

[395] A. Ibort and J. Marín-Solano. Coisotropic regularization of singular Lagrangians. Journal of Mathematical Physics, 36(10):5522–5539, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lozada:1995:AMS

[396] Abraham Lozada. Angular momentum and space–time hyperboloids. Journal of Mathematical Physics, 36(10):5540–5551, October 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nesterenko:1995:DRP

[397] V. V. Nesterenko, A. Feoli, and G. Scarpetta. Dynamics of relativistic particles with Lagrangians dependent on acceleration. Journal of Mathematical Physics, 36(10):5552–5564, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shaarawi:1995:EFC

[398] Amr M. Shaarawi, Richard W. Ziolkowski, and Ioannis M. Besieris. On the evanescent fields and the causality of the focus wave modes. Journal of Mathematical Physics, 36(10):5565–5587, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1147].

Simoni:1995:LTL

[399] A. Simoni, A. Stern, and I. Yakushin. Lorentz transformations as Lie– Poisson symmetries. Journal of Mathematical Physics, 36(10):5588–5597, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alli:1995:NMS

[400] Giovanni Alli and Geoffrey L. Sewell. New methods and structures in the theory of the multimode Dicke laser model. Journal of Mathematical Physics, 36(10):5598–5626, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golden:1995:SMC

[401] K. M. Golden. Statistical mechanics of conducting phase transitions. Journal of Mathematical Physics, 36(10):5627–5642, October 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://scitation.aip.org/content/aip/journal/jmp/
36/10/10.1063/1.531280.

Miyazawa:1995:ASD

[402] Toru Miyazawa. Algebraic structure of diffusion and scattering in one dimension. Journal of Mathematical Physics, 36(10):5643–5669, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cieslinski:1995:AMC

[403] Jan Cieśliński. An algebraic method to construct the Darboux matrix. Journal of Mathematical Physics, 36(10):5670–5706, October 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finley:1995:IDE

[404] J. D. Finley, III and John K. McIver. Infinite-dimensional Estabrook– Wahlquist prolongations for the sine-Gordon equation. Journal of Mathematical Physics, 36(10):5707–5734, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gafiychuk:1995:VRP

[405] V. V. Gafiychuk and I. A. Lubashevskiĭ. Variational representation of the projection dynamics and random motion of highly dissipative systems. Journal of Mathematical Physics, 36(10):5735–5752, October 1995. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Scoufis:1995:IVP

[406] George Scoufis and Christopher M. Cosgrove. On the initial value problem for the modified Benjamin–Ono equation. Journal of Mathematical Physics, 36(10):5753–5759, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhivkov:1995:DFD

[407] Angel Zhivkov and Milena Stanislavova. On the dynamics of the fourdimensional rigid body in a quadratic potential field. Journal of Mathematical Physics, 36(10):5760–5788, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barreto:1995:HMR

[408] W. Barreto and L. Castillo. Homothetic motion in radiating and dissipative spheres. Journal of Mathematical Physics, 36(10):5789–5799, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barros:1995:CVS

[409] A. Barros and C. Romero. Cosmic vacuum strings and domain walls in Brans–Dicke theory of gravity. Journal of Mathematical Physics, 36(10): 5800–5804, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bengtsson:1995:NMS

[410] Ingemar Bengtsson. Note on massive spin 2 in curved space. Journal of Mathematical Physics, 36(10):5805–5811, October 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Farid:1995:CSP

[411] Taha Bin Farid, Asghar Qadir, and M. Ziad. The classification of static plane symmetric space–times according to their Ricci collineations. Journal of Mathematical Physics, 36(10):5812–5828, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gavrilov:1995:IPE

[412] V. R. Gavrilov, V. D. Ivashchuk, and V. N. Mel′nikov. Integrable pseudo-Euclidean Toda-like systems in multidimensional cosmology with multicomponent perfect fluid. Journal of Mathematical Physics, 36(10):5829– 5847, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haddow:1995:PMS

[413] Barry M. Haddow. Purely magnetic space–times. Journal of Mathematical Physics, 36(10):5848–5854, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kehagias:1995:LCE

[414] A. Kehagias, J. Madore, J. Mourad, and G. Zoupanos. Linear connections on extended space–time. Journal of Mathematical Physics, 36(10):5855– 5867, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Macias:1995:TWC

[415] Alfredo Macías, Eckehard W. Mielke, Hugo A. Morales-Técotl, and Romualdo Tresguerres. Torsion and Weyl convector [covector] in metricaffine models of gravity. Journal of Mathematical Physics, 36(10):5868– 5876, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Magli:1995:KSG

[416] Giulio Magli. Kerr–Schild gravitational fields in matter and the Kerrinterior problem in general relativity. Journal of Mathematical Physics, 36(10):5877–5896, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moffat:1995:RTN

[417] J. W. Moffat. Regularity theorems in the nonsymmetric gravitational theory. Journal of Mathematical Physics, 36(10):5897–5915, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ballesteros:1995:LBC

[418] A. Ballesteros, N. A. Gromov, F. J. Herranz, M. A. del Olmo, and M. Santander. Lie bialgebra contractions and quantum deformations of quasi-orthogonal algebras. Journal of Mathematical Physics, 36(10): 5916–5937, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:1995:SRM

[419] G. S. Hall and Barry M. Haddow. Some remarks on metric and Weyl connections. Journal of Mathematical Physics, 36(10):5938–5948, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harnett:1995:SLA

[420] G. Harnett. The space of Lie algebra 2-cocycles of an affine transformation group. Journal of Mathematical Physics, 36(10):5949–5967, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

PfefferJohnston:1995:RRH

[421] Carolyn Pfeffer Johnston. Remarks on resolutions of Hilbert–Schmidt operators. Journal of Mathematical Physics, 36(10):5968–5978, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ky:1995:FDR

[422] Nguyen Anh Ky and Nedialka I. Stoilova. Finite-dimensional representations of the quantum superalgebra U_{q}[gl(2/2)]. II. Nontypical representations at generic q. Journal of Mathematical Physics, 36(10):5979–6003, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Milson:1995:CQE

[423] Robert Milson. On the construction of quasi-exactly solvable Schrödinger operators on homogeneous spaces. Journal of Mathematical Physics, 36(10):6004–6027, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nilsson:1995:NRA

[424] Bengt E. W. Nilsson and Per Sundell. A new relation among Catalan numbers. Journal of Mathematical Physics, 36(10):6028–6042, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:1995:CSV

[425] Xiao Dong Sun, Shi Kun Wang, and Ke Wu. Classification of six-vertex-type solutions of the colored Yang–Baxter equation. Journal of Mathematical Physics, 36(10):6043–6063, October 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Atiyah:1995:QTG

[426] Michael Atiyah. Quantum theory and geometry. Journal of Mathematical Physics, 36(11):6069–6072, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baez:1995:HDA

[427] John C. Baez and James Dolan. Higher-dimensional algebra and topological quantum field theory. Journal of Mathematical Physics, 36(11): 6073–6105, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lawrence:1995:AEW

[428] R. J. Lawrence. Asymptotic expansions of Witten–Reshetikhin–Turaev invariants for some simple 3-manifolds. Journal of Mathematical Physics, 36(11):6106–6129, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sawin:1995:PIT

[429] Stephen Sawin. Path integration in two-dimensional topological quantum field theory. Journal of Mathematical Physics, 36(11):6130–6136, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cattaneo:1995:TTD

[430] Alberto S. Cattaneo, Paolo Cotta-Ramusino, Jürg Fröhlich, and Maurizio Martellini. Topological BF theories in 3 and 4 dimensions. Journal of Mathematical Physics, 36(11):6137–6160, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barrett:1995:QGT

[431] John W. Barrett. Quantum gravity as topological quantum field theory. Journal of Mathematical Physics, 36(11):6161–6179, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crane:1995:CCQ

[432] Louis Crane. Clock and category: is quantum gravity algebraic. Journal of Mathematical Physics, 36(11):6180–6193, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Connes:1995:NGR

[433] Alain Connes. Noncommutative geometry and reality. Journal of Mathematical Physics, 36(11):6194–6231, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iochum:1995:FMR

[434] Bruno Iochum, Daniel Kastler, and Thomas Schücker. Fuzzy mass relations for the Higgs. Journal of Mathematical Physics, 36(11):6232–6254, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chamseddine:1995:GSC

[435] A. H. Chamseddine, J. Fröhlich, and O. Grandjean. The gravitational sector in the Connes–Lott formulation of the standard model. Journal of Mathematical Physics, 36(11):6255–6275, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Williams:1995:SQG

[436] Ruth M. Williams. Simplicial quantum gravity in four dimensions and invariants of discretized manifolds. Journal of Mathematical Physics, 36 (11):6276–6287, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwasaki:1995:DPR

[437] Junichi Iwasaki. A definition of the Ponzano–Regge quantum gravity model in terms of surfaces. Journal of Mathematical Physics, 36(11): 6288–6298, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ambjorn:1995:DTG

[438] J. Ambjørn, J. Jurkiewicz, and Y. Watabiki. Dynamical triangulations, a gateway to quantum gravity. Journal of Mathematical Physics, 36(11): 6299–6339, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brugmann:1995:MCS

[439] Bernd Brügmann and Enzo Marinari. Monte Carlo simulations of 4d simplicial quantum gravity. Journal of Mathematical Physics, 36(11): 6340–6352, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carfora:1995:HEE

[440] M. Carfora and A. Marzuoli. Holonomy and entropy estimates for dynamically triangulated manifolds. Journal of Mathematical Physics, 36 (11):6353–6376, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Susskind:1995:WH

[441] Leonard Susskind. The world as a hologram. Journal of Mathematical Physics, 36(11):6377–6396, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:1995:DCS

[442] Simonetta Frittelli, Ezra T. Newman, and Carlos N. Kozameh. On the dynamics of characteristic surfaces. Journal of Mathematical Physics, 36 (11):6397–6416, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1885].

Smolin:1995:LTQ

[443] Lee Smolin. Linking topological quantum field theory and nonperturbative quantum gravity. Journal of Mathematical Physics, 36(11):6417– 6455, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashtekar:1995:QDI

[444] Abhay Ashtekar, Jerzy Lewandowski, Donald Marolf, José Mourão, and Thomas Thiemann. Quantization of diffeomorphism invariant theories of connections with local degrees of freedom. Journal of Mathematical Physics, 36(11):6456–6493, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loll:1995:QAG

[445] R. Loll. Quantum aspects of 2 + 1 gravity. Journal of Mathematical Physics, 36(11):6494–6509, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DiBartolo:1995:KPS

[446] Cayetano Di Bartolo, Rodolfo Gambini, Jorge Griego, and Jorge Pullin. Knot polynomial states of quantum gravity in terms of loops and extended loops: some remarks. Journal of Mathematical Physics, 36(11): 6510–6528, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rovelli:1995:OGC

[447] Carlo Rovelli. Outline of a generally covariant quantum field theory and a quantum theory of gravity. Journal of Mathematical Physics, 36(11): 6529–6547, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rovelli:1995:OPV

[448] Carlo Rovelli and Lee Smolin. Overview. Journal of Mathematical Physics, 36(11):viii–ix, November 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:1995:NSO

[449] Stephen C. Anco. New spin-one gauge theory in three dimensions. Journal of Mathematical Physics, 36(12):6553–6565, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arthur:1995:VSC

[450] K. Arthur and D. H. Tchrakian. Vortex solutions in a class of 2-d symmetry breaking Skyrme-like models. Journal of Mathematical Physics, 36(12):6566–6582, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1995:JCM

[451] Xin Chen, Le Man Kuang, and Mo Lin Ge. Jaynes–Cummings model governed by the Milburn equations without diffusion approximation. Journal of Mathematical Physics, 36(12):6583–6595, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Combescure:1995:SSR

[452] M. Combescure and D. Robert. Semiclassical sum rules and generalized coherent states. Journal of Mathematical Physics, 36(12):6596–6610, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ilinski:1995:BPS

[453] Kirill N. Ilinski, Gleb V. Kalinin, and Vladislav V. Melezhik. Berry phase and supersymmetric topological index. Journal of Mathematical Physics, 36(12):6611–6624, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khrennikov:1995:SIA

[454] Andrew Khrennikov. Statistical interpretation of p-adic quantum theories with p-adic valued wave functions. Journal of Mathematical Physics, 36(12):6625–6632, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levay:1995:MSG

[455] Péter Lévay and Barnabás Apagyi. Modified symmetry generators related to solvable scattering problems. Journal of Mathematical Physics, 36(12):6633–6646, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Motovilov:1995:RRL

[456] A. K. Motovilov. Removal of the resolvent-like energy dependence from interactions and invariant subspaces of a total Hamiltonian. Journal of Mathematical Physics, 36(12):6647–6664, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Navarro:1995:CGP

[457] Miguel Navarro. Comments on Good’s proposal for new rules of quantization. Journal of Mathematical Physics, 36(12):6665–6672, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [13].

Sawin:1995:DSD

[458] Stephen Sawin. Direct sum decompositions and indecomposable TQFTs. Journal of Mathematical Physics, 36(12):6673–6680, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shiekh:1995:QCT

[459] A. Y. Shiekh. Quantum canonical transformations revisited. Journal of Mathematical Physics, 36(12):6681–6687, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beals:1995:IBP

[460] Richard Beals, G. M. Henkin, and N. N. Novikova. The inverse boundary problem for the Rayleigh system. Journal of Mathematical Physics, 36 (12):6688–6708, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fecko:1995:FCC

[461] Marián Fecko. “falling cat” connections and the momentum map. Journal of Mathematical Physics, 36(12):6709–6719, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pokrovski:1995:LPT

[462] A. A. Pokrovski and A. V. Badanin. Lax–Phillips theory for scattering by thin elastic bodies. Journal of Mathematical Physics, 36(12):6720–6736, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sawicz:1995:HIR

[463] Romuald Sawicz. Homogeneity in integral representations for effective properties of composite materials. Journal of Mathematical Physics, 36 (12):6737–6745, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://scitation.

aip.org/content/aip/journal/jmp/36/12/10.1063/1.531185.

Broidioi:1995:LAA

[464] M. Broidioi, B. Momont, and A. Verbeure. Lie algebra of anomalously scaled fluctuations. Journal of Mathematical Physics, 36(12):6746–6757, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mahan:1995:DTS

[465] G. D. Mahan. Diffusion through a slab. Journal of Mathematical Physics, 36(12):6758–6773, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1105].

Pavon:1995:HPS

[466] Michele Pavon. Hamilton’s principle in stochastic mechanics. Journal of Mathematical Physics, 36(12):6774–6800, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carbonaro:1995:HER

[467] P. Carbonaro and S. Giambò. Hodograph equations in relativistic gas dynamics admitting an infinite number of symmetries. Journal of Mathematical Physics, 36(12):6801–6808, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurel:1995:BVP

[468] Burak Gürel, Metin Gürses, and Ismagil Habibullin. Boundary value problems for integrable equations compatible with the symmetry algebra. Journal of Mathematical Physics, 36(12):6809–6821, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kraenkel:1995:BSW

[469] R. A. Kraenkel, M. A. Manna, J. C. Montero, and J. G. Pereira. Boussinesq solitary-wave as a multiple-time solution of the Korteweg–de Vries hierarchy. Journal of Mathematical Physics, 36(12):6822–6828, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pigli:1995:AIS

[470] Mircea Pigli. Adelic integrable systems. Journal of Mathematical Physics, 36(12):6829–6845, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:1995:LFF

[471] Manuel F. Rañada. Lax formalism for a family of integrable Toda-related n-particle systems. Journal of Mathematical Physics, 36(12):6846–6856, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:1995:ABT

[472] Huan Xiong Yang, Kang Li, and Yi Xin Chen. Auto-Bäcklund transformation as a gauge transformation preserving the form of Lax connection. Journal of Mathematical Physics, 36(12):6857–6861, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhu:1995:RCA

[473] Z. Zhu and D. G. Caldi. Reality of complex affine Toda solitons. Journal of Mathematical Physics, 36(12):6862–6871, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cataldo:1995:SSP

[474] Mauricio Cataldo and Patricio Salgado. Superposition of a static perfect fluid and a radial electric field. Journal of Mathematical Physics, 36(12): 6872–6878, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deAzcarraga:1995:RNS

[475] J. A. de Azcárraga and J. C. Pérez Bueno. Relativistic and Newtonian κ-space–times. Journal of Mathematical Physics, 36(12):6879–6896, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Husain:1995:ASS

[476] Viqar Husain. The affine symmetry and self-dual gravity. Journal of Mathematical Physics, 36(12):6897–6906, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lim:1995:LDK

[477] Paul H. Lim, J. M. Overduin, and Paul S. Wesson. Light deflection in Kaluza–Klein gravity. Journal of Mathematical Physics, 36(12):6907– 6914, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Perlick:1995:IDM

[478] Volker Perlick. Infinite-dimensional Morse theory and Fermat’s principle in general relativity. I. Journal of Mathematical Physics, 36(12):6915– 6928, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silva-Ortigoza:1995:KSS

[479] G. Silva-Ortigoza. Killing spinors and separability of Rarita–Schwinger’s equation in type {2, 2} backgrounds. Journal of Mathematical Physics, 36(12):6929–6936, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wallner:1995:NFE

[480] R. P. Wallner. A new form of Einstein’s equations. Journal of Mathematical Physics, 36(12):6937–6969, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zannias:1995:BHC

[481] T. Zannias. Black holes cannot support conformal scalar hair. Journal of Mathematical Physics, 36(12):6970–6980, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arutyunov:1995:PLS

[482] G. È. Arutyunov and P. B. Medvedev. On Poisson–Lie structure on the external algebra of the classical Lie groups. Journal of Mathematical Physics, 36(12):6981–6990, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGuise:1995:ACG

[483] Hubert de Guise and David J. Rowe. Asymptotic Clebsch–Gordan coefficients. Journal of Mathematical Physics, 36(12):6991–7008, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doktorov:1995:MNS

[484] E. V. Doktorov and V. S. Shchesnovich. Modified nonlinear Schrödinger equation: spectral transform and N-soliton solution. Journal of Mathematical Physics, 36(12):7009–7023, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Floreanini:1995:LSF

[485] Roberto Floreanini and Luc Vinet. Lie symmetries of finite-difference equations. Journal of Mathematical Physics, 36(12):7024–7042, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garajeu:1995:GST

[486] Daniela Garajeu, Richard Grimm, and Serge Lazzarini. W-gauge structures and their anomalies: an algebraic approach. Journal of Mathematical Physics, 36(12):7043–7072, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jing:1995:BFC

[487] Nai Huan Jing. Boson-fermion correspondence for Hall–Littlewood polynomials. Journal of Mathematical Physics, 36(12):7073–7080, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:1995:NQG

[488] Shahn Majid and M. J. Rodríguez-Plaza. Nonstandard quantum groups and superization. Journal of Mathematical Physics, 36(12):7081–7097, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pallotta:1995:CID

[489] M. Pallotta. On the convergence of the inverse diffraction transform kernel using Cesàro summability. Journal of Mathematical Physics, 36 (12):7098–7102, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shariati:1995:SMC

[490] A. Shariati and A. Aghamohammadi. A simple method for constructing the inhomogeneous quantum group IGL_{q}(n) and its universal enveloping algebra U_{q}(igl(n)). Journal of Mathematical Physics, 36(12):7103–7108, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhdanov:1995:GSD

[491] R. Z. Zhdanov, I. V. Revenko, and W. I. Fushchych. On the general solution of the d’Alembert equation with a nonlinear eikonal constraint and its applications. Journal of Mathematical Physics, 36(12):7109–7127, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moffat:1995:ENG

[492] J. W. Moffat. Erratum: “Nonsymmetric gravitational theory”. Journal of Mathematical Physics, 36(12):7128, December 1995. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [273].

Anonymous:1995:SPA

[493] Anonymous. Summary of the Physics and Astronomy Classification Scheme — 1995. Journal of Mathematical Physics, 36(12):7129, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1995:SPH

[494] Anonymous. Summary of PACS headings in current index. Journal of Mathematical Physics, 36(12):7130–7132, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1995:SIV

[495] Anonymous. Subject index to volume 36. Journal of Mathematical Physics, 36(12):7133–7161, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1995:AIV

[496] Anonymous. Author index to volume 36. Journal of Mathematical Physics, 36(12):7162–7183, December 1995. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alimohammadi:1996:QCS

[497] Masoud Alimohammadi. Quantum chains with GL_{q}(2) symmetry. Journal of Mathematical Physics, 37(1):1–5, January 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:1996:QES

[498] Carl M. Bender and Gerald V. Dunne. Quasi-exactly solvable systems and orthogonal polynomials. Journal of Mathematical Physics, 37(1):6– 11, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bovier:1996:RTC

[499] Pierre-Alain Bovier. A rigorous treatment of conformal blocks in a model of bosonic conformal field theory. Journal of Mathematical Physics, 37 (1):12–38, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dane:1996:ISG

[500] C. Dane and Y. A. Verdiyev. Integrable systems of group SO(1, 2) and Green’s functions. Journal of Mathematical Physics, 37(1):39–60, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:1996:TDS

[501] A. Das and C. Wotzasek. Two-dimensional supersymmetric harmonic oscillator carrying a representation of the GL(2|1) graded Lie algebra. Journal of Mathematical Physics, 37(1):61–71, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fauser:1996:VNO

[502] Bertfried Fauser. Vertex normal ordering as a consequence of nonsymmetric bilinear forms in Clifford algebras. Journal of Mathematical Physics, 37(1):72–83, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foussats:1996:FMC

[503] A. Foussats, E. Manavella, C. Repetto, O. P. Zandron, and O. S. Zandron. Fermionic matter coupled to higher derivative Chern–Simons theories. II. Journal of Mathematical Physics, 37(1):84–99, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herdegen:1996:AAC

[504] Andrzej Herdegen. Asymptotic algebra for charged particles and radiation. Journal of Mathematical Physics, 37(1):100–120, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hirokawa:1996:GPB

[505] Masao Hirokawa. General properties between the canonical correlation and the independent-oscillator model on a partial ^{∗}-algebra. Journal of Mathematical Physics, 37(1):121–146, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Inozemtsev:1996:STM

[506] V. I. Inozemtsev. Solution to three-magnon problem for S = 1/2 periodic quantum spin chains with elliptic exchange. Journal of Mathematical Physics, 37(1):147–159, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lledo:1996:QDH

[507] M. A. Lledó and M. García-Sucre. On the quantum deformations of Hamiltonian systems. Journal of Mathematical Physics, 37(1):160–173, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marques:1996:LRA

[508] G. C. Marques and D. Spehler. Left-right asymmetry and minimal coupling. Journal of Mathematical Physics, 37(1):174–195, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

MenaMarugan:1996:IAP

[509] Guillermo A. Mena Marugán. Involutions on the algebra of physical observables from reality conditions. Journal of Mathematical Physics, 37 (1):196–205, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Navarro:1996:GQC

[510] Miguel Navarro, Víctor Aldaya, and Manuel Calixto. Group quantization on configuration space. Journal of Mathematical Physics, 37(1):206–218, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seger:1996:FPP

[511] J. E. Seger and A. B. Balantekin. Fermion pair production from an electric field varying in two dimensions. Journal of Mathematical Physics, 37 (1):219–226, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:1996:HLG

[512] Huan-Qiang Zhou and Xi-Wen Guan. Hidden local gauge invariance in the one-dimensional Heisenberg XXZ model with the general boundary terms. Journal of Mathematical Physics, 37(1):227–232, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borghardt:1996:NSW

[513] Alexander A. Borghardt and Darya Ya. Karpenko. New solution of the wave equation with a uniformly moving point source. Journal of Mathematical Physics, 37(1):233–239, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grigore:1996:TSM

[514] D. R. Grigore. Transitive symplectic manifolds in 1 + 2 dimensions. Journal of Mathematical Physics, 37(1):240–253, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shivamoggi:1996:DSP

[515] Bhimsen K. Shivamoggi. Dielectric screening in a plasma: some rigorous results. Journal of Mathematical Physics, 37(1):254–262, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Honegger:1996:GCD

[516] Reinhard Honegger. Global cocycle dynamics for infinite mean field quantum systems interacting with the boson gas. Journal of Mathematical Physics, 37(1):263–282, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Truman:1996:SDE

[517] A. Truman and H. Z. Zhao. On stochastic diffusion equations and stochastic Burgers’ equations. Journal of Mathematical Physics, 37(1): 283–307, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaup:1996:SEM

[518] D. J. Kaup and T. I. Lakoba. The squared eigenfunctions of the massive Thirring model in laboratory coordinates. Journal of Mathematical Physics, 37(1):308–323, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:1996:EIS

[519] Q. M. Liu and A. S. Fokas. Exact interaction of solitary waves for certain nonintegrable equations. Journal of Mathematical Physics, 37(1):324– 345, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szydlowski:1996:GSJ

[520] Marek Szydłowski, Michael Heller, and Wiesław Sasin. Geometry of spaces with the Jacobi metric. Journal of Mathematical Physics, 37(1): 346–360, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abolghasem:1996:IMT

[521] G. Abolghasem, A. A. Coley, and D. J. Mc Manus. Induced matter theory and embeddings in Riemann flat space–times. Journal of Mathematical Physics, 37(1):361–373, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramidi:1996:NAA

[522] I. G. Avramidi. A new algebraic approach for calculating the heat kernel in quantum gravity. Journal of Mathematical Physics, 37(1):374–394, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clayton:1996:MNS

[523] M. A. Clayton. Massive NGT and spherically symmetric systems. Journal of Mathematical Physics, 37(1):395–420, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Glass:1996:AMK

[524] E. N. Glass. Angular momentum and Killing potentials. Journal of Mathematical Physics, 37(1):421–429, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maharaj:1996:EST

[525] S. D. Maharaj and P. G. L. Leach. Exact solutions for the Tikekar superdense star. Journal of Mathematical Physics, 37(1):430–437, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rendall:1996:ISS

[526] Alan D. Rendall. The initial singularity in solutions of the Einstein– Vlasov system of Bianchi type I. Journal of Mathematical Physics, 37 (1):438–451, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunne:1996:MP

[527] Robert S. Dunne. Majorana parafermions. Journal of Mathematical Physics, 37(1):452–459, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grigore:1996:PUI

[528] D. R. Grigore. The projective unitary irreducible representations of the Galilei group in 1 + 2 dimensions. Journal of Mathematical Physics, 37 (1):460–473, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

LeRoy:1996:GGS

[529] Bertrand Le Roy. A Z_{3}-graded generalization of supermatrices. Journal of Mathematical Physics, 37(1):474–483, January 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Links:1996:SFI

[530] J. R. Links and M. D. Gould. A q-superdimension formula for irreps of type I quantum superalgebras. Journal of Mathematical Physics, 37 (1):484–492, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McInnes:1996:CYC

[531] Brett McInnes. Calabi–Yau compactifications and the global structure of the standard group. Journal of Mathematical Physics, 37(1):493–498, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sachse:1996:XXX

[532] Sebastian Sachse and Ralf Weixler. ISO_{q}(3) and ISO_{q}(2, 1). Journal of Mathematical Physics, 37(1):499–507, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vacaru:1996:SSN

[533] Sergiu I. Vacaru. Spinor structures and nonlinear connections in vector bundles, generalized Lagrange and Finsler spaces. Journal of Mathematical Physics, 37(1):508–523, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandenHijligenberg:1996:DHA

[534] N. van den Hijligenberg and R. Martini. Differential Hopf algebra structures on the universal enveloping algebra of a Lie algebra. Journal of Mathematical Physics, 37(1):524–532, January 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ercolessi:1996:PIS

[535] E. Ercolessi, G. Morandi, F. Napoli, and P. Pieri. Path integrals for spinning particles, stationary phase and the Duistermaat–Heckmann [Heckman] theorem. Journal of Mathematical Physics, 37(2):535–553, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fuertes:1996:SCS

[536] Wifredo Garcia Fuertes and Juan Mateos Guilarte. Semilocal Chern– Simons defects. Journal of Mathematical Physics, 37(2):554–566, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujii:1996:CSG

[537] Kazuyuki Fujii, Taro Kashiwa, and Seiji Sakoda. Coherent states over Grassmann manifolds and the WKB exactness in path integral. Journal of Mathematical Physics, 37(2):567–602, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giller:1996:BPM

[538] S. Giller, C. Gonera, P. Kosiński, and P. Maślanka. Berry’s phase on SU(n)/S(U(1)×U(n−1)) manifolds. Journal of Mathematical Physics, 37 (2):603–607, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwai:1996:QMK

[539] Toshihiro Iwai, Yoshio Uwano, and Noriaki Katayama. Quantization of the multifold Kepler system. Journal of Mathematical Physics, 37(2): 608–624, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levay:1996:CCQ

[540] Péter Lévay, David McMullan, and Izumi Tsutsui. The canonical connection in quantum mechanics. Journal of Mathematical Physics, 37(2): 625–636, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rajeev:1996:PBW

[541] S. G. Rajeev and O. T. Turgut. Poisson brackets of Wilson loops and derivations of free algebras. Journal of Mathematical Physics, 37(2):

637–649, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burzlaff:1996:ZMR

[542] J. Burzlaff and D. H. Tchrakian. Zero modes of rotationally symmetric generalized vortices and vortex scattering. Journal of Mathematical Physics, 37(2):650–657, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heyman:1996:TDP

[543] Ehud Heyman. Time-dependent plane-wave spectrum representations for radiation from volume source distributions. Journal of Mathematical Physics, 37(2):658–681, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heyman:1996:TDM

[544] Ehud Heyman and A. J. Devaney. Time-dependent multipoles and their application for radiation from volume source distributions. Journal of Mathematical Physics, 37(2):682–692, February 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hwang:1996:TYM

[545] Dae Sung Hwang and Chang-Yeong Lee. Topological Yang–Mills theory with scalar and vector fields in two and three dimensions. Journal of Mathematical Physics, 37(2):693–701, February 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lhallabi:1996:SA

[546] T. Lhallabi. (2, 0) superconformal anomaly. Journal of Mathematical Physics, 37(2):702–718, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Song:1996:IVS

[547] Shougen Song, Chaoshun Qu, and Jishan He. Imaging of velocity singularities with multiscale operators. Journal of Mathematical Physics, 37 (2):719–731, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garbaczewski:1996:FKK

[548] Piotr Garbaczewski and Robert Olkiewicz. Feynman–Kac kernels in Markovian representations of the Schrödinger interpolating dynamics.

Journal of Mathematical Physics, 37(2):732–751, February 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:1996:UDL

[549] Jeong-Hoon Kim. A uniform diffusion limit for random wave propagation with turning point. Journal of Mathematical Physics, 37(2):752–768, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levy:1996:SMR

[550] Bernard C. Levy and Arthur J. Krener. Stochastic mechanics of reciprocal diffusions. Journal of Mathematical Physics, 37(2):769–802, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vlad:1996:NFD

[551] Marcel Ovidiu Vlad, John Ross, and Michael C. Mackey. Nonequilibrium fluctuation-dissipation relations for independent random rate processes with dynamical disorder. Journal of Mathematical Physics, 37(2):803– 835, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandes:1996:GIP

[552] Joe F. Q. Fernandes and Anthony W. C. Lun. Gauge invariant perturbations of black holes. I. Schwarzschild space–time. Journal of Mathematical Physics, 37(2):836–857, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [928].

Reboucas:1996:CME

[553] M. J. Rebouças, J. E. F. Skea, and R. K. Tavakol. Cosmological models expressible as gradient vector fields. Journal of Mathematical Physics, 37 (2):858–873, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zecca:1996:DER

[554] Antonio Zecca. The Dirac equation in the Robertson–Walker space–time. Journal of Mathematical Physics, 37(2):874–879, February 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ano:1996:GAT

[555] Noriaki Ano. Geometrical aspect of topologically twisted twodimensional conformal superalgebra. Journal of Mathematical Physics, 37(2):880–894, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bordag:1996:HKC

[556] M. Bordag, E. Elizalde, and K. Kirsten. Heat kernel coefficients of the Laplace operator on the D-dimensional ball. Journal of Mathematical Physics, 37(2):895–916, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:1996:PMO

[557] Won-Sang Chung and Anatoli U. Klimyk. On position and momentum operators in the q-oscillator algebra. Journal of Mathematical Physics, 37 (2):917–932, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciulli:1996:ACE

[558] M. Ciulli and T. D. Spearman. Analytic continuation from empirical data: a direct approach to the stabilization problem. Journal of Mathematical Physics, 37(2):933–941, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Donkov:1996:SRC

[559] Sava Donkov and Orlin Stoytchev. On some representations of current algebras in two dimensions and their central extensions. Journal of Mathematical Physics, 37(2):942–952, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkelstein:1996:DLG

[560] Robert J. Finkelstein. q-deformation of the Lorentz group. Journal of Mathematical Physics, 37(2):953–964, February 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foda:1996:PPI

[561] Omar Foda, Masato Okado, and S. Ole Warnaar. A proof of polynomial identities of type ŝl(n)_{1} ⊗ ŝl(n)_{1}/ŝl(n)_{2}. Journal of Mathematical Physics, 37(2):965–986, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gould:1996:TQS

[562] Mark D. Gould, Jon R. Links, and Yao-Zhong Zhang. Type-I quantum superalgebras, q-supertrace, and two-variable link polynomials. Journal of Mathematical Physics, 37(2):987–1003, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kocik:1996:SMM

[563] Jerzy Kocik and Jan Rzewuski. Structure of matrix manifolds and a particle model. Journal of Mathematical Physics, 37(2):1004–1028, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pleasants:1996:PCF

[564] Peter A. B. Pleasants, Michael Baake, and Johannes Roth. Planar coincidences for N-fold symmetry. Journal of Mathematical Physics, 37(2): 1029–1058, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sneddon:1996:IAI

[565] G. E. Sneddon. The identities of the algebraic invariants of the fourdimensional Riemann tensor. Journal of Mathematical Physics, 37(2): 1059–1075, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teleman:1996:CLG

[566] Kostake Teleman. Conformal Lorentz geometry revisited. Journal of Mathematical Physics, 37(2):1076–1085, February 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertolotti:1996:CRC

[567] R. Bertolotti, G. Contreras, L. A. Núñez, U. Percoco, and J. Carot. Comment on: “Ricci collineations of static spherically symmetric spacetimes”, [J. Math. Phys. 35 (1994), no. 6, 3005–3012; MR 95d:83017] by M. J. Amir, A. H. Bokhari and A. Qadir. Journal of Mathematical Physics, 37(2):1086–1088, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [11].

Amir:1996:ERC

[568] M. Jamil Amir, Ashfaque H. Bokhari, and Asghar Qadir. Erratum: “Ricci collineations of static spherically symmetric spacetimes”, [J. Math. Phys. 35 (1994), no. 6, 3005–3012; MR 95d:83017]. Journal of Mathematical Physics, 37(2):1089, February 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [11].

Adler:1996:NSI

[569] Ronald J. Adler and Ovid C. Jacob. Null surfaces, initial values, and evolution operators for spinor fields. Journal of Mathematical Physics, 37 (3):1091–1099, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brooks:1996:MET

[570] Roger Brooks and Arthur Lue. The monopole equations in topological Yang–Mills. Journal of Mathematical Physics, 37(3):1100–1105, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chadan:1996:GCC

[571] K. Chadan, R. Kobayashi, A. Martin, and J. Stubbe. Generalization of the Calogero–Cohn bound on the number of bound states. Journal of Mathematical Physics, 37(3):1106–1114, March 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deAzcarraga:1996:GTF

[572] J. A. de Azcárraga and A. J. Macfarlane. Group theoretical foundations of fractional supersymmetry. Journal of Mathematical Physics, 37(3): 1115–1127, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Elizalde:1996:ZRN

[573] Emilio Elizalde, Sergei D. Odintsov, and August Romeo. Zetaregularization of the O(N) nonlinear sigma model in D dimensions. Journal of Mathematical Physics, 37(3):1128–1147, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Epifanio:1996:CSN

[574] Giuseppina Epifanio, Todor S. Todorov, and Camillo Trapani. Complete sets of non-self-adjoint observables: an unbounded approach. Journal of Mathematical Physics, 37(3):1148–1160, March 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fine:1996:TSM

[575] Dana Stanley Fine. Topological sectors and measures on moduli space in quantum Yang–Mills on a Riemann surface. Journal of Mathematical Physics, 37(3):1161–1170, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Firsova:1996:TDW

[576] N. E. Firsova. On the time decay of a wave packet in a one-dimensional finite band periodic lattice. Journal of Mathematical Physics, 37(3):

1171–1181, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Handy:1996:VFM

[577] Carlos R. Handy, John Maweu, and Leticia Soto Atterberry. Variational formulation of a moment problem quantization method. Journal of Mathematical Physics, 37(3):1182–1196, March 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kar:1996:GPI

[578] T. K. Kar and Gautam Ghosh. Geometric phase for isotopic spin coherent states. Journal of Mathematical Physics, 37(3):1197–1203, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konkov:1996:QCG

[579] L. E. Kon′kov and S. V. Prants. Quantum chaos in the group-theoretical picture. Journal of Mathematical Physics, 37(3):1204–1217, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:1996:GPB

[580] Ali Mostafazadeh. Geometric phase, bundle classification, and group representation. Journal of Mathematical Physics, 37(3):1218–1233, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2063] and erratum [2064].

Ogurisu:1996:ESD

[581] Osamu Ogurisu. Essential spectrum of the Dirac Hamiltonian for a spin 1/2 neutral particle with an anomalous magnetic moment in an asymptotically constant magnetic field. Journal of Mathematical Physics, 37 (3):1234–1243, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmidt:1996:CML

[582] Hans-Jürgen Schmidt. Classical mechanics with lapse. Journal of Mathematical Physics, 37(3):1244–1252, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1996:SNC

[583] F. Calogero. Solvable (nonrelativistic, classical) n-body problems on the line. II. Journal of Mathematical Physics, 37(3):1253–1258, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Percus:1996:STN

[584] J. K. Percus. The stress tensor for nonlocal field equations. Journal of Mathematical Physics, 37(3):1259–1267, March 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gielerak:1996:GPF

[585] Roman Gielerak and Robert Olkiewicz. Gentle perturbations of the free Bose gas. The critical regime. Journal of Mathematical Physics, 37(3): 1268–1286, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kholodenko:1996:SMDa

[586] Arkady L. Kholodenko. Statistical mechanics of the deformable droplets on flat surfaces. Journal of Mathematical Physics, 37(3):1287–1313, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kholodenko:1996:SMDb

[587] Arkady L. Kholodenko. Statistical mechanics of the deformable droplets on Riemannian surfaces: applications to reptation and related problems. Journal of Mathematical Physics, 37(3):1314–1335, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Latrach:1996:NSP

[588] Khalid Latrach. On a nonlinear stationary problem arising in transport theory. Journal of Mathematical Physics, 37(3):1336–1348, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anderson:1996:ESB

[589] Arlen Anderson. An elegant solution of the n-body Toda problem. Journal of Mathematical Physics, 37(3):1349–1355, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delduc:1996:SKH

[590] F. Delduc, E. Ivanov, and S. Krivonos. N = 4 super KdV hierarchy in N = 4 and N = 2 superspaces. Journal of Mathematical Physics, 37(3): 1356–1381, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1009].

Lu:1996:FON

[591] Bao Qun Lu. A formula for obtaining new hereditary symmetries and new integrable equations. Journal of Mathematical Physics, 37(3):1382–1392, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vaklev:1996:SSG

[592] Y. Vaklev. Soliton solutions and gauge equivalence for the problem of Zakharov–Shabat and its generalizations. Journal of Mathematical Physics, 37(3):1393–1413, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brodbeck:1996:IPE

[593] Othmar Brodbeck and Norbert Straumann. Instability proof for Einstein–Yang–Mills solitons and black holes with arbitrary gauge groups. Journal of Mathematical Physics, 37(3):1414–1433, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burnett:1996:CRB

[594] Gregory A. Burnett. Collapsing regions and black hole formation. Journal of Mathematical Physics, 37(3):1434–1448, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lemos:1996:RDQ

[595] Nivaldo A. Lemos. Radiation-dominated quantum Friedmann models. Journal of Mathematical Physics, 37(3):1449–1460, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smoller:1996:UZS

[596] J. A. Smoller and A. G. Wasserman. Uniqueness of zero surface gravity SU(2) Einstein–Yang/Mills black holes. Journal of Mathematical Physics, 37(3):1461–1484, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stedile:1996:CGT

[597] E. Stedile and R. Duarte. Conformal gauge theory of (2+1)-dimensional gravity. Journal of Mathematical Physics, 37(3):1485–1500, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundland:1996:SMD

[598] A. M. Grundland, P. Winternitz, and W. J. Zakrzewski. On the solutions of the CP^{1} model in (2 + 1) dimensions. Journal of Mathematical Physics, 37(3):1501–1520, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jorgensen:1996:DEN

[599] Palle E. T. Jorgensen. A duality for endomorphisms of von Neumann algebras. Journal of Mathematical Physics, 37(3):1521–1538, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lafortune:1996:SFP

[600] Stéphane Lafortune and Pavel Winternitz. Superposition formulas for pseudounitary matrix Riccati equations. Journal of Mathematical Physics, 37(3):1539–1550, March 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bokanowski:1996:DDF

[601] O. Bokanowski and B. Grébert. Deformations of density functions in molecular quantum chemistry. Journal of Mathematical Physics, 37(4): 1553–1573, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorlas:1996:LSL

[602] T. C. Dorlas, N. Macris, and J. V. Pulé. Localization in single Landau bands. Journal of Mathematical Physics, 37(4):1574–1595, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Etesi:1996:SSB

[603] Gábor Etesi. Spontaneous symmetry breaking in the SO(3) gauge theory to discrete subgroups. Journal of Mathematical Physics, 37(4):1596– 1601, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feigenbaum:1996:DCP

[604] James Feigenbaum and Peter G. O. Freund. A q-deformation of the Coulomb problem. Journal of Mathematical Physics, 37(4):1602–1616, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fulop:1996:RWQ

[605] T. Fülöp. Reduced SL(2, R) WZNW quantum mechanics. Journal of Mathematical Physics, 37(4):1617–1631, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hostler:1996:CEI

[606] Levere Hostler. Commutator expansion. II. Relativistic reduced Green’s functions and the Lamb shift calculation. Journal of Mathematical Physics, 37(4):1632–1641, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivashchuk:1996:DCM

[607] V. D. Ivashchuk and V. N. Melnikov. Dually charged mesoatom on the space of constant negative curvature. Journal of Mathematical Physics, 37(4):1642–1649, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katriel:1996:RMD

[608] J. Katriel and C. Quesne. Recursively minimally-deformed oscillators. Journal of Mathematical Physics, 37(4):1650–1661, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lipinski:1996:QLE

[609] Z. Lipiński. Quantization of a loop extended SU(2) affine Kac–Moody algebra. Journal of Mathematical Physics, 37(4):1662–1667, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Varma:1996:QST

[610] S. Varma and E. C. G. Sudarshan. Quantum scattering theory in light of an exactly solvable model with rearrangement collisions. Journal of Mathematical Physics, 37(4):1668–1712, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vathsan:1996:RQG

[611] Radhika Vathsan. Remarks on the quantization of gauge theories. Journal of Mathematical Physics, 37(4):1713–1723, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [929].

Bette:1996:DIM

[612] Andreas Bette. Directly interacting massless particles—a twistor approach. Journal of Mathematical Physics, 37(4):1724–1734, April 1996.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1996:SBP

[613] F. Calogero. A solvable N-body problem in the plane. I. Journal of Mathematical Physics, 37(4):1735–1759, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gogilidze:1996:AFC

[614] S. A. Gogilidze, A. M. Khvedelidze, and V. N. Pervushin. On abelianization of first class constraints. Journal of Mathematical Physics, 37 (4):1760–1771, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nucci:1996:CKG

[615] M. C. Nucci. The complete Kepler group can be derived by Lie group analysis. Journal of Mathematical Physics, 37(4):1772–1775, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerberoff:1996:STS

[616] G. R. Guerberoff and G. A. Raggio. Standard thermal statistics with q-entropies. Journal of Mathematical Physics, 37(4):1776–1789, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerberoff:1996:NST

[617] Gustavo R. Guerberoff, Pedro A. Pury, and Guido A. Raggio. Nonstandard thermal statistics with q-entropies. Journal of Mathematical Physics, 37(4):1790–1811, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holevo:1996:CQM

[618] A. S. Holevo. Covariant quantum Markovian evolutions. Journal of Mathematical Physics, 37(4):1812–1832, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aden:1996:RSK

[619] H. Aden and B. Carl. On realizations of solutions of the KdV equation by determinants on operator ideals. Journal of Mathematical Physics, 37 (4):1833–1857, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorren:1996:SUS

[620] H. J. S. Dorren. Singular and unstable solutions of the Korteweg–de Vries hierarchy. Journal of Mathematical Physics, 37(4):1858–1870, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goriely:1996:IPI

[621] Alain Goriely. Integrability, partial integrability, and nonintegrability for systems of ordinary differential equations. Journal of Mathematical Physics, 37(4):1871–1893, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pickering:1996:SMM

[622] Andrew Pickering. The singular manifold method revisited. Journal of Mathematical Physics, 37(4):1894–1927, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schiff:1996:ZCF

[623] Jeremy Schiff. Zero curvature formulations of dual hierarchies. Journal of Mathematical Physics, 37(4):1928–1938, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beig:1996:KVA

[624] Robert Beig and Piotr T. Chruściel. Killing vectors in asymptotically flat space–times. I. Asymptotically translational Killing vectors and the rigid positive energy theorem. Journal of Mathematical Physics, 37(4): 1939–1961, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gupta:1996:SST

[625] Y. K. Gupta and J. R. Sharma. Similarity solutions for the type D fluid plates in 5-D flat space. Journal of Mathematical Physics, 37(4):1962– 1971, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Monte:1996:TCS

[626] E. M. Monte and M. D. Maia. The twisting connection of space–time. Journal of Mathematical Physics, 37(4):1972–1981, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Opher:1996:GMC

[627] Reuven Opher, Nilton O. Santos, and Anzhong Wang. Geodesic motion and confinement in van Stockum space–time. Journal of Mathematical Physics, 37(4):1982–1990, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

SanMiguel:1996:ORM

[628] A. San Miguel and F. Vicente. Overall rotation of a many-body system in Dixon’s theory. Journal of Mathematical Physics, 37(4):1991–2000, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alty:1996:OPS

[629] L. J. Alty and A. Chamblin. Obstructions to Pin structures on Kleinian manifolds. Journal of Mathematical Physics, 37(4):2001–2012, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Labastida:1996:RBK

[630] J. M. F. Labastida and E. Pérez. A relation between the Kauffman and the HOMFLY polynomials for torus knots. Journal of Mathematical Physics, 37(4):2013–2042, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nicorovici:1996:ARC

[631] N. A. Nicorovici, C. G. Poulton, and R. C. McPhedran. Analytical results for a class of sums involving Bessel functions and square arrays. Journal of Mathematical Physics, 37(4):2043–2052, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stern:1996:LPD

[632] A. Stern and I. Yakushin. Lie–Poisson deformation of the Poincaré algebra. Journal of Mathematical Physics, 37(4):2053–2070, April 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wasilkowski:1996:TPI

[633] Grzegorz W. Wasilkowski and Henryk Woźniakowski. On tractability of path integration. Journal of Mathematical Physics, 37(4):2071–2086, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2273].

TorresdelCastillo:1996:ADP

[634] G. F. Torres del Castillo. Addendum: “Debye potentials and Hamiltonians for magnetic field line flow”, [J. Math. Phys. 36 (1995), no. 7, 3413–3418; MR 96c:78004]. Journal of Mathematical Physics, 37(4):2089, April 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [247].

Baumann:1996:WTM

[635] Klaus Baumann. Why the Thirring model cannot fulfill canonical anticommutation relations. Journal of Mathematical Physics, 37(5):2093– 2106, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castagnino:1996:DRN

[636] M. Castagnino, G. Domenech, M. Levinas, and N. Umérez. The doublet representation of non-Hilbert eigenstates of the Hamiltonian. Journal of Mathematical Physics, 37(5):2107–2120, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hinrichsen:1996:MLP

[637] Haye Hinrichsen and Achim Kempf. Maximal localization in the presence of minimal uncertainties in positions and in momenta. Journal of Mathematical Physics, 37(5):2121–2137, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kanter:1996:DDC

[638] Marek Kanter. Discrete decay and continuous measurement. Journal of Mathematical Physics, 37(5):2138–2156, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimek:1996:QRS

[639] Slawomir Klimek and Andrzej Lesniewski. Quantum Riemann surfaces for arbitrary Planck’s constant. Journal of Mathematical Physics, 37 (5):2157–2165, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kostelecky:1996:RCO

[640] V. Alan Kostelecký and Neil Russell. Radial Coulomb and oscillator systems in arbitrary dimensions. Journal of Mathematical Physics, 37 (5):2166–2181, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saenz:1996:LTA

[641] A. W. Sáenz. Long-time approximation to the evolution of resonant and nonresonant anharmonic oscillators in quantum mechanics. Journal of Mathematical Physics, 37(5):2182–2205, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1218].

Tinebra:1996:EPB

[642] Fabrizio Tinebra. An equivalence proof of the background gauge field quantization and the conventional one. Journal of Mathematical Physics, 37(5):2206–2228, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aaberg:1996:TWN

[643] Ingegerd Åberg, Gerhard Kristensson, and David J. N. Wall. Transient waves in nonstationary media. Journal of Mathematical Physics, 37(5): 2229–2252, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:1996:OBB

[644] A. Das. An ongoing big bang model in the special relativistic Maxwell– Dirac equations. Journal of Mathematical Physics, 37(5):2253–2259, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenczewski:1996:QRW

[645] Romuald Lenczewski. Quantum random walk for U_{q}(su(2)) and a new example of quantum noise. Journal of Mathematical Physics, 37(5):2260– 2278, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vlad:1996:UCA

[646] Marcel Ovidiu Vlad, Ralf Metzler, Theo F. Nonnenmacher, and Michael C. Mackey. Universality classes for asymptotic behavior of relaxation processes in systems with dynamical disorder: dynamical generalizations of stretched exponential. Journal of Mathematical Physics, 37 (5):2279–2306, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:1996:BHS

[647] Q. P. Liu. Bi-Hamiltonian structures of the coupled AKNS hierarchy and the coupled Yajima–Oikawa hierarchy. Journal of Mathematical Physics, 37(5):2307–2314, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandeLeur:1996:SCK

[648] Johan van de Leur. The W_{1+∞}(gl_{s})-symmetries of the s-component KP hierarchy. Journal of Mathematical Physics, 37(5):2315–2337, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:1996:NIS

[649] Yongtang Wu and Xianguo Geng. A new integrable symplectic map associated with lattice soliton equations. Journal of Mathematical Physics, 37(5):2338–2345, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saa:1996:NNS

[650] Alberto Saa. New no-scalar-hair theorem for black holes. Journal of Mathematical Physics, 37(5):2346–2351, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:1996:PGR

[651] Stephen L. Adler. Projective group representations in quaternionic Hilbert space. Journal of Mathematical Physics, 37(5):2352–2360, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comments [926, 880] and response [927].

Anco:1996:DCL

[652] Stephen C. Anco and George Bluman. Derivation of conservation laws from nonlocal symmetries of differential equations. Journal of Mathematical Physics, 37(5):2361–2375, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bose:1996:PRD

[653] S. K. Bose. Projective representations of the (1+1)-dimensional Poincaré group. Journal of Mathematical Physics, 37(5):2376–2387, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brzezinski:1996:QHS

[654] Tomasz Brzeziński. Quantum homogeneous spaces as quantum quotient spaces. Journal of Mathematical Physics, 37(5):2388–2399, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1996:PGS

[655] Jin-Quan Chen, A. Klein, and Jia-Lun Ping. Point-group symmetrized boson representation. Algebraic solution for symmetry-adapted bases of O_{h}. Journal of Mathematical Physics, 37(5):2400–2425, May 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gould:1996:GEF

[656] M. D. Gould and J. R. Links. General eigenvalue formula for Casimir invariants of type I quantum superalgebras. Journal of Mathematical Physics, 37(5):2426–2456, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gjaja:1996:CLT

[657] Ivan Gjaja. Computation of Lie transformations from a power series: bounds and optimum truncation. Journal of Mathematical Physics, 37 (5):2457–2471, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lulek:1996:GIF

[658] T. Lulek and R. Olchawa. Geometric interpretation of fractal symmetries of a finite linear chain. Journal of Mathematical Physics, 37(5):2472– 2483, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masson:1996:SQM

[659] Thierry Masson. Submanifolds and quotient manifolds in noncommutative geometry. Journal of Mathematical Physics, 37(5):2484–2497, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1996:RWA

[660] D. J. Rowe and J. Repka. The Racah–Wigner algebra and coherent tensors. Journal of Mathematical Physics, 37(5):2498–2509, May 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sandoval:1996:EIS

[661] Leonidas Sandoval, Jr. Eichler integrals and string theory. Journal of Mathematical Physics, 37(5):2510–2526, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Swank:1996:DQF

[662] L. J. Swank. (α, β)-derivations on quantum function algebras. Journal of Mathematical Physics, 37(5):2527–2552, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vainerman:1996:SCA

[663] L. Vainerman and R. Kerner. On special classes of n-algebras. Journal of Mathematical Physics, 37(5):2553–2565, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marcinek:1996:EWA

[664] Władysław Marcinek and Robert Rałowski. Erratum: “On Wick algebras with braid relations”. Journal of Mathematical Physics, 37(5):2566, May 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [202].

Arthur:1996:SMS

[665] K. Arthur, G. Roche, D. H. Tchrakian, and Yisong Yang. Skyrme models with self-dual limits: d = 2, 3. Journal of Mathematical Physics, 37(6): 2569–2584, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Busch:1996:CPQ

[666] Paul Busch and Pekka J. Lahti. Correlation properties of quantum measurements. Journal of Mathematical Physics, 37(6):2585–2601, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1996:CQS

[667] Lusheng Chen, Gerald Dunne, Kurt Haller, and Edwin Lim-Lombridas. Canonical quantization of spontaneously broken topologically massive gauge theory. Journal of Mathematical Physics, 37(6):2602–2627, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkelstein:1996:CP

[668] Robert J. Finkelstein. The q-Coulomb problem. Journal of Mathematical Physics, 37(6):2628–2636, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gudder:1996:LPQ

[669] Stanley Gudder. Lattice properties of quantum effects. Journal of Mathematical Physics, 37(6):2637–2642, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nikolaev:1996:DQB

[670] A. S. Nikolaev. On the diagonalization of quantum Birkhoff–Gustavson normal form. Journal of Mathematical Physics, 37(6):2643–2661, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petz:1996:GQS

[671] Dénes Petz and Csaba Sudár. Geometries of quantum states. Journal of Mathematical Physics, 37(6):2662–2673, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Santopinto:1996:TBB

[672] Elena Santopinto, Roelof Bijker, and Francesco Iachello. Transformation brackets between U(ν + 1) ⊃ U(ν) ⊃ SO(ν) and U(ν + 1) ⊃ SO(ν + 1) ⊃ SO(ν). Journal of Mathematical Physics, 37(6):2674–2681, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Slater:1996:AQC

[673] Paul B. Slater. Applications of quantum and classical Fisher information to two-level complex and quaternionic and three-level complex systems. Journal of Mathematical Physics, 37(6):2682–2693, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Toller:1996:FQF

[674] M. Toller. Free quantum fields on the Poincaré group. Journal of Mathematical Physics, 37(6):2694–2730, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wiedemann:1996:SKM

[675] U. A. Wiedemann and N. P. Landsman. The Stückelberg-Kibble model as an example of quantized symplectic reduction. Journal of Mathematical Physics, 37(6):2731–2747, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delgado:1996:GFM

[676] Joaquín Delgado, Florin N. Diacu, Ernesto A. Lacomba, Angelo Mingarelli, Vasile Mioc, Ernesto Perez, and Cristina Stoica. The global flow of the Manev problem. Journal of Mathematical Physics, 37(6):2748– 2761, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Piccolo:1996:THE

[677] Santa Piccolo, Maria Cristina Recchioni, and Francesco Zirilli. The time harmonic electromagnetic field in a disturbed half-space: an existence theorem and a computational method. Journal of Mathematical Physics, 37(6):2762–2786, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bobylev:1996:GIS

[678] A. V. Bobylev, G. L. Caraffini, and G. Spiga. On group invariant solutions of the Boltzmann equation. Journal of Mathematical Physics, 37 (6):2787–2795, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Regge:1996:ESI

[679] Tullio Regge and Riccardo Zecchina. Exact solution of the Ising model on group lattices of genus g > 1. Journal of Mathematical Physics, 37 (6):2796–2814, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belleni-Morante:1996:SRM

[680] Aldo Belleni-Morante and Silvia Totaro. The successive reflection method in three-dimensional particle transport. Journal of Mathematical Physics, 37(6):2815–2823, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boiti:1996:NDS

[681] M. Boiti, J. Leon, and F. Pempinelli. Nonlinear discrete systems with nonanalytic dispersion relations. Journal of Mathematical Physics, 37 (6):2824–2850, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bueken:1996:MHS

[682] Peter Bueken. Multi-Hamiltonian structures for a class of degenerate completely integrable systems. Journal of Mathematical Physics, 37(6):

2851–2862, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1996:CIH

[683] F. Calogero and J.-P. Françoise. A completely integrable Hamiltonian system. Journal of Mathematical Physics, 37(6):2863–2871, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konstein:1996:SAO

[684] S. E. Konstein and M. A. Vasiliev. Supertraces on the algebras of observables of the rational Calogero model with harmonic potential. Journal of Mathematical Physics, 37(6):2872–2891, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hellaby:1996:NNS

[685] Charles Hellaby. The nonsimultaneous nature of the Schwarzschild R = 0 singularity. Journal of Mathematical Physics, 37(6):2892–2905, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendez:1996:QAC

[686] Vicenç Méndez and Josep Triginer. Qualitative analysis of causal cosmological models. Journal of Mathematical Physics, 37(6):2906–2919, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petters:1996:BNC

[687] A. O. Petters and H. J. Witt. Bounds on number of cusps due to point mass gravitational lenses. Journal of Mathematical Physics, 37(6):2920– 2933, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cornwell:1996:ACS

[688] J. F. Cornwell. Abstract carrier space formalism for the irreducible tensor operators of compact quantum group algebras. Journal of Mathematical Physics, 37(6):2934–2954, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeLeo:1996:QSR

[689] Stefano De Leo. Quaternions and special relativity. Journal of Mathematical Physics, 37(6):2955–2968, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doyle:1996:SWI

[690] P. W. Doyle and A. M. Grundland. Simple waves and invariant solutions of quasilinear systems. Journal of Mathematical Physics, 37(6):2969– 2978, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frochaux:1996:TRB

[691] E. Frochaux. Two relativistic boson models in the Schrödinger picture in three space–time dimensions. Journal of Mathematical Physics, 37 (6):2979–3000, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gazeau:1996:TDW

[692] Jean Pierre Gazeau and Vyacheslav Spiridonov. Toward discrete wavelets with irrational scaling factor. Journal of Mathematical Physics, 37(6): 3001–3013, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grot:1996:MSS

[693] Norbert Grot and Carlo Rovelli. Moduli-space structure of knots with intersections. Journal of Mathematical Physics, 37(6):3014–3021, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hambli:1996:CSG

[694] N. Hambli, J. Michelson, and R. T. Sharp. Character states and generator matrix elements for Sp(4) ⊃ SU(2) × U(1). Journal of Mathematical Physics, 37(6):3022–3031, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hinterleitner:1996:ESC

[695] F. Hinterleitner. Examples of separating coordinates for the Klein– Gordon equation in (2+1)-dimensional flat space–time. Journal of Mathematical Physics, 37(6):3032–3040, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kosinski:1996:BHS

[696] P. Kosiński, J. Lukierski, P. Maślanka, and J. Sobczyk. Bicrossproduct Hopf superalgebras and D = 4 κ-deformed Poincaré supergroup. Journal of Mathematical Physics, 37(6):3041–3049, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Satir:1996:PC

[697] Ahmet Satir. Preliminary classification of q_{t} = f(q, q_{x}, q_{xx}, q_{xxx}). Journal of Mathematical Physics, 37(6):3050–3061, June 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmidt:1996:RRF

[698] Jeffrey R. Schmidt. A Rodriguez [Rodrigues] formula and integration measure for the quantum deformed Legendre functions. Journal of Mathematical Physics, 37(6):3062–3072, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smith:1996:AST

[699] Jonathan D. H. Smith and Abraham A. Ungar. Abstract space–times and their Lorentz groups. Journal of Mathematical Physics, 37(6):3073– 3098, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guhr:1996:CIZ

[700] Thomas Guhr. Comment on: “The Itzykson–Zuber integral for U(m|n)”, [J. Math. Phys. 36 (1995), no. 6, 3085–3093; 1331309] by J. Alfaro, R. Medina and L. F. Urrutia. With a reply by Alfaro, Medina and Urrutia. Journal of Mathematical Physics, 37(6):3099–3100, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [220, 701].

Alfaro:1996:RCI

[701] Jorge Alfaro, Ricardo Medina, and Luis F. Urrutia. Reply to the “Comment on: The Itzykson–Zuber integral for U(m|n)”. Journal of Mathematical Physics, 37(6):3100–3101, June 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [700, 220].

Elizalde:1996:ATF

[702] E. Elizalde, M. Lygren, and D. V. Vassilevich. Antisymmetric tensor fields on spheres: functional determinants and non-local counterterms. Journal of Mathematical Physics, 37(7):3105–3117, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gavrilov:1996:PTP

[703] S. P. Gavrilov and D. M. Gitman. Proper time and path integral representations for the commutation function. Journal of Mathematical Physics, 37(7):3118–3130, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leitenberger:1996:QLP

[704] Frank Leitenberger. Quantum Lobachevsky planes. Journal of Mathematical Physics, 37(7):3131–3140, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miettinen:1996:LR

[705] Mauri Miettinen. On localization and regularization. Journal of Mathematical Physics, 37(7):3141–3152, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nenciu:1996:ODP

[706] Gheorghe Nenciu and Radu Purice. One-dimensional periodic Dirac Hamiltonians: semiclassical and high-energy asymptotics for gaps. Journal of Mathematical Physics, 37(7):3153–3167, July 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Renaud:1996:CDS

[707] Jacques Renaud. The contraction of the SU(1, 1) discrete series of representations by means of coherent states. Journal of Mathematical Physics, 37(7):3168–3179, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Truman:1996:QMC

[708] A. Truman and H. Z. Zhao. Quantum mechanics of charged particles in random electromagnetic fields. Journal of Mathematical Physics, 37 (7):3180–3197, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhdanov:1996:CSS

[709] R. Z. Zhdanov. Conditional symmetry and spectrum of the onedimensional Schrödinger equation. Journal of Mathematical Physics, 37 (7):3198–3217, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:1996:IEM

[710] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. Integral equation methods for the inverse problem with discontinuous wave speed. Journal of Mathematical Physics, 37(7):3218–3245, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deHoop:1996:GBC

[711] Maarten V. de Hoop. Generalization of the Bremmer coupling series. Journal of Mathematical Physics, 37(7):3246–3282, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Patch:1996:DTM

[712] S. K. Patch. Diffuse tomography modulo Grassmann and Laplace. Journal of Mathematical Physics, 37(7):3283–3305, July 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

BenAbdallah:1996:HMM

[713] N. Ben Abdallah and P. Degond. On a hierarchy of macroscopic models for semiconductors. Journal of Mathematical Physics, 37(7):3306–3333, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Douthett:1996:EES

[714] Jack Douthett and Richard Krantz. Energy extremes and spin configurations for the one-dimensional antiferromagnetic Ising model with arbitrary-range interaction. Journal of Mathematical Physics, 37(7): 3334–3353, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gielerak:1996:PIR

[715] R. Gielerak and A. L. Rebenko. On the Poisson integrals representation in the classical statistical mechanics of continuous systems. Journal of Mathematical Physics, 37(7):3354–3374, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pavon:1996:LDC

[716] Michele Pavon. Lagrangian dynamics for classical, Brownian, and quantum mechanical particles. Journal of Mathematical Physics, 37(7):3375– 3388, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:1996:GNH

[717] Manuel de León and David M. de Diego. On the geometry of nonholonomic Lagrangian systems. Journal of Mathematical Physics, 37 (7):3389–3414, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fokas:1996:AIH

[718] A. S. Fokas, R. H. J. Grimshaw, and D. E. Pelinovsky. On the asymptotic integrability of a higher-order evolution equation describing internal waves in a deep fluid. Journal of Mathematical Physics, 37(7):3415–3421, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioannidou:1996:SSN

[719] T. Ioannidou. Soliton solutions and nontrivial scattering in an integrable chiral model in (2 + 1) dimensions. Journal of Mathematical Physics, 37 (7):3422–3441, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kaup:1996:VMH

[720] D. J. Kaup and T. I. Lakoba. Variational method: how it can generate false instabilities. Journal of Mathematical Physics, 37(7):3442–3462, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kulish:1996:SPE

[721] P. P. Kulish, S. Rauch-Wojciechowski, and A. V. Tsiganov. Stationary problems for equation of the KdV type and dynamical r-matrices. Journal of Mathematical Physics, 37(7):3463–3482, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sedra:1996:HLS

[722] M. B. Sedra. On the huge Lie superalgebra of pseudo-superdifferential operators and super KP-hierarchies. Journal of Mathematical Physics, 37(7):3483–3490, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:1996:PAB

[723] Huan-Xiong Yang and You-Quan Li. Prolongation approach to Bäcklund transformation of Zhiber–Mikhailov–Shabat equation. Journal of Mathematical Physics, 37(7):3491–3497, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bokhari:1996:CCS

[724] Ashfaque H. Bokhari and A. R. Kashif. Curvature collineations of some static spherically symmetric space–times. Journal of Mathematical Physics, 37(7):3498–3504, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hajicek:1996:SGP

[725] P. Hájíček and C. J. Isham. The symplectic geometry of a parametrized scalar field on a curved background. Journal of Mathematical Physics, 37 (7):3505–3521, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hajicek:1996:PGT

[726] P. Hájíček and C. J. Isham. Perennials and the group-theoretical quantization of a parametrized scalar field on a curved background. Journal of Mathematical Physics, 37(7):3522–3538, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zecca:1996:SMF

[727] Antonio Zecca. Separation of the massless field equations for arbitrary spin in the Robertson–Walker space–time. Journal of Mathematical Physics, 37(7):3539–3547, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bougourzi:1996:QAA

[728] A. Hamid Bougourzi and Luc Vinet. A quantum analog of the Z algebra. Journal of Mathematical Physics, 37(7):3548–3567, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:1996:TPD

[729] Salih Çelïk. Two parameter deformation of Grassmann matrix group and supergroup. Journal of Mathematical Physics, 37(7):3568–3575, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daboul:1996:APS

[730] C. Daboul. Algebraic proof of the symmetric space theorem. Journal of Mathematical Physics, 37(7):3576–3586, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeVos:1996:KLC

[731] Koos De Vos and Peter van Driel. The Kazhdan–Lusztig conjecture for W algebras. Journal of Mathematical Physics, 37(7):3587–3610, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martin:1996:NHS

[732] Christiane Martin and Mohamed Zouagui. A noncommutative Hopf structure on C^{∞}[SL(2, C)] as a quantum Lorentz group. Journal of Mathematical Physics, 37(7):3611–3629, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reifler:1996:IGB

[733] Frank Reifler and Randall Morris. Inclusion of gauge bosons in the tensor formulation of the Dirac theory. Journal of Mathematical Physics, 37(7): 3630–3640, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salazar-Anaya:1996:RPI

[734] Gelasio Salazar-Anaya and Jesús Urías. Random perturbations of iterated maps. Journal of Mathematical Physics, 37(7):3641–3643, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cairo:1996:CHS

[735] Laurent Cairó and Marc R. Feix. Comments on: “Hamiltonian structures for the n-dimensional Lotka–Volterra equations”, [J. Math. Phys. 36 (1995), no. 7, 3520–3534; MR 96g:34010] by M. Plank. Journal of Mathematical Physics, 37(7):3644–3645, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [257, 1008].

Calogero:1996:ESO

[736] F. Calogero. Erratum: “Solution of the one-dimensional N-body problems with quadratic and/or inversely quadratic pair potentials”, [J. Math. Phys. 12 (1971), 419–436; MR 43 #5824]. Journal of Mathematical Physics, 37(7):3646, July 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [3].

Balakrishnan:1996:SEA

[737] Radha Balakrishnan and Partha Guha. Surface-embeddability approach to the dynamics of the inhomogeneous Heisenberg spin chain. Journal of Mathematical Physics, 37(8):3651–3661, August 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bayin:1996:CET

[738] Selçuk Ş. Bayin, J. P. Krisch, and Mustafa Özcan. The Casimir energy of the twisted string loop: uniform and two segment loops. Journal of Mathematical Physics, 37(8):3662–3674, August 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cadavid:1996:CPC

[739] A. C. Cadavid and R. J. Finkelstein. The q-Coulomb problem in configuration space. Journal of Mathematical Physics, 37(8):3675–3683, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cattaneo:1996:CWL

[740] Alberto S. Cattaneo. Cabled Wilson loops in BF theories. Journal of Mathematical Physics, 37(8):3684–3703, August 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chau:1996:KEQ

[741] Ling-Lie Chau and Itaru Yamanaka. The KZ equation and the quantumgroup difference equation in quantum self-dual Yang–Mills theory. Journal of Mathematical Physics, 37(8):3704–3712, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khudaverdian:1996:BVF

[742] O. M. Khudaverdian and A. Nersessian. Batalin–Vilkovisky formalism and integration theory on manifolds. Journal of Mathematical Physics, 37(8):3713–3724, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1996:BQG

[743] Chang-Yeong Lee, Dae Sung Hwang, and Yuval Ne’eman. BRST quantization of gauge theory in noncommutative geometry: matrix derivative approach. Journal of Mathematical Physics, 37(8):3725–3738, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsumoto:1996:NEC

[744] Masao Matsumoto. A natural extension of coherent state path integrals. Journal of Mathematical Physics, 37(8):3739–3752, August 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nergiz:1996:LVK

[745] Serdar Nergiz and Cihan Saçlıoğlu. Liouville vortex and ϕ^{4} kink solutions of the Seiberg–Witten equations. Journal of Mathematical Physics, 37 (8):3753–3759, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruan:1996:TBD

[746] W. Y. Ruan. Transformation bracket for 2D harmonic oscillator functions and its application to few-electron quantum dots. Journal of Mathematical Physics, 37(8):3760–3768, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tarantello:1996:MCS

[747] Gabriella Tarantello. Multiple condensate solutions for the Chern– Simons–Higgs theory. Journal of Mathematical Physics, 37(8):3769–3796, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wulkenhaar:1996:DSM

[748] R. Wulkenhaar. Deriving the standard model from the simplest twopoint K cycle. Journal of Mathematical Physics, 37(8):3797–3814, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zak:1996:DWH

[749] J. Zak. Discrete Weyl–Heisenberg transforms. Journal of Mathematical Physics, 37(8):3815–3823, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ahmad:1996:GFT

[750] Falih Ahmad and Mohsen Razzaghi. On the Green-functions technique and phase velocity approximation of axially symmetric fields in stratified media. Journal of Mathematical Physics, 37(8):3824–3832, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dotti:1996:SDM

[751] Gustavo Dotti and Carlos N. Kozameh. Self-dual Maxwell fields on curved space–times. Journal of Mathematical Physics, 37(8):3833–3853, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haines:1996:IIA

[752] A. John Haines and Maarten V. de Hoop. An invariant imbedding analysis of general wave scattering problems. Journal of Mathematical Physics, 37(8):3854–3881, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schray:1996:CSF

[753] Jörg Schray, Tevian Dray, Corinne A. Manogue, Robin W. Tucker, and Charles Wang. The construction of spinor fields on manifolds with smooth degenerate metrics. Journal of Mathematical Physics, 37(8): 3882–3896, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaveau:1996:MEB

[754] Bernard Gaveau and L. S. Schulman. Master equation based formulation of nonequilibrium statistical mechanics. Journal of Mathematical Physics, 37(8):3897–3932, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Langmann:1996:SME

[755] Edwin Langmann and Jouko Mickelsson. Scattering matrix in external field problems. Journal of Mathematical Physics, 37(8):3933–3953, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkel:1996:QES

[756] Federico Finkel, Artemio González-López, and Miguel A. Rodríguez. Quasi-exactly solvable potentials on the line and orthogonal polynomials. Journal of Mathematical Physics, 37(8):3954–3972, August 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garcia:1996:CBL

[757] Ariel O. García and Roberto C. Trinchero. Constructive building of the Lax pair in the non-linear sigma model. Journal of Mathematical Physics, 37(8):3973–3981, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suris:1996:IDB

[758] Yuri B. Suris. Integrable discretizations of the Bogoyavlensky lattices. Journal of Mathematical Physics, 37(8):3982–3996, August 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chicone:1996:IKB

[759] C. Chicone, B. Mashhoon, and D. G. Retzloff. On the ionization of a Keplerian binary system by periodic gravitational radiation. Journal of Mathematical Physics, 37(8):3997–4016, August 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See addendum [969].

Gil-Medrano:1996:RGR

[760] Olga Gil-Medrano. A Riemannian generalization of a result of DeWitt concerning Ricci-flat Lorentz metrics. Journal of Mathematical Physics, 37(8):4017–4024, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kleidis:1996:CSE

[761] K. Kleidis, H. Varvoglis, and D. B. Papadopoulos. Conditions on the stability of the external space solutions in a higher-dimensional quadratic theory of gravity. Journal of Mathematical Physics, 37(8):4025–4033, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Patel:1996:IKT

[762] L. K. Patel and Naresh Dadhich. Inhomogeneous Kasner-type cylindrically symmetric models in Kaluza–Klein space–time. Journal of Mathematical Physics, 37(8):4034–4040, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schilling:1996:NCG

[763] Troy A. Schilling. Non-covariance of the generalized holonomies: examples. Journal of Mathematical Physics, 37(8):4041–4052, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1996:PSE

[764] G. F. Torres del Castillo. Perturbations of solutions of the Einstein– Maxwell equations with a null background electromagnetic field. Journal of Mathematical Physics, 37(8):4053–4061, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bueken:1996:TDR

[765] Peter Bueken. Three-dimensional Riemannian manifolds with constant principal Ricci curvatures ρ_{1} = ρ_{2} ≠ ρ_{3}. Journal of Mathematical Physics, 37(8):4062–4075, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:1996:HMH

[766] Qing Ding and Bao-Qun Lu. On harmonic maps from R^{1,1} into Hilbert loop groups. Journal of Mathematical Physics, 37(8):4076–4088, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dubois-Violette:1996:CNG

[767] M. Dubois-Violette, J. Madore, T. Masson, and J. Mourad. On curvature in noncommutative geometry. Journal of Mathematical Physics, 37(8): 4089–4102, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:1996:SPS

[768] Carl M. Bender and Jade P. Vinson. Summation of power series by continued exponentials. Journal of Mathematical Physics, 37(8):4103– 4119, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figueroa-OFarrill:1996:SSS

[769] José M. Figueroa-O’Farrill and Sonia Stanciu. On the structure of symmetric self-dual Lie algebras. Journal of Mathematical Physics, 37(8): 4121–4134, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Floreanini:1996:DCJ

[770] Roberto Floreanini, Jean LeTourneux, and Luc Vinet. A q-deformed e(4) and continuous q-Jacobi polynomials. Journal of Mathematical Physics, 37(8):4135–4149, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gervois:1996:STI

[771] A. Gervois and M. L. Mehta. Some trigonometric identities. II. Journal of Mathematical Physics, 37(8):4150–4155, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holschneider:1996:CWT

[772] Matthias Holschneider. Continuous wavelet transforms on the sphere. Journal of Mathematical Physics, 37(8):4156–4165, August 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

vandenHijligenberg:1996:QDC

[773] N. van den Hijligenberg, R. Martini, and G. F. Post. Quantization of differential calculi on universal enveloping algebras. Journal of Mathematical Physics, 37(8):4166–4175, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanderJeugt:1996:LSE

[774] J. Van der Jeugt. On the Lie superalgebra embedding C(n+1) ⊃ B(0, n) and dimension formulas. Journal of Mathematical Physics, 37(8):4176– 4186, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Williams:1996:IF

[775] H. T. Williams. SU3 isoscalar factors. Journal of Mathematical Physics, 37(8):4187–4198, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Enginer:1996:EFS

[776] Y. Enginer, M. Hortaçsu, and N. Özdemir. Errata: “Fluctuations of a spherical gravitational impulsive wave”, [J. Math. Phys. 34 (1993), no. 2, 690–699; MR 93m:83024] by M. Hortaçsu and “Fluctuations of a rotating impulsive wave”, [ibid. 35 (1994), no. 6, 3043–3050; MR 95g:83036] by M. Hortaçsu and R. Kaya. Journal of Mathematical Physics, 37(8):4199, August 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [5, 12].

Arai:1996:CCR

[777] Asao Arai. Canonical commutation relations, the Weierstrass zeta function, and infinite-dimensional Hilbert space representations of the quantum group U_{q}(∫↕_{∈}). Journal of Mathematical Physics, 37(9):4203–4218, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:1996:HDS

[778] F. Bagarello and C. Trapani. The Heisenberg dynamics of spin systems: a quasi ^{∗}-algebras approach. Journal of Mathematical Physics, 37(9): 4219–4234, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bollini:1996:GSC

[779] C. G. Bollini, O. Civitarese, A. L. De Paoli, and M. C. Rocca. Gamow states as continuous linear functionals over analytical test functions.

Journal of Mathematical Physics, 37(9):4235–4242, September 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1996:SQV

[780] F. Calogero and J. F. van Diejen. Solvable quantum version of an integrable Hamiltonian system. Journal of Mathematical Physics, 37(9): 4243–4251, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chang:1996:CSP

[781] Zhe Chang. Coherent state path-integral representation of supersymmetric lattice models. Journal of Mathematical Physics, 37(9):4252–4273, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Droz-Vincent:1996:TBW

[782] Philippe Droz-Vincent. Two-body wave equations in curved space–time. Journal of Mathematical Physics, 37(9):4274–4291, September 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Honegger:1996:SBT

[783] Reinhard Honegger and Alfred Rieckers. Squeezing Bogoliubov transformations on the infinite mode CCR-algebra. Journal of Mathematical Physics, 37(9):4292–4309, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

LaChapelle:1996:PIS

[784] J. LaChapelle. Path integral solution of the Schrödinger equation in curvilinear coordinates: a straightforward procedure. Journal of Mathematical Physics, 37(9):4310–4319, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1106].

Liu:1996:SMF

[785] Yu-Fen Liu and Jaime Keller. A symmetry of massless fields. Journal of Mathematical Physics, 37(9):4320–4332, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mahan:1996:QBE

[786] G. D. Mahan. Quantum Boltzmann equation for photons. Journal of Mathematical Physics, 37(9):4333–4351, September 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Malbouisson:1996:VSE

[787] A. P. C. Malbouisson and N. F. Svaiter. On the vacuum stability in the Efimov–Fradkin model at finite temperature. Journal of Mathematical Physics, 37(9):4352–4367, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pashaev:1996:LPD

[788] Oktay K. Pashaev. The Lax pair by dimensional reduction of Chern– Simons gauge theory. Journal of Mathematical Physics, 37(9):4368–4387, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ciulli:1996:ATE

[789] S. Ciulli, S. Ispas, and M. K. Pidcock. Anomalous thresholds and edge singularities in electrical impedance tomography. Journal of Mathematical Physics, 37(9):4388–4417, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Radford:1996:LSD

[790] C. J. Radford. Localized solutions of the Dirac–Maxwell equations. Journal of Mathematical Physics, 37(9):4418–4433, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wurmser:1996:MRT

[791] Daniel Wurmser. A manifestly reciprocal theory of scattering in the presence of elastic media. Journal of Mathematical Physics, 37(9):4434– 4479, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meson:1996:IDS

[792] Alejandro M. Mesón and Fernando Vericat. Invariant of dynamical systems: a generalized entropy. Journal of Mathematical Physics, 37(9): 4480–4483, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morosi:1996:MTF

[793] Carlo Morosi and Livio Pizzocchero. r-matrix theory, formal Casimirs and the periodic Toda lattice. Journal of Mathematical Physics, 37(9): 4484–4513, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rangarajan:1996:SCS

[794] Govindan Rangarajan. Symplectic completion of symplectic jets. Journal of Mathematical Physics, 37(9):4514–4542, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruzzo:1996:SNG

[795] Ugo Bruzzo and Francesca Venzano. Solution of Nester’s gauge conditions. Journal of Mathematical Physics, 37(9):4543–4548, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hajac:1996:EAA

[796] Piotr M. Hajac. The Einstein action for algebras of matrix valued functions—toy models. Journal of Mathematical Physics, 37(9):4549– 4556, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roberts:1996:ISF

[797] M. D. Roberts. Imploding scalar fields. Journal of Mathematical Physics, 37(9):4557–4573, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Singleton:1996:ASS

[798] D. Singleton. Axially symmetric solutions for SU(2) Yang–Mills theory. Journal of Mathematical Physics, 37(9):4574–4583, September 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1996:ESE

[799] G. F. Torres del Castillo. Exact solutions of the Einstein equations with sources from linearized solutions. Journal of Mathematical Physics, 37 (9):4584–4589, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1592].

Cornwell:1996:ITO

[800] J. F. Cornwell. Irreducible tensor operators in the regular coaction formalisms of compact quantum group algebras. Journal of Mathematical Physics, 37(9):4590–4634, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dabrowski:1996:MPL

[801] Ludwik Dabrowski, Piotr M. Hajac, Giovanni Landi, and Pasquale Siniscalco. Metrics and pairs of left and right connections on bimodules. Journal of Mathematical Physics, 37(9):4635–4646, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakis:1996:DCL

[802] Aristophanes Dimakis and J. Madore. Differential calculi and linear connections. Journal of Mathematical Physics, 37(9):4647–4661, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorlas:1996:NCC

[803] T. C. Dorlas. A non-commutative central limit theorem. Journal of Mathematical Physics, 37(9):4662–4682, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fichtmuller:1996:LTF

[804] M. Fichtmüller and W. Weich. The limit transition q → 1 of the q-Fourier transform. Journal of Mathematical Physics, 37(9):4683–4689, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gianotti:1996:IFH

[805] Ricardo D. Gianotti, Antonio E. Rodriguez, and Fernando Vericat. On the inversion formula of the half-Hartley transform. Journal of Mathematical Physics, 37(9):4690–4692, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kajiwara:1996:DSR

[806] Kenji Kajiwara and Yasuhiro Ohta. Determinant structure of the rational solutions for the Painlevé II equation. Journal of Mathematical Physics, 37(9):4693–4704, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenglart:1996:STT

[807] E. Lenglart and G. Gouesbet. The separability “theorem” in terms of distributions with discussion of electromagnetic scattering theory. Journal of Mathematical Physics, 37(9):4705–4710, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odundun:1996:RUC

[808] Wolé Òdúndún. Remarks on the universal central extension of sdiff(T^{2})_{loc}. Journal of Mathematical Physics, 37(9):4711–4715, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oliveira-Neto:1996:ICS

[809] G. Oliveira-Neto. Identifying conical singularities. Journal of Mathematical Physics, 37(9):4716–4723, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Podles:1996:QEA

[810] P. Podleś. Quasitriangularity and enveloping algebras for inhomogeneous quantum groups. Journal of Mathematical Physics, 37(9):4724–4737, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Steinacker:1996:IQE

[811] Harold Steinacker. Integration on quantum Euclidean space and sphere. Journal of Mathematical Physics, 37(9):4738–4749, September 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thoma:1996:OOBa

[812] M. Thoma and R. T. Sharp. Orbit-orbit branching rules for families of classical Lie algebra-subalgebra pairs. Journal of Mathematical Physics, 37(9):4750–4757, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Torrisi:1996:GAA

[813] M. Torrisi, R. Tracinà, and A. Valenti. A group analysis approach for a nonlinear differential system arising in diffusion phenomena. Journal of Mathematical Physics, 37(9):4758–4767, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daniel:1996:EII

[814] M. Daniel, K. Porsezian, and M. Lakshmanan. Erratum: “On the integrability of the inhomogeneous spherically symmetric Heisenberg ferromagnet in arbitrary dimensions”, [J. Math. Phys. 35 (1994), no. 12, 6498–6510; MR 95h:82022]. Journal of Mathematical Physics, 37(9):4768, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [16].

Morato:1996:EMD

[815] Laura M. Morato and Lorenza Viola. Erratum: “Markov diffusions in comoving coordinates and stochastic quantization of the free relativistic spinless particle”, [J. Math. Phys. 36 (1995), no. 9, 4691–4710; MR 96k:81122]. Journal of Mathematical Physics, 37(9):4769, September 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [339].

Capasso:1996:MPS

[816] Federico Capasso, Jerome Faist, and Carlo Sirtori. Mesoscopic phenomena in semiconductor nanostructures by quantum design. Journal of Mathematical Physics, 37(10):4775–4791, October 1996. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buttiker:1996:DCQ

[817] M. Büttiker. Dynamic conductance and quantum noise in mesoscopic conductors. Journal of Mathematical Physics, 37(10):4793–4815, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Presilla:1996:TPR

[818] Carlo Presilla and Johannes Sjöstrand. Transport properties in resonant tunneling heterostructures. Journal of Mathematical Physics, 37(10): 4816–4844, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levitov:1996:ECS

[819] Leonid S. Levitov, Hyunwoo Lee, and Gordey B. Lesovik. Electron counting statistics and coherent states of electric current. Journal of Mathematical Physics, 37(10):4845–4866, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:1996:BSS

[820] P. Exner, P. Šeba, M. Tater, and D. Vaněk. Bound states and scattering in quantum waveguides coupled laterally through a boundary window. Journal of Mathematical Physics, 37(10):4867–4887, October 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Albeverio:1996:MRW

[821] S. Albeverio, F. Haake, P. Kurasov, M. Kuś, and P. Šeba. S-matrix, resonances, and wave functions for transport through billiards with leads. Journal of Mathematical Physics, 37(10):4888–4903, October 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Brouwer:1996:DMI

[822] P. W. Brouwer and C. W. J. Beenakker. Diagrammatic method of integration over the unitary group, with applications to quantum transport in mesoscopic systems. Journal of Mathematical Physics, 37(10): 4904–4934, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Falko:1996:SWF

[823] Vladimir I. Fal′ko and K. B. Efetov. Statistics of wave functions in mesoscopic systems. Journal of Mathematical Physics, 37(10):4935–4967, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andreev:1996:ELC

[824] A. V. Andreev, B. D. Simons, and B. L. Altshuler. Energy level correlations in disordered metals: beyond universality. Journal of Mathematical Physics, 37(10):4968–4985, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zirnbauer:1996:RSS

[825] Martin R. Zirnbauer. Riemannian symmetric superspaces and their origin in random-matrix theory. Journal of Mathematical Physics, 37(10):4986– 5018, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Walker:1996:SSR

[826] Paul N. Walker, María José Sánchez, and Michael Wilkinson. Singularities in the spectra of random matrices. Journal of Mathematical Physics, 37(10):5019–5032, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khorunzhy:1996:APL

[827] Alexei M. Khorunzhy, Boris A. Khoruzhenko, and Leonid A. Pastur. Asymptotic properties of large random matrices with independent entries. Journal of Mathematical Physics, 37(10):5033–5060, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chalker:1996:FLD

[828] John T. Chalker, Igor V. Lerner, and Robert A. Smith. Fictitious level dynamics: a novel approach to spectral statistics in disordered conductors. Journal of Mathematical Physics, 37(10):5061–5086, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Richter:1996:IDM

[829] Klaus Richter, Denis Ullmo, and Rodolfo A. Jalabert. Integrability and disorder in mesoscopic systems: application to orbital magnetism. Journal of Mathematical Physics, 37(10):5087–5110, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Graffi:1996:EPI

[830] Sandro Graffi and André Martinez. Ergodic properties of infinite quantum harmonic crystals: an analytic approach. Journal of Mathematical Physics, 37(10):5111–5135, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lenci:1996:EPQ

[831] Marco Lenci. Ergodic properties of the quantum ideal gas in the Maxwell–Boltzmann statistics. Journal of Mathematical Physics, 37(10): 5136–5157, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Parisi:1996:MFT

[832] Giorgio Parisi. A mean field theory for arrays of Josephson junctions. Journal of Mathematical Physics, 37(10):5158–5170, October 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geyler:1996:SPC

[833] V. A. Geyler, B. S. Pavlov, and I. Yu. Popov. Spectral properties of a charged particle in antidot array: a limiting case of quantum billiard. Journal of Mathematical Physics, 37(10):5171–5194, October 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guarneri:1996:SCS

[834] Italo Guarneri. Singular continuous spectra and discrete wave packet dynamics. Journal of Mathematical Physics, 37(10):5195–5206, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kreft:1996:MHT

[835] Ch. Kreft and R. Seiler. Models of the Hofstadter-type. Journal of Mathematical Physics, 37(10):5207–5243, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benatti:1996:ESQ

[836] Fabio Benatti. Entropy of a subalgebra and quantum estimation. Journal of Mathematical Physics, 37(10):5244–5258, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Landauer:1996:CSV

[837] R. Landauer. Comment: “Spatial variation of currents and fields due to localized scatterers in metallic conduction” [IBM J. Res. Develop. 1 (1957), 223–231; MR 19, 805g]. With a reprint of the 1957 original. Journal of Mathematical Physics, 37(10):5259–5268, October 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1, 2].

Barnich:1996:IBB

[838] Glenn Barnich and Marc Henneaux. Isomorphisms between the Batalin– Vilkovisky antibracket and the Poisson bracket. Journal of Mathematical Physics, 37(11):5273–5296, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beckers:1996:GCA

[839] J. Beckers and N. Debergh. On Green–Cusson ansätze and deformed supersymmetric quantum mechanics. Journal of Mathematical Physics, 37(11):5297–5309, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carvalho:1996:BRSa

[840] M. Carvalho, L. C. Q. Vilar, C. A. G. Sasaki, and S. P. Sorella. Becchi– Rouet–Stora cohomology of zero curvature systems. I. The complete ladder case. Journal of Mathematical Physics, 37(11):5310–5324, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carvalho:1996:BRSb

[841] M. Carvalho, L. C. Q. Vilar, C. A. G. Sasaki, and S. P. Sorella. Becchi– Rouet–Stora cohomology of zero curvature systems. II. The noncomplete ladder case. Journal of Mathematical Physics, 37(11):5325–5336, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Havlicek:1996:DCQ

[842] Hans Havlicek and Karl Svozil. Density conditions for quantum propositions. Journal of Mathematical Physics, 37(11):5337–5341, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1996:GIB

[843] Chang-Yeong Lee. Geometrical interpretation of BRST symmetry in topological Yang–Mills–Higgs theory. Journal of Mathematical Physics, 37(11):5342–5350, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lunev:1996:RQL

[844] F. A. Lunev. Reformulation of QCD in the language of general relativity. Journal of Mathematical Physics, 37(11):5351–5367, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudolph:1996:CEH

[845] Oliver Rudolph. On the consistent effect histories approach to quantum mechanics. Journal of Mathematical Physics, 37(11):5368–5379, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Svozil:1996:GDN

[846] K. Svozil and J. Tkadlec. Greechie diagrams, nonexistence of measures in quantum logics, and Kochen–Specker-type constructions. Journal of Mathematical Physics, 37(11):5380–5401, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szopa:1996:CFR

[847] M. Szopa, G. Mys, and A. Ceulemans. The canonical form of the Rabi Hamiltonian. Journal of Mathematical Physics, 37(11):5402–5411, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1996:BFM

[848] Y. Wang and Fu-Cho Pu. Boson-fermion model with two-body interactions: exact results. Journal of Mathematical Physics, 37(11):5412–5423, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1996:ERD

[849] Y. Wang and Fu-Cho Pu. Exact results on a Dirac-like Lee model. Journal of Mathematical Physics, 37(11):5424–5428, November 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:1996:MEA

[850] Stephen L. Adler and L. P. Horwitz. Microcanonical ensemble and algebra of conserved generators for generalized quantum dynamics. Journal of Mathematical Physics, 37(11):5429–5443, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoniouk:1996:DCU

[851] Alexander Val. Antoniouk and Alexandra Vict. Antoniouk. Decay of correlations and uniqueness of Gibbs lattice systems with nonquadratic interaction. Journal of Mathematical Physics, 37(11):5444–5454, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chattopadhyay:1996:ITI

[852] Sankar Chattopadhyay and Dipankar Ray. Integration of thermodynamic identities for a relativistic general equation of state. Journal of Mathematical Physics, 37(11):5455–5457, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Contucci:1996:LAP

[853] Pierluigi Contucci and Andreas Knauf. The low activity phase of some Dirichlet series. Journal of Mathematical Physics, 37(11):5458–5475, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cornille:1996:CIT

[854] H. Cornille and A. d’Almeida. Criteria for inverted temperatures in evaporation-condensation processes. Journal of Mathematical Physics, 37(11):5476–5495, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouquet:1996:EGR

[855] Serge Bouquet and H. Ralph Lewis. Extension of Giacomini’s results concerning invariants for one-dimensional time-dependent potentials. Journal of Mathematical Physics, 37(11):5496–5508, November 1996. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bouquet:1996:SIO

[856] Serge Bouquet and H. Ralph Lewis. A second invariant for one-degree-of-freedom, time-dependent Hamiltonians given a first invariant. Journal of Mathematical Physics, 37(11):5509–5517, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chacon:1996:ICE

[857] Ricardo Chacón, F. Balibrea, and M. A. López. Inhibition of chaotic escape from a potential well using small parametric modulations. Journal of Mathematical Physics, 37(11):5518–5523, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1996:LAS

[858] Deng yuan Chen and Da jun Zhang. Lie algebraic structures of (1 + 1)- dimensional Lax integrable systems. Journal of Mathematical Physics, 37(11):5524–5538, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Labrunie:1996:PFI

[859] Simon Labrunie. On the polynomial first integrals of the (a, b, c) Lotka– Volterra system. Journal of Mathematical Physics, 37(11):5539–5550, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levi:1996:SDD

[860] D. Levi and P. Winternitz. Symmetries of discrete dynamical systems. Journal of Mathematical Physics, 37(11):5551–5576, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tasnadi:1996:BNT

[861] Tamás Tasnádi. The behavior of nearby trajectories in magnetic billiards. Journal of Mathematical Physics, 37(11):5577–5598, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagrov:1996:IED

[862] V. G. Bagrov, V. V. Obukhov, and A. G. Sakhapov. Integration of the Einstein–Dirac equations. Journal of Mathematical Physics, 37(11): 5599–5610, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:1996:FDQ

[863] Shuxue Ding, Yasushige Maeda, and Masaru Siino. Four-dimensional quantum topology changes of space–times. Journal of Mathematical Physics, 37(11):5611–5626, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dray:1996:EEP

[864] Tevian Dray. Einstein’s equations in the presence of signature change. Journal of Mathematical Physics, 37(11):5627–5636, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunn:1996:KGI

[865] K. A. Dunn, Tina A. Harriott, and J. G. Williams. Kinks and geodesic incompleteness. Journal of Mathematical Physics, 37(11):5637–5651, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gleiser:1996:TPS

[866] Reinaldo J. Gleiser, Alexander Garate, and Carlos O. Nicasio. Topological properties of single gravisolitons. Journal of Mathematical Physics, 37(11):5652–5664, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heller:1996:NSS

[867] Michael Heller and Wiesław Sasin. Noncommutative structure of singularities in general relativity. Journal of Mathematical Physics, 37(11): 5665–5671, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mansouri:1996:EDI

[868] R. Mansouri and M. Khorrami. The equivalence of Darmois–Israel and distributional method for thin shells in general relativity. Journal of Mathematical Physics, 37(11):5672–5683, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1996:GEP

[869] G. F. Torres del Castillo. Gravitational and electromagnetic perturbations of the Schwarzschild solution. Journal of Mathematical Physics, 37 (11):5684–5694, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

BrinisUdeschini:1996:NTA

[870] Elisa Brinis Udeschini and Giulio Magli. A null-tetrad approach to Kerr– Schild gravitational fields in matter. Journal of Mathematical Physics, 37(11):5695–5714, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alexandrov:1996:HKN

[871] Sergei Alexandrov and Dmitri Vassilevich. Heat kernel for nonminimal operators on a Kähler manifold. Journal of Mathematical Physics, 37 (11):5715–5718, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alisauskas:1996:TCT

[872] Sigitas Ališauskas. Towards the canonical tensor operators of u_{q}(3). I. The maximal null space case. Journal of Mathematical Physics, 37(11): 5719–5746, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bautista:1996:QCA

[873] R. Bautista, A. Criscuolo, M. – Dur– dević, M. Rosenbaum, and J. D. Vergara. Quantum Clifford algebras from spinor representations. Journal of Mathematical Physics, 37(11):5747–5775, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:1996:IPP

[874] Jiasu Cao, Sailing He, and Vaughan H. Weston. An ill-posed problem of the continuation of transient data for a hyperbolic equation in a threedimensional inhomogeneous half-space. Journal of Mathematical Physics, 37(11):5776–5791, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coll:1996:LAG

[875] Bartolomé Coll and Fernando San José Martínez. On the Lie algebras generated by two two-forms in Minkowski space–time. Journal of Mathematical Physics, 37(11):5792–5804, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Falomir:1996:DDO

[876] H. Falomir, R. E. Gamboa Saraví, M. A. Muschietti, E. M. Santangelo, and J. E. Solomin. Determinants of Dirac operators with local boundary conditions. Journal of Mathematical Physics, 37(11):5805–5819, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gonera:1996:DCQ

[877] Cezary Gonera, Piotr Kosiński, and Paweł Maślanka. Differential calculi on quantum Minkowski space. Journal of Mathematical Physics, 37(11): 5820–5827, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rabinovich:1996:SNR

[878] S. Rabinovich, G. Berkolaiko, and S. Havlin. Solving nonlinear recursions. Journal of Mathematical Physics, 37(11):5828–5836, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suchanecki:1996:RHS

[879] Z. Suchanecki, I. Antoniou, S. Tasaki, and O. F. Bandtlow. Rigged Hilbert spaces for chaotic dynamical systems. Journal of Mathematical Physics, 37(11):5837–5847, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tao:1996:SPG

[880] Terry Tao and Andrew C. Millard. On the structure of projective group representations in quaternionic Hilbert space. Journal of Mathematical Physics, 37(11):5848–5857, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comments [651, 927, 926].

vandenHijligenberg:1996:ARN

[881] N. W. van den Hijligenberg and Yu. Yu. Kotchetkov. The absolute rigidity of the Neveu–Schwarz and Ramond superalgebras. Journal of Mathematical Physics, 37(11):5858–5868, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Varzugin:1996:AOR

[882] G. G. Varzugin. Asymptotics of oscillatory Riemann–Hilbert problems. Journal of Mathematical Physics, 37(11):5869–5892, November 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:1996:FSM

[883] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. Factorization of scattering matrices due to partitioning of potentials in one-dimensional Schrödinger-type equations. Journal of Mathematical Physics, 37(12): 5897–5915, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1256].

Basdevant:1996:IOB

[884] Jean-Louis Basdevant and André Martin. l^{P} interpolation and optimized bounds on pairwise interacting fermion systems. Journal of Mathematical Physics, 37(12):5916–5927, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baumgartner:1996:IRT

[885] B. Baumgartner. Interchannel resonances at a threshold. Journal of Mathematical Physics, 37(12):5928–5938, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bezak:1996:TCE

[886] V. Bezák. Transmission of conduction electrons through a symmetric pair of delta-barriers or delta-wells embedded in a semiconductor or a metal. Journal of Mathematical Physics, 37(12):5939–5957, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brooke:1996:LPP

[887] J. A. Brooke and F. E. Schroeck, Jr. Localization of the photon on phase space. Journal of Mathematical Physics, 37(12):5958–5986, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujii:1996:MPC

[888] Kazuyuki Fujii and Kunio Funahashi. Multi-periodic coherent states and the WKB exactness. Journal of Mathematical Physics, 37(12):5987–6011, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galvao:1996:HSD

[889] C. A. P. Galvão and Y. Nutku. Hamiltonian structure of Dubrovin’s equation of associativity in 2-d topological field theory. Journal of Mathematical Physics, 37(12):6012–6017, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gratus:1996:CQI

[890] Jonathan Gratus and Robin W. Tucker. Classical and quantum implications of the causality structure of two-dimensional space–times with degenerate metrics. Journal of Mathematical Physics, 37(12):6018–6032, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiers:1996:SOD

[891] K. A. Kiers and W. van Dijk. Scattering in one dimension: the coupled Schrödinger equation, threshold behaviour and Levinson’s theorem. Journal of Mathematical Physics, 37(12):6033–6059, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Navarro:1996:SRH

[892] D. J. Navarro and J. Navarro-Salas. Special-relativistic harmonic oscillator modeled by Klein–Gordon theory in anti-de Sitter space. Journal of Mathematical Physics, 37(12):6060–6073, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:1996:PYM

[893] Yüji Ohta. Prepotential of N = 2SU(2) Yang–Mills gauge theory coupled with a massive matter multiplet. Journal of Mathematical Physics, 37 (12):6074–6085, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schreckenberg:1996:SHQ

[894] S. Schreckenberg. Symmetry and history quantum theory: an analog of Wigner’s theorem. Journal of Mathematical Physics, 37(12):6086–6105, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:1996:LAC

[895] Hyun Seok Yang and Bum-Hoon Lee. Lie algebra cohomology and group structure of gauge theories. Journal of Mathematical Physics, 37(12): 6106–6120, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colatto:1996:DSC

[896] L. P. Colatto and J. L. Matheus-Valle. On q-deformed supersymmetric classical mechanical models. Journal of Mathematical Physics, 37(12): 6121–6129, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shevtsov:1996:IPN

[897] Boris M. Shevtsov. Inverse problem in nonstationary multidimensional medium. Journal of Mathematical Physics, 37(12):6130–6138, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saveliev:1996:TMD

[898] Vladimir Saveliev. A temperature and mass dependence of the linear Boltzmann collision operator from group theory point of view. Journal of Mathematical Physics, 37(12):6139–6151, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Constantin:1996:RES

[899] Peter Constantin, Charles R. Doering, and Edriss S. Titi. Rigorous estimates of small scales in turbulent flows. Journal of Mathematical Physics, 37(12):6152–6156, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ibort:1996:ESS

[900] A. Ibort, L. Martínez Alonso, and E. Medina Reus. Explicit solutions of supersymmetric KP hierarchies: supersolitons and solitinos. Journal of Mathematical Physics, 37(12):6157–6172, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kosmann-Schwarzbach:1996:LNO

[901] Y. Kosmann-Schwarzbach and F. Magri. Lax–Nijenhuis operators for integrable systems. Journal of Mathematical Physics, 37(12):6173–6197, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Labrunie:1996:GMT

[902] Simon Labrunie and Robert Conte. A geometrical method towards first integrals for dynamical systems. Journal of Mathematical Physics, 37 (12):6198–6206, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maccari:1996:KPE

[903] Attilio Maccari. The Kadomtsev–Petviashvili equation as a source of integrable model equations. Journal of Mathematical Physics, 37(12): 6207–6212, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oevel:1996:WSC

[904] Walter Oevel and Walter Strampp. Wronskian solutions of the constrained Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 37(12):6213–6219, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soto-Trevino:1996:HOM

[905] Cristina Soto-Treviño and Tasso J. Kaper. Higher-order Melnikov theory for adiabatic systems. Journal of Mathematical Physics, 37(12):6220– 6249, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashtekar:1996:PQG

[906] A. Ashtekar and M. Pierri. Probing quantum gravity through exactly soluble midi-superspaces. I. Journal of Mathematical Physics, 37(12): 6250–6270, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bezerra:1996:LVH

[907] V. B. Bezerra and P. S. Letelier. Loop variables and holonomies for a class of conical space–times. Journal of Mathematical Physics, 37(12): 6271–6282, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ibanez:1996:ABC

[908] J. Ibáñez and I. Olasagasti. Asymptotic behavior of a class of inhomogeneous scalar field cosmologies. Journal of Mathematical Physics, 37(12): 6283–6292, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maluf:1996:DGE

[909] José W. Maluf. On the distribution of gravitational energy in the de Sitter space. Journal of Mathematical Physics, 37(12):6293–6301, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maluf:1996:GER

[910] J. W. Maluf, E. F. Martins, and A. Kneip. Gravitational energy of rotating black holes. Journal of Mathematical Physics, 37(12):6302–6310, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bordemann:1996:SCS

[911] M. Bordemann, M. Brischle, C. Emmrich, and S. Waldmann. Subalgebras with converging star products in deformation quantization: an algebraic construction for CP^{n}. Journal of Mathematical Physics, 37 (12):6311–6323, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsko:1996:AMS

[912] A. G. Bytsko and L. D. Faddeev. (T^{∗}B)_{q}, q-analog of model space and the Clebsch–Gordan coefficients generating matrices. Journal of Mathematical Physics, 37(12):6324–6348, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daumens:1996:RBC

[913] Michel Daumens and Pierre Minnaert. Relations between the Casimir operators of sl(1|2) and osp(2|2) superalgebras. Journal of Mathematical Physics, 37(12):6349–6352, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fuhr:1996:WFA

[914] Hartmut Führ. Wavelet frames and admissibility in higher dimensions. Journal of Mathematical Physics, 37(12):6353–6366, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fukushima:1996:AAS

[915] Nobuhisa Fukushima. An algebra arising from 2-state chiral Potts model and Sklyanin algebra. Journal of Mathematical Physics, 37(12):6367– 6374, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gross:1996:CCF

[916] Christian Gross. Cohomology and connections on fiber bundles and applications to field theories. Journal of Mathematical Physics, 37(12): 6375–6394, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guhr:1996:IZL

[917] Thomas Guhr and Tilo Wettig. An Itzykson–Zuber-like integral and diffusion for complex ordinary and supermatrices. Journal of Mathematical Physics, 37(12):6395–6413, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heredero:1996:CIW

[918] Rafael Hernández Heredero and Peter J. Olver. Classification of invariant wave equations. Journal of Mathematical Physics, 37(12):6414–6438, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:1996:SAP

[919] E. G. Kalnins, W. Miller, Jr., and G. S. Pogosyan. Superintegrability and associated polynomial solutions: Euclidean space and the sphere in two dimensions. Journal of Mathematical Physics, 37(12):6439–6467, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klink:1996:MIT

[920] W. H. Klink and T. Ton-That. Multiplicity, invariants, and tensor product decompositions of compact groups. Journal of Mathematical Physics, 37(12):6468–6485, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Latrach:1996:EST

[921] Khalid Latrach and Aref Jeribi. On the essential spectrum of transport operators on L_{1}-spaces. Journal of Mathematical Physics, 37(12):6486– 6494, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:1996:BGC

[922] Shahn Majid. Braided geometry of the conformal algebra. Journal of Mathematical Physics, 37(12):6495–6509, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mas:1996:APS

[923] Javier Mas and Marcos Seco. The algebra of q-pseudodifferential symbols and the q-W_{KP}^{(n)} algebra. Journal of Mathematical Physics, 37(12):6510– 6529, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prakash:1996:CLA

[924] J. S. Prakash and H. S. Sharatchandra. A calculus for SU(3) leading to an algebraic formula for the Clebsch–Gordan coefficients. Journal of Mathematical Physics, 37(12):6530–6569, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thoma:1996:OOBb

[925] M. Thoma and R. T. Sharp. Orbit-orbit branching rules between classical simple Lie algebras and maximal reductive subalgebras. Journal of Mathematical Physics, 37(12):6570–6581, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Emch:1996:CRP

[926] Gérard G. Emch. Comments on a recent paper by S. L. Adler: “Projective group representations in quaternionic Hilbert spaces”, [J. Math. Phys. 37 (1996), no. 5, 2352–2360; 1385834]. Journal of Mathematical Physics, 37(12):6582–6585, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [651, 927, 880].

Adler:1996:RCR

[927] Stephen L. Adler. Response to: “Comments on a recent paper by S. L. Adler: ‘Projective group representations in quaternionic Hilbert spaces’ ”, [J. Math. Phys. 37 (1996), no. 12, 6582–6585; 1419186] by G. Emch. Journal of Mathematical Physics, 37(12):6586–6589, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [651, 926, 880].

Fernandes:1996:EGI

[928] Joe F. Q. Fernandes and Anthony W. C. Lun. Erratum: “Gauge invariant perturbations of black holes. I. Schwarzschild space–time”, [J. Math. Phys. 37 (1996), no. 2, 836–857; MR 96k:83073]. Journal of Mathematical Physics, 37(12):6590, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [552].

Vathsan:1996:ERQ

[929] Radhika Vathsan. Erratum: “Remarks on the quantization of gauge theories”. Journal of Mathematical Physics, 37(12):6590, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [611].

Anonymous:1996:SPA

[930] Anonymous. Summary of the Physics and Astronomy Classification Scheme — 1996. Journal of Mathematical Physics, 37(12):6591, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1996:SPH

[931] Anonymous. Summary of PACS headings in current index. Journal of Mathematical Physics, 37(12):6592–6595, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1996:SIV

[932] Anonymous. Subject index to volume 37. Journal of Mathematical Physics, 37(12):6596–6626, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1996:AIV

[933] Anonymous. Author index to volume 37. Journal of Mathematical Physics, 37(12):6627–6649, December 1996. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blanchard:1997:NNF

[934] Philippe Blanchard, Piotr Garbaczewski, and Robert Olkiewicz. Nonnegative Feynman–Kac kernels in Schrödinger’s interpolation problem. Journal of Mathematical Physics, 38(1):1–15, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dolinszky:1997:SSP

[935] T. Dolinszky. Scattering by singular potentials. B. Varying the linear parameters. Journal of Mathematical Physics, 38(1):16–29, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hwang:1997:NAT

[936] Dae Sung Hwang and Chang-Yeong Lee. Non-abelian topological mass generation in four dimensions. Journal of Mathematical Physics, 38(1): 30–38, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jung:1997:GAI

[937] Wolf Jung. Geometrical approach to inverse scattering for the Dirac equation. Journal of Mathematical Physics, 38(1):39–48, January 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karimipour:1997:LTF

[938] V. Karimipour and A. Mostafazadeh. Lattice topological field theory on nonorientable surfaces. Journal of Mathematical Physics, 38(1):49–66, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimek:1997:EPQ

[939] Sławomir Klimek, Andrzej Leśniewski, Neepa Maitra, and Ron Rubin. Ergodic properties of quantized toral automorphisms. Journal of Mathematical Physics, 38(1):67–83, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nieto:1997:DOSa

[940] Michael Martin Nieto and D. Rodney Truax. Displacement-operator squeezed states. I. Time-dependent systems having isomorphic symmetry algebras. Journal of Mathematical Physics, 38(1):84–97, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nieto:1997:DOSb

[941] Michael Martin Nieto and D. Rodney Truax. Displacement-operator squeezed states. II. Examples of time-dependent systems having isomorphic symmetry algebras. Journal of Mathematical Physics, 38(1):98–114, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pelc:1997:CCS

[942] Oskar Pelc and L. P. Horwitz. Construction of a complete set of states in relativistic scattering theory. Journal of Mathematical Physics, 38(1): 115–138, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pelc:1997:GCM

[943] Oskar Pelc and L. P. Horwitz. Generalization of the Coleman–Mandula theorem to higher dimension. Journal of Mathematical Physics, 38(1): 139–172, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zastawniak:1997:GSC

[944] Tomasz Zastawniak. Geometrical stochastic control and quantization. Journal of Mathematical Physics, 38(1):173–181, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

He:1997:TDE

[945] Sailing He and Vaughan H. Weston. Three-dimensional electromagnetic inverse scattering for bi-isotropic dispersive media. Journal of Mathematical Physics, 38(1):182–195, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiganov:1997:KTN

[946] A. V. Tsiganov. The Kowalewski top: a new Lax representation. Journal of Mathematical Physics, 38(1):196–211, January 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Corona-Corona:1997:HSD

[947] Gulmaro Corona-Corona. Hamiltonian structure for degenerate AKNS systems. Journal of Mathematical Physics, 38(1):212–225, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:1997:IST

[948] Nian-Ning Huang, Zong-Yun Chen, and Xiang-Jun Chen. An inverse scattering transform for the MKdV equation with non-vanishing boundary value. Journal of Mathematical Physics, 38(1):226–246, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kasman:1997:OPF

[949] Alex Kasman. Orthogonal polynomials and the finite Toda lattice. Journal of Mathematical Physics, 38(1):247–254, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kawahara:1997:PLC

[950] Yasuo Kawahara and Hyen Yeal Lee. Period lengths of cellular automata cam-90 with memory. Journal of Mathematical Physics, 38(1):255–266, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Knessl:1997:ADPa

[951] Charles Knessl and Joseph B. Keller. Advection diffusion past a strip. I. Normal incidence. Journal of Mathematical Physics, 38(1):267–282, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loris:1997:SCKa

[952] Ignace Loris and Ralph Willox. On solutions of constrained KP equations. Journal of Mathematical Physics, 38(1):283–291, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Radha:1997:ECS

[953] R. Radha and M. Lakshmanan. Exotic coherent structures in the (2 + 1)-dimensional long dispersive wave equation. Journal of Mathematical Physics, 38(1):292–299, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thacker:1997:PIT

[954] W. D. Thacker. A path integral for turbulence in incompressible fluids. Journal of Mathematical Physics, 38(1):300–320, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:1997:SVC

[955] Yunbo Zeng. Separation of variables for the constrained flows. Journal of Mathematical Physics, 38(1):321–329, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandes:1997:GIP

[956] Joe F. Q. Fernandes and Anthony W. C. Lun. Gauge invariant perturbations of black holes. II. Kerr space–time. Journal of Mathematical Physics, 38(1):330–349, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanimoto:1997:DCH

[957] Masayuki Tanimoto, Tatsuhiko Koike, and Akio Hosoya. Dynamics of compact homogeneous universes. Journal of Mathematical Physics, 38 (1):350–368, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bonatsos:1997:NDA

[958] D. Bonatsos, C. Daskaloyannis, P. Kolokotronis, A. Ludu, and C. Quesne. A nonlinear deformed su(2) algebra with a two-color quasitriangular Hopf structure. Journal of Mathematical Physics, 38(1):369–386, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1997:AEI

[959] Jin-Quan Chen and Jia-Lun Ping. Algebraic expressions for irreducible bases of icosahedral group. Journal of Mathematical Physics, 38(1):387– 410, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:1997:DCQ

[960] Heng Fan, Bo yu Hou, and Kang jie Shi. Drinfeld constructions of the quantum affine superalgebra U_{q}(ĝl(m/n)). Journal of Mathematical Physics, 38(1):411–433, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konopelchenko:1997:SCS

[961] B. G. Konopelchenko. Soliton curvatures of surfaces and spaces. Journal of Mathematical Physics, 38(1):434–457, January 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maluf:1997:GEC

[962] José W. Maluf and Andreas Kneip. Gravitational energy of conical defects. Journal of Mathematical Physics, 38(1):458–465, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mishra:1997:FSR

[963] A. K. Mishra and G. Rajasekaran. Fock space representation of differential calculus on the noncommutative quantum space. Journal of Mathematical Physics, 38(1):466–475, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mudrov:1997:QGA

[964] A. I. Mudrov. On quantization of Z_{2}-graded algebras. Journal of Mathematical Physics, 38(1):476–483, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1776].

Schwartz:1997:NOT

[965] Charles Schwartz. Nonlinear operators and their propagators. Journal of Mathematical Physics, 38(1):484–500, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tolar:1997:GCS

[966] J. Tolar and P. Trávníček. Graded contractions of symplectic Lie algebras in collective models. Journal of Mathematical Physics, 38(1):501–523, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1997:KDC

[967] Shuzhou Wang. Krein duality for compact quantum groups. Journal of Mathematical Physics, 38(1):524–534, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:1997:SQA

[968] R. B. Zhang. Symmetrizable quantum affine superalgebras and their representations. Journal of Mathematical Physics, 38(1):535–543, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chicone:1997:AIK

[969] C. Chicone, B. Mashhoon, and D. G. Retzloff. Addendum: “On the ionization of a Keplerian binary system by periodic gravitational radiation”, [J. Math. Phys. 37 (1996), no. 8, 3997–4016; MR 97d:70012]. Journal of Mathematical Physics, 38(1):544, January 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [759].

Abe:1997:OOI

[970] Mitsuo Abe, Noriaki Ikeda, and Noboru Nakanishi. Operator ordering index method for multiple commutators and Suzuki’s quantum analysis. Journal of Mathematical Physics, 38(2):547–555, February 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benvegnu:1997:RPI

[971] Stefan Benvegnù. Relativistic point interaction with Coulomb potential in one dimension. Journal of Mathematical Physics, 38(2):556–570, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Corley:1997:BSG

[972] S. R. Corley and O. W. Greenberg. Bound states in Galilean-invariant quantum field theory. Journal of Mathematical Physics, 38(2):571–581, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeLeo:1997:OR

[973] Stefano De Leo and Khaled Abdel-Khalek. Octonionic representations of GL(8, R) and GL(4, C). Journal of Mathematical Physics, 38(2):582–598, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferraro:1997:CGP

[974] Rafael Ferraro and Claudio Simeone. Canonical gauges in the path integral for parametrized systems. Journal of Mathematical Physics, 38(2): 599–610, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fulop:1997:RQM

[975] T. Fülöp. Relativistic quantum mechanics on the SL(2, R) space–time. Journal of Mathematical Physics, 38(2):611–621, February 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gabriel:1997:MSP

[976] Cl. Gabriel and Ph. Spindel. Massive spin-2 propagators on de Sitter space. Journal of Mathematical Physics, 38(2):622–638, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lay:1997:QO

[977] Wolfgang Lay. The quartic oscillator. Journal of Mathematical Physics, 38(2):639–647, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Metsaev:1997:MFP

[978] R. R. Metsaev. Massless fields in plane wave geometry. Journal of Mathematical Physics, 38(2):648–667, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mukku:1997:FDE

[979] C. Mukku. Fermion determinants and effective actions. Journal of Mathematical Physics, 38(2):668–681, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:1997:PYM

[980] Yűji Ohta. Prepotentials of N = 2SU(2) Yang–Mills theories coupled with massive matter multiplets. Journal of Mathematical Physics, 38(2): 682–696, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1294].

Omnes:1997:QCC

[981] Roland Omnès. Quantum-classical correspondence using projection operators. Journal of Mathematical Physics, 38(2):697–707, February 1997.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Omnes:1997:LRP

[982] Roland Omnès. Localization of relativistic particles. Journal of Mathematical Physics, 38(2):708–715, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sakoda:1997:ABS

[983] Seiji Sakoda and Minoru Omote. Aharonov–Bohm scattering: the role of the incident wave. Journal of Mathematical Physics, 38(2):716–737, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schnetz:1997:NR

[984] Oliver Schnetz. Natural renormalization. Journal of Mathematical Physics, 38(2):738–758, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schreckenberg:1997:SDF

[985] S. Schreckenberg. Symmetries of decoherence functionals. Journal of Mathematical Physics, 38(2):759–769, February 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sordoni:1997:IS

[986] V. Sordoni. Instantons and splitting. Journal of Mathematical Physics, 38(2):770–795, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Williams:1997:TCE

[987] Floyd L. Williams. Topological Casimir energy for a general class of Clifford–Klein space–times. Journal of Mathematical Physics, 38(2):796– 808, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:1997:ESL

[988] Bang-Rong Zhou. Exact solutions of linearized Schwinger–Dyson equation of fermion self-energy. Journal of Mathematical Physics, 38(2):809– 820, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1997:MSP

[989] F. Calogero. Motion of strings in the plane: a solvable model. I. Journal of Mathematical Physics, 38(2):821–829, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kramer:1997:KES

[990] Eric M. Kramer. The von Karman equations, the stress function, and elastic ridges in high dimensions. Journal of Mathematical Physics, 38 (2):830–846, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sardanashvily:1997:SEM

[991] Gennadi A. Sardanashvily. Stress-energy-momentum tensors in constraint field theories. Journal of Mathematical Physics, 38(2):847–866, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gough:1997:IFQ

[992] John Gough. The interacting-free quantum stochastic limit of quantum field theory. Journal of Mathematical Physics, 38(2):867–881, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferreira:1997:TFD

[993] Luiz A. Ferreira, J. Luis Miramontes, and Joaquín Sánchez Guillén. Tau-functions and dressing transformations for zero-curvature affine integrable equations. Journal of Mathematical Physics, 38(2):882–901, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Knessl:1997:ADPb

[994] Charles Knessl and Joseph B. Keller. Advection-diffusion past a strip. II. Oblique incidence. Journal of Mathematical Physics, 38(2):902–925, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McDaniel:1997:FPT

[995] Andrew McDaniel and Lawrence Smolinsky. The flow of the G_{2} periodic Toda lattice. Journal of Mathematical Physics, 38(2):926–945, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Katanaev:1997:ETD

[996] M. O. Katanaev. Euclidean two-dimensional gravity with torsion. Journal of Mathematical Physics, 38(2):946–980, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zha:1997:UAC

[997] Chao zheng Zha and Wei zhong Zhao. The universal anomalies and condition for anomaly-free high-order Virasoro algebra. Journal of Mathematical Physics, 38(2):981–988, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:1997:LOI

[998] Y. Brihaye, Stefan Giller, and Piotr Kosinski. Linear operators with invariant polynomial space and graded algebra. Journal of Mathematical Physics, 38(2):989–999, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bueken:1997:TDL

[999] Peter Bueken. Three-dimensional Lorentzian manifolds with constant principal Ricci curvatures ρ_{1} = ρ_{2} ≠ ρ_{3}. Journal of Mathematical Physics, 38(2):1000–1013, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

daCosta:1997:FIS

[1000] G. A. T. F. da Costa. Feynman identity: a special case. I. Journal of Mathematical Physics, 38(2):1014–1034, February 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duplij:1997:NSC

[1001] Steven Duplij. Noninvertible N = 1 superanalog of complex structure. Journal of Mathematical Physics, 38(2):1035–1040, February 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lascoux:1997:RTH

[1002] Alain Lascoux, Bernard Leclerc, and Jean-Yves Thibon. Ribbon tableaux, Hall–Littlewood functions, quantum affine algebras, and unipotent varieties. Journal of Mathematical Physics, 38(2):1041–1068, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruijsenaars:1997:FOA

[1003] S. N. M. Ruijsenaars. First order analytic difference equations and integrable quantum systems. Journal of Mathematical Physics, 38(2):1069– 1146, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scolarici:1997:QRM

[1004] G. Scolarici and L. Solombrino. Quaternionic representations of magnetic groups. Journal of Mathematical Physics, 38(2):1147–1160, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stovicek:1997:RDQ

[1005] Pavel Šťovíček and Reidun Twarock. Representations of U_{⟨}(∫⊓(N)) derived from quantum flag manifolds. Journal of Mathematical Physics, 38(2):1161–1182, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suzuki:1997:CEP

[1006] Masuo Suzuki. Compact exponential product formulas and operator functional derivative. Journal of Mathematical Physics, 38(2):1183–1196, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhdanov:1997:SVD

[1007] R. Z. Zhdanov. Separation of variables in (1+2)-dimensional Schrödinger equations. Journal of Mathematical Physics, 38(2):1197–1217, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kerner:1997:CHS

[1008] Edward H. Kerner. Comment on Hamiltonian structures for the ndimensional Lotka–Volterra equations. Journal of Mathematical Physics, 38(2):1218–1223, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [257, 735].

Delduc:1997:ESK

[1009] F. Delduc, E. Ivanov, and S. Krivonos. Erratum: “N = 4 super KdV hierarchy in N = 4 and N = 2 superspaces”, [J. Math. Phys. 37 (1996), no. 3, 1356–1381; MR 97e:58106]. Journal of Mathematical Physics, 38 (2):1224, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [590].

Bockenhauer:1997:HLW

[1010] Jens Böckenhauer and Jürgen Fuchs. Higher level WZW sectors from free fermions. Journal of Mathematical Physics, 38(3):1227–1256, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Booth:1997:DME

[1011] H. S. Booth and C. J. Radford. The Dirac–Maxwell equations with cylindrical symmetry. Journal of Mathematical Physics, 38(3):1257–1268, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carminati:1997:NCS

[1012] Lionel Carminati, Bruno Iochum, and Thomas Schücker. The noncommutative constraints on the standard model à la Connes. Journal of Mathematical Physics, 38(3):1269–1280, March 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:1997:NAB

[1013] Ho-Meoyng Choi, Chueng-Ryong Ji, and Hyun Kyu Lee. Non-abelian Berry’s phase in a slowly rotating electric field. Journal of Mathematical Physics, 38(3):1281–1288, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Erdos:1997:DPN

[1014] László Erdős. Dia- and paramagnetism for nonhomogeneous magnetic fields. Journal of Mathematical Physics, 38(3):1289–1317, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Florig:1997:SIA

[1015] Martin Florig and Stephen J. Summers. On the statistical independence of algebras of observables. Journal of Mathematical Physics, 38(3):1318– 1328, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jensen:1997:GRS

[1016] Bjørn Jensen. Geometries of relativistic strings and other p-branes. Journal of Mathematical Physics, 38(3):1329–1346, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kempf:1997:QFT

[1017] Achim Kempf. On quantum field theory with nonzero minimal uncertainties in positions and momenta. Journal of Mathematical Physics, 38 (3):1347–1372, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klein:1997:BOA

[1018] Markus Klein, André Martinez, and Xue Ping Wang. On the Born– Oppenheimer approximation of diatomic wave operators. II. Singular potentials. Journal of Mathematical Physics, 38(3):1373–1396, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martina:1997:TFT

[1019] L. Martina, O. K. Pashaev, and G. Soliani. Topological field theories and integrable models. Journal of Mathematical Physics, 38(3):1397–1412, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naito:1997:IVF

[1020] Seichi Naito. Inlaying vertex function and scattering amplitude. Journal of Mathematical Physics, 38(3):1413–1453, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Navarro:1997:GQC

[1021] Miguel Navarro, Victor Aldaya, and Manuel Calixto. Group quantization on configuration space: gauge symmetries and linear fields. Journal of Mathematical Physics, 38(3):1454–1476, March 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chacon:1997:CGR

[1022] Ricardo Chacón. Chaos and geometrical resonance in the damped pendulum subjected to periodic pulses. Journal of Mathematical Physics, 38 (3):1477–1483, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Donnat:1997:DNG

[1023] Phillipe Donnat and Jeffrey Rauch. Dispersive nonlinear geometric optics. Journal of Mathematical Physics, 38(3):1484–1523, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ray:1997:BAS

[1024] Subhankar Ray. Bethe ansatz study for ground state of Fateev Zamolodchikov model. Journal of Mathematical Physics, 38(3):1524–1558, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aratyn:1997:CKM

[1025] H. Aratyn, L. A. Ferreira, J. F. Gomes, and A. H. Zimerman. Constrained KP models as integrable matrix hierarchies. Journal of Mathematical Physics, 38(3):1559–1576, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karimipour:1997:SHS

[1026] V. Karimipour. A solvable Hamiltonian system: integrability and actionangle variables. Journal of Mathematical Physics, 38(3):1577–1582, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nunez:1997:SRR

[1027] Manuel Núñez. Some rigorous results for the kinematic dynamo problem with general boundary conditions. Journal of Mathematical Physics, 38 (3):1583–1592, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kriele:1997:OTS

[1028] Marcus Kriele and Sean A. Hayward. Outer trapped surfaces and their apparent horizon. Journal of Mathematical Physics, 38(3):1593–1604, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petters:1997:MGL

[1029] A. O. Petters. Multiplane gravitational lensing. III. Upper bound on number of images. Journal of Mathematical Physics, 38(3):1605–1613, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Steinbauer:1997:URN

[1030] R. Steinbauer. The ultrarelativistic Reissner–Nordstrøm field in the Colombeau algebra. Journal of Mathematical Physics, 38(3):1614–1622, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abawi:1997:ELS

[1031] Ahmad T. Abawi, Roger F. Dashen, and Herbert Levine. The eigenvalues of the Laplacian on a sphere with boundary conditions specified on a segment of a great circle. Journal of Mathematical Physics, 38(3):1623– 1649, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abramov:1997:HGG

[1032] Viktor Abramov, Richard Kerner, and Bertrand Le Roy. Hypersymmetry: a Z_{3}-graded generalization of supersymmetry. Journal of Mathematical Physics, 38(3):1650–1669, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baguis:1997:MWR

[1033] P. Baguis and T. Stavracou. Marsden–Weinstein reduction on graded symplectic manifolds. Journal of Mathematical Physics, 38(3):1670–1684, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doliwa:1997:HMP

[1034] Adam Doliwa. Harmonic maps and periodic Toda systems. Journal of Mathematical Physics, 38(3):1685–1691, March 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gorbar:1997:HKE

[1035] E. V. Gorbar. Heat kernel expansion for operators containing a root of the Laplace operator. Journal of Mathematical Physics, 38(3):1692–1699, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kent:1997:MOM

[1036] Robert D. Kent and Mordechay Schlesinger. Multibody operator matrix elements and subduction coefficients in U(n). II. Journal of Mathematical Physics, 38(3):1700–1709, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salcedo:1997:RCP

[1037] L. L. Salcedo. Representation of complex probabilities. Journal of Mathematical Physics, 38(3):1710–1722, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wall:1997:IIH

[1038] David J. N. Wall and Peter Olsson. Invariant imbedding and hyperbolic heat waves. Journal of Mathematical Physics, 38(3):1723–1749, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bellissard:1997:ISI

[1039] J. Bellissard. Introduction [to special issue on condensed matter physics]. Journal of Mathematical Physics, 38(4):1753–1756, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jaksic:1997:STT

[1040] Vojkan Jakšić and Claude-Alain Pillet. Spectral theory of thermal relaxation. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1757–1780, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Akkermans:1997:TBC

[1041] E. Akkermans. Twisted boundary conditions and transport in disordered systems. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1781–1793, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Roche:1997:ETP

[1042] S. Roche, G. Trambly de Laissardière, and D. Mayou. Electronic transport properties of quasicrystals. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1794–1822, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kellendonk:1997:TET

[1043] Johannes Kellendonk. Topological equivalence of tilings. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38 (4):1823–1842, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Allouche:1997:SOR

[1044] J.-P. Allouche. Schrödinger operators with Rudin–Shapiro potentials are not palindromic. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1843–1848, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Destainville:1997:EOD

[1045] N. Destainville and J. F. Sadoc. Excitations in one dimension: a geometrical view of the transfer matrix method. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1849– 1863, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Varga:1997:TWG

[1046] Peter Varga. Taming of the wild group of magnetic translations. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1864–1869, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guhr:1997:SCC

[1047] Thomas Guhr and Axel Müller-Groeling. Spectral correlations in the crossover between GUE and Poisson regularity: on the identification of scales. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1870–1887, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mirlin:1997:SSA

[1048] Alexander D. Mirlin. Spatial structure of anomalously localized states in disordered conductors. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1888–1917, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fyodorov:1997:SRP

[1049] Yan V. Fyodorov and Hans-Jürgen Sommers. Statistics of resonance poles, phase shifts and time delays in quantum chaotic scattering: random matrix approach for systems with broken time-reversal invariance. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1918–1981, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simons:1997:TRS

[1050] B. D. Simons, O. Agam, and A. V. Andreev. Time-reversal symmetry breaking and the field theory of quantum chaos. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):1982– 2006, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zirnbauer:1997:TTI

[1051] Martin R. Zirnbauer. Toward a theory of the integer quantum Hall transition: continuum limit of the Chalker–Coddington model. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38 (4):2007–2036, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1959].

March:1997:DMI

[1052] N. H. March. Density matrices for itinerant and localized electrons with and without external fields. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):2037–2052, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sewell:1997:DLR

[1053] Geoffrey L. Sewell. Off-diagonal long range order and superconductive electrodynamics. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):2053–2071, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bach:1997:AHF

[1054] Volker Bach and Jurij Poelchau. Accuracy of the Hartree–Fock approximation for the Hubbard model. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):2072–2083, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Macris:1997:GSL

[1055] N. Macris and J. L. Lebowitz. Ground states and low-temperature phases of itinerant electrons interacting with classical fields: a review of rigorous results. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):2084–2103, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baesens:1997:FCL

[1056] C. Baesens and R. S. MacKay. Finite coherence length for equilibrium states of generalized adiabatic Holstein models. Quantum problems in condensed matter physics. Journal of Mathematical Physics, 38(4):2104– 2114, April 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:1997:CSQ

[1057] Stephen L. Adler and Andrew C. Millard. Coherent states in quaternionic quantum mechanics. Journal of Mathematical Physics, 38(5):2117–2126, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bay:1997:SQO

[1058] Karlheinz Bay and Wolfgang Lay. The spectrum of the quartic oscillator. Journal of Mathematical Physics, 38(5):2127–2131, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chan:1997:IO

[1059] Gordon Chan, Robert Finkelstein, and Vadim Oganesyan. The qisotropic oscillator. Journal of Mathematical Physics, 38(5):2132–2147, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:1997:COD

[1060] W.-S. Chung. The gl\sqrt{q}(n)-covariant oscillators and q-deformed quantum mechanics in n dimensions. Journal of Mathematical Physics, 38(5): 2148–2153, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fu:1997:HST

[1061] Hong-Chen Fu and Ryu Sasaki. Hypergeometric states and their nonclassical properties. Journal of Mathematical Physics, 38(5):2154–2166, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gitman:1997:PMW

[1062] D. M. Gitman and A. E. Gonçalves. Pseudoclassical model for Weyl particle in 10 dimensions. Journal of Mathematical Physics, 38(5):2167– 2170, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Henderson:1997:RPI

[1063] R. J. Henderson and S. G. Rajeev. Renormalized path integral in quantum mechanics. Journal of Mathematical Physics, 38(5):2171–2180, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jallouli:1997:IAM

[1064] H. Jallouli and H. Sazdjian. Incorporation of anomalous magnetic moments in the two-body relativistic wave equations of constraint theory. Journal of Mathematical Physics, 38(5):2181–2196, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Low:1997:CRQ

[1065] Stephen G. Low. Canonically relativistic quantum mechanics: representations of the unitary semidirect Heisenberg group, U(1, 3) ⊗_{s} H(1, 3). Journal of Mathematical Physics, 38(5):2197–2209, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Malbouisson:1997:ARY

[1066] A. P. C. Malbouisson, B. F. Svaiter, and N. F. Svaiter. Analytic regularization of the Yukawa model at finite temperature. Journal of Mathematical Physics, 38(5):2210–2218, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manoliu:1997:QST

[1067] Mihaela Manoliu. Quantization of symplectic tori in a real polarization. Journal of Mathematical Physics, 38(5):2219–2254, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pris:1997:NGB

[1068] Igor Pris and Thomas Schücker. Noncommutative geometry beyond the standard model. Journal of Mathematical Physics, 38(5):2255–2265, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pushnitski:1997:SLO

[1069] A. B. Pushnitski. The spectrum of Liouville operators and multiparticle Hamiltonians associated to one-particle Hamiltonians with singular continuous spectrum. Journal of Mathematical Physics, 38(5):2266–2273, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Slater:1997:IGE

[1070] Paul B. Slater. Information gains expected from separate and joint measurements of N identical spin-1/2 systems: noninformative Bayesian analyses. Journal of Mathematical Physics, 38(5):2274–2280, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spohn:1997:ACR

[1071] Herbert Spohn. Asymptotic completeness for Rayleigh scattering. Journal of Mathematical Physics, 38(5):2281–2296, May 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cardoso:1997:WEK

[1072] J. G. Cardoso. Wave equations and kinematical integrals for complexified Maxwell–Klein–Gordon fields. Journal of Mathematical Physics, 38(5): 2297–2307, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goldberg:1997:UHF

[1073] J. Goldberg, Thomas Berger, and Roger Dashen. Uniform high frequency approximation to scattering from infinite strip. Journal of Mathematical Physics, 38(5):2308–2331, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ibanez:1997:DGP

[1074] Raúl Ibáñez, Manuel de León, Juan C. Marrero, and David Martín de Diego. Dynamics of generalized Poisson and Nambu–Poisson brackets. Journal of Mathematical Physics, 38(5):2332–2344, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:1997:ISS

[1075] E. G. Kalnins, S. Benenti, and Willard Miller, Jr. Integrability, Stäckel spaces, and rational potentials. Journal of Mathematical Physics, 38 (5):2345–2365, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Magnoli:1997:SIP

[1076] N. Magnoli and G. A. Viano. The source identification problem in electromagnetic theory. Journal of Mathematical Physics, 38(5):2366–2388, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thacker:1997:NFC

[1077] W. D. Thacker. New formulation of the classical path integral with reparametrization invariance. Journal of Mathematical Physics, 38(5): 2389–2416, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Uehara:1997:MSR

[1078] Masahiro Uehara, Ken’ichiro Yashiro, and Sumio Ohkawa. A method for solving Riemann–Hilbert boundary value problems in nonreciprocal wave propagation. Journal of Mathematical Physics, 38(5):2417–2434, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Woodcock:1997:MTC

[1079] Harry W. Woodcock. Multiple-time constants of the motion for manifestly Poincaré-invariant Newtonian-like equations of motion for point particles. Journal of Mathematical Physics, 38(5):2435–2452, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Uchaikin:1997:LFA

[1080] V. Uchaikin and G. Gusarov. Lévy flight applied to random media problems. Journal of Mathematical Physics, 38(5):2453–2464, May 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duan:1997:CKS

[1081] Jinqiao Duan, Charles Bu, Hongjun Gao, and Mario Taboada. On a coupled Kuramoto–Sivashinsky and Ginzburg–Landau-type model for the Marangoni convection. Journal of Mathematical Physics, 38(5):2465– 2474, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ginibre:1997:LEC

[1082] J. Ginibre and G. Velo. Localized estimates and Cauchy problem for the logarithmic complex Ginzburg–Landau equation. Journal of Mathematical Physics, 38(5):2475–2482, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kowalski:1997:NDS

[1083] K. Kowalski. Nonlinear dynamical systems and classical orthogonal polynomials. Journal of Mathematical Physics, 38(5):2483–2505, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palev:1997:MBW

[1084] T. D. Palev and N. I. Stoilova. Many-body Wigner quantum systems. Journal of Mathematical Physics, 38(5):2506–2523, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1997:ADS

[1085] Bixiang Wang and Boling Guo. Attractors for the Davey–Stewartson systems on R^{2}. Journal of Mathematical Physics, 38(5):2524–2534, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:1997:FBS

[1086] Ruguang Zhou. The finite-band solution of the Jaulent–Miodek equation. Journal of Mathematical Physics, 38(5):2535–2546, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ahmad:1997:HMS

[1087] Daud Ahmad and M. Ziad. Homothetic motions of spherically symmetric space–times. Journal of Mathematical Physics, 38(5):2547–2552, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bezerra:1997:GED

[1088] Valdir B. Bezerra. Global effects due to a chiral cone. Journal of Mathematical Physics, 38(5):2553–2564, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chimento:1997:FID

[1089] Luis P. Chimento. Form invariance of differential equations in general relativity. Journal of Mathematical Physics, 38(5):2565–2576, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hartle:1997:SRE

[1090] James B. Hartle and Zoltan Perjés. Solutions of the Regge equations on some triangulations of CP^{2}. Journal of Mathematical Physics, 38 (5):2577–2586, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kopeikin:1997:PLS

[1091] S. M. Kopeikin. Propagation of light in the stationary field of multipole gravitational lens. Journal of Mathematical Physics, 38(5):2587–2601, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasinski:1997:TIP

[1092] Andrzej Krasiński, Hernando Quevedo, and Roberto A. Sussman. On the thermodynamical interpretation of perfect fluid solutions of the Einstein equations with no symmetry. Journal of Mathematical Physics, 38(5): 2602–2610, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nilsson:1997:SBT

[1093] Ulf S. Nilsson and Claes Uggla. Stationary Bianchi type II perfect fluid models. Journal of Mathematical Physics, 38(5):2611–2615, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:1997:GSA

[1094] Dahl Park and Jae Kwan Kim. General static axially symmetric solutions of (2 + 1)-dimensional Einstein–Maxwell-dilaton theory. Journal of Mathematical Physics, 38(5):2616–2625, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruce:1997:ARS

[1095] S. Bruce and P. Salgado. Affine realizations of sphere algebras. Journal of Mathematical Physics, 38(5):2626–2630, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrev:1997:IRR

[1096] V. K. Dobrev and P. Truini. Irregular U_{q}(sl(3)) representations at roots of unity via Gel′fand-(Weyl)-Zetlin basis. Journal of Mathematical Physics, 38(5):2631–2651, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Drabant:1997:DHA

[1097] Bernhard Drabant. Differential Hopf algebra structure of the quantum standard complex. Journal of Mathematical Physics, 38(5):2652–2659, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kim:1997:ALL

[1098] Jeong-Hoon Kim. An asymptotic limit law with a singularly perturbed drift and a random noise. Journal of Mathematical Physics, 38(5):2660– 2675, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Minnaert:1997:AFR

[1099] Pierre Minnaert and Marek Mozrzymas. Analytical formulas for Racah coefficients and 6-j symbols of the quantum superalgebra U_{q}(osp(1|2)).

Journal of Mathematical Physics, 38(5):2676–2693, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Norris:1997:SNB

[1100] L. K. Norris. Schouten–Nijenhuis brackets. Journal of Mathematical Physics, 38(5):2694–2709, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rangarajan:1997:RCH

[1101] Govindan Rangarajan. Representations of Sp(6, R) and SU(3) carried by homogeneous polynomials. Journal of Mathematical Physics, 38(5):2710– 2719, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strasburger:1997:WCO

[1102] Aleksander Strasburger and Arkadiusz Orłowski. Weighted class operators: a new approach to tensor operators for compact groups. Journal of Mathematical Physics, 38(5):2720–2727, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanderJeugt:1997:CCL

[1103] J. Van der Jeugt. Coupling coefficients for Lie algebra representations and addition formulas for special functions. Journal of Mathematical Physics, 38(5):2728–2740, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zapatrin:1997:PRD

[1104] Roman R. Zapatrin. Polyhedral representations of discrete differential manifolds. Journal of Mathematical Physics, 38(5):2741–2750, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gieseler:1997:CDT

[1105] U. D. J. Gieseler and J. G. Kirk. Comment on: “Diffusion through a slab”, [J. Math. Phys. 36 (1995), no. 12, 6758–6773; MR 96j:82040] by G. D. Mahan. Journal of Mathematical Physics, 38(5):2751–2754, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [465].

Wurm:1997:CPI

[1106] Alexander Wurm and John LaChapelle. Comment on: “Path integral solution of the Schrödinger equation in curvilinear coordinates: a straightforward procedure”, [J. Math. Phys. 37 (1996), no. 9, 4310–4319; MR 97i:81064] by LaChapelle. Journal of Mathematical Physics, 38(5):2755– 2757, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [784].

Anonymous:1997:AOAa

[1107] Anonymous. Announcement: Online access to Journal of Mathematical Physics. Journal of Mathematical Physics, 38(5):i, May 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arians:1997:GAI

[1108] Silke Arians. Geometric approach to inverse scattering for the Schrödinger equation with magnetic and electric potentials. Journal of Mathematical Physics, 38(6):2761–2773, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brevik:1997:CEP

[1109] I. Brevik and R. Sollie. On the Casimir energy for a 2N-piece relativistic string. Journal of Mathematical Physics, 38(6):2774–2785, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Constantinescu:1997:GKD

[1110] Radu Constantinescu. The generalized Koszul differential in the BRST quantization. Journal of Mathematical Physics, 38(6):2786–2794, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkel:1997:QES

[1111] Federico Finkel, Artemio González-López, and Miguel A. Rodríguez. Quasi-exactly solvable spin 1/2 Schrödinger operators. Journal of Mathematical Physics, 38(6):2795–2811, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujii:1997:MCS

[1112] Kazuyuki Fujii and Kunio Funahashi. Multiperiodic coherent states and the Wentzel–Kramers–Brillouin exactness. II. Noncompact case and classical theories revisited. Journal of Mathematical Physics, 38(6):2812– 2831, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:1997:STF

[1113] M. N. Hounkonnou, M. Hounkpe, and J. Shabani. Scattering theory for finitely many sphere interactions supported by concentric spheres.

Journal of Mathematical Physics, 38(6):2832–2850, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jorjadze:1997:CQS

[1114] G. Jorjadze. Constrained quantization on symplectic manifolds and quantum distribution functions. Journal of Mathematical Physics, 38 (6):2851–2879, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krutov:1997:FST

[1115] A. F. Krutov, D. I. Muravyev, and V. E. Troitsky. A factorization of a special type S matrix into Jost matrices. Journal of Mathematical Physics, 38(6):2880–2887, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liguori:1997:FRE

[1116] A. Liguori, M. Mintchev, and M. Rossi. Fock representations of exchange algebras with involution. Journal of Mathematical Physics, 38(6):2888– 2898, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McRae:1997:PTC

[1117] S. M. McRae and E. R. Vrscay. Perturbation theory and the classical limit of quantum mechanics. Journal of Mathematical Physics, 38(6): 2899–2921, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moretti:1997:CQP

[1118] Valter Moretti. Canonical quantization of photons in a Rindler wedge. Journal of Mathematical Physics, 38(6):2922–2953, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:1997:LAA

[1119] Yűji Ohta. Langevin approach for abelian topological gauge theory. Journal of Mathematical Physics, 38(6):2954–2962, June 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarel:1997:CST

[1120] B. Sarel and L. P. Horwitz. A chiral spin theory in the framework of an invariant evolution parameter formalism. Journal of Mathematical Physics, 38(6):2963–2995, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1997:SDE

[1121] G. F. Torres del Castillo and L. C. Cortés-Cuautli. Solution of the Dirac equation in the field of a magnetic monopole. Journal of Mathematical Physics, 38(6):2996–3006, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leznov:1997:IPS

[1122] A. N. Leznov, B. Piette, and W. J. Zakrzewski. On the integrability of pure Skyrme models in two dimensions. Journal of Mathematical Physics, 38(6):3007–3011, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mackrodt:1997:SI

[1123] Christof Mackrodt and Helmut Reeh. Summational invariants. Journal of Mathematical Physics, 38(6):3012–3019, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beltrametti:1997:EAS

[1124] E. G. Beltrametti and S. Bugajski. Effect algebras and statistical physical theories. Journal of Mathematical Physics, 38(6):3020–3030, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Constantin:1997:SSN

[1125] Peter Constantin and Jiahong Wu. Statistical solutions of the Navier– Stokes equations on the phase space of vorticity and the inviscid limits. Journal of Mathematical Physics, 38(6):3031–3045, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsutsumi:1997:MEG

[1126] Masayoshi Tsutsumi, Hironori Kasai, and Takeshi Ōishi. The Meissner effect and the Ginzburg–Landau equations in the presence of an applied magnetic field. Journal of Mathematical Physics, 38(6):3046–3054, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:1997:CAS

[1127] Manuel de León and David Martín de Diego. A constraint algorithm for singular Lagrangians subjected to nonholonomic constraints. Journal of Mathematical Physics, 38(6):3055–3062, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dragovic:1997:IPP

[1128] V. Dragović and B. Jovanović. On integrable potential perturbations of the billiard system within an ellipsoid. Journal of Mathematical Physics, 38(6):3063–3068, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geng:1997:NFD

[1129] Xianguo Geng and Yongtang Wu. New finite-dimensional integrable systems and explicit solutions of Hirota–Satsuma coupled Korteweg– de Vries equation. Journal of Mathematical Physics, 38(6):3069–3080, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamm:1997:UDS

[1130] Andreas Hamm. Uncertain dynamical systems defined by pseudomeasures. Journal of Mathematical Physics, 38(6):3081–3109, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moroz:1997:KPE

[1131] Irene M. Moroz. The Kadomtsev–Petviashvili equation under rapid forcing. Journal of Mathematical Physics, 38(6):3110–3122, June 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruan:1997:HDD

[1132] Hang yu Ruan and Sen yue Lou. Higher-dimensional dromion structures: Jimbo–Miwa–Kadomtsev–Petviashvili system. Journal of Mathematical Physics, 38(6):3123–3136, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1997:DGA

[1133] Guanxiang Wang and Shu Zhu. On the dimension of the global attractor for the damped sine-Gordon equation. Journal of Mathematical Physics, 38(6):3137–3141, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergqvist:1997:LPK

[1134] Göran Bergqvist. A Lanczos potential in Kerr geometry. Journal of Mathematical Physics, 38(6):3142–3154, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:1997:EDR

[1135] Ya-Jun Gao, Zai-Zhe Zhong, and Yuan-Xing Gui. Extended double Riemann–Hilbert problem and multiple representations of the infinitedimensional symmetric group for the stationary axisymmetric vacuum field equations. Journal of Mathematical Physics, 38(6):3155–3165, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kleidis:1997:HDM

[1136] K. Kleidis, A. Kuiroukidis, D. B. Papadopoulos, and H. Varvoglis. Higher-dimensional models in gravitational theories of quartic Lagrangians. Journal of Mathematical Physics, 38(6):3166–3188, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kleidis:1997:CSK

[1137] K. Kleidis and D. B. Papadopoulos. Cosmological solutions in Kaluza– Klein theories of quadratic Lagrangians. Journal of Mathematical Physics, 38(6):3189–3208, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bal:1997:UTC

[1138] Subrata Bal, K. V. Shajesh, and Debabrata Basu. A unified treatment of the characters of SU(2) and SU(1, 1). Journal of Mathematical Physics, 38(6):3209–3229, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chung:1997:RWR

[1139] W.-S. Chung and Takashi Suzuki. On representations of U_{q}osp(1|2) when q is a root of unity. Journal of Mathematical Physics, 38(6):3230–3237, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dubin:1997:IRU

[1140] D. A. Dubin, M. A. Hennings, and A. I. Solomon. Integrable representations of the ultra-commutation relations. Journal of Mathematical Physics, 38(6):3238–3262, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Georgelin:1997:NRG

[1141] Y. Georgelin, J. Madore, T. Masson, and J. Mourad. On the noncommutative Riemannian geometry of GL_{q}(n). Journal of Mathematical Physics, 38(6):3263–3277, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grothaus:1997:FIT

[1142] Martin Grothaus, Dinkar C. Khandekar, José L. da Silva, and Ludwig Streit. The Feynman integral for time-dependent anharmonic oscillators. Journal of Mathematical Physics, 38(6):3278–3299, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khorunzhy:1997:EDL

[1143] A. Khorunzhy and G. J. Rodgers. Eigenvalue distribution of large dilute random matrices. Journal of Mathematical Physics, 38(6):3300–3320, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:1997:DOL

[1144] Yong Moon Park and Hyun Jae Yoo. Dirichlet operators on loop spaces: essential self-adjointness and log-Sobolev inequality. Journal of Mathematical Physics, 38(6):3321–3346, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Staahl:1997:LVT

[1145] Fredrik Ståhl. A local variational theory for the Schmidt metric. Journal of Mathematical Physics, 38(6):3347–3357, June 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wulkenhaar:1997:NGG

[1146] Raimar Wulkenhaar. Noncommutative geometry with graded differential Lie algebras. Journal of Mathematical Physics, 38(6):3358–3390, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Overfelt:1997:CEF

[1147] P. L. Overfelt. Comment on: “On the evanescent fields and the causality of the focus wave modes”, [J. Math. Phys. 36 (1995), no. 10, 5565–5587; MR 96g:78017] by A. M. Shaarawi, R. W. Ziolkowski and I. M. Besieris. Journal of Mathematical Physics, 38(6):3391–3395, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [398].

Anonymous:1997:AOAb

[1148] Anonymous. Announcement: Online access to Journal of Mathematical Physics. Journal of Mathematical Physics, 38(6):i, June 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:1997:NGT

[1149] Stephen C. Anco. Novel generalization of three-dimensional Yang–Mills theory. Journal of Mathematical Physics, 38(7):3399–3413, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bellinger:1997:SPI

[1150] James N. Bellinger. Symmetries of preon interactions modeled as a finite group. Journal of Mathematical Physics, 38(7):3414–3426, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bordner:1997:LCT

[1151] Andrew J. Bordner. Linear canonical transformations in quantum mechanics. Journal of Mathematical Physics, 38(7):3427–3434, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Casini:1997:APS

[1152] Roberto Casini. Algebraic proof of a sum rule occurring in Stark broadening of hydrogen lines. Journal of Mathematical Physics, 38(7):3435– 3445, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1335].

Fan:1997:PES

[1153] Heng Fan, Bo yu Hou, and Kang jie Shi. Properties of eigenstates of the six-vertex model with twisted and open boundary conditions. Journal of Mathematical Physics, 38(7):3446–3456, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frick:1997:NHO

[1154] Rudolf A. Frick. A new Hamiltonian operator for a massive relativistic particle with spin one in a generalized Heisenberg/Schrödinger picture. Journal of Mathematical Physics, 38(7):3457–3465, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1521].

Lavrov:1997:EAC

[1155] P. M. Lavrov, S. D. Odintsov, and A. A. Reshetnyak. Effective action of composite fields for general gauge theories in Batalin, Lavrov, and Tyutin covariant formalism. Journal of Mathematical Physics, 38(7): 3466–3477, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Luo:1997:CSA

[1156] Shunlong Luo. SU(1, 1) coherent states and associated Wick symbol calculus. Journal of Mathematical Physics, 38(7):3478–3488, July 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:1997:QCT

[1157] Ali Mostafazadeh. Quantum canonical transformations and exact solution of the Schrödinger equation. Journal of Mathematical Physics, 38 (7):3489–3496, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sacks:1997:RSA

[1158] Paul E. Sacks. Recovery of singularities from amplitude information. Journal of Mathematical Physics, 38(7):3497–3507, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anco:1997:NSN

[1159] Stephen C. Anco and George Bluman. Nonlocal symmetries and nonlocal conservation laws of Maxwell’s equations. Journal of Mathematical Physics, 38(7):3508–3532, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Elipe:1997:PFA

[1160] A. Elipe and V. Lanchares. Phase flow of an axially symmetrical gyrostat with one constant rotor. Journal of Mathematical Physics, 38(7):3533– 3544, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tip:1997:NEM

[1161] A. Tip. Nonconducting electromagnetic media with rotational invariance: transition operators and Green’s functions. Journal of Mathematical Physics, 38(7):3545–3570, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arlotti:1997:ENS

[1162] Luisa Arlotti and Miroslaw Lachowicz. Euler and Navier–Stokes limits of the Uehling–Uhlenbeck quantum kinetic equations. Journal of Mathematical Physics, 38(7):3571–3588, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1997:MMI

[1163] M. Chen. On the moment methods and irreversible thermodynamics. Journal of Mathematical Physics, 38(7):3589–3602, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hietarinta:1997:ETE

[1164] Jarmo Hietarinta and Frank Nijhoff. The eight tetrahedron equations. Journal of Mathematical Physics, 38(7):3603–3615, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karasu:1997:PCC

[1165] Ayşe Karasu. Painlevé classification of coupled Korteweg–de Vries systems. Journal of Mathematical Physics, 38(7):3616–3622, July 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bartnik:1997:SSS

[1166] Robert Bartnik. The structure of spherically symmetric su(n) Yang–Mills fields. Journal of Mathematical Physics, 38(7):3623–3638, July 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bokhari:1997:CSS

[1167] Ashfaque H. Bokhari, Asghar Qadir, M. Shahan Ahmed, and Mohammad Asghar. Classification of spherically symmetric static space–times by their curvature collineations. Journal of Mathematical Physics, 38(7): 3639–3649, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Malec:1997:SGN

[1168] Edward Malec. Self-gravitating nonlinear scalar fields. Journal of Mathematical Physics, 38(7):3650–3668, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Masuda:1997:IEF

[1169] Takahiro Masuda and Hisao Suzuki. Integral equations of fields on a rotating black hole. Journal of Mathematical Physics, 38(7):3669–3678, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bates:1997:TGF

[1170] Jason W. Bates. On toroidal Green’s functions. Journal of Mathematical Physics, 38(7):3679–3691, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Becker:1997:NIO

[1171] Peter A. Becker. Normalization integrals of orthogonal Heun functions. Journal of Mathematical Physics, 38(7):3692–3699, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:1997:CLL

[1172] Carl M. Bender and Alexander Tovbis. Continuum limit of lattice approximation schemes. Journal of Mathematical Physics, 38(7):3700–3717, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruhn:1997:FCC

[1173] B. Bruhn. Factorization of complex canonical transformations. Journal of Mathematical Physics, 38(7):3718–3734, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deAzcarraga:1997:GSN

[1174] J. A. de Azcárraga, J. M. Izquierdo, A. M. Perelomov, and J. C. Pérez-Bueno. The Z_{2}-graded Schouten–Nijenhuis bracket and generalized super-Poisson structures. Journal of Mathematical Physics, 38(7):3735– 3749, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dobrev:1997:PRG

[1175] V. K. Dobrev and P. Truini. Polynomial realization of the U_{q}(sl(3)) Gel′fand-(Weyl)-Zetlin basis. Journal of Mathematical Physics, 38(7): 3750–3767, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:1997:LAE

[1176] Javier Fernandez and Marcela Zuccalli. On Lie algebra extensions in a symplectic framework. Journal of Mathematical Physics, 38(7):3768– 3771, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mann:1997:PKV

[1177] E. Mann. The perturbed Korteweg–de Vries equation considered anew. Journal of Mathematical Physics, 38(7):3772–3785, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Negro:1997:NCGa

[1178] J. Negro, M. A. del Olmo, and A. Rodríguez-Marco. Nonrelativistic conformal groups. Journal of Mathematical Physics, 38(7):3786–3809, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Negro:1997:NCGb

[1179] J. Negro, M. A. del Olmo, and A. Rodríguez-Marco. Nonrelativistic conformal groups. II. Further developments and physical applications. Journal of Mathematical Physics, 38(7):3810–3831, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pelc:1997:NFS

[1180] Oskar Pelc. A new family of solvable self-dual Lie algebras. Journal of Mathematical Physics, 38(7):3832–3840, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schwartz:1997:NOI

[1181] Charles Schwartz. Nonlinear operators. II. Journal of Mathematical Physics, 38(7):3841–3862, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:1997:BBW

[1182] R. B. Zhang. Bott–Borel–Weil construction for quantum supergroup U_{q}(gl(m|n)). Journal of Mathematical Physics, 38(7):3863–3884, July 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1997:AAMa

[1183] Anonymous. Announcement: Accepted manuscript status inquiry system. Journal of Mathematical Physics, 38(7):i, July 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ahmedov:1997:OGF

[1184] H. Ahmedov and I. H. Duru. A q-oscillator Green’s function. Journal of Mathematical Physics, 38(8):3889–3894, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bosanac:1997:RCT

[1185] S. Danko Bosanac. Relativistic classical theory of a free particle. Journal of Mathematical Physics, 38(8):3895–3907, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Darewych:1997:NFR

[1186] J. W. Darewych, D. V. Shapoval, I. V. Simenog, and A. G. Sitenko. Nonperturbative formulation of relativistic two-particle states in the scalar Yukawa model. Journal of Mathematical Physics, 38(8):3908–3924, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Derezinski:1997:LRS

[1187] J. Dereziński and C. Gérard. Long-range scattering in the position representation. Journal of Mathematical Physics, 38(8):3925–3942, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doebner:1997:VBC

[1188] H.-D. Doebner and H.-J. Mann. Vector bundles over configuration spaces of nonidentical particles: topological potentials and internal degrees of freedom. Journal of Mathematical Physics, 38(8):3943–3952, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fatibene:1997:NFC

[1189] L. Fatibene, M. Ferraris, and M. Francaviglia. Nöther formalism for conserved quantities in classical gauge field theories. II. The arbitrary bosonic matter case. Journal of Mathematical Physics, 38(8):3953–3967, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fu:1997:NBM

[1190] Hong-Chen Fu and Ryu Sasaki. Negative binomial and multinomial states: probability distributions and coherent states. Journal of Mathematical Physics, 38(8):3968–3987, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Han:1997:GMB

[1191] Kyoung Ho Han and H. J. Shin. Grassmann manifold bosonization of QCD in two dimensions. Journal of Mathematical Physics, 38(8):3988– 3995, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

LeYaouanc:1997:HNC

[1192] A. Le Yaouanc, L. Oliver, and J.-C. Raynal. The Hamiltonian (p^{2} + m^{2})^{1/2} − α/r near the critical value α_{c} = 2/π. Journal of Mathematical Physics, 38(8):3997–4012, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Markvardsen:1997:SME

[1193] Anders J. Markvardsen and Neil F. Johnson. A solvable model for excitonic complexes in one dimension. Journal of Mathematical Physics, 38 (8):4013–4022, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mohamed:1997:ADS

[1194] Abderemane Mohamed. Asymptotic of the density of states for the Schrödinger operator with periodic electromagnetic potential. Journal of Mathematical Physics, 38(8):4023–4051, August 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nakamura:1997:PSM

[1195] Toru Nakamura. Path space measures for Dirac and Schrödinger equations: nonstandard analytical approach. Journal of Mathematical Physics, 38(8):4052–4072, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scotti:1997:NQI

[1196] Antonio Scotti and Alexander Ushveridze. Nonlinear quantization of integrable classical systems. Journal of Mathematical Physics, 38(8): 4073–4085, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lyakhovich:1997:SPS

[1197] S. L. Lyakhovich, A. A. Sharapov, and K. M. Shekhter. Spinning particle in six dimensions. Journal of Mathematical Physics, 38(8):4086–4103, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

dosSantos:1997:GST

[1198] Roberto J. V. dos Santos. Generalization of Shannon’s theorem for Tsallis entropy. Journal of Mathematical Physics, 38(8):4104–4107, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belich:1997:TSD

[1199] H. Belich and R. Paunov. A_{n}^{(1)} Toda solitons and the dressing symmetry. Journal of Mathematical Physics, 38(8):4108–4127, August 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chau:1997:SCK

[1200] Ling-Lie Chau, Jiin-Chang Shaw, and Ming-Hsien Tu. Solving the constrained KP hierarchy by gauge transformations. Journal of Mathematical Physics, 38(8):4128–4137, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doktorov:1997:STS

[1201] E. V. Doktorov. Spectral transform and solitons for the three-wave coupling model with nontrivial boundary conditions. Journal of Mathematical Physics, 38(8):4138–4150, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1448].

Maccari:1997:UIN

[1202] Attilio Maccari. Universal and integrable nonlinear evolution systems of equations in 2 + 1 dimensions. Journal of Mathematical Physics, 38(8): 4151–4164, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:1997:SSQ

[1203] Manuel F. Rañada. Superintegrable n = 2 systems, quadratic constants of motion, and potentials of Drach. Journal of Mathematical Physics, 38 (8):4165–4178, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Suris:1997:NQP

[1204] Yuri B. Suris. Nonlocal quadratic Poisson algebras, monodromy map, and Bogoyavlensky lattices. Journal of Mathematical Physics, 38(8): 4179–4201, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Das:1997:SSC

[1205] A. Das, N. Tariq, D. Aruliah, and T. Biech. Spherically symmetric collapse of an anisotropic fluid body into an exotic black hole. Journal of Mathematical Physics, 38(8):4202–4227, August 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paiva:1997:LST

[1206] F. M. Paiva, M. J. Rebouças, and A. F. F. Teixeira. Limits of space– times in five dimensions and their relation to the Segrè types. Journal of Mathematical Physics, 38(8):4228–4236, August 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Racz:1997:EES

[1207] István Rácz. On Einstein’s equations for space–times admitting a nonnull Killing field. Journal of Mathematical Physics, 38(8):4237–4254, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cieslinski:1997:GFI

[1208] Jan Cieśliński. A generalized formula for integrable classes of surfaces in Lie algebras. Journal of Mathematical Physics, 38(8):4255–4272, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dahlqvist:1997:EPS

[1209] Per Dahlqvist. On the effect of pruning on the singularity structure of zeta functions. Journal of Mathematical Physics, 38(8):4273–4282, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gratus:1997:NBS

[1210] J. Gratus. A natural basis of states for the noncommutative sphere and its Moyal bracket. Journal of Mathematical Physics, 38(8):4283–4300, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Horuzhy:1997:RBA

[1211] S. S. Horuzhy and A. V. Voronin. Representations of the BRST algebra and unsolvable algebraic problems. Journal of Mathematical Physics, 38 (8):4301–4322, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jakobsen:1997:TSQ

[1212] Hans Plesner Jakobsen. Tensoring with small quantized representations. Journal of Mathematical Physics, 38(8):4323–4335, August 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kachkachi:1997:PPC

[1213] M. Kachkachi and S. Kouadik. Proof of Polyakov conjecture on supercomplex plane. Journal of Mathematical Physics, 38(8):4336–4342, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Keller:1997:BIB

[1214] Joseph B. Keller and Hans F. Weinberger. Boundary and initialboundary value problems for separable backward-forward parabolic problems. Journal of Mathematical Physics, 38(8):4343–4353, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McInnes:1997:DFS

[1215] Brett McInnes. Disconnected forms of the standard group. Journal of Mathematical Physics, 38(8):4354–4362, August 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1997:AAC

[1216] D. J. Rowe and J. Repka. An algebraic algorithm for calculating Clebsch– Gordan coefficients; application to SU(2) and SU(3). Journal of Mathematical Physics, 38(8):4363–4388, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simon:1997:SDS

[1217] David S. Simon. Semiclassical determinants of Schrödinger operators on bundles with curvature of low rank. Journal of Mathematical Physics, 38 (8):4389–4397, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saenz:1997:ELT

[1218] A. W. Sáenz. Erratum: “Long-time approximation to the evolution of resonant and nonresonant anharmonic oscillators in quantum mechanics”, [J. Math. Phys. 37 (1996), no. 5, 2182–2205; 1 385 824]. Journal of Mathematical Physics, 38(8):4398, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [641].

Anonymous:1997:AAMb

[1219] Anonymous. Announcement: Accepted manuscript status inquiry system. Journal of Mathematical Physics, 38(8):i, August 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arthur:1997:TCG

[1220] K. Arthur, G. M. O’Brien, and D. H. Tchrakian. Towards a Coulomb gas of instantons of the SO(4) × U(1) Higgs model on R_{4}. Journal of Mathematical Physics, 38(9):4403–4421, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujii:1997:EBG

[1221] Kazuyuki Fujii and Kunio Funahashi. Extension of the Barut–Girardello coherent state and path integral. Journal of Mathematical Physics, 38 (9):4422–4434, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gebert:1997:ASV

[1222] R. W. Gebert and H. Nicolai. An affine string vertex operator construction at an arbitrary level. Journal of Mathematical Physics, 38(9): 4435–4450, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Goswami:1997:FDL

[1223] G. Goswami and P. Bandyopadhyay. Fermion doubling on a lattice and topological aspects of chiral anomaly. Journal of Mathematical Physics, 38(9):4451–4461, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hammou:1997:BOQ

[1224] A. B. Hammou and M. Lagraa. The BRST operator of quantum symmetries: the quantum analogs of Donaldson invariants. Journal of Mathematical Physics, 38(9):4462–4473, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Podles:1997:DOG

[1225] P. Podleś. The Dirac operator and gamma matrices for quantum Minkowski spaces. Journal of Mathematical Physics, 38(9):4474–4491, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Samsonov:1997:SSS

[1226] Boris F. Samsonov. Supersymmetry and supercoherent states of a nonrelativistic free particle. Journal of Mathematical Physics, 38(9):4492–4503, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benn:1997:DPM

[1227] I. M. Benn, Philip Charlton, and Jonathan Kress. Debye potentials for Maxwell and Dirac fields from a generalization of the Killing–Yano equation. Journal of Mathematical Physics, 38(9):4504–4527, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clark:1997:CFR

[1228] Karen E. Clark. A continued fraction representation for the effective conductivity of a two-dimensional polycrystal. Journal of Mathematical Physics, 38(9):4528–4541, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fecko:1997:DRO

[1229] Marián Fecko. On 3 + 1 decompositions with respect to an observer field via differential forms. Journal of Mathematical Physics, 38(9):4542–4560, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vivarelli:1997:FIK

[1230] Maria Dina Vivarelli. On a first integral of the Kepler problem. Journal of Mathematical Physics, 38(9):4561–4569, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Streater:1997:DBP

[1231] R. F. Streater. Dynamics of Brownian particles in a potential. Journal of Mathematical Physics, 38(9):4570–4575, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Auberson:1997:CHN

[1232] G. Auberson. A class of homogeneous nonlinear evolution equations with stable, localized solutions in any dimension. Journal of Mathematical Physics, 38(9):4576–4593, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Birol:1997:SAM

[1233] İnanç Birol and Avadis Hacinliyan. Signs and approximate magnitudes of Lyapunov exponents in continuous time dynamical systems. Journal of Mathematical Physics, 38(9):4594–4605, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Casati:1997:NFK

[1234] Paolo Casati, Gregorio Falqui, Franco Magri, and Marco Pedroni. A note on fractional KdV hierarchies. Journal of Mathematical Physics, 38(9):4606–4628, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGroot:1997:ESB

[1235] E. H. de Groot and C. Dalitz. Exact solution of a boundary value problem in semiconductor kinetic theory. Journal of Mathematical Physics, 38(9): 4629–4643, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Orlov:1997:APO

[1236] A. Yu. Orlov and P. Winternitz. Algebra of pseudodifferential operators and symmetries of equations in the Kadomtsev–Petviashvili hierarchy. Journal of Mathematical Physics, 38(9):4644–4674, September 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Porsezian:1997:PAN

[1237] K. Porsezian. Painlevé analysis of new higher-dimensional soliton equation. Journal of Mathematical Physics, 38(9):4675–4679, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ragnisco:1997:IDS

[1238] Orlando Ragnisco and Yuri B. Suris. Integrable discretizations of the spin Ruijsenaars–Schneider models. Journal of Mathematical Physics, 38(9):4680–4691, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Steudel:1997:VLD

[1239] H. Steudel, R. Meinel, and G. Neugebauer. Vandermonde-like determinants and N-fold Darboux/Bäcklund transformations. Journal of Mathematical Physics, 38(9):4692–4695, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Darian:1997:CEY

[1240] B. K. Darian and H. P. Künzle. Cosmological Einstein–Yang–Mills equations. Journal of Mathematical Physics, 38(9):4696–4713, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iriondo:1997:NSB

[1241] Mirta Iriondo, Carlos N. Kozameh, and Alejandra Rojas. Null surfaces and the Bach equations. Journal of Mathematical Physics, 38(9):4714– 4729, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marolf:1997:SDS

[1242] Donald Marolf, José Mourão, and Thomas Thiemann. The status of diffeomorphism superselection in Euclidean 2 + 1 gravity. Journal of Mathematical Physics, 38(9):4730–4740, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mars:1997:STA

[1243] Marc Mars and Roustam M. Zalaletdinov. Space–time averages in macroscopic gravity and volume-preserving coordinates. Journal of Mathematical Physics, 38(9):4741–4757, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:1997:RQP

[1244] S. L. Adler and G. G. Emch. A rejoinder on quaternionic projective representations. Journal of Mathematical Physics, 38(9):4758–4762, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertini:1997:RFM

[1245] L. Bertini and N. Cancrini. Reduction formula for moments of stochastic integrals. Journal of Mathematical Physics, 38(9):4763–4770, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chodos:1997:ZFR

[1246] Alan Chodos and András Kaiser. Zeta function regularization in de Sitter space: the Minkowski limit. Journal of Mathematical Physics, 38 (9):4771–4782, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gould:1997:CIC

[1247] M. D. Gould and N. I. Stoilova. Casimir invariants and characteristic identities for gl(∞). Journal of Mathematical Physics, 38(9):4783–4793, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gudkov:1997:FET

[1248] V. V. Gudkov. A family of exact travelling wave solutions to nonlinear evolution and wave equations. Journal of Mathematical Physics, 38(9): 4794–4803, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bilge:1997:GSD

[1249] Ayşe Hümeyra Bilge, Tekin Dereli, and ¸ Sahin Koçak. The geometry of self-dual two-forms. Journal of Mathematical Physics, 38(9):4804–4814, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karabulut:1997:DGP

[1250] Hasan Karabulut and Edwin L. Sibert, III. Distributed Gaussian polynomials and associated Gaussian quadratures. Journal of Mathematical Physics, 38(9):4815–4831, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konya:1997:GMJ

[1251] B. Kónya, G. Lévai, and Z. Papp. Green’s matrix from Jacobimatrix Hamiltonian. Journal of Mathematical Physics, 38(9):4832–4844, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:1997:FFS

[1252] S. Majid. Fermionic q-Fock space and braided geometry. Journal of Mathematical Physics, 38(9):4845–4853, September 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prakash:1997:NBC

[1253] J. S. Prakash and H. S. Sharatchandra. A new boson calculus for SU(3). Journal of Mathematical Physics, 38(9):4854–4862, September 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rebollo-Neira:1997:FME

[1254] L. Rebollo-Neira, J. Fernandez-Rubio, and A. Plastino. Frames: a maximum entropy statistical estimate of the inverse problem. Journal of Mathematical Physics, 38(9):4863–4871, September 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seiler:1997:NUC

[1255] Erhard Seiler and Karim Yildirim. New universality classes in onedimensional O(N)-invariant spin-models with an n-parametric action. Journal of Mathematical Physics, 38(9):4872–4881, September 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

SassolideBianchi:1997:CFS

[1256] M. Sassoli de Bianchi. Comment on: “Factorization of scattering matrices due to partitioning of potentials in one-dimensional Schrödinger-type equations”, [J. Math. Phys. 37 (1996), no. 12, 5897–5915; MR 97j:34111] by T. Aktosun, M. Klaus and C. van der Mee. Journal of Mathematical Physics, 38(9):4882–4883, September 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [883].

Anonymous:1997:ASEa

[1257] Anonymous. Announcement: Submitting electronic graphics files. Journal of Mathematical Physics, 38(9):v–vi, September 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Allen:1997:QAO

[1258] T. Allen, C. Anastassiou, and W. H. Klink. The quartic anharmonic oscillator and its associated nonconstant magnetic field. Journal of Mathematical Physics, 38(10):4887–4899, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chadan:1997:SCE

[1259] K. Chadan and R. Kobayashi. A sufficient condition for the existence of bound states in a potential. Journal of Mathematical Physics, 38(10): 4900–4908, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2272].

Hall:1997:BSE

[1260] Richard L. Hall and Nasser Saad. Bounds on Schrödinger eigenvalues for polynomial potentials in N dimensions. Journal of Mathematical Physics, 38(10):4909–4913, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hamilton:1997:HND

[1261] J. J. Hamilton. Hypercomplex numbers and the description of spin states. Journal of Mathematical Physics, 38(10):4914–4928, October 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Iochum:1997:UCC

[1262] Bruno Iochum, Daniel Kastler, and Thomas Schücker. On the universal Chamseddine–Connes action. I. Details of the action computation. Journal of Mathematical Physics, 38(10):4929–4950, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jallouli:1997:GTR

[1263] H. Jallouli and H. Sazdjian. Gauge transformations in relativistic twoparticle constraint theory. Journal of Mathematical Physics, 38(10): 4951–4970, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kopf:1997:TFD

[1264] Tomáš Kopf, A. E. Santana, and F. C. Khanna. Thermal field dynamics and bialgebras. Journal of Mathematical Physics, 38(10):4971–4979, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naito:1997:GFP

[1265] Seichi Naito. Generating functionals of physical vertex operators in superstring. Journal of Mathematical Physics, 38(10):4980–5009, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Odzijewicz:1997:CSM

[1266] A. Odzijewicz and M. Świetochowski. Coherent states map for MIC-Kepler system. Journal of Mathematical Physics, 38(10):5010–5030, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sanchez-Ruiz:1997:LPH

[1267] Jorge Sánchez-Ruiz. Logarithmic potential of Hermite polynomials and information entropies of the harmonic oscillator eigenstates. Journal of Mathematical Physics, 38(10):5031–5043, October 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thomas:1997:CSM

[1268] Launey J. Thomas, III and Eyvind H. Wichmann. On the causal structure of Minkowski spacetime. Journal of Mathematical Physics, 38(10): 5044–5086, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Znojil:1997:OAP

[1269] Miloslav Znojil. r^{D} oscillators with arbitrary D > 0 and perturbation expansions with Sturmians. Journal of Mathematical Physics, 38(10): 5087–5097, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krupkova:1997:MSN

[1270] Olga Krupková. Mechanical systems with nonholonomic constraints. Journal of Mathematical Physics, 38(10):5098–5126, October 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brodsky:1997:CFE

[1271] Anatol M. Brodsky. Casimir free energy of a spherical cavity in a dielectric medium. Journal of Mathematical Physics, 38(10):5127–5142, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buffet:1997:MCP

[1272] E. Buffet and J. V. Pulé. A model of continuous polymers with random charges. Journal of Mathematical Physics, 38(10):5143–5152, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1997:SLT

[1273] M. Chen. On the second law of thermodynamics for a dissipative system. Journal of Mathematical Physics, 38(10):5153–5166, October 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guha:1997:AKS

[1274] Partha Guha. Adler–Kostant–Symes construction, bi-Hamiltonian manifolds, and KdV equations. Journal of Mathematical Physics, 38(10): 5167–5182, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiso:1997:HEE

[1275] Kazuhiro Kiso, Eiji Kaneda, and Ei ichiro Kawai. Hierarchies of evolution equations associated with certain symmetric spaces. Journal of Mathematical Physics, 38(10):5183–5189, October 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loris:1997:SCKb

[1276] Ignace Loris and Ralph Willox. On solutions of constrained Kadomtsev– Petviashvili equations: Grammians. Journal of Mathematical Physics, 38(10):5190–5197, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matsuno:1997:IVP

[1277] Y. Matsuno and D. J. Kaup. Initial value problem of the linearized Benjamin–Ono equation and its applications. Journal of Mathematical Physics, 38(10):5198–5224, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:1997:BBG

[1278] Cangtao Zhou, C. H. Lai, and M. Y. Yu. Bifurcation behavior of the generalized Lorenz equations at large rotation numbers. Journal of Mathematical Physics, 38(10):5225–5239, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bayin:1997:CEM

[1279] Selçuk Ş. Bayın and Mustafa Özcan. Casimir energy of the massless conformal scalar field on S-2 by the point-splitting method. Journal of Mathematical Physics, 38(10):5240–5255, October 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coley:1997:HSF

[1280] A. A. Coley, J. Ibáñez, and R. J. van den Hoogen. Homogeneous scalar field cosmologies with an exponential potential. Journal of Mathematical Physics, 38(10):5256–5271, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kreiss:1997:GEE

[1281] Heinz-Otto Kreiss, Gabriel B. Nagy, Omar E. Ortiz, and Oscar A. Reula. Global existence and exponential decay for hyperbolic dissipative relativistic fluid theories. Journal of Mathematical Physics, 38(10):5272– 5279, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marklund:1997:SRM

[1282] Mattias Marklund and Zoltán Perjés. Stationary rotating matter in general relativity. Journal of Mathematical Physics, 38(10):5280–5292, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendoza:1997:FMO

[1283] Ramón Mendoza, Pedro Gómez, and Fernando Moraes. Fluctuating metrics in one-dimensional manifolds. Journal of Mathematical Physics, 38(10):5293–5300, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Orlyansky:1997:SFS

[1284] Oleg Yu. Orlyansky. Singularity-free static fluid spheres in general relativity. Journal of Mathematical Physics, 38(10):5301–5304, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saha:1997:NSF

[1285] B. Saha and G. N. Shikin. Nonlinear spinor field in Bianchi type-I universe filled with perfect fluid: exact self-consistent solutions. Journal of Mathematical Physics, 38(10):5305–5318, October 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alfaro:1997:ORS

[1286] Jorge Alfaro, Ricardo Medina, and Luis F. Urrutia. Orthogonality relations and supercharacter formulas of U(m|n) representations. Journal of Mathematical Physics, 38(10):5319–5349, October 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1997:AEE

[1287] Goong Chen, Stephen A. Fulling, and Jianxin Zhou. Asymptotic equipartition of energy by nodal points of an eigenfunction. Journal of Mathematical Physics, 38(10):5350–5360, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dickenstein:1997:ISF

[1288] Alicia Dickenstein, Mirta Susana Iriondo, and Teresita Alejandra Rojas. Integrating singular functions on the sphere. Journal of Mathematical Physics, 38(10):5361–5370, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haring-Oldenburg:1997:TCC

[1289] Reinhard Häring-Oldenburg. Tensor categories of Coxeter type B and QFT on the half plane. Journal of Mathematical Physics, 38(10):5371– 5382, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jacobs:1997:TCC

[1290] Andrew D. Jacobs and J. F. Cornwell. Twisting 2-cocycles for the construction of new non-standard quantum groups. Journal of Mathematical Physics, 38(10):5383–5401, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Janeczko:1997:EIS

[1291] S. Janeczko. Equivalence of isotropic submanifolds and symmetry. Journal of Mathematical Physics, 38(10):5402–5415, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:1997:STD

[1292] E. G. Kalnins, W. Miller, Jr., and G. S. Pogosyan. Superintegrability on the two-dimensional hyperboloid. Journal of Mathematical Physics, 38(10):5416–5433, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Radko:1997:ASD

[1293] O. V. Radko and A. A. Vladimirov. On the algebraic structure of differential calculus on quantum groups. Journal of Mathematical Physics, 38(10):5434–5446, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:1997:EPY

[1294] Yűji Ohta. Erratum: “Prepotentials of N = 2SU(2) Yang–Mills theories coupled with massive matter multiplets”, [J. Math. Phys. 38 (1997), no. 2, 682–696; 1 434 203]. Journal of Mathematical Physics, 38(10):5447, October 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [980].

Anonymous:1997:ASEb

[1295] Anonymous. Announcement: Submitting electronic graphics files. Journal of Mathematical Physics, 38(10):ix–x, October 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashtekar:1997:QYM

[1296] Abhay Ashtekar, Jerzy Lewandowski, Donald Marolf, Jose Mourao, and Thomas Thiemann. SU(N) quantum Yang–Mills theory in two dimensions: A complete solution. Journal of Mathematical Physics, 38(11): 5453–5483, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bessis:1997:OPR

[1297] N. Bessis and G. Bessis. Open perturbation and the Riccati equation: Algebraic determination of the quartic anharmonic oscillator energies and eigenfunctions. Journal of Mathematical Physics, 38(11):5483–5492, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bracken:1997:RSP

[1298] Paul Bracken and Jiři Čížek. Reconstruction of secular polynomials for Hubbard model from energy perturbation series for weak and strong coupling cases. Journal of Mathematical Physics, 38(11):5493–5504, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cotaescu:1997:OAQ

[1299] Ion I. Cotăescu and Gheorghe Drăgănescu. The operator algebra of the quantum relativistic oscillator. Journal of Mathematical Physics, 38(11): 5505–5514, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dereli:1997:CWW

[1300] T. Dereli and A. Vercin. W_{∞}-covariance of the Weyl–Wigner– Groenewold–Moyal quantization. Journal of Mathematical Physics, 38 (11):5515–5530, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Drechsler:1997:GSQ

[1301] W. Drechsler. Geometro-stochastically quantized fields with internal spin variables. Journal of Mathematical Physics, 38(11):5531–5558, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figueroa-OFarrill:1997:SST

[1302] Jose M. Figueroa-O’Farrill. N = 2 structures in string theories. Journal of Mathematical Physics, 38(11):5559–5575, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Houghton:1997:MPS

[1303] Conor J. Houghton and Paul M. Sutcliffe. SU(N) monopoles and Platonic symmetry. Journal of Mathematical Physics, 38(11):5576–5589, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kanter:1997:CQP

[1304] Marek Kanter. Can a quantum process be oblivious to observation? Journal of Mathematical Physics, 38(11):5590–5604, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimek:1997:QKF

[1305] Stawomir Klimek and Andrzej Lesniewski. Quantized Kronecker flows and almost periodic quantum field theory. Journal of Mathematical Physics, 38(11):5605–5625, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pupyshev:1997:KBV

[1306] V. I. Pupyshev, A. V. Scherbinin, and N. F. Stepanov. The Kirkwood– Buckingham variational method and the boundary value problems for the molecular Schrödinger equation. Journal of Mathematical Physics, 38(11):5626–5633, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruan:1997:TBH

[1307] W. Y. Ruan and C. G. Bao. Transformation bracket for 2D hyperspherical harmonics and its applications to few-anyon problems. Journal of Mathematical Physics, 38(11):5634–5642, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudolph:1997:TOR

[1308] Oliver Rudolph and J. D. Maitland Wright. On tracial operator representations of quantum decoherence functionals. Journal of Mathematical Physics, 38(11):5643–5652, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schaefer:1997:CPI

[1309] Joe Schaefer. Covariant path integrals on hyperbolic surfaces. Journal of Mathematical Physics, 38(11):5653–5662, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zapata:1997:CAD

[1310] Jose A. Zapata. A combinatorial approach to diffeomorphism invariant quantum gauge theories. Journal of Mathematical Physics, 38(11):5663– 5681, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arutyunov:1997:ERS

[1311] G. E. Arutyunov, S. A. Frolov, and P. B. Medvedev. Elliptic Ruijsenaars– Schneider model from the cotangent bundle over the two-dimensional current group. Journal of Mathematical Physics, 38(11):5682–5689, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rivera:1997:EST

[1312] R. Rivera and D. Villarroel. An exactly solvable two-body problem with retarded interactions and radiation reaction in classical electrodynamics. Journal of Mathematical Physics, 38(11):5690–5706, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1729] and response [1730].

Das:1997:DSP

[1313] Shawpawn Kumar Das, Mylavarapu Sita Janaki, and Brahmananda Dasgupta. Divergences in the solutions of the plasma screening equation. Journal of Mathematical Physics, 38(11):5707–5710, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1997:CIH

[1314] F. Calogero. A class of integrable Hamiltonian systems whose solutions are (perhaps) all completely periodic. Journal of Mathematical Physics, 38(11):5711–5719, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gonzalez-Lopez:1997:IDE

[1315] Artemio Gonzalez-Lopez, Rafael Hernandez Heredero, and Gloria Mari Beffa. Invariant differential equations and the Adler–Gel’fand–Dikii bracket. Journal of Mathematical Physics, 38(11):5720–5738, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marciniak:1997:CKV

[1316] Krzysztof Marciniak. Coupled Korteweg–de Vries hierarchy with sources and its Newton decomposition. Journal of Mathematical Physics, 38(11): 5739–5755, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shaw:1997:MAB

[1317] Jiin-Chang Shaw and Ming-Hsien Tu. Miura and auto-Bácklund transformations for the cKP and cmKP hierarchies. Journal of Mathematical Physics, 38(11):5756–5773, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bartnik:1997:SFN

[1318] Robert Bartnik. Shear-free null quasi-spherical space–times. Journal of Mathematical Physics, 38(11):5774–5791, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chaudhuri:1997:ASM

[1319] S. Chaudhuri and K. C. Das. Axially symmetric metrics from Laplace’s seed by inverse scattering method. Journal of Mathematical Physics, 38 (11):5792–5806, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clement:1997:FWC

[1320] Gerard Clement. Flat wormholes from cosmic strings. Journal of Mathematical Physics, 38(11):5807–5819, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Denisova:1997:DMP

[1321] Irene P. Denisova and Mahavir Dalal. Development of the method of potentials for the problems of gravitation-electromagnetic conversion. Journal of Mathematical Physics, 38(11):5820–5832, November 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hashimoto:1997:AAG

[1322] Yoshitake Hashimoto, Yukinori Yasui, Sayuri Miyagi, and Takayoshi Ootsuka. Applications of the Ashtekar gravity to four-dimensional hyperkahler geometry and Yang–Mills instantons. Journal of Mathematical Physics, 38(11):5833–5839, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heller:1997:GAN

[1323] Michael Heller, Wiešlaw Sasin, and Dominique Lambert. Groupoid approach to noncommutative quantization of gravity. Journal of Mathematical Physics, 38(11):5840–5853, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zafiris:1997:IST

[1324] Elias Zafiris. Incorporation of space–time symmetries in Einstein’s field equations. Journal of Mathematical Physics, 38(11):5854–5877, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Balakrishnan:1997:TGC

[1325] Radha Balakrishnan and Raphael Blumenfeld. Transformation of general curve evolution to a modified Belavin–Polyakov equation. Journal of Mathematical Physics, 38(11):5878–5888, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cho:1997:LCG

[1326] Sunggoo Cho and Kwang Sung Park. Linear connections on graphs. Journal of Mathematical Physics, 38(11):5889–5904, November 1997. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DiFrancesco:1997:MD

[1327] P. Di Francesco. SU(N) meander determinants. Journal of Mathematical Physics, 38(11):5905–5943, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fushchych:1997:HSE

[1328] W. I. Fushchych and A. G. Nikitin. Higher symmetries and exact solutions of linear and nonlinear Schrödinger equation. Journal of Mathematical Physics, 38(11):5944–5959, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:1997:QBL

[1329] Ya-Jun Gao and Yuan-Xing Gui. Quantized braided linear algebras relating to quantized braided groups. Journal of Mathematical Physics, 38 (11):5960–5967, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guil:1997:DSF

[1330] Francisco Guil and Manuel Manas. The Darboux system: Finite-rank constraints and Darboux transformations. Journal of Mathematical Physics, 38(11):5968–5975, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karadayi:1997:ECCa

[1331] H. R. Karadayi and M. Gungormez. An explicit construction of Casimir operators and eigenvalues. I. Journal of Mathematical Physics, 38(11): 5976–5990, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karadayi:1997:ECCb

[1332] H. R. Karadayi and M. Gungormez. An explicit construction of Casimir operators and eigenvalues. Il. Journal of Mathematical Physics, 38(11): 5991–6007, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nash:1997:CDS

[1333] Patrick L. Nash. On the construction of a Dirac spinor that generates a specified tetrad. Journal of Mathematical Physics, 38(11):6008–6017, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quesne:1997:CQU

[1334] C. Quesne. Colored quantum universal enveloping algebras. Journal of Mathematical Physics, 38(11):6018–6039, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Casini:1997:EAP

[1335] Roberto Casini. Erratum: “Algebraic proof of a sum rule occurring in Stark broadening of hydrogen lines” (J. Math. Phys. 38, 3435-3445 (1997)). Journal of Mathematical Physics, 38(11):6040, November 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1152].

Abdesselam:1997:THA

[1336] Boucif Abdesselam. The twisted Heisenberg algebra ⊓_{h,w}(⟨(4)). Journal of Mathematical Physics, 38(12):6045–6060, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaive:1997:ULD

[1337] Bruno Blaive and Michel Cadilhac. Upper limit of the discrete hydrogenlike wave functions: Expansion in the inverse principal quantum number n^{−1}. Journal of Mathematical Physics, 38(12):6061–6071, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bohm:1997:GJV

[1338] A. Bohm, Mark Loewe, S. Maxson, P. Patuleanu, C. Puntmann, and M. Gadella. Gamow–Jordan vectors and non-reducible density operators from higher-order S-matrix poles. Journal of Mathematical Physics, 38 (12):6072–6100, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cassinelli:1997:IQA

[1339] G. Cassinelli, E. De Vito, and A. Levrero. Integrability of the quantum adiabatic evolution and geometric phases. Journal of Mathematical Physics, 38(12):6101–6118, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cheng:1997:CSB

[1340] G. Cheng and T. K. Kuo. Condition on the symmetry-breaking solution of the Schwinger–Dyson equation. Journal of Mathematical Physics, 38 (12):6119–6125, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delabaere:1997:ESE

[1341] Eric Delabaere, Herve Dillinger, and Frederic Pham. Exact semiclassical expansions for one-dimensional quantum oscillators. Journal of Mathematical Physics, 38(12):6126–6184, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:1997:GQJ

[1342] Manuel de Leon, Juan C. Marrero, and Edith Padron. On the geometric quantization of Jacobi manifolds. Journal of Mathematical Physics, 38 (12):6185–6213, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fuertes:1997:SDS

[1343] Wifredo Garcia Fuertes and Juan Mateos Guilarte. Self-dual solitons in N = 2 supersymmetric Chern–Simons gauge theory. Journal of Mathematical Physics, 38(12):6214–6229, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Millard:1997:QMC

[1344] Andrew C. Millard. Quantum mechanics in classical dynamics. Journal of Mathematical Physics, 38(12):6230–6248, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Myung:1997:TDB

[1345] Yun Soo Myung. Two-dimensional boson and W -symmetry in the quantum Hall effect. Journal of Mathematical Physics, 38(12):6249–6264, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nesterenko:1997:JZF

[1346] V. V. Nesterenko and I. G. Pirozhenko. Justification of the zeta function renormalization in rigid string model. Journal of Mathematical Physics, 38(12):6265–6280, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oshima:1997:RCS

[1347] Kazuto Oshima. de Rham cohomology of SO(n) by supersymmetric quantum mechanics. Journal of Mathematical Physics, 38(12):6281– 6286, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Park:1997:DEG

[1348] Yong Moon Park and Hyun Hye Shin. Dynamical entropy of generalized quantum Markov chains over infinite dimensional algebras. Journal of Mathematical Physics, 38(12):6287–6303, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pule:1997:IDL

[1349] J. V. Pule and M. Scrowston. Infinite degeneracy for a Landau Hamiltonian with Poisson impurities. Journal of Mathematical Physics, 38(12): 6304–6314, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:1997:PDO

[1350] Yaping Yang, Weiguo Feng, and Xiang Wu. The q-phase-difference operator and two-mode q-coherent states. Journal of Mathematical Physics, 38(12):6315–6327, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borzdov:1997:FDW

[1351] George N. Borzdov. Frequency domain wave-splitting techniques for plane stratified bianisotropic media. Journal of Mathematical Physics, 38(12):6328–6366, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giannoni:1997:ARB

[1352] Fabio Giannoni, Paolo Piccione, and Jose A. Verderesi. An approach to the relativistic brachistochrone problem by sub-Riemannian geometry. Journal of Mathematical Physics, 38(12):6367–6381, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dubrovsky:1997:DMS

[1353] V. G. Dubrovsky. The dressing method and the solutions with constant asymptotic values at infinity of DS-II equation. Journal of Mathematical Physics, 38(12):6382–6400, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:1997:IML

[1354] Sen-Yue Lou and Xing-Biao Hu. Infinitely many Lax pairs and symmetry constraints of the KP equation. Journal of Mathematical Physics, 38(12): 6401–6427, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meshkov:1997:IIS

[1355] A. G. Meshkov. Integrability and integrodifferential substitutions. Journal of Mathematical Physics, 38(12):6428–6443, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:1997:MMG

[1356] Roman G. Smirnov. Magri–Morosi–Gel’fand-Dorfman’s bi-Hamiltonian constructions in the action-angle variables. Journal of Mathematical Physics, 38(12):6444–6454, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Willox:1997:DBD

[1357] R. Willox, T. Tokihiro, and J. Satsuma. Darboux and binary Darboux transformations for the nonautonomous discrete KP equation. Journal of Mathematical Physics, 38(12):6455–6469, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunn:1997:ECS

[1358] K. A. Dunn, Tina A. Harriott, and J. G. Williams. Energy conditions for a spherically symmetric kink space–time. Journal of Mathematical Physics, 38(12):6470–6474, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ichinose:1997:GCG

[1359] Shoichi Ichinose and Noriaki Ikeda. Graphical classification of global SO(n) invariants and independent general invariants. Journal of Mathematical Physics, 38(12):6475–6521, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smoller:1997:RNL

[1360] J. A. Smoller and A. G. Wasserman. Reissner–Nordstrom-like solutions of the SU(2) Einstein–Yang/Mills equations. Journal of Mathematical Physics, 38(12):6522–6559, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanimoto:1997:HSC

[1361] Masayuki Tanimoto, Tatsuhiko Koike, and Akio Hosoya. Hamiltonian structures for compact homogeneous universes. Journal of Mathematical Physics, 38(12):6560–6577, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benenti:1997:ICV

[1362] Sergio Benenti. Intrinsic characterization of the variable separation in the Hamilton–Jacobi equation. Journal of Mathematical Physics, 38(12): 6578–6602, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Colavecchia:1997:HIA

[1363] F. D. Colavecchia, G. Gasaneo, and C. R. Garibotti. Hypergeometric integrals arising in atomic collisions physics. Journal of Mathematical Physics, 38(12):6603–6612, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jinzenji:1997:QCR

[1364] Masao Jinzenji. On quantum cohomology rings for hypersurfaces in CP^{N−1}. Journal of Mathematical Physics, 38(12):6613–6638, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kota:1997:TBB

[1365] V. K. B. Kota. Transformation brackets between U(N) ⊃ SO(N) ⊃ SO(N_{a})⊕SO(N_{b}) and U(N) ⊃ U(N_{a})⊕U(N_{b}) ⊃ SO(N_{a})⊕SO(N_{b}). Journal of Mathematical Physics, 38(12):6639–6647, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levi:1997:CEH

[1366] D. Levi and R. Yamilov. Conditions for the existence of higher symmetries of evolutionary equations on the lattice. Journal of Mathematical Physics, 38(12):6648–6674, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Plastino:1997:RIM

[1367] A. R. Plastino, H. G. Miller, A. Plastino, and G. D. Yen. The role of information measures in the determination of the maximum entropyminimum norm solution of the generalized inverse problem. Journal of Mathematical Physics, 38(12):6675–6682, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sachse:1997:QUE

[1368] Sebastian Sachse and Ralf Weixler. The quantized universal enveloping algebras Uq(iso(N)), Uq(e(3, 1)) and Uq(e(N)) and the representation theory of Uq(e(3)). Journal of Mathematical Physics, 38(12):6683–6691, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1997:SPA

[1369] Anonymous. Summary of the Physics and Astronomy Classification Scheme — 1996–1997. Journal of Mathematical Physics, 38(12):6692, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1997:SPH

[1370] Anonymous. Summary of PACS headings in current index. Journal of Mathematical Physics, 38(12):6693–6697, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1997:SIV

[1371] Anonymous. Subject index to volume 38. Journal of Mathematical Physics, 38(12):6698–6729, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1997:AIV

[1372] Anonymous. Author index to volume 38. Journal of Mathematical Physics, 38(12):6730–6751, December 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beals:1998:DST

[1373] Richard Beals, David Sattinger, and Eric Williams. A Dirac sea and thermodynamic equilibrium for the quantized three-wave interaction. Journal of Mathematical Physics, 39(1):1–29, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergere:1998:PSC

[1374] M. C. Bergère. Proof of Serban’s conjecture. Journal of Mathematical Physics, 39(1):30–46, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dabrowski:1998:ABE

[1375] L. Dabrowski and P. Šťovíček. Aharonov–Bohm effect with d-type interaction. Journal of Mathematical Physics, 39(1):47–62, January 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hempel:1998:EGP

[1376] Rainer Hempel and Serge Z. Levendorskiĭ. On eigenvalues in gaps for perturbed magnetic Schrödinger operators. Journal of Mathematical Physics, 39(1):63–78, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivanova:1998:IDA

[1377] T. A. Ivanova. On infinite-dimensional algebras of symmetries of the self-dual Yang–Mills equations. Journal of Mathematical Physics, 39(1): 79–87, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

King:1998:QSQ

[1378] Christopher King and Andrezj Leśniewski. Quantum sources and a quantum coding theorem. Journal of Mathematical Physics, 39(1):88–101, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1998:MFH

[1379] Jyh-Hao Lee and Oktay K. Pashaev. Moving frames hierarchy and BF theory. Journal of Mathematical Physics, 39(1):102–123, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Long:1998:TBD

[1380] Peter Long and Horace W. Crater. Two-body Dirac equations for general covariant interactions and their coupled Schrödinger-like forms. Journal of Mathematical Physics, 39(1):124–160, January 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maamache:1998:UTA

[1381] M. Maamache. Unitary transformation approach to the exact solution for a class of time-dependent nonlinear Hamiltonian systems. Journal of Mathematical Physics, 39(1):161–169, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manoliu:1998:ACSa

[1382] Mihaela Manoliu. Abelian Chern–Simons theory. I. A topological quantum field theory. Journal of Mathematical Physics, 39(1):170–206, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manoliu:1998:ACSb

[1383] Mihaela Manoliu. Abelian Chern–Simons theory. II. A functional integral approach. Journal of Mathematical Physics, 39(1):207–217, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pitowsky:1998:IFG

[1384] Itamar Pitowsky. Infinite and finite Gleason’s theorems and the logic of indeterminacy. Journal of Mathematical Physics, 39(1):218–228, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prosser:1998:ESH

[1385] Reese T. Prosser. On the energy spectrum of the hydrogen atom in a photon field I. Journal of Mathematical Physics, 39(1):229–277, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Robinett:1998:POT

[1386] R. W. Robinett. Periodic orbit theory analysis of a continuous family of quasi-circular billiards. Journal of Mathematical Physics, 39(1):278–298, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wheeler:1998:NCG

[1387] James T. Wheeler. New conformal gauging and the electromagnetic theory of Weyl. Journal of Mathematical Physics, 39(1):299–328, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cnen:1998:SLT

[1388] M. Cnen and H. C. Tseng. On the second law of thermodynamics and contact geometry. Journal of Mathematical Physics, 39(1):329–344, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fort:1998:ITD

[1389] J. Fort and J. E. Llebot. Information-theoretical derivation of an extended thermodynamical description of radiative systems. Journal of Mathematical Physics, 39(1):345–354, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dragovic:1998:CCT

[1390] Vladimir Dragović and Milena Radnović. Conditions of Cayley’s type for ellipsoidal billiard. Journal of Mathematical Physics, 39(1):355–362, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1998:HCS

[1391] You-De Wang. Heisenberg chain systems from compact manifolds into S2. Journal of Mathematical Physics, 39(1):363–371, January 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

GarciadeAndrade:1998:SDD

[1392] L. C. Garcia de Andrade. Spacetime defects: Domain walls and torsion. Journal of Mathematical Physics, 39(1):372–379, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasinski:1998:RDSa

[1393] Andrzej Krasiński. Rotating dust solutions of Einstein’s equations with 3dimensional symmetry groups. I. two Killing fields spanned on ua and wa. Journal of Mathematical Physics, 39(1):380–400, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasinski:1998:RDSb

[1394] Andrzej Krasiński. Rotating dust solutions of Einstein’s equations with 3dimensional symmetry groups. II. one Killing field spanned on ua and wa. Journal of Mathematical Physics, 39(1):401–422, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Madore:1998:QST

[1395] J. Madore and J. Mourad. Quantum space–time and classical gravity. Journal of Mathematical Physics, 39(1):423–442, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oliveira-Neto:1998:CRS

[1396] G. Oliveira-Neto. Conditions for regular space — times with compact spatial sections. Journal of Mathematical Physics, 39(1):443–453, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shishkin:1998:DPP

[1397] G. V. Shishkin and W. H. Abdel Salam. The Dirac particle in the presence of the tensor wave on the background of gravitation. Journal of Mathematical Physics, 39(1):454–466, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1998:ASPa

[1398] Jin-Quan Chen and Peng-Dong Fan. Algebraic solutions for point groups: The tetrahedral group for the group chain T ⊃ C3. Journal of Mathematical Physics, 39(1):467–488, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dayi:1998:DSP

[1399] Ömer F. Dayi. q-deformed star products and Moyal brackets. Journal of Mathematical Physics, 39(1):489–496, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Degasperis:1998:SRF

[1400] A. Degasperis, S. De Lillo, and F. Tinebra. Solutions of 1 + 1 relativistic field models. Journal of Mathematical Physics, 39(1):497–531, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Falomir:1998:DOD

[1401] H. Falomir, R. E. Gamboa Saraví, and E. M. Santangelo. Dirac operator on a disk with global boundary conditions. Journal of Mathematical Physics, 39(1):532–544, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gorski:1998:PDO

[1402] Andrzej Z. Górski and Jacek Szmigielski. On pairs of difference operators satisfying (D, X) = Id. Journal of Mathematical Physics, 39(1):545–568, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

John:1998:LFW

[1403] Sarah John. Lattice functions, wavelet aliasing, and SO(3) mappings of orthonormal filters. Journal of Mathematical Physics, 39(1):569–593, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Milhorat:1998:SDO

[1404] Jean-Louis Milhorat. Spectrum of the Dirac operator on Gr_{2}(C^{m+2}). Journal of Mathematical Physics, 39(1):594–609, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Minnaert:1998:RPR

[1405] Pierre Minnaert and Marek Mozrzymas. Representation properties, Racah sum rule, and Biedenharn–Elliott identity for uq(osp(1/2)). Journal of Mathematical Physics, 39(1):610–624, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ran:1998:UCY

[1406] Z. Ran. Unobstructedness of Calabi–Yau orbi-Kleinfolds. Journal of Mathematical Physics, 39(1):625–629, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sun:1998:AER

[1407] Hong zhou Sun and Dong Ruan. Algebraic expressions for O(N) ⊂ O(N − 1) reduction factors for the three-rowed irreducible representations. Journal of Mathematical Physics, 39(1):630–649, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiganov:1998:ADM

[1408] A. V. Tsiganov. Automorphisms of sl(2) and dynamical r-matrices. Journal of Mathematical Physics, 39(1):650–664, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Walton:1998:DCW

[1409] Mark A. Walton. Demazure characters and WZW fusion rules. Journal of Mathematical Physics, 39(1):665–681, January 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bordemann:1998:DII

[1410] Martin Bordemann and Jens Hoppe. Diffeomorphism invariant integrable field theories and hypersurface motions in Riemannian manifolds. Journal of Mathematical Physics, 39(2):683–694, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chalykh:1998:NIG

[1411] O. Chalykh, M. Feigin, and A. Veselov. New integrable generalizations of Calogero–Moser quantum problem. Journal of Mathematical Physics, 39 (2):695–703, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dass:1998:ATL

[1412] Tulsi Dass and Yogesh Joglekar. Augmented temporal logic formalism for histories-based generalized quantum mechanics. Journal of Mathematical Physics, 39(2):704–729, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dubois-Violette:1998:SN

[1413] Michel Dubois-Violette, John Madore, and Richard Kerner. Shadow of noncommutativity. Journal of Mathematical Physics, 39(2):730–738, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Florek:1998:RFV

[1414] Wojciech Florek and Stanisław Wałcerz. The role of a form of vector potential — normalization of the antisymmetric gauge. Journal of Mathematical Physics, 39(2):739–748, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Henderson:1998:RCP

[1415] R. J. Henderson and S. G. Rajeev. Renormalized contact potential in two dimensions. Journal of Mathematical Physics, 39(2):749–759, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Helffer:1998:SLD

[1416] B. Helffer. Splitting in large dimension and infrared estimates. II. moment inequalities. Journal of Mathematical Physics, 39(2):760–776, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Honegger:1998:UIO

[1417] Reinhard Honegger and Alfred Rieckers. Unitary implementations of one-parameter squeezing groups. Journal of Mathematical Physics, 39 (2):777–801, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leitenberger:1998:QTR

[1418] Frank Leitenberger. Quantum theory of the real and the complexified projective line. Journal of Mathematical Physics, 39(2):802–817, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marcinek:1998:PSM

[1419] Władysław Marcinek. Particles in singular magnetic field. Journal of Mathematical Physics, 39(2):818–830, February 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1731].

Mickelsson:1998:VPG

[1420] Jouko Mickelsson. Vacuum polarization and the geometric phase: Gauge invariance. Journal of Mathematical Physics, 39(2):831–837, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Seshadri:1998:LOI

[1421] S. Seshadri, V. Balakrishnan, and S. Lakshmibala. Ladder operators for isospectral oscillators. Journal of Mathematical Physics, 39(2):838–847, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sorella:1998:ACV

[1422] S. P. Sorella, L. C. Q. Vilar, O. S. Ventura, and C. A. G. Sasaki. Algebraic characterization of vector supersymmetry in topological field theories. Journal of Mathematical Physics, 39(2):848–866, February 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Wu:1998:QPB

[1423] Yihren Wu. Quantization of a particle in a background Yang–Mills field. Journal of Mathematical Physics, 39(2):867–875, February 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deLeon:1998:HSC

[1424] Manuel de León, Eugenio Merino, José A. Oubiña, Paulo R. Rodrigues, and Modesto R. Salgado. Hamiltonian systems on k-cosymplectic manifolds. Journal of Mathematical Physics, 39(2):876–893, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

DeSanto:1998:DBI

[1425] J. A. DeSanto and P. A. Martin. On the derivation of boundary integral equations for scattering by an infinite two-dimensional rough surface. Journal of Mathematical Physics, 39(2):894–912, February 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karimipour:1998:ISN

[1426] V. Karimipour. Integrable structure of the new Calogero models. Journal of Mathematical Physics, 39(2):913–920, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joussot:1998:QCF

[1427] Patrice Joussot and Valentin A. Zagrebnov. Quantum critical fluctuations in a ferroelectric model: Quasi-average approach. Journal of Mathematical Physics, 39(2):921–930, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klosek:1998:CNC

[1428] M. M. Kłosek and P. S. Hagan. Colored noise and a characteristic level crossing problem. Journal of Mathematical Physics, 39(2):931–953, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mendez:1998:VPN

[1429] Vicenç Méndez. Variational principles for nonlinear dynamical systems. Journal of Mathematical Physics, 39(2):954–966, February 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Samsonov:1998:DPS

[1430] Boris F. Samsonov. Distortion of a phase space under the Darboux transformation. Journal of Mathematical Physics, 39(2):967–975, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ustinov:1998:RSD

[1431] N. V. Ustinov. The reduced self-dual Yang–Mills equation, binary and infinitesimal Darboux transformations. Journal of Mathematical Physics, 39(2):976–985, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:1998:NCS

[1432] Zixiang Zhou. Nonlinear constraints and soliton solutions of 1 + 2dimensional three-wave equation. Journal of Mathematical Physics, 39 (2):986–997, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mignemi:1998:CMI

[1433] S. Mignemi and H.-J. Schmidt. Classification of multidimensional inflationary models. Journal of Mathematical Physics, 39(2):998–1010, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petters:1998:FPD

[1434] A. O. Petters and F. J. Wicklin. Fixed points due to gravitational lenses. Journal of Mathematical Physics, 39(2):1011–1023, February 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sopuerta:1998:SGK

[1435] Carlos F. Sopuerta. Stationary generalized Kerr–Schild spacetimes. Journal of Mathematical Physics, 39(2):1024–1039, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Branson:1998:VEV

[1436] Thomas P. Branson, Peter B. Gilkey, and Dmitri V. Vassilevich. Vacuum expectation value asymptotics for second order differential operators on manifolds with boundary. Journal of Mathematical Physics, 39(2):1040– 1049, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2263].

Brauer:1998:BSS

[1437] Uwe Brauer. Breakdown of smooth solutions of the three-dimensional Euler–Poisson system. Journal of Mathematical Physics, 39(2):1050– 1074, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsenko:1998:ZFP

[1438] Andrei A. Bytsenko and Floyd L. Williams. Zeta functions on a product of Einstein manifolds, and the multiplicative anomaly. Journal of Mathematical Physics, 39(2):1075–1086, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deGuise:1998:CGC

[1439] Hubert de Guise and David J. Rowe. Clebsch–Gordan coefficients in the asymptotic limit. Journal of Mathematical Physics, 39(2):1087–1106, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Forger:1998:IPM

[1440] Michael Forger. Invariant polynomials and Molien functions. Journal of Mathematical Physics, 39(2):1107–1141, February 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gawrylczyk:1998:GAM

[1441] Jarosław Gawrylczyk. Generalization of the algebra of the matrix pseudodifferential operators. Journal of Mathematical Physics, 39(2):1142– 1152, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Honegger:1998:ETF

[1442] Reinhard Honegger. Enlarged test-function spaces for the global free folia dynamics on the CCR-algebra. Journal of Mathematical Physics, 39(2): 1153–1169, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kovalev:1998:GAR

[1443] Vladimir F. Kovalev, Veniamin V. Pustovalov, and Dmitrii V. Shirkov. Group analysis and renormgroup symmetries. Journal of Mathematical Physics, 39(2):1170–1188, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Patrangenaru:1998:TDM

[1444] Victor Patrangenaru. Three-dimensional metrics with a spherical homogeneous model. Journal of Mathematical Physics, 39(2):1189–1198, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Quesne:1998:DCQ

[1445] C. Quesne. Duals of colored quantum universal enveloping algebras and colored universal I-matrices. Journal of Mathematical Physics, 39(2): 1199–1222, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reinsch:1998:YMF

[1446] Matthias Reinsch. Yang–Mills field equations for the Coriolis tensor in the n-body problem. Journal of Mathematical Physics, 39(2):1223–1235, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thiemann:1998:ILT

[1447] T. Thiemann. The inverse loop transform. Journal of Mathematical Physics, 39(2):1236–1248, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doktorov:1998:EST

[1448] E. V. Doktorov. Erratum: “Spectral transform and solitons for the threewave coupling model with nontrivial boundary conditions”, (J. Math. Phys. 38, 4138–4150 (1997)). Journal of Mathematical Physics, 39(2): 1249, February 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1201].

Baez:1998:QDI

[1449] John C. Baez and Kirill V. Krasnov. Quantization of diffeomorphisminvariant theories with fermions. Journal of Mathematical Physics, 39 (3):1251–1271, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernoff:1998:OSD

[1450] Andrew Bernoff and Peter Sternberg. Onset of superconductivity in decreasing fields for general domains. Journal of Mathematical Physics, 39 (3):1272–1284, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brink:1998:HAS

[1451] Lars Brink, Alexander Turbiner, and Niclas Wyllard. Hidden algebras of the (super) Calogero and Sutherland models. Journal of Mathematical Physics, 39(3):1285–1315, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cattaneo:1998:BST

[1452] Alberto S. Cattaneo, Paolo Cotta-Ramusino, and Maurizio Rinaldi. BRST symmetries for the tangent gauge group. Journal of Mathematical Physics, 39(3):1316–1339, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dranov:1998:DTA

[1453] A. Dranov, J. Kellendonk, and R. Seiler. Discrete time adiabatic theorems for quantum mechanical systems. Journal of Mathematical Physics, 39(3):1340–1349, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Filippov:1998:MIT

[1454] G. F. Filippov, A. F. Lisetskyi, and J. P. Draayer. Microscopic SU_{π} ⊗ SU_{ν}(3) interpretation of a two triaxial rotor model. Journal of Mathematical Physics, 39(3):1350–1365, March 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ghosh:1998:GAO

[1455] Gautam Ghosh. Generalized annihilation operator coherent states. Journal of Mathematical Physics, 39(3):1366–1372, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Holevo:1998:RND

[1456] A. S. Holevo. Radon–Nikodym derivatives of quantum instruments. Journal of Mathematical Physics, 39(3):1373–1387, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khrennikov:1998:PAS

[1457] Andrew Khrennikov. p-adic stochastic hidden variable model. Journal of Mathematical Physics, 39(3):1388–1401, March 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naito:1998:PVO

[1458] Seichi Naito. Physical vertex operators in the Ramond sector. Journal of Mathematical Physics, 39(3):1402–1440, March 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schlesinger:1998:TQM

[1459] Karl-Georg Schlesinger. On topology and quantum mechanics. Journal of Mathematical Physics, 39(3):1441–1450, March 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Willis:1998:SID

[1460] Barton L. Willis. Similarity and intertwining of Dirac operators. Journal of Mathematical Physics, 39(3):1451–1457, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eidus:1998:RPI

[1461] D. Eidus. Radiation principle for inhomogeneous media. Journal of Mathematical Physics, 39(3):1458–1477, March 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prieto:1998:EGR

[1462] Angel Prieto. On the electrodynamic group: The relativistic radiation reaction force. Journal of Mathematical Physics, 39(3):1478–1494, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tveter:1998:DHE

[1463] Frank Thomas Tveter. Deriving the Hamilton equations of motion for a nonconservative system using a variational principle. Journal of Mathematical Physics, 39(3):1495–1500, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

FariadaVeiga:1998:CLH

[1464] P. A. Faria da Veiga, M. O’Carroll, and R. Schor. A classical large N hierarchical vector model in three dimensions: A nonzero fixed point and canonical decay of correlation functions. Journal of Mathematical Physics, 39(3):1501–1516, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gaveau:1998:TNF

[1465] Bernard Gaveau and L. S. Schulman. Theory of nonequilibrium firstorder phase transitions for stochastic dynamics. Journal of Mathematical Physics, 39(3):1517–1533, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miller:1998:SMB

[1466] Steven Miller. A stochastic method for Brownian-like optical transport calculations in anisotropic biosuspensions and blood. Journal of Mathematical Physics, 39(3):1534–1550, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Endean:1998:NTE

[1467] Geoffrey Endean. A note on Tauber’s expanding universe in conformally flat coordinates. Journal of Mathematical Physics, 39(3):1551– 1554, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iriondo:1998:FSS

[1468] Mirta S. Iriondo, Enzo O. Leguizamón, and Oscar A. Reula. Fast and slow solutions in General Relativity: The initialization procedure. Journal of Mathematical Physics, 39(3):1555–1565, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ashdown:1998:MRR

[1469] M. A. J. Ashdown, S. S. Somaroo, S. F. Gull, C. J. L. Doran, and A. N. Lasenby. Multilinear representations of rotation groups within geometric algebra. Journal of Mathematical Physics, 39(3):1566–1588, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1998:ADS

[1470] Cheng Chen and G. Cheng. Anomalous dimension in the solution of the Barenblatt’s equation. Journal of Mathematical Physics, 39(3):1589– 1600, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foda:1998:DMV

[1471] Omar Foda, Kailash C. Misra, and Masato Okado. Damazure modules and vertex models: The sl(2) case. Journal of Mathematical Physics, 39 (3):1601–1622, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hayaishi:1998:ODG

[1472] Norifumi Hayaishi and Kei Miki. L operators and Drinfeld’s generators. Journal of Mathematical Physics, 39(3):1623–1636, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:1998:TSC

[1473] Wen ling Huang. Transformations of strongly causal space–times preserving null geodesics. Journal of Mathematical Physics, 39(3):1637–1641, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

MacLeod:1998:SCT

[1474] Malcolm A. MacLeod. The spherical curl transform of a linear forcefree magnetic field. Journal of Mathematical Physics, 39(3):1642–1658, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sneddon:1998:IAI

[1475] G. E. Sneddon. The identities of the algebraic invariants of the fourdimensional Riemann tensor. II. Journal of Mathematical Physics, 39 (3):1659–1679, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thomas:1998:CDI

[1476] Launey J. Thomas, III and Eyvind H. Wichmann. On a class of distributions of interest in quantum field theory ac. Journal of Mathematical Physics, 39(3):1680–1719, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramidi:1998:ECA

[1477] I. G. Avramidi. Erratum: “Covariant algebraic method for calculation of the low-energy heat kernel”, (J. Math. Phys. 36, 5055–5070 (1995)). Journal of Mathematical Physics, 39(3):1720, March 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [364].

Adler:1998:GFP

[1478] Stephen L. Adler. Gauge fixing in the partition function for generalized quantum dynamics. Journal of Mathematical Physics, 39(4):1723–1729, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arians:1998:GAI

[1479] Silke Arians. Geometric approach to inverse scattering for hydrogenlike systems in a homogeneous magnetic field. Journal of Mathematical Physics, 39(4):1730–1743, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ban:1998:PSR

[1480] Masashi Ban. Phase-space representation of quantum state vectors. Journal of Mathematical Physics, 39(4):1744–1765, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [1982].

Haba:1998:FIR

[1481] Z. Haba. Feynman integral in regularized nonrelativistic quantum electrodynamics. Journal of Mathematical Physics, 39(4):1766–1787, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herdegen:1998:SPC

[1482] Andrzej Herdegen. Semidirect product of CCR and CAR algebras and asymptotic states in quantum electrodynamics. Journal of Mathematical Physics, 39(4):1788–1817, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isham:1998:CTC

[1483] C. J. Isham, N. Linden, K. Sawidou, and S. Schreckenberg. Continuous time and consistent histories. Journal of Mathematical Physics, 39(4): 1818–1834, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lesniewski:1998:CLQ

[1484] Andrzej Lesniewski, Ron Rubin, and Nathan Salwen. Classical limits for quantum maps on the torus. Journal of Mathematical Physics, 39 (4):1835–1847, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naito:1998:OPE

[1485] Seichi Naito. Operator product expansions and stress operators. Journal of Mathematical Physics, 39(4):1848–1886, April 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Saenz:1998:LSA

[1486] A. W. Sáenz. Lie-series approach to the evolution of resonant and nonresonant anharmonic oscillators in quantum mechanics. Journal of Mathematical Physics, 39(4):1887–1909, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2266].

Sezgin:1998:GPT

[1487] M. Sezgin, A. Y. Verdiyev, and Y. A. Verdiyev. Generalized Poschl– Teller, Toda, Morse potentials and SL(2, R) group. Journal of Mathematical Physics, 39(4):1910–1918, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spector:1998:DPS

[1488] Donald Spector. Duality, partial supersymmetry, and arithmetic number theory. Journal of Mathematical Physics, 39(4):1919–1927, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Volkel:1998:SSB

[1489] A. H. Völkel. Symmetry and symmetry breaking in quantumchromo(flavor)-dynamics. Journal of Mathematical Physics, 39(4):1928– 1956, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:1998:WSO

[1490] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. Wave scattering in one dimension with absorption. Journal of Mathematical Physics, 39(4):1957–1992, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meson:1998:GFH

[1491] Alejandro M. Mesón and Femando Vericat. Geodesic flow on an hyperboloid model and the two-body motion under repulsive Coulomb forces. Journal of Mathematical Physics, 39(4):1993–1996, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Panayotaros:1998:DCS

[1492] Panayotis Panayotaros. Diagrams in classical and semiclassical perturbation theory. Journal of Mathematical Physics, 39(4):1997–2003, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cornille:1998:ESN

[1493] H. Cornille. Exact solutions for nonconservative two-velocity discrete Bolkmann models. Journal of Mathematical Physics, 39(4):2004–2018, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Coyle:1998:IMP

[1494] Lester N. Coyle. Infinite moments of the partition function for random walks in a random potential. Journal of Mathematical Physics, 39(4): 2019–2034, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyazawa:1998:AAS

[1495] Toru Miyazawa. Algebraic approach to scattering theory of the Fokker– Planck equation. Journal of Mathematical Physics, 39(4):2035–2076, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Okumura:1998:RSL

[1496] Ko Okumura. A reformulation of simple liquids theory — renormalization by one-, two-, and three-particle densities. Journal of Mathematical Physics, 39(4):2077–2102, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurses:1998:ICK

[1497] Metin Gürses and Atalay Karasu. Integrable coupled KdV systems. Journal of Mathematical Physics, 39(4):2103–2111, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:1998:KEP

[1498] Sen yue Lou. KdV extensions with Painlevé property. Journal of Mathematical Physics, 39(4):2112–2121, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Myrzakulov:1998:SDI

[1499] R. Myrzakulov, S. Vijayalakshmi, R. N. Syzdykova, and M. Lakshmanan. On the simplest (2 + 1) dimensional integrable spin systems and their equivalent nonlinear Schrödinger equations. Journal of Mathematical Physics, 39(4):2122–2140, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bergqvist:1998:PPB

[1500] Göran Bergqvist. Positivity properties of the Bel–Robinson tensor. Journal of Mathematical Physics, 39(4):2141–2147, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krasinski:1998:RDSc

[1501] Andrzej Krasiński. Rotating dust solutions of Einstein’s equations with three-dimensional symmetry groups. III. all Killing fields linearly independent of ua and wa. Journal of Mathematical Physics, 39(4):2148– 2179, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reboucas:1998:RST

[1502] M. J. Reboučas and A. F. F. Teixeira. Riemannian space–times of Gödel type in five dimensions. Journal of Mathematical Physics, 39(4):2180– 2192, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sajko:1998:GCM

[1503] William N. Sajko, Paul S. Wesson, and Hongya Liu. Gauge conditions in modern Kaluza–Klein theory. Journal of Mathematical Physics, 39 (4):2193–2200, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Steinbauer:1998:GGD

[1504] R. Steinbauer. Geodesics and geodesic deviation for impulsive gravitational waves. Journal of Mathematical Physics, 39(4):2201–2212, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tolksdorf:1998:EHY

[1505] Júrgen Tolksdorf. The Einstein–Hilbert–Yang–Mills–Higgs action and the Dirac–Yukawa operator. Journal of Mathematical Physics, 39(4): 2213–2241, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Braham:1998:YMF

[1506] S. P. Braham and J. Gegenberg. Yang–Mills flow and uniformization theorems. Journal of Mathematical Physics, 39(4):2242–2253, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brau:1998:IEF

[1507] F. Brau. Integral equation formulation of the spinless Salpeter equation. Journal of Mathematical Physics, 39(4):2254–2263, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buchwalter:1998:RPL

[1508] S. A. Buchwalter and P. Moylan. Representations of the Poincaré Lie algebra from unitary representations of SO(1, 4). Journal of Mathematical Physics, 39(4):2264–2272, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:1998:FBR

[1509] X. M. Ding, B. Y. Hou, B. Yuan Hou, and L. Zhao. Free boson representation of DYh(glN)k and DYh(slN)k. Journal of Mathematical Physics, 39(4):2273–2289, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golenishcheva-Kutuzova:1998:CA

[1510] Maria I. Golenishcheva-Kutuzova and Victor G. Kac. G-conformal algebras. Journal of Mathematical Physics, 39(4):2290–2305, April 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gratus:1998:NBS

[1511] J. Gratus. A natural basis for spinor and vector fields on the noncommutative sphere. Journal of Mathematical Physics, 39(4):2306–2324, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karaa:1998:EEG

[1512] Samir Karaa. Extremal eigenvalue gaps for the Schrödinger operator with Dirichlet boundary conditions. Journal of Mathematical Physics, 39(4): 2325–2332, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kiselev:1998:PTD

[1513] O. M. Kiselev. Perturbation theory for the Dirac equation in twodimensional space. Journal of Mathematical Physics, 39(4):2333–2345, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ludu:1998:WQA

[1514] A. Ludu, M. Greiner, and J. P. Draayer. Wavelets and quantum algebras. Journal of Mathematical Physics, 39(4):2346–2361, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McInnes:1998:EPS

[1515] Brett McInnes. Existence of parallel spinors on nonsimply connected Riemannian manifolds. Journal of Mathematical Physics, 39(4):2362– 2366, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sartori:1998:OSC

[1516] G. Sartori and V. Talamini. Orbit spaces of compact coregular simple Lie groups with 2, 3 and 4 basic polynomial invariants: Effective tools in the analysis of invariant potentials. Journal of Mathematical Physics, 39 (4):2367–2401, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sontz:1998:ESB

[1517] Stephen Bruce Sontz. Entropy and the Segal–Bargmann transform. Journal of Mathematical Physics, 39(4):2402–2417, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Vercin:1998:OPA

[1518] A. Vercin. Ordered products, W (∞)-algebra, and two-variable, definiteparity, orthogonal polynomials. Journal of Mathematical Physics, 39(4): 2418–2427, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wagner:1998:ENO

[1519] Peter Wagner. Evaluation of nonrelativistic one-loop Feynman integrals by distributional methods. Journal of Mathematical Physics, 39(4):2428– 2436, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeiner:1998:BWF

[1520] P. Zeiner, R. Dirl, and B. L. Davies. Bloch and Wannier functions composed of Gaussian orbitals — analyticity, localization, and expectation values. Journal of Mathematical Physics, 39(4):2437–2454, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frick:1998:ENH

[1521] Rudolf A. Frick. Erratum: “A new Hamiltonian operator for a massive relativistic particle with spin one in a generalized Heisenberg/ Schrödinger picture” (J. Math. Phys. 38, 3457–3465 (1997)). Journal of Mathematical Physics, 39(4):2455, April 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1154].

Antoniou:1998:GVD

[1522] I. E. Antoniou, M. Gadella, and G. P. Pronko. Gamow vectors for degenerate scattering resonances. Journal of Mathematical Physics, 39(5): 2459–2475, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arai:1998:RTA

[1523] Asao Arai. Representation-theoretic aspects of two-dimensional quantum systems in singular vector potentials: Canonical commutation relations, quantum algebras. Journal of Mathematical Physics, 39(5):2476–2498, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Babujian:1998:GKZ

[1524] H. M. Babujian and A. V. Kitaev. Generalized Knizhnik–Zamolodchikov equations and isomonodromy quantization of the equations integrable via the Inverse Scattering Transform: Maxwell–Bloch system with pumping. Journal of Mathematical Physics, 39(5):2499–2506, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bering:1998:GBF

[1525] K. Bering. Geometry of the Batalin–Fradkin–Vilkovisky theorem. Journal of Mathematical Physics, 39(5):2507–2519, May 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Buslaev:1998:IPS

[1526] Vladimir Buslaev and Alain Grigis. Imaginary parts of Stark–Wannier resonances. Journal of Mathematical Physics, 39(5):2520–2550, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dekar:1998:ESS

[1527] Liès Dekar, Lyazid Chetouani, and Theophile F. Hammann. An exactly soluble Schrödinger equation with smooth position-dependent mass.

Journal of Mathematical Physics, 39(5):2551–2563, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hasegawa:1998:ILS

[1528] Hiroshi Hasegawa and Jian-Zhong Ma. Intermediate level statistics with one-parameter random matrix ensembles. Journal of Mathematical Physics, 39(5):2564–2583, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mangano:1998:PIA

[1529] Gianpiero Mangano. Path integral approach to noncommutative space– times. Journal of Mathematical Physics, 39(5):2584–2591, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mavromatis:1998:TQE

[1530] Harry A. Mavromatis. From Turbiner’s quasi-exactly soluble potentials in N dimensions to analytic solutions of the combined Coulomb plus oscillator system. Journal of Mathematical Physics, 39(5):2592–2596, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scherer:1998:QAI

[1531] Wolfgang Scherer. Quantum averaging. III. time-dependent systems. Journal of Mathematical Physics, 39(5):2597–2634, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Soloviev:1998:UTD

[1532] M. A. Soloviev. A uniqueness theorem for distributions and its application to nonlocal quantum field theory. Journal of Mathematical Physics, 39(5):2635–2642, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thomas:1998:SFL

[1533] Launey J. Thomas III and Eyvind H. Wichmann. Standard forms of local nets in quantum field theory. Journal of Mathematical Physics, 39 (5):2643–2681, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yamashita:1998:HDR

[1534] Hideyasu Yamashita. Hyperfinite-dimensional representations of canonical commutation relation. Journal of Mathematical Physics, 39(5):2682– 2692, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioannidou:1998:SMC

[1535] T. Ioannidou and W. J. Zakrzewski. Solutions of the modified chiral model in (2 + 1) dimensions. Journal of Mathematical Physics, 39(5): 2693–2701, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mosley:1998:APE

[1536] Shaun N. Mosley. Alternative potentials for the electromagnetic field. Journal of Mathematical Physics, 39(5):2702–2713, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sardanashvily:1998:HTD

[1537] Gennadi A. Sardanashvily. Hamiltonian time-dependent mechanics. Journal of Mathematical Physics, 39(5):2714–2729, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bagarello:1998:NST

[1538] Fabio Bagarello and Geoffrey L. Sewell. New structures in the theory of the laser model. II. microscopic dynamics and a nonequilibrium entropy principle. Journal of Mathematical Physics, 39(5):2730–2747, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Franz:1998:DBS

[1539] U. Franz and R. Schott. Diffusion on braided spaces. Journal of Mathematical Physics, 39(5):2748–2762, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lindblad:1998:BMQ

[1540] Göran Lindblad. Brownian motion of quantum harmonic oscillators: Existence of a subdynamics. Journal of Mathematical Physics, 39(5): 2763–2780, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Montarnal:1998:SHL

[1541] Philippe Montarnal. Study of hydrodynamical limits in a multicollision scale Boltzmann equation for semiconductors. Journal of Mathematical Physics, 39(5):2781–2799, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cervero:1998:MTB

[1542] J. M. Cerveró and P. G. Estévez. Miura transformation between two nonlinear equations in 2+1 dimensions. Journal of Mathematical Physics, 39 (5):2800–2807, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Corona:1998:CCT

[1543] Gulmaro Corona Corona. Couple coefficients for the two-block AKNS system. Journal of Mathematical Physics, 39(5):2808–2812, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ju:1998:ASC

[1544] Guo xing Ju, Shi kun Wang, and Ke Wu. The algebraic structure of the gl(n.m) color Calogero–Sutherland models. Journal of Mathematical Physics, 39(5):2813–2820, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rerikh:1998:NIO

[1545] K. V. Rerikh. Nonalgebraic integrability of one reversible dynamical system of the Cremona type. Journal of Mathematical Physics, 39(5): 2821–2832, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vandeLeur:1998:SBT

[1546] Johan van de Leur. Schlesinger–Bäcklund transformations for the Ncomponent KP. Journal of Mathematical Physics, 39(5):2833–2847, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhou:1998:LRM

[1547] Ruguang Zhou. Lax representation, r-matrix method, and separation of variables for the Neumann-type restricted flow. Journal of Mathematical Physics, 39(5):2848–2858, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Andersson:1998:CFA

[1548] F. Andersson and S. Brian Edgar. Curvature-free asymmetric metric connections and Lanczos potentials in Kerr–Schild spacetimes. Journal of Mathematical Physics, 39(5):2859–2861, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gao:1998:CRS

[1549] Sijie Gao, Zhiquan Kuang, and Canbin Liang. “clock rate synchronizable” reference frames in curved space–times. Journal of Mathematical Physics, 39(5):2862–2865, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivashchuk:1998:MCQ

[1550] V. D. Ivashchuk and V. N. Melnikov. Multidimensional classical and quantum cosmology with intersecting p-branes. Journal of Mathematical Physics, 39(5):2866–2888, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Avramidi:1998:GFH

[1551] Ivan G. Avramidi. Green functions of higher-order differential operators. Journal of Mathematical Physics, 39(5):2889–2909, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boyallian:1998:QHW

[1552] Carina Boyallian, Victor G. Kac, José I. Liberati, and Catherine H. Yan. Quasifinite highest weight modules over the Lie algebra of matrix differential operators on the circle. Journal of Mathematical Physics, 39 (5):2910–2928, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Figueiredo:1998:ASI

[1553] A. Figueiredo, T. M. Rocha Filho, and L. Brenig. Algebraic structures and invariant manifolds of differential systems. Journal of Mathematical Physics, 39(5):2929–2946, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Folguera:1998:DST

[1554] Alejandra Folguera. Davey–Stewartson type equations of the second and third order: Derivation and classification. Journal of Mathematical Physics, 39(5):2947–2968, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mnatsakanova:1998:IRH

[1555] M. Mnatsakanova, G. Morchio, F. Strocchi, and Yu Vernov. Irreducible representations of the Heisenberg algebra in Krein spaces. Journal of Mathematical Physics, 39(5):2969–2982, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tahri:1998:QGS

[1556] E. H. Tahri. Quantum group structure on the quantum plane. Journal of Mathematical Physics, 39(5):2983–2992, May 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoniou:1998:RGV

[1557] I. Antoniou, M. Gadella, I. Prigogine, and G. P. Pronko. Relativistic Gamow vectors. Journal of Mathematical Physics, 39(6):2995–3018, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Applebaum:1998:QMM

[1558] David Applebaum. Quantum martingale measures and stochastic partial differential equations in Fock space. Journal of Mathematical Physics, 39 (6):3019–3030, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arthur:1998:IEG

[1559] K. Arthur, Y. Brihaye, and D. H. Tchrakian. Interaction energy of generalized Abelian Higgs vortices. Journal of Mathematical Physics, 39 (6):3031–3042, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cotaescu:1998:RSS

[1560] Ion I. Cotăescu. A relativistic supersymmetry and some algebraic consequences. Journal of Mathematical Physics, 39(6):3043–3049, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dehesa:1998:SAL

[1561] J. S. Dehesa, R. J. Yáñez, A. I. Aptekarev, and V. Buyarov. Strong asymptotics of Laguerre polynomials and information entropies of twodimensional harmonic oscillator and one-dimensional Coulomb potentials. Journal of Mathematical Physics, 39(6):3050–3060, June 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eberspacher:1998:MNS

[1562] Matthias Eberspächer and Werner Scheid. The modified newton-Sabatier method for the coupled channel inverse scattering problem with charged particles at fixed energy. Journal of Mathematical Physics, 39(6):3061– 3072, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frieler:1998:NAS

[1563] K. Frieler and K.-H. Rehren. A non-Abelian square root of Abelian vertex operators. Journal of Mathematical Physics, 39(6):3073–3090, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grundling:1998:CDQ

[1564] Hendrik Grundling and C. A. Hurst. Constrained dynamics for quantum mechanics. I. restricting a particle to a surface. Journal of Mathematical Physics, 39(6):3091–3119, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oshima:1998:RCS

[1565] Kazuto Oshima. de Rham cohomology of SO(n) and some related manifolds by supersymmetric quantum mechanics. Journal of Mathematical Physics, 39(6):3120–3130, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simeone:1998:QMO

[1566] Claudio Simeone. Quantization of minisuperspaces as ordinary gauge systems. Journal of Mathematical Physics, 39(6):3131–3137, June 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cendra:1998:MVE

[1567] Hernán Cendra, Darryl D. Holm, Mark J. W. Hoyle, and Jerrold E. Marsden. The Maxwell–Vlasov equations in Euler–Poincaré form. Journal of Mathematical Physics, 39(6):3138–3157, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sonego:1998:MEO

[1568] Sebastiano Sonego and Marek A. Abramowicz. Maxwell equations and the optical geometry. Journal of Mathematical Physics, 39(6):3158–3166, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez:1998:MDC

[1569] Ariel Fernández and Andrés Colubri. Microscopic dynamics from a coarsely defined solution to the protein folding problem. Journal of Mathematical Physics, 39(6):3167–3187, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerra:1998:FEC

[1570] Francesco Guerra and Andreas Knauf. Free energy and correlations of the number-theoretical spin chain. Journal of Mathematical Physics, 39 (6):3188–3202, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bhutani:1998:GSS

[1571] O. P. Bhutani and K. Singh. Generalized similarity solutions for the type D fluid in five-dimensional flat space. Journal of Mathematical Physics, 39(6):3203–3212, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaszak:1998:SBH

[1572] Maciej Błaszak. On separability of bi-Hamiltonian chain with degenerated Poisson structures. Journal of Mathematical Physics, 39(6):3213– 3235, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cerkaski:1998:ALF

[1573] Marcin Cerkaski. An application of Lax’s formalism to a class of curvilinear axially symmetric coordinates in three-dimensional space. Journal of Mathematical Physics, 39(6):3236–3245, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1998:PCC

[1574] Deng yuan Chen. The potential constraints of the complex (1 + 2)- dimensional soliton systems and related Hamiltonian equations. Journal of Mathematical Physics, 39(6):3246–3259, June 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:1998:TCB

[1575] L. Herrera and J. Martínez. Thermal conduction before relaxation in slowly rotating fluids. Journal of Mathematical Physics, 39(6):3260–3270, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qiao:1998:DIS

[1576] Zhijun Qiao and Walter Strampp. On different integrable systems sharing the same nondynamical r-matrix. Journal of Mathematical Physics, 39 (6):3271–3279, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Abdussattar:1998:FST

[1577] Abdussattar and Babita Dwivedi. Fluid space–times and conharmonic symmetries. Journal of Mathematical Physics, 39(6):3280–3295, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barrett:1998:RSN

[1578] John W. Barrett and Louis Crane. Relativistic spin networks and quantum gravity. Journal of Mathematical Physics, 39(6):3296–3302, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doran:1998:EST

[1579] Chris Doran, Anthony Lasenby, Anthony Challinor, and Stephen Gull. Effects of spin-torsion in gauge theory gravity. Journal of Mathematical Physics, 39(6):3303–3321, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feoli:1998:DSB

[1580] A. Feoli, W. R. Wood, and G. Papini. A dynamical symmetry breaking model in Weyl space. Journal of Mathematical Physics, 39(6):3322–3331, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Low:1998:SSW

[1581] Robert Low. Stable singularities of wave-fronts in general relativity. Journal of Mathematical Physics, 39(6):3332–3335, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nilsson:1998:SPF

[1582] Ulf S. Nilsson, Claes Uggla, and Mattias Marklund. Static perfect fluid cylinders. Journal of Mathematical Physics, 39(6):3336–3346, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thiemann:1998:CFM

[1583] T. Thiemann. Closed formula for the matrix elements of the volume operator in canonical quantum gravity. Journal of Mathematical Physics, 39(6):3347–3371, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thiemann:1998:LOC

[1584] T. Thiemann. A length operator for canonical quantum gravity. Journal of Mathematical Physics, 39(6):3372–3392, June 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bott:1998:NHF

[1585] E. Bott, M. Methfessel, W. Krabs, and P. C. Schmidt. Nonsingular Hankel functions as a new basis for electronic structure calculations. Journal of Mathematical Physics, 39(6):3393–3425, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:1998:DCS

[1586] Salih Çelik, Sultan A. Çelik, and Metin Arik. Differential calculus on the h-superplane. Journal of Mathematical Physics, 39(6):3426–3436, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fiore:1998:DML

[1587] Gaetano Fiore. Deforming maps for Lie group covariant creation and annihilation operators. Journal of Mathematical Physics, 39(6):3437– 3452, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gieres:1998:SCO

[1588] François Gieres and Stéphane Gourmelen. d = 2, N = 2 superconformally covariant operators and super W -algebras. Journal of Mathematical Physics, 39(6):3453–3475, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:1998:AIT

[1589] Avinash Khare and Bhabani Prasad Mandal. Anti-isospectral transformations, orthogonal polynomials, and quasi-exactly solvable problems. Journal of Mathematical Physics, 39(6):3476–3486, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lyubashenko:1998:GLA

[1590] Volodymyr Lyubashenko and Anthony Sudbery. Generalized Lie algebras of type A_{n}. Journal of Mathematical Physics, 39(6):3487–3504, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Spichak:1998:SCE

[1591] Stanislav Spichak and Valerii Stognii. Symmetry classification and exact solutions of the Kramers equation. Journal of Mathematical Physics, 39 (6):3505–3510, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1998:EES

[1592] G. F. Torres del Castillo. Erratum: “Exact solutions of the Einstein equations with sources from linearized solutions” (J. Math. Phys. 37, 4584–4589 (1996)). Journal of Mathematical Physics, 39(6):3511, June 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [799].

Ananos:1998:ARM

[1593] G. N. J. Añaños, A. P. C. Malbouisson, and N. F. Svaiter. Analytic regularization of the (ψ̄_{i}, ψ_{i})_{D}^{2} model at finite temperature and density. Journal of Mathematical Physics, 39(7):3515–3521, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castagnino:1998:DRN

[1594] M. Castagnino, G. Domenech, M. Levinas, and N. Umérez. The doublet representation of non-Hilbert eigenstates of the Hamiltonian. II. Journal of Mathematical Physics, 39(7):3522–3529, July 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Emmrich:1998:MWK

[1595] C. Emmrich and H. Römer. Multicomponent Wentzel–Kramers–Brillouin approximation on arbitrary symplectic manifolds: A star product approach. Journal of Mathematical Physics, 39(7):3530–3546, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gavrilov:1998:QEF

[1596] S. P. Gavrilov, D. M. Gitman, and A. E. Gonçalves. QED in external field with space–time uniform invariants: Exact solutions. Journal of Mathematical Physics, 39(7):3547–3567, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hennig:1998:ESE

[1597] D. Hennig. Exponential stability for the energy exchange dynamics of the Holstein model. Journal of Mathematical Physics, 39(7):3568–3576, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jurco:1998:AKS

[1598] Branislav Jurčo and Peter Schupp. Adler–Kostant–Symes scheme for face and Calogero–Moser–Sutherland-type models. Journal of Mathematical Physics, 39(7):3577–3588, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kunhardt:1998:IST

[1599] Walter Kunhardt. On infravacua and superselection theory. Journal of Mathematical Physics, 39(7):3589–3599, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanWinter:1998:IEF

[1600] Clasine van Winter. An increasing entropy for a free quantum particle. Journal of Mathematical Physics, 39(7):3600–3618, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bandelloni:1998:KSD

[1601] Giuseppe Bandelloni and Serge Lazzarini. Kodaira–Spencer deformation of complex structures and Lagrangian field theory. Journal of Mathematical Physics, 39(7):3619–3642, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marengo:1998:TDP

[1602] Edwin A. Marengo and Anthony J. Devaney. Time-dependent plane wave and multipole expansions of the electromagnetic field. Journal of Mathematical Physics, 39(7):3643–3660, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stump:1998:ATE

[1603] Daniel R. Stump. Arnol’d’s transformation — An example of a generalized canonical transformation. Journal of Mathematical Physics, 39 (7):3661–3669, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lukkarinen:1998:LAQ

[1604] J. Lukkarinen. Lattice approximation of quantum statistical traces at a complex temperature. Journal of Mathematical Physics, 39(7):3670– 3677, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Percival:1998:GPD

[1605] D. J. Percival and P. A. Robinson. Generalized plasma dispersion functions. Journal of Mathematical Physics, 39(7):3678–3693, July 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ahluwalia:1998:ADA

[1606] Daljit S. Ahluwalia, Joseph B. Keller, and Charles Knessl. Advection– diffusion around a curved obstacle. Journal of Mathematical Physics, 39 (7):3694–3710, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorren:1998:LTK

[1607] H. J. S. Dorren. A linearizing transformation for the Korteweg–de Vries equation; generalizations to higher-dimensional nonlinear partial differential equations. Journal of Mathematical Physics, 39(7):3711–3729, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Felipe:1998:ESI

[1608] Raúl Felipe and Fausto Ongay. N-extended superelliptic integrable systems. Journal of Mathematical Physics, 39(7):3730–3737, July 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:1998:MFH

[1609] Wen-Jui Huang, Jiin-Chang Shaw, and Ming-Hsien Tu. Matrix formulation of Hamiltonian structures of constrained KP hierarchy. Journal of Mathematical Physics, 39(7):3738–3753, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khusnutdinova:1998:SHG

[1610] K. R. Khusnutdinova and H. Steudel. Second harmonic generation: Hamiltonian structures and particular solutions. Journal of Mathematical Physics, 39(7):3754–3764, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lakshmanan:1998:MCS

[1611] M. Lakshmanan, R. Myrzakulov, S. Vijayalakshmi, and A. K. Danlybaeva. Motion of curves and surfaces and nonlinear evolution equations in (2+1) dimensions. Journal of Mathematical Physics, 39(7):3765–3771, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leblond:1998:SSL

[1612] H. Leblond. The secular solutions of the linearized Korteweg–de Vries equation. Journal of Mathematical Physics, 39(7):3772–3782, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tasnadi:1998:HCM

[1613] Tamás Tasnádi. Hard chaos in magnetic billiards (on the hyperbolic plane). Journal of Mathematical Physics, 39(7):3783–3804, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clayton:1998:CGR

[1614] M. A. Clayton. Canonical general relativity: The diffeomorphism constraints and spatial frame transformations. Journal of Mathematical Physics, 39(7):3805–3816, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:1998:GLS

[1615] L. Herrera and N. O. Santos. Geodesics in Lewis space–time. Journal of Mathematical Physics, 39(7):3817–3827, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bautista:1998:PBW

[1616] César Bautista. A Poincaré–Birkhoff–Witt theorem for generalized Lie color algebras. Journal of Mathematical Physics, 39(7):3828–3843, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Billig:1998:PVO

[1617] Yuly Billig. Principal vertex operator representations for toroidal Lie algebras. Journal of Mathematical Physics, 39(7):3844–3864, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feigin:1998:EBC

[1618] B. L. Feigin, A. M. Semikhatov, and I. Yu. Tipunin. Equivalence between chain categories of representations of affine sl(2 ) and N = 2 superconformal algebras. Journal of Mathematical Physics, 39(7):3865–3905, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flude:1998:EAF

[1619] James P. M. Flude. The Edmonds asymptotic formulas for the 3-j and 6-j symbols. Journal of Mathematical Physics, 39(7):3906–3915, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ritter:1998:CRB

[1620] Stefan Ritter. On a class of Robin boundary value problems in physical geodesy. Journal of Mathematical Physics, 39(7):3916–3926, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Thomova:1998:SDS

[1621] Z. Thomova, P. Winternitz, and W. J. Zakrzewski. Solutions of (2 + 1)- dimensional spin systems. Journal of Mathematical Physics, 39(7):3927– 3944, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1998:ICP

[1622] Anonymous. Information for contributors. Journal of Mathematical Physics, 39(7):v–??, July 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Torresani:1998:ISI

[1623] Bruno Torrésani. Introduction to the special issue on wavelet and timefrequency analysis. Journal of Mathematical Physics, 39(8):3949–3953, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TwarequeAli:1998:GTS

[1624] S. Twareque Ali. A general theorem on square-integrability: Vector coherent states. Journal of Mathematical Physics, 39(8):3954–3964, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aniello:1998:WTD

[1625] Paolo Aniello, Gianni Cassinelli, Ernesto De Vito, and Alberto Levrero. Wavelet transforms and discrete frames associated to semidirect products. Journal of Mathematical Physics, 39(8):3965–3973, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fuhr:1998:CWT

[1626] Hartmut Führ. Continuous wavelet transforms with Abelian dilation groups. Journal of Mathematical Physics, 39(8):3974–3986, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoine:1998:WSR

[1627] J.-P. Antoine and P. Vandergheynst. Wavelets on the n-sphere and related manifolds. Journal of Mathematical Physics, 39(8):3987–4008, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paul:1998:WTF

[1628] T. Paul and A. Uribe. Weighted trace formula near a hyperbolic trajectory and complex orbits. Journal of Mathematical Physics, 39(8): 4009–4015, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Flandrin:1998:SPM

[1629] Patrick Flandrin. Separability, positivity, and minimum uncertainty in time–frequency energy distributions. Journal of Mathematical Physics, 39(8):4016–4040, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matz:1998:TFT

[1630] Gerald Matz and Franz Hlawatsch. Time-frequency transfer function calculus (symbolic calculus) of linear time-varying systems (linear operators) based on a generalized underspread theory. Journal of Mathematical Physics, 39(8):4041–4070, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bertrand:1998:SCT

[1631] J. Bertrand and P. Bertrand. Symbolic calculus on the time-frequency half-plane. Journal of Mathematical Physics, 39(8):4071–4090, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frick:1998:WAS

[1632] Peter Frick, Alex Grossmann, and Philippe Tchamitchian. Wavelet analysis of signals with gaps. Journal of Mathematical Physics, 39(8):4091– 4107, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lemarie-Rieusset:1998:SRN

[1633] Pierre Gilles Lemarié-Rieusset. Some remarks on the Navier–Stokes equations in R^{3}. Journal of Mathematical Physics, 39(8):4108–4118, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lewalle:1998:FIS

[1634] Jacques Lewalle. Formal improvements in the solution of the wavelettransformed Poisson and diffusion equations. Journal of Mathematical Physics, 39(8):4119–4128, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jaffard:1998:OSP

[1635] Stéphane Jaffard. Oscillation spaces: Properties and applications to fractal and multifractal functions. Journal of Mathematical Physics, 39 (8):4129–4141, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Arneodo:1998:RCW

[1636] A. Arneodo, E. Bacry, and J. F. Muzy. Random cascades on wavelet dyadic trees. Journal of Mathematical Physics, 39(8):4142–4164, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerin:1998:TDS

[1637] Charles-Antoine Guerin and Matthias Holschneider. Time-dependent scattering on fractal measures. Journal of Mathematical Physics, 39(8): 4165–4194, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zak:1998:OSL

[1638] J. Zak. Orthonormal sets of localized functions for a Landau level. Journal of Mathematical Physics, 39(8):4195–4200, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gazeau:1998:TWP

[1639] Jean Pierre Gazeau, Jiří Patera, and Edita Pelantová. Tau-wavelets in the plane. Journal of Mathematical Physics, 39(8):4201–4212, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bernuau:1998:WBA

[1640] Guillaume Bernuau. Wavelets bases adapted to a self-similar quasicrystal. Journal of Mathematical Physics, 39(8):4213–4225, August 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:1998:SWA

[1641] Yuchuan Wei and Nanxian Chen. Square wave analysis. Journal of Mathematical Physics, 39(8):4226–4245, August 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aktosun:1998:NBS

[1642] Tuncay Aktosun, Martin Klaus, and Cornelis van der Mee. On the number of bound states for the one-dimensional Schrödinger equation. Journal of Mathematical Physics, 39(9):4249–4256, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brandao:1998:SSG

[1643] Augusto Brandão. Symplectic structure for Gaussian diffusions. Journal of Mathematical Physics, 39(9):4257–4283, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Casini:1998:ECM

[1644] Roberto Casini. The effect of configuration mixing on the first-order moments of polarized hydrogen lines. Journal of Mathematical Physics, 39(9):4284–4298, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Davydychev:1998:GAF

[1645] A. I. Davydychev and R. Delbourgo. A geometrical angle on Feynman integrals. Journal of Mathematical Physics, 39(9):4299–4334, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

deOliveira:1998:PDV

[1646] César R. de Oliveira. A proof of the dynamical version of the Bombieri– Taylor conjecture. Journal of Mathematical Physics, 39(9):4335–4342, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duan:1998:GDT

[1647] Yishi Duan and Libin Fu. The general decomposition theory of SU (2 ) gauge potential, topological structure and bifurcation of SU (2 ) Chern density. Journal of Mathematical Physics, 39(9):4343–4355, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fan:1998:EQA

[1648] Heng Fan, Bo yu Hou, Kang jie Shi, and Wen li Yang. The elliptic quantum algebra A_{q,p}(ŝl_{n}) and its bosonization at level one. Journal of Mathematical Physics, 39(9):4356–4368, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fassari:1998:CCT

[1649] Silvestro Fassari and Martin Klaus. Coupling constant thresholds of perturbed periodic Hamiltonians. Journal of Mathematical Physics, 39 (9):4369–4416, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Froman:1998:JCF

[1650] Nanny Fröman and Per Olof Fröman. JWKB connection-formula problem revisited again via Borel summation. Journal of Mathematical Physics, 39(9):4417–4429, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gargett:1998:DEO

[1651] T. D. Gargett and I. N. McArthur. Derivative expansion of one loop effective actions for Yang–Mills fields. Journal of Mathematical Physics, 39(9):4430–4448, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hattori:1998:MDC

[1652] Tetsuya Hattori and Hiroshi Watanabe. Mathematical derivation of chiral anomaly in lattice gauge theory with Wilson’s action. Journal of Mathematical Physics, 39(9):4449–4475, September 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kuna:1998:FIC

[1653] Tobias Kuna, Ludwig Streit, and Werner Westerkamp. Feynman integrals for a class of exponentially growing potentials. Journal of Mathematical Physics, 39(9):4476–4491, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mitra:1998:SLF

[1654] A. N. Mitra and R. Ramanathan. On simulating Liouvillian flow from quantum mechanics via Wigner functions. Journal of Mathematical Physics, 39(9):4492–4498, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:1998:TCF

[1655] Ali Mostafazadeh. Two-component formulation of the Wheeler–DeWitt equation. Journal of Mathematical Physics, 39(9):4499–4512, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Peter:1998:QAF

[1656] Ingo J. Peter and Gérard G. Emch. Quantum Anosov flows: A new family of examples. Journal of Mathematical Physics, 39(9):4513–4539, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sato:1998:TFL

[1657] Haru-Tada Sato. Thermodynamic fermion loop in a constant magnetic field. Journal of Mathematical Physics, 39(9):4540–4545, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Crawford:1998:AEE

[1658] John David Crawford and Anandhan Jayaraman. Amplitude equations for electrostatic waves: Multiple species. Journal of Mathematical Physics, 39(9):4546–4577, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Echeverria-Enriquez:1998:MFC

[1659] Arturo Echeverría-Enríquez, Miguel C. Muñoz-Lecanda, and Narciso Román-Roy. Multivector fields and connections: Setting Lagrangian equations in field theories. Journal of Mathematical Physics, 39(9):4578– 4603, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Felbacq:1998:LAD

[1660] D. Felbacq, B. Guizal, and F. Zolla. Limit analysis of the diffraction of a plane wave by a one-dimensional periodic medium. Journal of Mathematical Physics, 39(9):4604–4607, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belgiorno:1998:BHT

[1661] F. Belgiorno. Black hole thermodynamics and spectral analysis. Journal of Mathematical Physics, 39(9):4608–4633, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Degond:1998:MMS

[1662] P. Degond and C. Schmeiser. Macroscopic models for semiconductor heterostructures. Journal of Mathematical Physics, 39(9):4634–4663, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gozani:1998:PWT

[1663] Joseph Gozani. Propagation of waves through non-Gaussian longitudinally correlated random media. Journal of Mathematical Physics, 39(9): 4664–4674, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krantz:1998:MES

[1664] Richard J. Krantz, Jack Douthett, and Steven D. Doty. Maximally even sets and the devil’s-staircase phase diagram for the one-dimensional Ising antiferromagnet with arbitrary-range interaction. Journal of Mathematical Physics, 39(9):4675–4682, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bogdanov:1998:ABAa

[1665] L. V. Bogdanov and B. G. Konopelchenko. Analytic-bilinear approach to integrable hierarchies. I. Generalized KP hierarchy. Journal of Mathematical Physics, 39(9):4683–4700, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bogdanov:1998:ABAb

[1666] L. V. Bogdanov and B. G. Konopelchenko. Analytic-bilinear approach to integrable hierarchies. II. Multicomponent KP and 2D Toda lattice hierarchies. Journal of Mathematical Physics, 39(9):4701–4728, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Delduc:1998:SDS

[1667] F. Delduc and L. Gallot. Supersymmetric Drinfeld–Sokolov reduction. Journal of Mathematical Physics, 39(9):4729–4745, September 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Fan:1998:ODM

[1668] Heng Fan, Bo-Yu Hou, Guang-Liang Li, Kang-Jie Shi, and Yan-Shen Wang. A one-dimensional many-body integrable model from Z_{n} Belavin model with open boundary conditions. Journal of Mathematical Physics, 39(9):4746–4758, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gonera:1998:DPI

[1669] Cezary Gonera. Direct proof of integrability of Calogero–Moser model. Journal of Mathematical Physics, 39(9):4759–4765, September 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:1998:SNR

[1670] Xing-Biao Hu and Zuo-Nong Zhu. Some new results on the Blaszak– Marciniak lattice: Bäcklund transformation and nonlinear superposition formula. Journal of Mathematical Physics, 39(9):4766–4772, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shaw:1998:BDB

[1671] Jiin-Chang Shaw and Ming-Hsien Tu. Binary Darboux–Bäcklund transformations for the Manin–Radul super KdV hierarchy. Journal of Mathematical Physics, 39(9):4773–4784, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsuchida:1998:ISD

[1672] Takayuki Tsuchida, Hideaki Ujino, and Miki Wadati. Integrable semidiscretization of the coupled modified KdV equations. Journal of Mathematical Physics, 39(9):4785–4813, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Canarutto:1998:PDT

[1673] Daniel Canarutto. Possibly degenerate tetrad gravity and Maxwell–Dirac fields. Journal of Mathematical Physics, 39(9):4814–4823, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hajicek:1998:GTQ

[1674] P. Hájíček. Group-theoretical quantization of 2+1 gravity in the metrictorus sector. Journal of Mathematical Physics, 39(9):4824–4848, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mavromatos:1998:ETH

[1675] N. E. Mavromatos and E. Winstanley. Existence theorems for hairy black holes in ∫⊓(N) Einstein–Yang–Mills theories. Journal of Mathematical Physics, 39(9):4849–4873, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sardanashvily:1998:CSS

[1676] Gennadi A. Sardanashvily. Covariant spin structure. Journal of Mathematical Physics, 39(9):4874–4890, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanimoto:1998:NVG

[1677] Masayuki Tanimoto. New varieties of Gowdy space–times. Journal of Mathematical Physics, 39(9):4891–4898, September 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barannyk:1998:CSC

[1678] L. F. Barannyk, P. Basarab-Horwath, and W. I. Fushchych. On the classification of subalgebras of the conformal algebra with respect to inner automorphisms. Journal of Mathematical Physics, 39(9):4899–4922, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:1998:QDS

[1679] Y. Brihaye. Quommutator deformations of spl(N, 1). Journal of Mathematical Physics, 39(9):4923–4927, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fauser:1998:ETO

[1680] Bertfried Fauser. On an easy transition from operator dynamics to generating functionals by Clifford algebras. Journal of Mathematical Physics, 39(9):4928–4947, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Galaktionov:1998:MSI

[1681] Victor A. Galaktionov and Sergey A. Posashkov. Maximal sign-invariants of quasilinear parabolic equations with gradient diffusivity. Journal of Mathematical Physics, 39(9):4948–4964, September 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guichardet:1998:ROA

[1682] A. Guichardet. On the representations of a Q-oscillator algebra. Journal of Mathematical Physics, 39(9):4965–4969, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

John:1998:ACG

[1683] P. John, O. Moritsch, M. Schweda, and S. P. Sorella. Algebraic characterization of gauge anomalies on a nontrivial bundle. Journal of Mathematical Physics, 39(9):4970–4981, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Juszczak:1998:CLD

[1684] C. Juszczak and J. T. Sobczyk. Classification of low-dimensional Lie super-bialgebras. Journal of Mathematical Physics, 39(9):4982–4992, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kakei:1998:IOD

[1685] Saburo Kakei. Intertwining operators for a degenerate double affine Hecke algebra and multivariable orthogonal polynomials. Journal of Mathematical Physics, 39(9):4993–5006, September 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lagraa:1998:LAS

[1686] M. Lagraa and N. Touhami. Lie-algebraic structure from inhomogeneous Hopf algebras. Journal of Mathematical Physics, 39(9):5007–5014, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pati:1998:IDA

[1687] K. C. Pati and D. Parashar. Iwasawa decomposition of affine Kac–Moody algebras using Satake diagrams. Journal of Mathematical Physics, 39(9): 5015–5023, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scheunert:1998:CLS

[1688] M. Scheunert and R. B. Zhang. Cohomology of Lie superalgebras and their generalizations. Journal of Mathematical Physics, 39(9):5024–5061, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanderJeugt:1998:RGF

[1689] J. Van der Jeugt and R. Jagannathan. Realizations of su(1, 1) and U_{q} (su(1, 1)) and generating functions for orthogonal polynomials. Journal of Mathematical Physics, 39(9):5062–5078, September 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Adler:1998:CEC

[1690] Stephen L. Adler and Achim Kempf. Corrections to the emergent canonical commutation relations arising in the statistical mechanics of matrix models. Journal of Mathematical Physics, 39(10):5083–5097, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruning:1998:CFO

[1691] E. Brüning and S. Nagamachi. Closure of field operators, asymptotic Abelianness, and vacuum structure in hyperfunction quantum field theory. Journal of Mathematical Physics, 39(10):5098–5111, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Elout:1998:ICE

[1692] Melchior O. Elout, David Z. Goodson, Carl D. Elliston, Shi-Wei Huang, Alexei V. Sergeev, and Deborah K. Watson. Improving the convergence and estimating the accuracy of summation approximants of 1/D expansions for Coulombic systems. Journal of Mathematical Physics, 39(10): 5112–5122, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Escher:1998:FRN

[1693] Jutta Escher and J. P. Draayer. Fermion realization of the nuclear Sp(6 , R) model. Journal of Mathematical Physics, 39(10):5123–5147, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ferrari:1998:OFB

[1694] Franco Ferrari and Jan T. Sobczyk. Operator formalism for bosonic beta–gamma fields on general algebraic curves. Journal of Mathematical Physics, 39(10):5148–5165, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garg:1998:ADW

[1695] Anupam Garg. Application of the discrete Wentzel–Kramers–Brillouin method to spin tunneling. Journal of Mathematical Physics, 39(10):

5166–5179, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2261].

Giannitrapani:1998:QCE

[1696] R. Giannitrapani. Quantum coordinates of an event in local quantum physics. Journal of Mathematical Physics, 39(10):5180–5182, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Labastida:1998:KIV

[1697] J. M. F. Labastida and Esther Pérez. Kontsevich integral for Vassiliev invariants from Chern–Simons perturbation theory in the light-cone gauge. Journal of Mathematical Physics, 39(10):5183–5198, October 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1998:LAC

[1698] C.-W. H. Lee and S. G. Rajeev. A Lie algebra for closed strings, spin chains, and gauge theories. Journal of Mathematical Physics, 39(10): 5199–5230, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szmytkowski:1998:OFW

[1699] Radosław Szmytkowski. Operator formulation of Wigner’s R-matrix theories for the Schrödinger and Dirac equations. Journal of Mathematical Physics, 39(10):5231–5252, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2104].

Wu:1998:EBT

[1700] Zuo-Bing Wu and Jin-Yan Zeng. Extension of Bertrand’s theorem and factorization of the radial Schrödinger equation. Journal of Mathematical Physics, 39(10):5253–5259, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yu:1998:CIT

[1701] Sixia Yu and Yongde Zhang. A constructive index theorem, phase operators, and phase-difference operators. Journal of Mathematical Physics, 39(10):5260–5267, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1998:ISM

[1702] F. Calogero. Integrable and solvable many-body problems in the plane via complexification. Journal of Mathematical Physics, 39(10):5268– 5291, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marciniak:1998:TFN

[1703] Krzysztof Marciniak and Stefan Rauch-Wojciechowski. Two families of nonstandard Poisson structures for Newton equations. Journal of Mathematical Physics, 39(10):5292–5306, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miller:1998:SCF

[1704] Steven D. Miller. Stochastic construction of a Feynman path integral representation for Green’s functions in radiative transfer. Journal of Mathematical Physics, 39(10):5307–5315, October 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Meyer-Ortmanns:1998:MCF

[1705] Hildegard Meyer-Ortmanns and Thomas Reisz. The monotony criterion for a finite size scaling analysis of phase transitions. Journal of Mathematical Physics, 39(10):5316–5323, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aldrovandi:1998:CID

[1706] R. Aldrovandi and L. P. Freitas. Continuous iteration of dynamical maps. Journal of Mathematical Physics, 39(10):5324–5336, October 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Belich:1998:VOR

[1707] H. Belich, G. Cuba, and R. Paunov. Vertex operator representation of the soliton tau functions in the A_{n}^{(1)} Toda models by dressing transformations. Journal of Mathematical Physics, 39(10):5337–5363, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:1998:HDP

[1708] Sen yue Lou and Jian jun Xu. Higher dimensional Painlevé integrable models from the Kadomtsev–Petviashvili equation. Journal of Mathematical Physics, 39(10):5364–5376, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pelinovsky:1998:RSK

[1709] Dmitry Pelinovsky. Rational solutions of the KP hierarchy and the dynamics of their poles. II. Construction of the degenerate polynomial solutions. Journal of Mathematical Physics, 39(10):5377–5395, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sabatier:1998:LIT

[1710] Pierre C. Sabatier. On local inverse transforms. Journal of Mathematical Physics, 39(10):5396–5405, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dolan:1998:LPV

[1711] P. Dolan and B. D. Muratori. The Lanczos potential for vacuum space– times with an Ernst potential. Journal of Mathematical Physics, 39(10): 5406–5420, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gravel:1998:SSS

[1712] Pierre Gravel. Static stringlike solutions in five-dimensional relativity. Journal of Mathematical Physics, 39(10):5421–5426, October 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Harris:1998:UFC

[1713] Steven G. Harris. Universality of the future chronological boundary. Journal of Mathematical Physics, 39(10):5427–5445, October 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kechkin:1998:SGM

[1714] Oleg Kechkin and Maria Yurova. Symplectic gravity models in four, three and two dimensions. Journal of Mathematical Physics, 39(10):5446–5457, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mak:1998:FCB

[1715] M. K. Mak and T. Harko. Full causal bulk-viscous cosmological models. Journal of Mathematical Physics, 39(10):5458–5476, October 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nurowski:1998:CFE

[1716] Paweł Nurowski. On a certain formulation of the Einstein equations. Journal of Mathematical Physics, 39(10):5477–5480, October 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Nurowski:1998:EFA

[1717] Paweł Nurowski. Elliptic fibrations associated with the Einstein space– times. Journal of Mathematical Physics, 39(10):5481–5490, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tilley:1998:PIC

[1718] Daniel Tilley and Roy Maartens. Perturbations in inflationary cosmologies with smooth exit. Journal of Mathematical Physics, 39(10):5491– 5501, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1998:ASP

[1719] Jin-Quan Chen and Peng-Dong Fan. Algebraic solutions for point groups: The tetrahedral group for the group chain T ⊃ D_{2} ⊃ C_{2}. Journal of Mathematical Physics, 39(10):5502–5518, October 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1998:AEC

[1720] Jin-Quan Chen and Peng-Dong Fan. Algebraic expressions of the Clebsch–Gordan coefficients of the point group T^{†}. Journal of Mathematical Physics, 39(10):5519–5535, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Foong:1998:AOD

[1721] S. K. Foong. Addendum: “Overcoming dissipative distortions by a waveform restorer and designer”, [J. Math. Phys. 36, 2324 (1995)]. Journal of Mathematical Physics, 39(10):5536–5542, October 1998. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [173].

Lam:1998:DTO

[1722] C. S. Lam. Decomposition of time-ordered products and path-ordered exponentials. Journal of Mathematical Physics, 39(10):5543–5558, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molev:1998:FDI

[1723] A. I. Molev. Finite-dimensional irreducible representations of twisted Yangians. Journal of Mathematical Physics, 39(10):5559–5600, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mountain:1998:ITC

[1724] Arthur J. Mountain. Invariant tensors and Casimir operators for simple compact Lie groups. Journal of Mathematical Physics, 39(10):5601–5607, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mudrov:1998:TCF

[1725] A. I. Mudrov. Twisting cocycles in fundamental representation and triangular bi-crossproduct Hopf algebras. Journal of Mathematical Physics, 39(10):5608–5616, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Musette:1998:BTP

[1726] Micheline Musette and Robert Conte. Bäcklund transformation of partial differential equations from the Painlevé–Gambier classification. I. Kaup– Kupershmidt equation. Journal of Mathematical Physics, 39(10):5617– 5630, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:1998:CGRa

[1727] Feng Pan and J. P. Draayer. Complementary group resolution of the SU (n) outer multiplicity problem. I. the Littlewood rules and a complementary U(∈, \ − ∈) group structure. Journal of Mathematical Physics, 39(10):5631–5641, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pan:1998:CGRb

[1728] Feng Pan and J. P. Draayer. Complementary group resolution of the SU (n) outer multiplicity problem. II. recoupling approach for SU (3 ) ⊃ U (2 ) reduced Wigner coefficients. Journal of Mathematical Physics, 39(10):5642–5662, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hnizdo:1998:CES

[1729] V. Hnizdo. Comment on “An exactly solvable two-body problem with retarded interactions and radiation reaction in classical electrodynamics”, [J. Math. Phys. 38, 5690 (1997)]. Journal of Mathematical Physics, 39(10):5663, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1312, 1730].

Rivera:1998:RCE

[1730] R. Rivera and D. Villarroel. Response to “Comment on ‘An exactly solvable two-body problem with retarded interactions and radiation reaction in classical electrodynamics’ ” (J. Math. Phys. 39, 5663 (1998)). Journal of Mathematical Physics, 39(10):5664, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1312, 1729].

Marcinek:1998:EPS

[1731] Władysław Marcinek. Erratum: “Particles in singular magnetic field”, [J. Math. Phys. 39, 818 (1998)]. Journal of Mathematical Physics, 39 (10):5665, October 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1419].

Barcelos-Neto:1998:TFG

[1732] J. Barcelos-Neto and R. Thibes. Two-form gauge fields in the Becchi– Rouet–Stora–Tyutin superspace. Journal of Mathematical Physics, 39 (11):5669–5692, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaive:1998:EHL

[1733] Bruno Blaive and Michel Cadilhac. Expansion of the hydrogen-like wave functions in the inverse principal quantum number, for the discrete and continuous spectra. Journal of Mathematical Physics, 39(11):5693–5699, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bodurov:1998:CHE

[1734] Theodore Bodurov. Complex Hamiltonian evolution equations and field theory. Journal of Mathematical Physics, 39(11):5700–5715, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daumens:1998:CQR

[1735] Michel Daumens and Pierre Minnaert. Canonical and quasicanonical realizations of the osp(D/d) superalgebra and the Coulomb problem in superspaces. Journal of Mathematical Physics, 39(11):5716–5725, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Evans:1998:WTN

[1736] T. S. Evans, T. W. B. Kibble, and D. A. Steer. Wick’s theorem for nonsymmetric normal ordered products and contractions. Journal of Mathematical Physics, 39(11):5726–5738, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Floreanini:1998:ADI

[1737] Roberto Floreanini, Luc Lapointe, and Luc Vinet. Algebraic description of S_{N}-invariant oscillator states. Journal of Mathematical Physics, 39 (11):5739–5761, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Georgiev:1998:RCF

[1738] L. S. Georgiev and I. T. Todorov. Rational conformal field theory extensions of W_{1+∞} in terms of bilocal fields. Journal of Mathematical Physics, 39(11):5762–5771, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gudder:1998:HAQ

[1739] Stanley Gudder. A histories approach to quantum mechanics. Journal of Mathematical Physics, 39(11):5772–5788, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:1998:NQS

[1740] Avinash Khare and Bhabani Prasad Mandal. A number of quasiexactly solvable N-body problems. Journal of Mathematical Physics, 39(11): 5789–5797, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Krause:1998:QKR

[1741] J. Krause. SU(2) quantum kinematics: Rotation-observable versus angular-momentum generalized commutation relations. Journal of Mathematical Physics, 39(11):5798–5810, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Levai:1998:UTC

[1742] G. Lévai, B. Kónya, and Z. Papp. Unified treatment of the Coulomb and harmonic oscillator potentials in D dimensions. Journal of Mathematical Physics, 39(11):5811–5823, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

OKeeffe:1998:SMC

[1743] D. O’Keeffe, D. H. Tchrakian, and Yisong Yang. Sigma models coupled with Abelian gauge fields. Journal of Mathematical Physics, 39(11):5824– 5831, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palev:1998:HWI

[1744] T. D. Palev and N. I. Stoilova. Highest weight irreducible representations of the quantum algebra U_{h}(A_{∞}). Journal of Mathematical Physics, 39 (11):5832–5849, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudolph:1998:SHQ

[1745] Oliver Rudolph. Symmetries of history quantum theories and decoherence functionals. Journal of Mathematical Physics, 39(11):5850–5855, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chandre:1998:VTA

[1746] C. Chandre and H. R. Jauslin. A version of Thirring’s approach to the Kolmogorov–Arnold–Moser theorem for quadratic Hamiltonians with degenerate twist. Journal of Mathematical Physics, 39(11):5856–5865, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dragovic:1998:PTB

[1747] Vladimir Dragović and Milena Radnović. On periodical trajectories of the billiard systems within an ellipsoid in R^{d} and generalized Cayley’s condition. Journal of Mathematical Physics, 39(11):5866–5869, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Margetis:1998:EFA

[1748] Dionisios Margetis. Electromagnetic fields in air of traveling-wave currents above the earth. Journal of Mathematical Physics, 39(11):5870– 5893, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pekarsky:1998:PVS

[1749] Sergey Pekarsky and Jerrold E. Marsden. Point vortices on a sphere: Stability of relative equilibria. Journal of Mathematical Physics, 39(11):

5894–5907, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sarlet:1998:SSV

[1750] W. Sarlet, R. Sahadevan, and F. Cantrijn. Symmetries, separability and volume forms. Journal of Mathematical Physics, 39(11):5908–5924, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teleman:1998:REA

[1751] Kostake Teleman. Relativistic eccentric anomalies. Journal of Mathematical Physics, 39(11):5925–5933, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wollensak:1998:MFA

[1752] Matthias Wollensak. Maxwell fields in anisotropic space–times. Journal of Mathematical Physics, 39(11):5934–5945, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsuchiya:1998:DFS

[1753] Osamu Tsuchiya. Determinant formula for the six-vertex model with reflecting end. Journal of Mathematical Physics, 39(11):5946–5951, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dixon:1998:TCC

[1754] Tony W. Dixon and Brian G. Kenny. Transition to criticality in circle maps at the golden mean. Journal of Mathematical Physics, 39(11): 5952–5963, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:1998:FDM

[1755] Yunbo Zeng and Runliang Lin. Families of dynamical r-matrices and Jacobi inversion problem for nonlinear evolution equations. Journal of Mathematical Physics, 39(11):5964–5983, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ansorg:1998:TGM

[1756] M. Ansorg. Timelike geodesic motions within the general relativistic gravitational field of the rigidly rotating disk of dust. Journal of Mathematical Physics, 39(11):5984–6000, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Beem:1998:CHE

[1757] John K. Beem and Andrzej Królak. Cauchy horizon end points and differentiability. Journal of Mathematical Physics, 39(11):6001–6010, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bicak:1998:SAF

[1758] J. Bičák and A. Pravdová. Symmetries of asymptotically flat electrovacuum space–times and radiation. Journal of Mathematical Physics, 39 (11):6011–6039, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bombelli:1998:CRW

[1759] Luca Bombelli, Fernando Lombardo, and Mario Castagnino. Chaos in Robertson–Walker cosmology. Journal of Mathematical Physics, 39(11): 6040–6050, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Breton:1998:CPW

[1760] Nora Bretón, Alberto García, Alfredo Macías, and Gustavo Yáñez. Colliding plane waves in terms of Jacobi functions. Journal of Mathematical Physics, 39(11):6051–6065, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hollmann:1998:HSA

[1761] Helia Hollmann. A harmonic space approach to quantum gravity of stationary space–times with SO(3 ) symmetry. Journal of Mathematical Physics, 39(11):6066–6085, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Liu:1998:CSI

[1762] Quan-Hui Liu. The complete Schwarzschild interior and exterior solution in the harmonic coordinate system. Journal of Mathematical Physics, 39 (11):6086–6090, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2101] and response [2102].

Bagarello:1998:ATA

[1763] F. Bagarello. Applications of topological ^{∗}-algebras of unbounded operators. Journal of Mathematical Physics, 39(11):6091–6105, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blake:1998:CMM

[1764] Stephen Blake. Calculus on the multivector manifold spin ^{+}(p, q). Journal of Mathematical Physics, 39(11):6106–6117, November 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blazic:1998:FDH

[1765] Novica Blažić, Neda Bokan, and Zoran Rakić. Foliation of a dynamically homogeneous neutral manifold. Journal of Mathematical Physics, 39(11): 6118–6124, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Byrd:1998:DGA

[1766] Mark Byrd. Differential geometry on SU (3 ) with applications to three state systems. Journal of Mathematical Physics, 39(11):6125–6136, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2262].

Giannoni:1998:ETR

[1767] Fabio Giannoni and Paolo Piccione. An existence theory for relativistic brachistochrones in stationary space–times. Journal of Mathematical Physics, 39(11):6137–6152, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grensing:1998:MRS

[1768] G. Grensing. Modified real structures for quantum groups and applications. Journal of Mathematical Physics, 39(11):6153–6161, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hernandez-Bermejo:1998:HSD

[1769] Benito Hernández-Bermejo and Víctor Fairén. Hamiltonian structure and Darboux theorem for families of generalized Lotka–Volterra systems. Journal of Mathematical Physics, 39(11):6162–6174, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Komori:1998:COG

[1770] Yasushi Komori and Kazuhiro Hikami. Conserved operators of the generalized elliptic Ruijsenaars models. Journal of Mathematical Physics, 39(11):6175–6190, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Paschke:1998:DST

[1771] Mario Paschke and Andrzej Sitarz. Discrete spectral triples and their symmetries. Journal of Mathematical Physics, 39(11):6191–6205, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Silva:1998:TFY

[1772] Ronaldo Rodrigues Silva. The trace formulas yield the inverse metric formula. Journal of Mathematical Physics, 39(11):6206–6213, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1998:BRR

[1773] D. J. Rowe and J. Repka. Branching rules for restriction of the Weil representations of Sp(n, R) to its maximal parabolic subgroup CM (n). Journal of Mathematical Physics, 39(11):6214–6224, November 1998. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruan:1998:AES

[1774] Dong Ruan and Hong zhou Sun. Algebraic expressions for some SO(N ) Racah coefficients. Journal of Mathematical Physics, 39(11):6225–6240, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kota:1998:EAS

[1775] J. Kóta and Z. Perjés. Erratum: “All stationary vacuum metrics with shearing geodesic eigenrays”, [J. Math. Phys. 13, 1695 (1972)]. Journal of Mathematical Physics, 39(11):6241, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [4].

Mudrov:1998:EQG

[1776] A. I. Mudrov. Erratum: “On quantization of Z_{2}-graded algebras”, [J. Math. Phys. 38, 476–483 (1997)]. Journal of Mathematical Physics, 39 (11):6242, November 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [964].

Atakishiyev:1998:WDF

[1777] Natig M. Atakishiyev, Sergey M. Chumakov, and Kurt Bernardo Wolf. Wigner distribution function for finite systems. Journal of Mathematical Physics, 39(12):6247–6261, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feldmann:1998:RSM

[1778] D. Matthew Feldmann, Paul J. Pelzl, and Frederick W. King. Resolution of some mathematical problems arising in the relativistic treatment of the S states of three-electron systems. Journal of Mathematical Physics, 39(12):6262–6275, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finster:1998:LSR

[1779] Felix Finster. Local U (2,2) symmetry in relativistic quantum mechanics. Journal of Mathematical Physics, 39(12):6276–6290, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gielerak:1998:SPS

[1780] R. Gielerak, L. Jakóbczyk, and R. Olkiewicz. Stochastically positive structures on Weyl algebras. The case of quasi-free states. Journal of Mathematical Physics, 39(12):6291–6328, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Golodets:1998:GSA

[1781] Valentin Ya. Golodets and Sergey V. Neshveyev. Gibbs states for AF algebras. Journal of Mathematical Physics, 39(12):6329–6344, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:1998:MEG

[1782] Richard L. Hall, Nasser Saad, and Attila B. von Keviczky. Matrix elements for a generalized spiked harmonic oscillator. Journal of Mathematical Physics, 39(12):6345–6352, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kunhardt:1998:ILS

[1783] Walter Kunhardt. On infravacua and the localization of sectors. Journal of Mathematical Physics, 39(12):6353–6363, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lahti:1998:COE

[1784] Pekka Lahti, Sylvia Pulmannova, and Kari Ylinen. Coexistent observables and effects in a convexity approach. Journal of Mathematical Physics, 39(12):6364–6371, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Landsman:1998:SQC

[1785] N. P. Landsman. Strict quantization of coadjoint orbits. Journal of Mathematical Physics, 39(12):6372–6383, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lesimple:1998:DMG

[1786] Marc Lesimple. Deformations of massless Gupta–Bleuler triplets in 3 + 2 de Sitter space. Journal of Mathematical Physics, 39(12):6384–6391, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nerses:1998:SET

[1787] A. Nerses and E. E. Kunhardt. Solitary excitations of a two-dimensional electron gas. Journal of Mathematical Physics, 39(12):6392–6402, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scutaru:1998:TPB

[1788] H. Scutaru. Transition probabilities between quasifree states. Journal of Mathematical Physics, 39(12):6403–6415, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tjernberg:1998:SAS

[1789] O. Tjernberg. Symmetry adapted states for Hubbard clusters. Journal of Mathematical Physics, 39(12):6416–6423, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanLeeuwen:1998:HOL

[1790] Hans van Leeuwen and Hans Maassen. The q-harmonic oscillator in a lattice model. Journal of Mathematical Physics, 39(12):6424–6440, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yang:1998:PCS

[1791] Yaping Yang, Zhixin Lin, Shuangyuan Xie, Weiguo Feng, and Xiang Wu. q-phase-coherent states and their squeezing properties. Journal of Mathematical Physics, 39(12):6441–6453, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhdanov:1998:SSM

[1792] Renat Zhdanov and Maxim Lutfullin. On separable Schrödinger–Maxwell equations. Journal of Mathematical Physics, 39(12):6454–6458, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1998:SRC

[1793] Goong Chen, Sze-Bi Hsu, and Jianxin Zhou. Snapback repellers as a cause of chaotic vibration of the wave equation with a van der Pol boundary condition and energy injection at the middle of the span. Journal of Mathematical Physics, 39(12):6459–6489, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borland:1998:IGW

[1794] Lisa Borland, Angel R. Plastino, and Constantino Tsallis. Information gain within nonextensive thermostatistics. Journal of Mathematical Physics, 39(12):6490–6501, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1958].

Brody:1998:QCE

[1795] Dorje C. Brody and Lane P. Hughston. The quantum canonical ensemble. Journal of Mathematical Physics, 39(12):6502–6508, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Choi:1998:DFD

[1796] Veni Choi, Yong Moon Park, and Hyun Jae Yoo. Dirichlet forms and Dirichlet operators for infinite particle systems: Essential selfadjointness. Journal of Mathematical Physics, 39(12):6509–6536, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diacu:1998:CGP

[1797] Florin Diacu and Dan Şelaru. Chaos in the Gyldén problem. Journal of Mathematical Physics, 39(12):6537–6546, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maccari:1998:GHE

[1798] Attilio Maccari. A generalized Hirota equation in 2+1 dimensions. Journal of Mathematical Physics, 39(12):6547–6551, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pelinovsky:1998:BNE

[1799] Dmitry E. Pelinovsky and Catherine Sulem. Bifurcations of new eigenvalues for the Benjamin–Ono equation. Journal of Mathematical Physics, 39(12):6552–6572, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gen:1998:TIA

[1800] Uchida Gen and Tetsuya Shiromizu. Timelike infinity and asymptotic symmetry. Journal of Mathematical Physics, 39(12):6573–6592, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gionti:1998:SQG

[1801] Gabriele Gionti. Simplicial quantum gravity in the elongated phase. Journal of Mathematical Physics, 39(12):6593–6602, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giulini:1998:TGA

[1802] Domenico Giulini. Is there a general area theorem for black holes? Journal of Mathematical Physics, 39(12):6603–6606, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Groah:1998:TIB

[1803] Jeffrey M. Groah. Time-independent bounds on solutions with shocks to the relativistic Euler equations on conformally flat space–times with curvature singularities. Journal of Mathematical Physics, 39(12):6607– 6630, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moreschi:1998:RFS

[1804] Osvaldo M. Moreschi and Sergio Dain. Rest frame system for asymptotically flat space–times. Journal of Mathematical Physics, 39(12):6631– 6650, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2225].

Zannias:1998:SBH

[1805] T. Zannias. On stationary black holes of the Einstein conformally invariant scalar system. Journal of Mathematical Physics, 39(12):6651–6667, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bobenko:1998:IDE

[1806] A. I. Bobenko, B. Lorbeer, and Yu. B. Suris. Integrable discretizations of the Euler top. Journal of Mathematical Physics, 39(12):6668–6683, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:1998:HMG

[1807] Qing Ding. On harmonic maps into gauge groups. Journal of Mathematical Physics, 39(12):6684–6695, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Duan:1998:TSS

[1808] Yishi Duan and Sheng Li. The topological structure of the space–time disclination. Journal of Mathematical Physics, 39(12):6696–6705, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frontini:1998:MEG

[1809] M. Frontini and A. Tagliani. Maximum entropy in the generalized moment problem. Journal of Mathematical Physics, 39(12):6706–6714, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Magnus:1998:UNI

[1810] W. Magnus and W. Schoenmaker. On the use of a new integral theorem for the quantum mechanical treatment of electric circuits. Journal of Mathematical Physics, 39(12):6715–6719, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McCarthy:1998:SPE

[1811] J. E. McCarthy, E. Yu. Backhaus, and J. Fajans. Solving Poisson’s equation with interior conditions. Journal of Mathematical Physics, 39 (12):6720–6729, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosengren:1998:TSF

[1812] Hjalmar Rosengren. On the triple sum formula for Wigner 9j-symbols. Journal of Mathematical Physics, 39(12):6730–6744, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhdanov:1998:LSI

[1813] R. Z. Zhdanov. Lie symmetry and integrability of ordinary differential equations. Journal of Mathematical Physics, 39(12):6745–6756, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1998:SPA

[1814] Anonymous. Summary of the Physics and Astronomy Classification Scheme — 1996 (as amended for 1998). Journal of Mathematical Physics, 39(12):6757–??, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1998:PHU

[1815] Anonymous. PACS headings used in the present index. Journal of Mathematical Physics, 39(12):6758–??, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1998:APT

[1816] Anonymous. Announcement: Probabilistic techniques in equilibrium and nonequilibrium statistical physics. Journal of Mathematical Physics, 39 (12):I–??, December 1998. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boillat:1999:BIE

[1817] Guy Boillat and Alberto Strumia. On the Born–Infeld electron: Spin effects. Journal of Mathematical Physics, 40(1):1–11, January 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brody:1999:TQS

[1818] Dorje C. Brody and Lane P. Hughston. Thermalization of quantum states. Journal of Mathematical Physics, 40(1):12–18, January 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chae:1999:ENC

[1819] Dongho Chae, Jongmin Han, and Oleg Yu. Imanuvilov. Existence and nonexistence in Chern–Simons–Higgs theory with a constant electric charge density. Journal of Mathematical Physics, 40(1):19–34, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Damak:1999:STD

[1820] Mondher Damak. On the spectral theory of dispersive N-body Hamiltonians. Journal of Mathematical Physics, 40(1):35–48, January 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Doebner:1999:GTQ

[1821] H.-D. Doebner, G. A. Goldin, and P. Nattermann. Gauge transformations in quantum mechanics and the unification of nonlinear Schrödinger equations. Journal of Mathematical Physics, 40(1):49–63, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loo:1999:RRT

[1822] Ken Loo. A rigorous real time Feynman path integral. Journal of Mathematical Physics, 40(1):64–70, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mattes:1999:SOM

[1823] M. Mattes and M. Sorg. Second-order mixtures in relativistic Schrödinger theory. Journal of Mathematical Physics, 40(1):71–92, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pletyukhov:1999:OKP

[1824] M. V. Pletyukhov and E. A. Tolkachev. 8 D oscillator and 5 D Kepler problem: The case of nontrivial constraints. Journal of Mathematical Physics, 40(1):93–100, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Robinett:1999:IVN

[1825] R. W. Robinett. Isolated versus nonisolated periodic orbits in variants of the two-dimensional square and circular billiards. Journal of Mathematical Physics, 40(1):101–122, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanWinter:1999:COF

[1826] Clasine van Winter. Chaotic observables for a free quantum particle. Journal of Mathematical Physics, 40(1):123–139, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Weber:1999:NPI

[1827] T. A. Weber. Nonlocal potentials, isolated states, and Levinson’s theorem. Journal of Mathematical Physics, 40(1):140–149, January 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

Xia:1999:EGS

[1828] Jingbo Xia. An estimate of the ground state energy of the fractional quantum Hall effect. Journal of Mathematical Physics, 40(1):150–155, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Skaropoulos:1999:RAS

[1829] Nikolaos C. Skaropoulos and Dimitrios P. Chrissoulidis. Rigorous application of the stochastic functional method to plane-wave scattering from a random cylindrical surface. Journal of Mathematical Physics, 40(1): 156–168, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teixeira:1999:LET

[1830] F. L. Teixeira and W. C. Chew. Lattice electromagnetic theory from a topological viewpoint. Journal of Mathematical Physics, 40(1):169–187, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calzada:1999:POG

[1831] Juan A. Calzada, Mariano A. del Olmo, and Miguel A. Rodríguez. Pseudo-orthogonal groups and integrable dynamical systems in two dimensions. Journal of Mathematical Physics, 40(1):188–209, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Damianou:1999:SHS

[1832] P. A. Damianou and C. Sophocleous. Symmetries of Hamiltonian systems with two degrees of freedom. Journal of Mathematical Physics, 40(1): 210–235, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:1999:SCM

[1833] Manuel F. Rañada. Superintegrability of the Calogero–Moser system: Constants of motion, master symmetries, and time-dependent symmetries. Journal of Mathematical Physics, 40(1):236–247, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruan:1999:SDI

[1834] Hang-Yu Ruan and Yi-Xin Chen. The study of dromion interactions of (2 + 1)-dimensional integrable systems. Journal of Mathematical Physics, 40(1):248–255, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rylov:1999:SWF

[1835] Yuri A. Rylov. Spin and wave function as attributes of ideal fluid. Journal of Mathematical Physics, 40(1):256–278, January 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tsiganov:1999:SSA

[1836] A. V. Tsiganov. The Stäckel systems and algebraic curves. Journal of Mathematical Physics, 40(1):279–298, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ward:1999:LPI

[1837] R. S. Ward. Lax pairs for integrable lattice systems. Journal of Mathematical Physics, 40(1):299–308, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Glass:1999:SPE

[1838] E. N. Glass and Jonathan Kress. Solutions of Penrose’s equation. Journal of Mathematical Physics, 40(1):309–317, January 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hajicek:1999:BHI

[1839] Petr Hájíček. Black hole interacting with matter as a simple dynamical system. Journal of Mathematical Physics, 40(1):318–339, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Isenberg:1999:ABP

[1840] James Isenberg and Satyanad Kichenassamy. Asymptotic behavior in polarized T^{2}-symmetric vacuum space–times. Journal of Mathematical Physics, 40(1):340–352, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jantzen:1999:KRA

[1841] R. T. Jantzen and C. Uggla. The kinematical role of automorphisms in the orthonormal frame approach to Bianchi cosmology. Journal of Mathematical Physics, 40(1):353–368, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Biswas:1999:ILT

[1842] T. Biswas and Satish D. Joglekar. Inverse Laplace transform and perturbation theory. Journal of Mathematical Physics, 40(1):369–382, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:1999:EEF

[1843] Simonetta Frittelli, Ezra T. Newman, and Gilberto Silva-Ortigoza. The Eikonal equation in flat space: Null surfaces and their singularities. I. Journal of Mathematical Physics, 40(1):383–407, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fuchs:1999:SSF

[1844] J. Fuchs, A. Ch. Ganchev, K. Szlachányi, and P. Vecsernyés. S_{4} symmetry of 6-j symbols and Frobenius–Schur indicators in rigid monoidal C^{∗} categories. Journal of Mathematical Physics, 40(1):408–426, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hadjiivanov:1999:HAP

[1845] L. K. Hadjiivanov, A. P. Isaev, O. V. Ogievetsky, P. N. Pyatov, and I. T. Todorov. Hecke algebraic properties of dynamical R-matrices. Application to related quantum matrix algebras. Journal of Mathematical Physics, 40(1):427–448, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lastaria:1999:LSR

[1846] Federico G. Lastaria. Lie subalgebras of real and complex orthogonal groups in dimension four. Journal of Mathematical Physics, 40(1):449– 453, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leandre:1999:SSB

[1847] Rémi Léandre. String structure over the Brownian bridge. Journal of Mathematical Physics, 40(1):454–479, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Longhi:1999:CRB

[1848] G. Longhi and M. Materassi. A canonical realization of the BMS algebra. Journal of Mathematical Physics, 40(1):480–500, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pati:1999:IAI

[1849] K. C. Pati, D. Parashar, and R. S. Kaushal. Involutive automorphisms and Iwasawa decomposition of some hyperbolic Kac–Moody algebras. Journal of Mathematical Physics, 40(1):501–510, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pravica:1999:PED

[1850] David W. Pravica. Propagation estimates for dispersive wave equations: Application to the stratified wave equation. Journal of Mathematical Physics, 40(1):511–527, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Stoyanov:1999:PLS

[1851] Ognyan Stoyanov. Poisson–Lie structures on infinite-dimensional jet groups and quantum groups related to them. Journal of Mathematical Physics, 40(1):528–582, January 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amorim:1999:FVC

[1852] R. Amorim and J. Barcelos-Neto. Functional versus canonical quantization of a nonlocal massive vector-gauge theory. Journal of Mathematical Physics, 40(2):585–600, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borchers:1999:MGQ

[1853] H. J. Borchers and J. Yngvason. Modular groups of quantum fields in thermal states. Journal of Mathematical Physics, 40(2):601–624, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Burgbacher:1999:TSH

[1854] Frank Burgbacher, Claus Lämmerzahl, and Alfredo Macias. Is there a stable hydrogen atom in higher dimensions? Journal of Mathematical Physics, 40(2):625–634, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chakrabarti:1999:IIM

[1855] A. Chakrabarti. Implementation of an iterative map in the construction of (quasi)periodic instantons: Chaotic aspects and discontinuous rotation numbers. Journal of Mathematical Physics, 40(2):635–673, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geyer:1999:OCLa

[1856] B. Geyer, P. M. Lavrov, and D. Mülsch. The osp(1,2)-covariant Lagrangian quantization of irreducible massive gauge theories. Journal of Mathematical Physics, 40(2):674–698, February 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:1999:CIE

[1857] Richard L. Hall. Constructive inversion of energy trajectories in quantum mechanics. Journal of Mathematical Physics, 40(2):699–707, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:1999:STDa

[1858] E. G. Kalnins, G. C. Williams, W. Miller, Jr., and G. S. Pogosyan. Superintegrability in three-dimensional Euclidean space. Journal of Mathematical Physics, 40(2):708–725, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Links:1999:BAS

[1859] Jon Links, Angela Foerster, and Michael Karowski. Bethe ansatz solution of a closed spin 1 XXZ Heisenberg chain with quantum algebra symmetry. Journal of Mathematical Physics, 40(2):726–735, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Manton:1999:DS

[1860] N. S. Manton. Deconstructing supersymmetry. Journal of Mathematical Physics, 40(2):736–750, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miller:1999:SQG

[1861] Steven D. Miller. Stochastic quantum geometry within a (1+1) manifold: A basic construction. Journal of Mathematical Physics, 40(2):751–765, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morales:1999:RCS

[1862] Jorge A. Morales, Erik Deumens, and Yngve Öhrn. On rotational coherent states in molecular quantum dynamics. Journal of Mathematical Physics, 40(2):766–786, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wulkenhaar:1999:GTG

[1863] Raimar Wulkenhaar. Gauge theories with graded differential Lie algebras. Journal of Mathematical Physics, 40(2):787–794, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cantrijn:1999:RCS

[1864] Frans Cantrijn, Manuel de León, Juan Carlos Marrero, and David Martín de Diego. Reduction of constrained systems with symmetries. Journal of Mathematical Physics, 40(2):795–820, February 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lopez:1999:GIV

[1865] M. Castrillón López and J. Muñoz Masqué. Gauge-invariant variationally trivial problems on t^{∗}m. Journal of Mathematical Physics, 40(2):821– 829, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chen:1999:GST

[1866] M. Chen. On the geometric structure of thermodynamics. Journal of Mathematical Physics, 40(2):830–837, February 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Miyazawa:1999:GFF

[1867] Toru Miyazawa. Green’s function of the Fokker–Planck equation: General formula of frequency expansion. Journal of Mathematical Physics, 40 (2):838–856, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kouranbaeva:1999:CHE

[1868] Shinar Kouranbaeva. The Camassa–Holm equation as a geodesic flow on the diffeomorphism group. Journal of Mathematical Physics, 40(2): 857–868, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marchioro:1999:LSR

[1869] Carlo Marchioro. Large smoke rings with concentrated vorticity. Journal of Mathematical Physics, 40(2):869–883, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Porubov:1999:EPS

[1870] A. V. Porubov and M. G. Velarde. Exact periodic solutions of the complex Ginzburg–Landau equation. Journal of Mathematical Physics, 40 (2):884–896, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2176] and response [2177].

Rainer:1999:SRC

[1871] A. Rainer and H. Stephani. Similarity reduction for a class of algebraically special perfect fluids. Journal of Mathematical Physics, 40(2): 897–908, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brodbeck:1999:EEA

[1872] Othmar Brodbeck, Simonetta Frittelli, Peter Hübner, and Oscar A. Reula. Einstein’s equations with asymptotically stable constraint propagation. Journal of Mathematical Physics, 40(2):909–923, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bronnikov:1999:GBS

[1873] K. A. Bronnikov. Gravitating brane systems: Some general theorems. Journal of Mathematical Physics, 40(2):924–938, February 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giachetta:1999:NRD

[1874] Giovanni Giachetta. Nonlinear realizations of the diffeomorphism group in metric-affine gauge theory of gravity. Journal of Mathematical Physics, 40(2):939–954, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gladush:1999:SSG

[1875] V. D. Gladush and R. A. Konoplya. Split structures in general relativity and the Kaluza–Klein theories. Journal of Mathematical Physics, 40(2): 955–979, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nollert:1999:QEQ

[1876] Hans-Peter Nollert and Richard H. Price. Quantifying excitations of quasinormal mode systems. Journal of Mathematical Physics, 40(2):980– 1010, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barnea:1999:RMC

[1877] Nir Barnea. A recursive method for the construction of irreducible representations of the orthogonal group O(n). Journal of Mathematical Physics, 40(2):1011–1022, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brandt:1999:DGS

[1878] Friedemann Brandt. Deformations of global symmetries in the extended antifield formalism. Journal of Mathematical Physics, 40(2):1023–1040, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:1999:EEA

[1879] Simonetta Frittelli, Ezra T. Newman, and Gilberto Silva-Ortigoza. The Eikonal equation in asymptotically flat space–times. Journal of Mathematical Physics, 40(2):1041–1056, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garbaczewski:1999:CNA

[1880] Piotr Garbaczewski and Robert Olkiewicz. Cauchy noise and affiliated stochastic processes. Journal of Mathematical Physics, 40(2):1057–1073, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Green:1999:HDQ

[1881] Gavin Green. Howe duality for the quantum groups U_{⨿}u(m, n), U_{⨿}u(M). Journal of Mathematical Physics, 40(2):1074–1086, February 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Keller:1999:STP

[1882] Joseph B. Keller. Singularities at the tip of a plane angular sector. Journal of Mathematical Physics, 40(2):1087–1092, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Newman:1999:CSD

[1883] Ezra T. Newman and Alejandro Perez. Characteristic surface data for the eikonal equation. Journal of Mathematical Physics, 40(2):1093–1102, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schleef:1999:AGP

[1884] David Schleef, Michelle Parry, Shu-Ju Tu, Brian Woodahl, and Ephraim Fischbach. Application of geometric probability techniques to the evaluation of interaction energies arising from a general radial potential. Journal of Mathematical Physics, 40(2):1103–1112, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:1999:EDC

[1885] Simonetta Frittelli, Carlos N. Kozameh, and Erza T. Newman. Erratum: “On the dynamics of characteristic surfaces”, [J. Math. Phys. 36, 6397–6416 (1995)]. Journal of Mathematical Physics, 40(2):1113, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [442].

Frittelli:1999:ELM

[1886] Simonetta Frittelli, Carlos N. Kozameh, and Ezra T. Newman. Erratum: “Lorentzian metrics from characteristic surfaces”, [J. Math. Phys. 36, 4975–4983 (1995)]. Journal of Mathematical Physics, 40(2):1114, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [359].

Frittelli:1999:EGC

[1887] Simonetta Frittelli, Carlos N. Kozameh, and Ezra T. Newman. Erratum: “GR via characteristic surfaces”, [J. Math. Phys. 36, 4984–5004 (1995)]. Journal of Mathematical Physics, 40(2):1115, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [360].

Frittelli:1999:ELE

[1888] Simonetta Frittelli, Carlos N. Kozameh, and Ezra T. Newman. Erratum: “Linearized Einstein theory via null surfaces”, [J. Math. Phys. 36, 5005–5022 (1995)]. Journal of Mathematical Physics, 40(2):1116, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [361].

Brau:1999:ASR

[1889] F. Brau. Analytical solution of the relativistic Coulomb problem with a hard core interaction for a one-dimensional spinless Salpeter equation. Journal of Mathematical Physics, 40(3):1119–1126, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2224].

Brevik:1999:NSP

[1890] I. Brevik, E. Elizalde, R. Sollie, and J. B. Aarseth. A new scaling property of the Casimir energy for a piecewise uniform string. Journal of Mathematical Physics, 40(3):1127–1135, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brihaye:1999:DSL

[1891] Y. Brihaye, B. Kleihaus, and D. H. Tchrakian. Dyon–Skyrmion lumps. Journal of Mathematical Physics, 40(3):1136–1152, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:1999:MVS

[1892] Shijin Ding, Boling Guo, and Fengqiu Su. Measure-valued solution to the strongly degenerate compressible Heisenberg chain equations. Journal of Mathematical Physics, 40(3):1153–1162, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dudek:1999:RWP

[1893] Mariusz Dudek, Stefan Giller, and Piotr Milczarski. Rectangular well as perturbation. Journal of Mathematical Physics, 40(3):1163–1179, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dunne:1999:BIF

[1894] R. S. Dunne. A braided interpretation of fractional supersymmetry in higher dimensions. Journal of Mathematical Physics, 40(3):1180–1196, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrmann:1999:SPH

[1895] D. J. L. Herrmann and T. Janssen. On spectral properties of Harperlike models. Journal of Mathematical Physics, 40(3):1197–1214, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiroshima:1999:WCL

[1896] Fumio Hiroshima. Weak coupling limit and removing an ultraviolet cutoff for a Hamiltonian of particles interacting with a quantized scalar field. Journal of Mathematical Physics, 40(3):1215–1236, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kudaka:1999:UPP

[1897] Shoju Kudaka and Shuichi Matsumoto. Uncertainty principle for proper time and mass. Journal of Mathematical Physics, 40(3):1237–1245, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comments [2268, 2269].

Lin:1999:PIR

[1898] De-Hone Lin. Path integral for the relativistic three-dimensional Aharonov–Bohm–Coulomb system. Journal of Mathematical Physics, 40 (3):1246–1254, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McInnes:1999:MSS

[1899] Brett McInnes. Metric symmetries and spin asymmetries of Ricci-flat Riemannian manifolds. Journal of Mathematical Physics, 40(3):1255– 1267, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Michoel:1999:NBE

[1900] T. Michoel and A. Verbeure. Nonextensive Bose–Einstein condensation model. Journal of Mathematical Physics, 40(3):1268–1279, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Niederle:1999:ESS

[1901] J. Niederle and A. G. Nikitin. Extended supersymmetries for the Schrödinger–Pauli equation. Journal of Mathematical Physics, 40(3): 1280–1293, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Nouri:1999:DWD

[1902] Saeid Nouri. Determination of Wigner distribution function for the ddimensional Coulomb problem. Journal of Mathematical Physics, 40(3): 1294–1299, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Olkiewicz:1999:DSI

[1903] R. Olkiewicz. Dynamical semigroups for interacting quantum and classical systems. Journal of Mathematical Physics, 40(3):1300–1316, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Radulescu:1999:APS

[1904] O. Radulescu and T. Janssen. Arithmetic properties of spectra produced by Farey hierarchies of approximants. Journal of Mathematical Physics, 40(3):1317–1333, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rybkin:1999:TFB

[1905] Alexei Rybkin. On a trace formula of the Buslaev–Faddeev type for a long-range potential. Journal of Mathematical Physics, 40(3):1334–1343, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schlesinger:1999:TQM

[1906] Karl-Georg Schlesinger. Toward quantum mathematics. I. from quantum set theory to universal quantum mechanics. Journal of Mathematical Physics, 40(3):1344–1358, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Craig:1999:AMP

[1907] Scott Craig, Florin Diacu, Ernesto A. Lacomba, and Ernesto Perez. On the anisotropic Manev problem. Journal of Mathematical Physics, 40(3): 1359–1375, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giachetta:1999:NCT

[1908] G. Giachetta, L. Mangiarotti, and G. Sardanashvily. Nonholonomic constraints in time-dependent mechanics. Journal of Mathematical Physics, 40(3):1376–1390, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Marin:1999:EET

[1909] Marin Marin. An evolutionary equation in thermoelasticity of dipolar bodies. Journal of Mathematical Physics, 40(3):1391–1399, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Prodan:1999:TMS

[1910] E. Prodan. Transfer matrices for scalar fields on curved spaces. Journal of Mathematical Physics, 40(3):1400–1405, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Estevez:1999:DTS

[1911] P. G. Estévez. Darboux transformation and solutions for an equation in 2 + 1 dimensions. Journal of Mathematical Physics, 40(3):1406–1419, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loris:1999:SRB

[1912] Ignace Loris and Ralph Willox. Symmetry reductions of the BKP hierarchy. Journal of Mathematical Physics, 40(3):1420–1431, March 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shengfan:1999:DGAa

[1913] Zhou Shengfan. On dimension of the global attractor for damped nonlinear wave equations. Journal of Mathematical Physics, 40(3):1432–1438, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shirasaki:1999:ASF

[1914] Ryōen Shirasaki and Kaoru Iwano. Analytical study of fractionally charged solitons in a one-dimensional trimerized electron–phonon system. Journal of Mathematical Physics, 40(3):1439–1465, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:1999:CCS

[1915] G. S. Hall and M. S. Capocci. Classification and conformal symmetry in three-dimensional space–times. Journal of Mathematical Physics, 40(3): 1466–1478, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kunzinger:1999:RSC

[1916] M. Kunzinger and R. Steinbauer. A rigorous solution concept for geodesic and geodesic deviation equations in impulsive gravitational waves. Journal of Mathematical Physics, 40(3):1479–1489, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maluf:1999:SBE

[1917] J. W. Maluf and J. F. da Rocha-Neto. Static Bondi energy in the teleparallel equivalent of general relativity. Journal of Mathematical Physics, 40 (3):1490–1503, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Montesinos:1999:RSN

[1918] Merced Montesinos, Hugo A. Morales-Técotl, Luis F. Urrutia, and J. David Vergara. Real sector of the nonminimally coupled scalar field to self-dual gravity. Journal of Mathematical Physics, 40(3):1504–1517, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dimakis:1999:DRG

[1919] Aristophanes Dimakis and Folkert Müller-Hoissen. Discrete Riemannian geometry. Journal of Mathematical Physics, 40(3):1518–1548, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Izmestev:1999:CLA

[1920] A. A. Izmest’ev, G. S. Pogosyan, A. N. Sissakian, and P. Winternitz. Contractions of Lie algebras and separation of variables. The n-dimensional sphere. Journal of Mathematical Physics, 40(3):1549–1573, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palev:1999:HWI

[1921] T. D. Palev and N. I. Stoilova. Highest weight irreducible representations of the Lie superalgebra gl(1 |∞). Journal of Mathematical Physics, 40 (3):1574–1594, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruijsenaars:1999:GLFa

[1922] S. N. M. Ruijsenaars. Generalized Lamé functions. I. the elliptic case. Journal of Mathematical Physics, 40(3):1595–1626, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ruijsenaars:1999:GLFb

[1923] S. N. M. Ruijsenaars. Generalized Lamé functions. II. hyperbolic and trigonometric specializations. Journal of Mathematical Physics, 40(3): 1627–1663, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sontz:1999:RKS

[1924] Stephen Bruce Sontz. On the reproducing kernel of the Segal–Bargmann space. Journal of Mathematical Physics, 40(3):1664–1676, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sontz:1999:RLS

[1925] Stephen Bruce Sontz. A reverse log-Sobolev inequality in the Segal– Bargmann space. Journal of Mathematical Physics, 40(3):1677–1695, March 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alonso:1999:NAS

[1926] M. A. Alonso and G. W. Forbes. New approach to semiclassical analysis in mechanics. Journal of Mathematical Physics, 40(4):1699–1718, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baumann:1999:BBF

[1927] Klaus Baumann. Bounded Bose fields in 1 + 1 dimensions commuting for space- and timelike distances. Journal of Mathematical Physics, 40 (4):1719–1737, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cervero:1999:FAN

[1928] José M. Cerveró and Juan D. Lejarreta. The Floquet analysis and noninteger higher harmonics generation. Journal of Mathematical Physics, 40 (4):1738–1755, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chadan:1999:GBS

[1929] Khosrow Chadan, Reido Kobayashi, and Monique Lassaut. Generalization of the Birman–Schwinger method for the number of bound states. Journal of Mathematical Physics, 40(4):1756–1763, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Froman:1999:ELQ

[1930] Per Olof Fröman, Kilian Larsson, and Anders Hökback. Energy levels of a quantal particle enclosed in N identical, mirror symmetric wells of a periodic potential. Numerical test of phase-integral formulas. Journal of Mathematical Physics, 40(4):1764–1779, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Griesemer:1999:IPR

[1931] Marcel Griesemer and Christian Tix. Instability of a pseudo-relativistic model of matter with self-generated magnetic field. Journal of Mathematical Physics, 40(4):1780–1791, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grigoriev:1999:GSM

[1932] M. A. Grigoriev, A. M. Semikhatov, and I. Yu. Tipunin. Gauge symmetries of the master action in the Batalin–Vilkovisky formalism. Journal of Mathematical Physics, 40(4):1792–1806, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karki:1999:PIL

[1933] Topi Kärki. On path integral localization and the Laplacian. Journal of Mathematical Physics, 40(4):1807–1830, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Korff:1999:TPS

[1934] C. Korff, G. Lang, and R. Schrader. Two-particle scattering theory for anyons. Journal of Mathematical Physics, 40(4):1831–1869, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lee:1999:PBN

[1935] C.-W. H. Lee and S. G. Rajeev. Poisson brackets of normal-ordered Wilson loops. Journal of Mathematical Physics, 40(4):1870–1890, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:1999:DES

[1936] Yūji Ohta. Differential equations for scaling relation in N = 2 supersymmetric SU (2 ) Yang–Mills theory coupled with massive hypermultiplet. Journal of Mathematical Physics, 40(4):1891–1900, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Teufel:1999:FAS

[1937] S. Teufel, D. Dürr, and K. Münch-Berndl. The flux-across-surfaces theorem for short range potentials and wave functions without energy cutoffs. Journal of Mathematical Physics, 40(4):1901–1922, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Perng:1999:CQS

[1938] Shyan-Ming Perng. On conserved quantities at spatial infinity. Journal of Mathematical Physics, 40(4):1923–1950, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ayari:1999:GIS

[1939] M. A. Ayari, V. Hussin, and P. Winternitz. Group invariant solutions for the N = 2 super Korteweg–de Vries equation. Journal of Mathematical Physics, 40(4):1951–1965, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dorren:1999:INP

[1940] H. J. S. Dorren. On the integrability of nonlinear partial differential equations. Journal of Mathematical Physics, 40(4):1966–1976, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Eleonsky:1999:IPS

[1941] V. M. Eleonsky and V. G. Korolev. Isospectral problem for Schrödinger operator: Evolutional viewpoint. Journal of Mathematical Physics, 40 (4):1977–1992, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ghosh:1999:SSL

[1942] Sasanka Ghosh, Anjan Kundu, and Sudipta Nandy. Soliton solutions, Liouville integrability and gauge equivalence of Sasa Satsuma equation. Journal of Mathematical Physics, 40(4):1993–2000, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:1999:HBA

[1943] Xing-Biao Hu, Yong-Tang Wu, and Xian-Guo Geng. Hirota bilinear approach to a new integrable differential-difference system. Journal of Mathematical Physics, 40(4):2001–2010, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2103].

Vakhnenko:1999:HFS

[1944] V. O. Vakhnenko. High-frequency soliton-like waves in a relaxing medium. Journal of Mathematical Physics, 40(4):2011–2020, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gen:1999:ASS

[1945] Uchida Gen and Tetsuya Shiromizu. Asymptotically Schwarzschild space–times. Journal of Mathematical Physics, 40(4):2021–2031, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guilfoyle:1999:WTF

[1946] Brendan S. Guilfoyle. Weyl-type fields with geodesic lines of force. Journal of Mathematical Physics, 40(4):2032–2045, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reisenberger:1999:RSN

[1947] Michael P. Reisenberger. On relativistic spin network vertices. Journal of Mathematical Physics, 40(4):2046–2054, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Agacy:1999:GKP

[1948] R. L. Agacy. Generalized Kronecker and permanent deltas, their spinor and tensor equivalents and applications. Journal of Mathematical Physics, 40(4):2055–2063, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bonanos:1999:ATS

[1949] S. Bonanos. The algebra of two symmetric matrices: Proving completeness and deriving syzygies for a set of invariants of the Riemann tensor. Journal of Mathematical Physics, 40(4):2064–2073, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cairo:1999:DMS

[1950] Laurent Cairó, Marc R. Feix, and Jaume Llibre. Darboux method and search of invariants for the Lotka–Volterra and complex quadratic systems. Journal of Mathematical Physics, 40(4):2074–2091, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Conte:1999:BTP

[1951] Robert Conte, Micheline Musette, and A. Michel Grundland. Bäcklund transformation of partial differential equations from the Painlevé– Gambier classification. II. Tzitzéica equation. Journal of Mathematical Physics, 40(4):2092–2106, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gotay:1999:GVH

[1952] Mark J. Gotay. On the Groenewold–Van Hove problem for R^{2}n. Journal of Mathematical Physics, 40(4):2107–2116, April 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Haine:1999:JPE

[1953] Luc Haine and Jean-Pierre Semengue. The Jacobi polynomial ensemble and the Painlevé VI equation. Journal of Mathematical Physics, 40(4): 2117–2134, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Havlicek:1999:RCS

[1954] M. Havlíček, A. U. Klimyk, and S. Pošta. Representations of the cyclically symmetric q-deformed algebra so_{q}(3). Journal of Mathematical Physics, 40(4):2135–2161, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Heim:1999:ASH

[1955] Thomas A. Heim and Dmitry Green. Alternative sets of hyperspherical harmonics: Satisfying cusp conditions through frame transformations. Journal of Mathematical Physics, 40(4):2162–2180, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lahti:1999:OIP

[1956] Pekka Lahti, Juha-Pekka Pellonpää, and Kari Ylinen. Operator integrals and phase space observables. Journal of Mathematical Physics, 40(4): 2181–2189, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Raghavan:1999:VBR

[1957] K. N. Raghavan and P. Sankaran. On Verma bases for representations of sl(n, C). Journal of Mathematical Physics, 40(4):2190–2195, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2264].

Borland:1999:EIG

[1958] L. Borland, A. R. Plastino, and C. Tsallis. Erratum: “Information gain within generalized thermostatistics”, [J. Math. Phys. 39, 6490 (1998)]. Journal of Mathematical Physics, 40(4):2196, April 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1794].

Zirnbauer:1999:ETT

[1959] Martin R. Zirnbauer. Erratum: “Toward a theory of the integer quantum Hall transition: Continuum limit of the Chalker–Coddington model”, [J. Math. Phys. 38, 2007 (1997)]. Journal of Mathematical Physics, 40(4): 2197, April 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1051].

Bender:1999:SQM

[1960] Carl M. Bender, Stefan Boettcher, and Peter N. Meisinger. PT - symmetric quantum mechanics. Journal of Mathematical Physics, 40 (5):2201–2229, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gorbunov:1999:HSB

[1961] I. V. Gorbunov and S. L. Lyakhovich. Hidden supersymmetry and Berezin quantization of N = 2, D = 3 spinning superparticles. Journal of Mathematical Physics, 40(5):2230–2253, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hall:1999:IIP

[1962] Richard L. Hall. An inversion inequality for potentials in quantum mechanics. Journal of Mathematical Physics, 40(5):2254–2258, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ichinose:1999:WAH

[1963] Shoichi Ichinose and Noriaki Ikeda. Weyl anomaly in higher dimensions and Feynman rules in coordinate space. Journal of Mathematical Physics, 40(5):2259–2290, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kalnins:1999:STDb

[1964] E. G. Kalnins, W. Miller, Jr., Ye. M. Hakobyan, and G. S. Pogosyan. Superintegrability on the two-dimensional hyperboloid. II. Journal of Mathematical Physics, 40(5):2291–2306, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Konya:1999:CSM

[1965] B. Kónya and Z. Papp. On the Coulomb Sturmian matrix elements of relativistic Coulomb Green’s operators. Journal of Mathematical Physics, 40(5):2307–2310, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Majid:1999:QGF

[1966] Shahn Majid. Quantum geometry of field extensions. Journal of Mathematical Physics, 40(5):2311–2323, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Martinez-y-Romero:1999:NRA

[1967] R. P. Martínez y Romero, A. L. Salas-Brito, and Jaime Saldaña-Vega. Nonunitary representations of the SU (2 ) algebra in the Dirac equation with a Coulomb potential. Journal of Mathematical Physics, 40(5):2324– 2336, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mourao:1999:PPQ

[1968] J. M. Mourão, T. Thiemann, and J. M. Velhinho. Physical properties of quantum field theory measures. Journal of Mathematical Physics, 40 (5):2337–2353, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Penson:1999:NGC

[1969] K. A. Penson and A. I. Solomon. New generalized coherent states. Journal of Mathematical Physics, 40(5):2354–2363, May 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sajko:1999:WPK

[1970] W. N. Sajko, P. S. Wesson, and H. Liu. Waves and particles in Kaluza– Klein theory. Journal of Mathematical Physics, 40(5):2364–2380, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iwai:1999:CQM

[1971] Toshihiro Iwai. Classical and quantum mechanics of jointed rigid bodies with vanishing total angular momentum. Journal of Mathematical Physics, 40(5):2381–2399, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:1999:ASD

[1972] Wen-Xiu Ma and Benno Fuchssteiner. Algebraic structure of discrete zero curvature equations and master symmetries of discrete evolution equations. Journal of Mathematical Physics, 40(5):2400–2418, May 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mishev:1999:FIK

[1973] Yordan P. Mishev. On the Fay identity for Korteweg–de Vries tau functions and the identity for the Wronskian of squared solutions of Sturm– Liouville equation. Journal of Mathematical Physics, 40(5):2419–2428, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schimming:1999:DPE

[1974] Rainer Schimming and Walter Strampp. Differential polynomial expressions related to the Kadomtsev–Petviashvili and Korteweg–de Vries hierarchies. Journal of Mathematical Physics, 40(5):2429–2444, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wang:1999:AKG

[1975] Bixiang Wang and Horst Lange. Attractors for the Klein–Gordon– Schrödinger equation. Journal of Mathematical Physics, 40(5):2445– 2457, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Benoit:1999:AFS

[1976] Patricia M. Benoit and Alan A. Coley. Anisotropic fluid spherically symmetric space–times admitting a kinematic self-similarity. Journal of Mathematical Physics, 40(5):2458–2469, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Giulini:1999:GTS

[1977] Domenico Giulini. The generalized thin-sandwich problem and its local solvability. Journal of Mathematical Physics, 40(5):2470–2482, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Iriondo:1999:NCL

[1978] Mirta Iriondo, Carlos N. Kozameh, and Alejandra T. Rojas. Null cones from I^{+} and Legendre submanifolds. Journal of Mathematical Physics, 40(5):2483–2493, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:1999:HAS

[1979] Salih Çelik. Hopf algebra structure of Gr_{q} (1—1) related to GL_{q} (1—1). Journal of Mathematical Physics, 40(5):2494–2499, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Matos:1999:ICF

[1980] Tonatiuh Matos, Ulises Nucamendi, and Petra Wiederhold. SU (N )- and SO(N )-invariant chiral fields: One- and two-dimensional subspaces. Journal of Mathematical Physics, 40(5):2500–2513, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Warnaar:1999:TI

[1981] S. Ole Warnaar. q-Trinomial identities. Journal of Mathematical Physics, 40(5):2514–2530, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moller:1999:CPS

[1982] Klaus B. Møller. Comment on phase-space representation of quantum state vectors. Journal of Mathematical Physics, 40(5):2531–2535, May 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [?].

Baker:1999:MNE

[1983] L. M. Baker and D. B. Fairlie. Moyal–Nahm equations. Journal of Mathematical Physics, 40(6):2539–2548, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v40/i6/p2539_s1.

Bodmann:1999:RPI

[1984] Bernhard Bodmann, Hajo Leschke, and Simone Warzel. A rigorous path integral for quantum spin using flat-space Wiener regularization. Journal of Mathematical Physics, 40(6):2549–2559, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2549_s1.

Daniel:1999:GEC

[1985] M. Daniel and K. Manivannan. On the geometric equivalence of certain discrete integrable Heisenberg ferromagnetic spin chains. Journal of Mathematical Physics, 40(6):2560–2567, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2560_s1.

Field:1999:GCS

[1986] Timothy R. Field and Lane P. Hughston. The geometry of coherent states. Journal of Mathematical Physics, 40(6):2568–2583, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/
p2568_s1.

Fleischhack:1999:NTL

[1987] Christian Fleischhack. A new type of loop independence and SU(N) quantum Yang–Mills theory in two dimensions. Journal of Mathematical Physics, 40(6):2584–2610, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p2584_s1. See erratum [2027].

Furlani:1999:QMV

[1988] Edward P. Furlani. Quantization of massive vector fields in curved space– time. Journal of Mathematical Physics, 40(6):2611–2626, June 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2611_
s1.

Kellendonk:1999:ESV

[1989] Johannes Kellendonk. Exact spectral values for discrete quantum pendulum-integrals. Journal of Mathematical Physics, 40(6):2627–2639, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v40/i6/p2627_s1.

Khare:1999:EGS

[1990] Avinash Khare. Exact ground state of several N-body problems with an N-body potential. Journal of Mathematical Physics, 40(6):2640– 2646, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p2640_s1.

Lu:1999:EPG

[1991] Kening Lu and Xing-Bin Pan. Eigenvalue problems of Ginzburg–Landau operator in bounded domains. Journal of Mathematical Physics, 40(6): 2647–2670, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p2647_s1.

March:1999:SSC

[1992] N. H. March, L. M. Nieto, C. Amovilli, L. C. Balbás, and M. L. Glasser. Slater sum for central field problems characterized by its s-wave component alone. Journal of Mathematical Physics, 40(6):2671–2679, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v40/i6/p2671_s1.

Reifler:1999:FST

[1993] Frank Reifler and Randall Morris. Flavor symmetry of the tensor Dirac theory. Journal of Mathematical Physics, 40(6):2680–2697, June 1999.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/
p2680_s1.

Roepstorff:1999:SHF

[1994] G. Roepstorff. Superconnections and the Higgs field. Journal of Mathematical Physics, 40(6):2698–2715, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v40/i6/p2698_s1.

vanWyk:1999:MSP

[1995] C. B. van Wyk. Momentum and spin of a particle with spin unity. Journal of Mathematical Physics, 40(6):2716–2734, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2716_s1.

Vilar:1999:ZCF

[1996] L. C. Q. Vilar, S. P. Sorella, and C. A. G. Sasaki. Zero curvature formalism of the four-dimensional Yang–Mills theory in superspace. Journal of Mathematical Physics, 40(6):2735–2756, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2735_s1.

Zeiner:1999:CSB

[1997] P. Zeiner, R. Dirl, and B. L. Davies. Complete sets of Bloch and Wannier functions composed of oscillator eigenfunctions. Journal of Mathematical Physics, 40(6):2757–2781, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p2757_s1.

Gomez-Ullate:1999:SDD

[1998] D. Gómez-Ullate, S. Lafortune, and P. Winternitz. Symmetries of discrete dynamical systems involving two species. Journal of Mathematical Physics, 40(6):2782–2804, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p2782_s1.

Gough:1999:DCF

[1999] John Gough. Dissipative canonical flows in classical and quantum mechanics. Journal of Mathematical Physics, 40(6):2805–2815, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/
p2805_s1.

Nash:1999:NFI

[2000] Patrick L. Nash. A new first integral for a binary rigid body collision of arbitrarily short duration. Journal of Mathematical Physics, 40(6):2816– 2829, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p2816_s1.

Nizette:1999:TCE

[2001] Michel Nizette and Alain Goriely. Towards a classification of Euler– Kirchhoff filaments. Journal of Mathematical Physics, 40(6):2830–2866, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v40/i6/p2830_s1.

Rivera:1999:ESS

[2002] R. Rivera and D. Villarroel. An exact solution for several charges in classical electrodynamics. Journal of Mathematical Physics, 40(6):2867– 2881, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p2867_s1.

delCastillo:1999:TDF

[2003] G. F. Torres del Castillo and J. Mercado-Pérez. Three-dimensional formulation of the Maxwell equations for stationary space–times. Journal of Mathematical Physics, 40(6):2882–2890, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2882_s1.

Barbachoux:1999:HBB

[2004] C. Barbachoux, F. Debbasch, and J. P. Rivet. Hydrodynamic behavior of Brownian particles in a position-dependent constant force field. Journal of Mathematical Physics, 40(6):2891–2908, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2891_s1.

Melnikov:1999:IDC

[2005] Yu. Melnikov and E. Yarevsky. Integral-difference collision operators: Analytical and numerical spectral analysis. Journal of Mathematical Physics, 40(6):2909–2921, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p2909_s1.

Aratyn:1999:SKP

[2006] H. Aratyn, E. Nissimov, and S. Pacheva. Supersymmetric Kadomtsev– Petviashvili hierarchy: “ghost” symmetry structure, reductions, and Darboux–Bäcklund solutions. Journal of Mathematical Physics, 40(6): 2922–2932, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p2922_s1.

Chen:1999:HFN

[2007] Zhi-De Chen, Xiang-Jun Chen, and Nian-Ning Huang. Hamiltonian formalism for the nonlinear Schrödinger equation in physical space–time. Journal of Mathematical Physics, 60(6):2933–2948, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/p2933_s1.

Gattobigio:1999:NSE

[2008] Mario Gattobigio, Antonio Liguori, and Mihail Mintchev. The nonlinear Schrödinger equation on the half line. Journal of Mathematical Physics, 40(6):2949–2970, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p2949_s1.

Geng:1999:FBS

[2009] Xianguo Geng and Yongtang Wu. Finite-band solutions of the classical Boussinesq–Burgers equations. Journal of Mathematical Physics, 40(6): 2971–2982, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p2971_s1.

Sabatier:1999:NDL

[2010] Pierre C. Sabatier. New direct linearizations for KdV and solutions of the other Cauchy problem. Journal of Mathematical Physics, 40(6):2983– 3020, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p2983_s1.

Tu:1999:HSG

[2011] Ming-Hsien Tu and Jiin-Chang Shaw. Hamiltonian structures of generalized Manin–Radul super-KdV and constrained super KP hierarchies. Journal of Mathematical Physics, 40(6):3021–3034, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/p3021_s1.

Ali:1999:LAG

[2012] G. Alì and John K. Hunter. Large amplitude gravitational waves. Journal of Mathematical Physics, 40(6):3035–3052, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p3035_s1.

Bonilla:1999:SET

[2013] Miguel Á. G. Bonilla and Carlos F. Sopuerta. Super-energy tensor for space–times with vanishing scalar curvature. Journal of Mathematical Physics, 40(6):3053–3062, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p3053_s1.

Cotsakis:1999:VCS

[2014] Spiros Cotsakis, John Miritzis, and Laurent Querella. Variational and conformal structure of nonlinear metric-connection gravitational Lagrangians. Journal of Mathematical Physics, 40(6):3063–3071, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v40/i6/p3063_s1.

Rudnicki:1999:RWM

[2015] Ryszard Rudnicki and Marek Wolf. Random walk with memory. Journal of Mathematical Physics, 40(6):3072–3083, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p3072_s1.

Rudnicki:1999:CCO

[2016] Wiešlaw Rudnicki and Paweł Zieba. Curvature conditions for the occurrence of a class of spacetime singularities. Journal of Mathematical Physics, 40(6):3084–3090, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p3084_s1.

Celik:1999:TPE

[2017] Sultan A. Çelik, Emanullah Hizel, and Salih Çelik. Two-parametric extension of h-deformation of Gr(1|1). Journal of Mathematical Physics, 40(6):3091–3098, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v40/i6/p3091_s1.

Gascon:1999:IVF

[2018] F. G. Gascón and D. Peralta Salas. Instability of vector fields induced by first integrals. Journal of Mathematical Physics, 40(6):3099–3103, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v40/i6/p3099_s1.

Havlicek:1999:NSF

[2019] M. Havlícek, S. Posta, and P. Winternitz. Nonlinear superposition formulas based on imprimitive group action. Journal of Mathematical Physics, 40(6):3104–3122, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v40/i6/p3104_s1.

Hong:1999:CUM

[2020] Jin Hong. Center and universal R-matrix for quantized Borcherds superalgebras. Journal of Mathematical Physics, 40(6):3123–3145, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v40/i6/p3123_s1.

Irac-Astaud:1999:URQ

[2021] M. Irac-Astaud and C. Quesne. Unitary representations of the quantum algebra su_{q}(2) on a real two-dimensional sphere for q ∈ R^{+} or generic q ∈ S^{1}. Journal of Mathematical Physics, 40(6):3146–3161, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v40/i6/p3146_s1.

Khristov:1999:ACK

[2022] E. Kh. Khristov. On an application of Crum–Krein transform to expansions in products of solutions of two Sturm–Liouville equations. Journal of Mathematical Physics, 40(6):3162–3174, June 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v40/i6/p3162_s1.

Lee:1999:GRQ

[2023] H. C. Lee and R. B. Zhang. Geometry and representations of the quantum supergroup OSP_{q}(1|2n). Journal of Mathematical Physics, 40(6): 3175–3190, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p3175_s1.

Miki:1999:TLA

[2024] Kei Miki. Toroidal and level 0 U′_{q}(s_{l_{n+1}}) actions on U_{q}(g_{l_{n+1}}) modules. Journal of Mathematical Physics, 40(6):3191–3210, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v40/i6/p3191_s1.

Ohta:1999:PFO

[2025] Yuji Ohta. Picard–Fuchs ordinary differential systems in N = 2 supersymmetric Yang–Mills theories. Journal of Mathematical Physics, 40(6): 3211–3226, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v40/i6/p3211_s1.

Xu:1999:SCI

[2026] Feng Xu. On the signature of certain intersection forms. Journal of Mathematical Physics, 40(6):3227–3234, June 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v40/i6/p3227_s1.

Fleischhack:2000:ENT

[2027] Christian Fleischhack. Erratum: “A new type of loop independence and SU(N) quantum Yang–Mills theory in two dimensions”, [J. Math. Phys. 40, 2584 (1999)]. Journal of Mathematical Physics, 41(1):614, January 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1987].

Bender:1999:QFT

[2028] Carl M. Bender, Dorje C. Brody, and Bernhard K. Meister. Quantum field theory of partitions. Journal of Mathematical Physics, 40(7):3239– 3245, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dittmann:1999:CMR

[2029] J. Dittmann and A. Uhlmann. Connections and metrics respecting purification of quantum states. Journal of Mathematical Physics, 40(7):3246– 3267, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Finkel:1999:FPT

[2030] Federico Finkel, Artemio González-López, Niky Kamran, and Miguel A. Rodríguez. On form-preserving transformations for the time-dependent Schrödinger equation. Journal of Mathematical Physics, 40(7):3268– 3274, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gyger:1999:LIS

[2031] Sébastien Gyger and Philippe A. Martin. Lifetimes of impurity states in crossed magnetic and electric fields. Journal of Mathematical Physics, 40 (7):3275–3282, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Keyl:1999:OCP

[2032] M. Keyl and R. F. Werner. Optimal cloning of pure states, testing single clones. Journal of Mathematical Physics, 40(7):3283–3299, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Modanese:1999:TMF

[2033] Giovanni Modanese. Tunneling of a massless field through a 3D Gaussian barrier. Journal of Mathematical Physics, 40(7):3300–3310, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mostafazadeh:1999:GAP

[2034] Ali Mostafazadeh. Generalized adiabatic product expansion: A nonperturbative method of solving the time-dependent Schrödinger equation. Journal of Mathematical Physics, 40(7):3311–3326, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Somaroo:1999:GAC

[2035] Shyamal Somaroo, Anthony Lasenby, and Chris Doran. Geometric algebra and the causal approach to multiparticle quantum mechanics. Journal of Mathematical Physics, 40(7):3327–3340, July 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hu:1999:DGI

[2036] Kai Hu. A differential-geometric interpretation of Kirchhoff’s elastic rods. Journal of Mathematical Physics, 40(7):3341–3352, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kane:1999:SEM

[2037] C. Kane, J. E. Marsden, and M. Ortiz. Symplectic-energy-momentum preserving variational integrators. Journal of Mathematical Physics, 40 (7):3353–3371, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shivamoggi:1999:MBL

[2038] Bhimsen K. Shivamoggi and David K. Rollins. Magnetohydrodynamic boundary layer on a flat plate: Further analytic results. Journal of Mathematical Physics, 40(7):3372–3378, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bracken:1999:WES

[2039] P. Bracken, A. M. Grundland, and L. Martina. The Weierstrass–Enneper system for constant mean curvature surfaces and the completely integrable sigma model. Journal of Mathematical Physics, 40(7):3379–3403, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Varga:1999:UDC

[2040] Péter Varga. Unitary deformations and complex soliton equations. Journal of Mathematical Physics, 40(7):3404–3408, July 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:1999:FDI

[2041] Yongtang Wu and Xianguo Geng. A finite-dimensional integrable system associated with the three-wave interaction equations. Journal of Mathematical Physics, 40(7):3409–3430, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zaalani:1999:PSR

[2042] Nadhem Zaalani. Phase space reduction and Poisson structure. Journal of Mathematical Physics, 40(7):3431–3438, July 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Baki:1999:CAL

[2043] P. Baki and J. O. Malo. Cauchy analysis of the linearized skew sector of the massive nonsymmetric gravitational theory. Journal of Mathematical Physics, 40(7):3439–3445, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Borowiec:1999:ACA

[2044] Andrzej Borowiec, Marco Ferraris, Mauro Francaviglia, and Igor Volovich. Almost-complex and almost-product Einstein manifolds from a variational principle. Journal of Mathematical Physics, 40(7):3446– 3464, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cai:1999:NPS

[2045] Rong-Gen Cai and J. B. Griffiths. Null particle solutions in threedimensional (anti-) de Sitter spaces. Journal of Mathematical Physics, 40(7):3465–3475, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Clayton:1999:CGR

[2046] M. A. Clayton. Canonical general relativity: Matter fields in a general linear frame. Journal of Mathematical Physics, 40(7):3476–3489, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Diaz:1999:RDE

[2047] Raúl Durán Díaz, Pedro M. Gadea, and José A. Oubiña. Reductive decompositions and Einstein–Yang–Mills equations associated to the oscillator group. Journal of Mathematical Physics, 40(7):3490–3498, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrero:1999:RCM

[2048] Alicia Herrero and Juan Antonio Morales. Radial conformal motions in Minkowski space–time. Journal of Mathematical Physics, 40(7):3499– 3508, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ilha:1999:DCO

[2049] Anderson Ilha, Antares Kleber, and José P. S. Lemos. Dimensionally continued Oppenheimer–Snyder gravitational collapse: Solutions in odd dimensions. Journal of Mathematical Physics, 40(7):3509–3518, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Szydlowski:1999:DJM

[2050] Marek Szydłowski. Desingularization of Jacobi metrics and chaos in general relativity. Journal of Mathematical Physics, 40(7):3519–3539, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:1999:PMC

[2051] Xiao Zhang. Positive mass conjecture for five-dimensional Lorentzian manifolds. Journal of Mathematical Physics, 40(7):3540–3552, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:1999:TPN

[2052] Salih Çelik and Sultan A. Çelik. Two-parameter nonstandard deformation of 2×2 matrices. Journal of Mathematical Physics, 40(7):3553–3560, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Leite:1999:CFE

[2053] F. Silva Leite and P. Crouch. Closed forms for the exponential mapping on matrix Lie groups based on Putzer’s method. Journal of Mathematical Physics, 40(7):3561–3568, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mugan:1999:SOS

[2054] U. Muğan and A. Sakka. Second-order second-degree Painlevé equations related with Painlevé I–VI equations and Fuchsian-type transformations. Journal of Mathematical Physics, 40(7):3569–3587, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Oeckl:1999:CDC

[2055] Robert Oeckl. Classification of differential calculi on U_{q}(mathfrakb_{+}), classical limits, and duality. Journal of Mathematical Physics, 40(7): 3588–3603, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1999:RWG

[2056] D. J. Rowe, B. C. Sanders, and H. de Guise. Representations of the Weyl group and Wigner functions for SU (3 ). Journal of Mathematical Physics, 40(7):3604–3615, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Scharnhorst:1999:SIM

[2057] K. Scharnhorst. A special irreducible matrix representation of the real Clifford algebra C (3,1). Journal of Mathematical Physics, 40(7):3616– 3631, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simon:1999:CCF

[2058] R. Simon, S. Chaturvedi, and V. Srinivasan. Congruences and canonical forms for a positive matrix: Application to the Schweinler–Wigner extremum principle. Journal of Mathematical Physics, 40(7):3632–3642, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Tanriver:1999:IPA

[2059] Ugur Tanriver and S. Roy Choudhury. Invariant Painlevé analysis and coherent structures of two families of reaction-diffusion equations. Journal of Mathematical Physics, 40(7):3643–3653, July 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wei:1999:FAB

[2060] Yuchuan Wei. Frequency analysis based on general periodic functions. Journal of Mathematical Physics, 40(7):3654–3684, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:1999:CGS

[2061] Wen-Xiu Ma. Comment on “Generalized W_{∞} symmetry algebra of the conditionally integrable nonlinear evolution equation” [J. Math. Phys. 36, 3492 (1995)]. Journal of Mathematical Physics, 40(7):3685–3690, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [254] and response [2062].

Lou:1999:RCG

[2062] Sen yue Lou and Jian ping Weng. Response to “Comment on ‘Generalized W_{∞} symmetry algebra of the conditionally integrable nonlinear evolution equation’ ”, [J. Math. Phys. 40, 3685 (1999)]. Journal of Mathematical Physics, 40(7):3691–3692, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [254, 2061].

Azad:1999:CGP

[2063] H. Azad, M. N. Qureshi, and M. Ziad. Comment on “Geometric phase, bundle classification, and group representation”, [J. Math. Phys. 37, 1218 (1996)]. Journal of Mathematical Physics, 40(7):3693–3696, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [580, 2064].

Mostafazadeh:1999:EGP

[2064] Ali Mostafazadeh. Erratum: “Geometric phase, bundle classification, and group representation”, [J. Math. Phys. 37, 1218 (1996)]. Journal of Mathematical Physics, 40(7):3697, July 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [580, 2063].

Aktosun:1999:ASC

[2065] Tuncay Aktosun and Martin Klaus. Asymptotics of the scattering coefficients for a generalized Schrödinger equation. Journal of Mathematical Physics, 40(8):3701–3709, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bambusi:1999:DHM

[2066] Dario Bambusi. On the dynamics of the Holstein model from the anticontinuous limit. Journal of Mathematical Physics, 40(8):3710–3717, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ban:1999:PSR

[2067] Masashi Ban. Phase-space representation of quantum state vectors: The relative-state approach and the displacement-operator approach. Journal of Mathematical Physics, 40(8):3718–3722, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bassi:1999:ODM

[2068] Zorawar S. Bassi and André LeClair. A one-dimensional model for nlevel atoms coupled to an electromagnetic field. Journal of Mathematical Physics, 40(8):3723–3731, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Berard:1999:MMF

[2069] A. Bérard, Y. Grandati, and H. Mohrbach. Magnetic monopole in the Feynman’s derivation of Maxwell equations. Journal of Mathematical Physics, 40(8):3732–3737, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Castro:1999:GMN

[2070] Carlos Castro and Jerzy Plebański. The generalized Moyal–Nahm and continuous Moyal–Toda equations. Journal of Mathematical Physics, 40 (8):3738–3760, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chim:1999:CCA

[2071] Leung Chim. Central charge and the Andrews–Bailey construction. Journal of Mathematical Physics, 40(8):3761–3772, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guerrero:1999:MIT

[2072] J. Guerrero, M. Calixto, and V. Aldaya. Modular invariance on the torus and Abelian Chern–Simons theory. Journal of Mathematical Physics, 40 (8):3773–3790, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hara:1999:FFA

[2073] Yuji Hara, Michio Jimbo, Hitoshi Konno, Satoru Odake, and Jun’ichi Shiraishi. Free field approach to the dilute A_{L} models. Journal of Mathematical Physics, 40(8):3791–3826, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:1999:ESDa

[2074] M. N. Hounkonnou and J. E. B. Mendy. Exact solutions of the Dirac equation in a nonfactorizable metric. Journal of Mathematical Physics, 40(8):3827–3842, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Moretti:1999:OLS

[2075] Valter Moretti. One-loop stress-tensor renormalization in curved background: The relation between ζ-function and point-splitting approaches, and an improved point-splitting procedure. Journal of Mathematical Physics, 40(8):3843–3875, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ramm:1999:IPI

[2076] A. G. Ramm. Inverse problem for an inhomogeneous Schrödinger equation. Journal of Mathematical Physics, 40(8):3876–3880, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Smirnov:1999:SMC

[2077] Yu. F. Smirnov and Anju Sharma. Spin-mass content of the Bhabha particle and the group chain SU(4) ⊃ Sp(4) ⊃ SU(2) × U(1). Journal of Mathematical Physics, 40(8):3881–3894, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lancellotti:1999:TAT

[2078] Carlo Lancellotti and J. J. Dorning. Time-asymptotic traveling-wave solutions to the nonlinear Vlasov–Poisson–Ampère equations. Journal of Mathematical Physics, 40(8):3895–3917, August 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sonego:1999:CBL

[2079] Sebastiano Sonego. Conformal behavior of the Lorentz–Dirac equation and Machian particle dynamics. Journal of Mathematical Physics, 40(8): 3918–3924, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Veerman:1999:SSF

[2080] J. J. P. Veerman, D. Bazeia, and Fernando Moraes. Soliton stability in a Z (2) field theory. Journal of Mathematical Physics, 40(8):3925–3929, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pastur:1999:RSE

[2081] L. Pastur, M. Shcherbina, and B. Tirozzi. On the replica symmetric equations for the Hopfield model. Journal of Mathematical Physics, 40 (8):3930–3947, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cao:1999:RBK

[2082] Cewen Cao, Yongtang Wu, and Xianguo Geng. Relation between the Kadometsev–Petviashvili equation and the confocal involutive system. Journal of Mathematical Physics, 40(8):3948–3970, August 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Maccari:1999:NID

[2083] Attilio Maccari. A new integrable Davey–Stewartson-type equation. Journal of Mathematical Physics, 40(8):3971–3977, August 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banach:1999:ECPa

[2084] Zbigniew Banach. Equivalence classes of perturbations in cosmologies of Bianchi types I and V: Formulation. Journal of Mathematical Physics, 40(8):3978–3994, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Banach:1999:ECPb

[2085] Zbigniew Banach. Equivalence classes of perturbations in cosmologies of Bianchi types I and V: Interpretation. Journal of Mathematical Physics, 40(8):3995–4010, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carrion:1999:GTS

[2086] H. L. Carrion, M. J. Rebouças, and A. F. F. Teixeira. Gödel-type space– times in induced matter gravity theory. Journal of Mathematical Physics, 40(8):4011–4027, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fernandez-Jambrina:1999:GCO

[2087] L. Fernández-Jambrina. Geodesic completeness of orthogonally transitive cylindrical space–times. Journal of Mathematical Physics, 40(8):4028– 4034, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garecki:1999:CAS

[2088] Janusz Garecki. Canonical angular supermomentum tensors in general relativity. Journal of Mathematical Physics, 40(8):4035–4055, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Glass:1999:SSS

[2089] E. N. Glass and J. P. Krisch. Scale symmetries of spherical string fluids. Journal of Mathematical Physics, 40(8):4056–4063, August 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Herrera:1999:LCS

[2090] L. Herrera, Filipe M. Paiva, and N. O. Santos. The Levi-Civita space– time as a limiting case of the γ space–time. Journal of Mathematical Physics, 40(8):4064–4071, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivashchuk:1999:HKM

[2091] V. D. Ivashchuk, S.-W. Kim, and V. N. Melnikov. Hyperbolic Kac– Moody algebra from intersecting p-branes. Journal of Mathematical Physics, 40(8):4072–4083, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2270].

Montesinos:1999:CTA

[2092] Merced Montesinos. Connection with torsion from Ashtekar–Barbero connection. Journal of Mathematical Physics, 40(8):4084–4088, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:1999:OIP

[2093] Yūji Ohta. One-instanton prepotentials from WDVV equations in N = 2 supersymmetric SU (4 ) Yang–Mills theory. Journal of Mathematical Physics, 40(8):4089–4098, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

delCastillo:1999:SDG

[2094] G. F. Torres del Castillo and H. G. Solís-Rodríguez. Self-duality and gravitational perturbations. Journal of Mathematical Physics, 40(8): 4099–4105, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoniou:1999:SLN

[2095] I. Antoniou, S. A. Shkarin, and Z. Suchanecki. The spectrum of the Liouville–von Neumann operator in the Hilbert–Schmidt space. Journal of Mathematical Physics, 40(8):4106–4118, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bytsenko:1999:AAT

[2096] A. A. Bytsenko, A. E. Gonçalves, M. Simões, and F. L. Williams. Anomalies and analytic torsion on hyperbolic manifolds. Journal of Mathematical Physics, 40(8):4119–4133, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Daboul:1999:MRO

[2097] Jamil Daboul and Robert Delbourgo. Matrix representation of octonions and generalizations. Journal of Mathematical Physics, 40(8):4134–4150, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dallmer:1999:LAA

[2098] Elke Dallmer. On Lie algebras all nilpotent subalgebras of which are Abelian. Journal of Mathematical Physics, 40(8):4151–4156, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ding:1999:HAE

[2099] Jintai Ding. Hopf algebra extension of a Zamolochikov algebra and its double. Journal of Mathematical Physics, 40(8):4157–4164, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klimek:1999:CLL

[2100] M. Klimek. Conservation laws for linear equations on braided linear spaces. Journal of Mathematical Physics, 40(8):4165–4176, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gergely:1999:CCS

[2101] László Á. Gergely. Comment on “The complete Schwarzschild interior and exterior solution in the harmonic coordinate system”, [J. Math. Phys. 39, 6086 (1998)]. Journal of Mathematical Physics, 40(8):4177–4178, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1762] and response [2102].

Liu:1999:RCC

[2102] Quan-Hui Liu. Response to “Comment on ‘The complete Schwarzschild interior and exterior solution in the harmonic coordinate system’ ”, [J. Math. Phys. 40, 4177 (1999)]. Journal of Mathematical Physics, 40(8): 4179, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1762, 2101].

Hu:1999:EHB

[2103] Xing-Biao Hu, Yong-Tang Wu, and Xian-Guo Geng. Erratum: “Hirota bilinear approach to a new integrable differential-difference system”, [J. Math. Phys. 40, 2001 (1999)]. Journal of Mathematical Physics, 40(8): 4180, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1943].

Szmytkowski:1999:EOF

[2104] Radosław Szmytkowski. Erratum: “Operator formulation of Wigner’s R-matrix theories for the Schrödinger and Dirac equations”, [J. Math. Phys. 39, 5231 (1998)]. Journal of Mathematical Physics, 40(8):4181, August 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1699].

Actor:1999:PLB

[2105] Alfred Actor. Point and line boundaries in scalar Casimir theory. Journal of Mathematical Physics, 40(9):4185–4207, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Calogero:1999:PSQ

[2106] F. Calogero. Partially solvable quantum many-body problems in Ddimensional space (D = 1, 2, 3, …). Journal of Mathematical Physics, 40(9):4208–4226, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fujiwara:1999:ETC

[2107] Akio Fujiwara and Hiroshi Nagaoka. An estimation theoretical characterization of coherent states. Journal of Mathematical Physics, 40(9): 4227–4239, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:1999:ESDb

[2108] M. N. Hounkonnou and J. E. B. Mendy. Exact solutions of Dirac equation for neutrinos in presence of external fields. Journal of Mathematical Physics, 40(9):4240–4254, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hounkonnou:1999:ESM

[2109] M. N. Hounkonnou, M. Hounkpe, and J. Shabani. Exactly solvable models of δ′-sphere interactions in nonrelativistic quantum mechanics. Journal of Mathematical Physics, 40(9):4255–4273, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jafarizadeh:1999:DPS

[2110] M. A. Jafarizadeh and S. K. Moayedi. Degeneracy and parasupersymmetry of Dirac Hamiltonian in (2 + 1)-space–time. Journal of Mathematical Physics, 40(9):4274–4289, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Melia:1999:PLE

[2111] Joseph Melia and Michael Redhead. On a purported local extension of the quantum formalism. Journal of Mathematical Physics, 40(9):4290– 4295, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Muravyev:1999:OGP

[2112] Dmitry I. Muravyev. On one generalization for the projection matrices method in the S-matrix factorization problem. Journal of Mathematical Physics, 40(9):4296–4302, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Turgut:1999:CMG

[2113] O. T. Turgut. Classical mechanics and geometric quantization on an infinite dimensional disc and Grassmannian. Journal of Mathematical Physics, 40(9):4303–4324, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhao:1999:GTS

[2114] Wei zhong Zhao. Gauge theory of w symmetry. Journal of Mathematical Physics, 40(9):4325–4330, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ardavan:1999:MHD

[2115] H. Ardavan. Method of handling the divergences in the radiation theory of sources that move faster than their waves. Journal of Mathematical Physics, 40(9):4331–4336, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See comment [2267].

Bahn:1999:NDI

[2116] Changsoo Bahn, Yong Moon Park, and Hyun Jae Yoo. Nonequilibrium dynamics of infinite particle systems with infinite range interactions. Journal of Mathematical Physics, 40(9):4337–4358, September 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shevtsov:1999:BIP

[2117] Boris M. Shevtsov. Backscattering and inverse problem in random media. Journal of Mathematical Physics, 40(9):4359–4373, September 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blohm:1999:SDI

[2118] Henning Blohm. Solution of a discrete inverse scattering problem and of the Cauchy problem of a class of discrete evolution equations. Journal of Mathematical Physics, 40(9):4374–4392, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Corona:1999:GGL

[2119] Gulmaro Corona Corona. Generalized Gel’fand–Levitan integral equation for two block Ablowitz–Kaup–Newell–Segur systems. Journal of Mathematical Physics, 40(9):4393–4399, September 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gentili:1999:IWL

[2120] F. Gentili and G. Morchio. Irreversible weak limits of classical dynamical systems. Journal of Mathematical Physics, 40(9):4400–4418, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:1999:CAK

[2121] Wen-Xiu Ma and Ruguang Zhou. A coupled AKNS–Kaup–Newell soliton hierarchy. Journal of Mathematical Physics, 40(9):4419–4428, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Salasnich:1999:ICS

[2122] Luca Salasnich. Instability and chaos in spatially homogeneous field theories. Journal of Mathematical Physics, 40(9):4429–4443, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Shengfan:1999:DGAb

[2123] Zhou Shengfan. Dimension of the global attractor for damped semilinear wave equations with critical exponent. Journal of Mathematical Physics, 40(9):4444–4451, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:1999:FQB

[2124] Yunbo B. Zeng and Wen-Xiu Ma. Families of quasi-bi-Hamiltonian systems and separability. Journal of Mathematical Physics, 40(9):4452– 4473, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anile:1999:GGW

[2125] A. M. Anile, John K. Hunter, and Binh Truong. The generation of gravitational waves by the resonant interaction of sound waves. Journal of Mathematical Physics, 40(9):4474–4494, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bicak:1999:GWVa

[2126] Jiří Bičák and Jiří Podolský. Gravitational waves in vacuum spacetimes with cosmological constant. I. classification and geometrical properties of nontwisting type N solutions. Journal of Mathematical Physics, 40 (9):4495–4505, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bicak:1999:GWVb

[2127] Jiří Bičák and Jiří Podolský. Gravitational waves in vacuum spacetimes with cosmological constant. II. deviation of geodesics and interpretation of nontwisting type N solutions. Journal of Mathematical Physics, 40 (9):4506–4517, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lu:1999:CKS

[2128] H. Lü, C. N. Pope, and J. Rahmfeld. A construction of Killing spinors on S^{n}. Journal of Mathematical Physics, 40(9):4518–4526, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Simeone:1999:GFG

[2129] Claudio Simeone. Gauge fixation and global phase time for minisuperspaces. Journal of Mathematical Physics, 40(9):4527–4537, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wu:1999:ESS

[2130] S. Q. Wu and X. Cai. Exact solutions to sourceless charged massive scalar field equation on Kerr–Newman background. Journal of Mathematical Physics, 40(9):4538–4548, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hasegawa:1999:CDO

[2131] Koji Hasegawa, Takeshi Ikeda, and Tetsuya Kikuchi. Commuting difference operators arising from the elliptic C_{2}^{(1)}-face model. Journal of Mathematical Physics, 40(9):4549–4568, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kulish:1999:EJT

[2132] P. P. Kulish, V. D. Lyakhovsky, and A. I. Mudrov. Extended Jordanian twists for Lie algebras. Journal of Mathematical Physics, 40(9):4569– 4586, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Petitjean:1999:RMS

[2133] Michel Petitjean. On the root mean square quantitative chirality and quantitative symmetry measures. Journal of Mathematical Physics, 40 (9):4587–4595, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Schmudgen:1999:ORD

[2134] Konrad Schmüdgen. Operator representations of a q-deformed Heisenberg algebra. Journal of Mathematical Physics, 40(9):4596–4605, September 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amorim:1999:EGT

[2135] R. Amorim and J. Barcelos-Neto. Extended gauge theories starting from the matter field Lagrangian. Journal of Mathematical Physics, 40(10): 4609–4615, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bender:1999:LOP

[2136] Carl M. Bender and Gerald V. Dunne. Large-order perturbation theory for a non-Hermitian PT -symmetric Hamiltonian. Journal of Mathematical Physics, 40(10):4616–4621, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Cartas-Fuentevilla:1999:CCF

[2137] R. Cartas-Fuentevilla. A conserved current for the field perturbations in the Einstein–Yang–Mills-dilaton–axion theory. Journal of Mathematical Physics, 40(10):4622–4629, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Exner:1999:NPC

[2138] Pavel Exner and Simeon A. Vugalter. On the number of particles that a curved quantum waveguide can bind. Journal of Mathematical Physics, 40(10):4630–4638, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hwang:1999:GBA

[2139] Stephen Hwang, Robert Marnelius, and Panagiotis Saltsidis. A general BRST approach to string theories with zeta function regularizations. Journal of Mathematical Physics, 40(10):4639–4657, October 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kahn:1999:TDO

[2140] Peter B. Kahn and Yair Zarmi. Time dependence of operators in anharmonic quantum oscillators: Explicit perturbative analysis. Journal of Mathematical Physics, 40(10):4658–4663, October 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kerner:1999:NSQ

[2141] Edward H. Kerner. Nonassociative structure of quantum mechanics in curved space–time. Journal of Mathematical Physics, 40(10):4664–4676, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kovacs:1999:ACA

[2142] Tamás G. Kovács and E. T. Tomboulis. Absence of confinement in the absence of vortices. Journal of Mathematical Physics, 40(10):4677–4687, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lahti:1999:CPO

[2143] Pekka Lahti and Juha-Pekka Pellonpää. Covariant phase observables in quantum mechanics. Journal of Mathematical Physics, 40(10):4688– 4698, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

McInnes:1999:SGS

[2144] Brett McInnes. The semispin groups in string theory. Journal of Mathematical Physics, 40(10):4699–4712, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Naito:1999:QSF

[2145] Seichi Naito. Quantum superstring field theory in the B_{′}-gauge and the physical scattering amplitudes. Journal of Mathematical Physics, 40(10): 4713–4781, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reinsch:1999:ACG

[2146] Matthias W. Reinsch and James J. Morehead. Asymptotics of Clebsch– Gordan coefficients. Journal of Mathematical Physics, 40(10):4782–4806, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

AresdeParga:1999:ESH

[2147] G. Ares de Parga and R. Mares. Exact solution of the Herrera equation of motion in classical electrodynamics. Journal of Mathematical Physics, 40(10):4807–4812, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bal:1999:DES

[2148] Guillaume Bal, George Papanicolaou, and Leonid Ryzhik. Diffusive energy scattering from weakly random surfaces. Journal of Mathematical Physics, 40(10):4813–4827, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Guo:1999:TEG

[2149] Maozheng Guo and Xiao-Ping Wang. Transport equations for a general class of evolution equations with random perturbations. Journal of Mathematical Physics, 40(10):4828–4858, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Fatone:1999:THE

[2150] Lorella Fatone, Cristina Pignotti, Maria Cristina Recchioni, and Francesco Zirilli. Time harmonic electromagnetic scattering from a bounded obstacle: An existence theorem and a computational method. Journal of Mathematical Physics, 40(10):4859–4887, October 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Podlipenko:1999:ESP

[2151] Yu. K. Podlipenko and Yu. V. Shestopalov. On the electromagnetic scattering problem for an infinite dielectric cylinder of an arbitrary cross section located in the wedge. Journal of Mathematical Physics, 40(10): 4888–4902, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

TorresdelCastillo:1999:JPM

[2152] G. F. Torres del Castillo. Jacobi’s principle for magnetic interactions. Journal of Mathematical Physics, 40(10):4903–4910, October 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Woodside:1999:UTC

[2153] Dale A. Woodside. Uniqueness theorems for classical four-vector fields in Euclidean and Minkowski spaces. Journal of Mathematical Physics, 40(10):4911–4943, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Charret:1999:GAF

[2154] I. C. Charret, E. V. Corrêa Silva, S. M. de Souza, O. Rojas Santos, and M. T. Thomaz. Grassmann algebra and fermions at finite temperature. Journal of Mathematical Physics, 40(10):4944–4955, October 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kerimov:1999:UGS

[2155] Azer Kerimov. Uniqueness of Gibbs states in one-dimensional antiferromagnetic model with long-range interaction. Journal of Mathematical Physics, 40(10):4956–4974, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Witte:1999:LAE

[2156] N. S. Witte and D. Bessis. The Lanczos algorithm for extensive manybody systems in the thermodynamic limit. Journal of Mathematical Physics, 40(10):4975–4994, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Doebner:1999:SKH

[2157] H.-D. Doebner and R. Z. Zhdanov. The stationary KdV hierarchy and so (2,1) as a spectrum generating algebra. Journal of Mathematical Physics, 40(10):4995–5003, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kamran:1999:AES

[2158] Niky Kamran and Robert Milson. Algebraic exact solvability of trigonometric-type Hamiltonians associated to root systems. Journal of Mathematical Physics, 40(10):5004–5013, October 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kresic-Juric:1999:LGA

[2159] Saša Krešić-Jurić. A loop group approach to the C. Neumann problem and Moser–Veselov factorization. Journal of Mathematical Physics, 40(10):5014–5025, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ranada:1999:SST

[2160] Manuel F. Rañada and Mariano Santander. Superintegrable systems on the two-dimensional sphere S^{2} and the hyperbolic plane H^{2}. Journal of Mathematical Physics, 40(10):5026–5057, October 1999. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Strachan:1999:DFM

[2161] I. A. B. Strachan. Degenerate Frobenius manifolds and the bi-Hamiltonian structure of rational Lax equations. Journal of Mathematical Physics, 40(10):5058–5079, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yoshida:1999:SBC

[2162] Z. Yoshida and S. M. Mahajan. Simultaneous Beltrami conditions in coupled vortex dynamics. Journal of Mathematical Physics, 40(10):5080– 5091, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Billyard:1999:QAE

[2163] Andrew P. Billyard, Alan A. Coley, and James E. Lidsey. Qualitative analysis of early universe cosmologies. Journal of Mathematical Physics, 40(10):5092–5105, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Binosi:1999:QSF

[2164] Daniele Binosi and Sergio Zerbini. Quantum scalar field in D-dimensional static black hole space–times. Journal of Mathematical Physics, 40(10): 5106–5116, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruna:1999:LSEa

[2165] Lluís Bruna and Joan Girbau. Linearization stability of the Einstein equation for Robertson–Walker models. I. Journal of Mathematical Physics, 40(10):5117–5130, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bruna:1999:LSEb

[2166] Lluís Bruna and Joan Girbau. Linearization stability of the Einstein equation for Robertson–Walker models. II. Journal of Mathematical Physics, 40(10):5131–5137, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Budzynski:1999:DCH

[2167] Robert J. Budzyński, Witold Kondracki, and Andrzej Królak. On the differentiability of Cauchy horizons. Journal of Mathematical Physics, 40(10):5138–5142, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Frittelli:1999:WPF

[2168] Simonetta Frittelli and Oscar A. Reula. Well-posed forms of the 3 + 1 conformally-decomposed Einstein equations. Journal of Mathematical Physics, 40(10):5143–5156, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rainer:1999:EMM

[2169] Martin Rainer. The effective σ-model of multidimensional gravity. Journal of Mathematical Physics, 40(10):5157–5183, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aniello:1999:FIS

[2170] Paolo Aniello, Gianni Cassinelli, Ernesto De Vito, and Alberto Levrero. Frames from imprimitivity systems. Journal of Mathematical Physics, 40(10):5184–5202, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bandos:1999:GSS

[2171] I. Bandos, E. Ivanov, A. A. Kapustnikov, and S. A. Ulanov. General solution of string inspired nonlinear equations. Journal of Mathematical Physics, 40(10):5203–5223, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Becker:1999:NPT

[2172] Peter A. Becker. A new perturbative technique for solving integro-partial differential equations. Journal of Mathematical Physics, 40(10):5224– 5239, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Huang:1999:IRH

[2173] Jingfang Huang. Integral representations of harmonic lattice sums. Journal of Mathematical Physics, 40(10):5240–5246, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Makai:1999:SBV

[2174] M. Makai and Y. Orechwa. Symmetries of boundary value problems in mathematical physics. Journal of Mathematical Physics, 40(10):5247– 5263, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:1999:QHS

[2175] Yao-Zhong Zhang and Mark D. Gould. Quasi-Hopf superalgebras and elliptic quantum supergroups. Journal of Mathematical Physics, 40(10): 5264–5282, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Conte:1999:CEP

[2176] Robert Conte and Micheline Musette. Comment on “Exact periodic solutions of the complex Ginzburg–Landau equation”, [J. Math. Phys. 40, 884 (1999)]. Journal of Mathematical Physics, 40(10):5283, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1870] and response [2177].

Porubov:1999:RCE

[2177] Alexei V. Porubov and Manuel G. Velarde. Response to “Comment on ‘Exact periodic solutions of the complex Ginzburg–Landau equation’ ”, [J. Math. Phys. 40, 5283 (1999)]. Journal of Mathematical Physics, 40 (10):5284, October 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2176].

Aktosun:1999:SES

[2178] Tuncay Aktosun. On the Schrödinger equation with steplike potentials. Journal of Mathematical Physics, 40(11):5289–5305, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amorim:1999:BFF

[2179] Ricardo Amorim and Ronaldo Thibes. On the Batalin, Fradkin, Fradkina, and Tyutin quantization of first order systems. Journal of Mathematical Physics, 40(11):5306–5317, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Challifour:1999:PSF

[2180] John L. Challifour and John P. Clancy. A path space formula for Gauss vectors in Chern–Simons quantum electrodynamics. Journal of Mathematical Physics, 40(11):5318–5340, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chalykh:1999:MIS

[2181] O. A. Chalykh, V. M. Goncharenko, and A. P. Veselov. Multidimensional integrable Schrödinger operators with matrix potential. Journal of Mathematical Physics, 40(11):5341–5355, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Feng:1999:EMD

[2182] Y. J. Feng and C. S. Lam. Evaluation of multiloop diagrams via lightcone integration. Journal of Mathematical Physics, 40(11):5356–5370, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gould:1999:QGF

[2183] Mark D. Gould and Yao-Zhong Zhang. Quasispin graded-fermion formalism and gl(m|n) ↓ osp(m|n) branching rules. Journal of Mathematical Physics, 40(11):5371–5386, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grothaus:1999:CRQ

[2184] Martin Grothaus and Ludwig Streit. Construction of relativistic quantum fields in the framework of white noise analysis. Journal of Mathematical Physics, 40(11):5387–5405, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gusynin:1999:DEE

[2185] V. P. Gusynin and I. A. Shovkovy. Derivative expansion of the effective action for quantum electrodynamics in 2+1 and 3+1 dimensions. Journal of Mathematical Physics, 40(11):5406–5439, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioannidou:1999:MHM

[2186] Theodora Ioannidou and Paul M. Sutcliffe. Monopoles and harmonic maps. Journal of Mathematical Physics, 40(11):5440–5455, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Joye:1999:AES

[2187] A. Joye, F. Monti, S. Guérin, and H. R. Jauslin. Adiabatic evolution for systems with infinitely many eigenvalue crossings. Journal of Mathematical Physics, 40(11):5456–5472, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Khare:1999:NSQ

[2188] Avinash Khare and Uday Sukhatme. New solvable and quasiexactly solvable periodic potentials. Journal of Mathematical Physics, 40(11): 5473–5494, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Knill:1999:RSO

[2189] Oliver Knill. Random Schrödinger operators arising from lattice gauge fields. I. existence and examples. Journal of Mathematical Physics, 40 (11):5495–5510, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Loo:1999:NFP

[2190] Ken Loo. Nonstandard Feynman path integral for the harmonic oscillator. Journal of Mathematical Physics, 40(11):5511–5521, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Margetis:1999:BEC

[2191] Dionisios Margetis. Bose–Einstein condensation in an external potential at zero temperature: Solitary-wave theory. Journal of Mathematical Physics, 40(11):5522–5543, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Molnar:1999:GWU

[2192] Lajos Molnár. A generalization of Wigner’s unitary–antiunitary theorem to Hilbert modules. Journal of Mathematical Physics, 40(11):5544–5554, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Morales:1999:TNP

[2193] J. Morales and J. J. Peña. Two new potentials for the free particle model. Journal of Mathematical Physics, 40(11):5555–5564, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pavon:1999:DWF

[2194] Michele Pavon. Derivation of the wave function collapse in the context of Nelson’s stochastic mechanics. Journal of Mathematical Physics, 40 (11):5565–5577, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pushnitski:1999:ESS

[2195] Alexander Pushnitski. Estimates for the spectral shift function of the polyharmonic operator. Journal of Mathematical Physics, 40(11):5578– 5592, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Reuter:1999:SDK

[2196] M. Reuter. Symplectic Dirac–Kähler fields. Journal of Mathematical Physics, 40(11):5593–5640, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sato:1999:WLG

[2197] Haru-Tada Sato. World-line Green functions with momentum and source conservations. Journal of Mathematical Physics, 40(11):5641– 5656, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Savvidou:1999:AOC

[2198] K. Savvidou. The action operator for continuous-time histories. Journal of Mathematical Physics, 40(11):5657–5674, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Yanez:1999:EIH

[2199] R. J. Yáñez, W. Van Assche, R. González-Férez, and Jesús S. Dehesa. Entropic integrals of hyperspherical harmonics and spatial entropy of Ddimensional central potentials. Journal of Mathematical Physics, 40(11): 5675–5686, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ambjorn:1999:BAT

[2200] J. Ambjørn, A. Avakyan, T. Hakobyan, and A. Sedrakyan. Bethe Ansatz and thermodynamic limit of affine quantum group invariant extensions of the t – J model. Journal of Mathematical Physics, 40(11):5687–5701, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lesniewski:1999:MRM

[2201] Andrew Lesniewski and Mary Beth Ruskai. Monotone Riemannian metrics and relative entropy on noncommutative probability spaces. Journal of Mathematical Physics, 40(11):5702–5724, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Blaszak:1999:TSM

[2202] Maciej Błaszak. Theory of separability of multi-Hamiltonian chains. Journal of Mathematical Physics, 40(11):5725–5748, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gordoa:1999:NSP

[2203] Pilar R. Gordoa and Andrew Pickering. Nonisospectral scattering problems: A key to integrable hierarchies. Journal of Mathematical Physics, 40(11):5749–5786, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Takasaki:1999:ECM

[2204] Kanehisa Takasaki. Elliptic Calogero–Moser systems and isomonodromic deformations. Journal of Mathematical Physics, 40(11):5787–5821, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

vanDiejen:1999:SFR

[2205] J. F. van Diejen. A Sato formula for reflectionless finite difference operators. Journal of Mathematical Physics, 40(11):5822–5834, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Germinario:1999:EMR

[2206] Anna Germinario and Fabio Giannoni. Existence and multiplicity results for massive particles trajectories in a universe with boundary. Journal of Mathematical Physics, 40(11):5835–5848, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivanov:1999:ASE

[2207] B. V. Ivanov. Algebraically special, expanding and twisting gravitational fields with vanishing Newman–Unti–Tamburino parameter. Journal of Mathematical Physics, 40(11):5849–5859, November 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Klauder:1999:NQG

[2208] John R. Klauder. Noncanonical quantization of gravity. I. foundations of affine quantum gravity. Journal of Mathematical Physics, 40(11):5860– 5882, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Qadir:1999:SSR

[2209] Asghar Qadir and Azad A. Siddiqui. K slicing the Schwarzschild and the Reissner–Nordstrom spacetimes. Journal of Mathematical Physics, 40(11):5883–5889, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rudolph:1999:ESE

[2210] Gerd Rudolph, Torsten Tok, and Igor Volobuev. Exact solutions in Einstein–Yang–Mills–Dirac systems. Journal of Mathematical Physics, 40(11):5890–5904, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sneddon:1999:IAI

[2211] G. E. Sneddon. The identities of the algebraic invariants of the fourdimensional Riemann tensor. III. Journal of Mathematical Physics, 40 (11):5905–5920, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Aizawa:1999:SR

[2212] N. Aizawa. Symplecton for U_{⟨}(sl(2 )) and representations of SL_{h}(2 ). Journal of Mathematical Physics, 40(11):5921–5938, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alisauskas:1999:TCT

[2213] Sigitas Ališauskas. Towards the canonical tensor operators of u_{q}(3). II. the denominator function problem. Journal of Mathematical Physics, 40 (11):5939–5955, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Antoine:1999:GWC

[2214] J.-P. Antoine and I. Mahara. Galilean wavelets: Coherent states of the affine Galilei group. Journal of Mathematical Physics, 40(11):5956–5971, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Carnovale:1999:BFT

[2215] Giovanna Carnovale. On the braided Fourier transform on the ndimensional quantum space. Journal of Mathematical Physics, 40(11): 5972–5997, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Celik:1999:TPD

[2216] Salih Çelik and Sultan A. Çelik. Two-parameter differential calculus on the h-superplane. Journal of Mathematical Physics, 40(11):5998–6008, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chryssomalakos:1999:HOB

[2217] C. Chryssomalakos, J. A. de Azcárraga, A. J. Macfarlane, and J. C. Pérez Bueno. Higher-order BRST and anti-BRST operators and cohomology for compact Lie algebras. Journal of Mathematical Physics, 40(11):6009– 6032, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Karp:1999:PIF

[2218] Robert L. Karp, Freydoon Mansouri, and Jung S. Rno. Product integral formalism and non-Abelian Stokes theorem. Journal of Mathematical Physics, 40(11):6033–6043, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kleinert:1999:FDW

[2219] H. Kleinert and A. Chervyakov. Functional determinants from Wronski Green functions. Journal of Mathematical Physics, 40(11):6044–6051, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lagraa:1999:IQG

[2220] M. Lagraa and N. Touhami. The inhomogeneous quantum groups from differential calculi with classical dimension. Journal of Mathematical Physics, 40(11):6052–6070, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ma:1999:BTS

[2221] Wen-Xiu Ma, Xing-Biao Hu, Si-Ming Zhu, and Yong-Tang Wu. Bäcklund transformation and its superposition principle of a Blaszak–Marciniak four-field lattice. Journal of Mathematical Physics, 40(11):6071–6086, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Repka:1999:SFD

[2222] Joe Repka and Hubert de Guise. Some finite dimensional indecomposable representations of E (2). Journal of Mathematical Physics, 40(11):6087– 6109, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhang:1999:LOR

[2223] Yao-Zhong Zhang. Level-one representations and vertex operators of quantum affine superalgebra U_{q}[gl(N̂ |N )]. Journal of Mathematical Physics, 40(11):6110–6124, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Brau:1999:EAS

[2224] F. Brau. Erratum: “Analytical solution of the relativistic Coulomb problem with a hard core interaction for a one-dimensional spinless Salpeter equation”, [J. Math. Phys. 40, 1119 (1999)]. Journal of Mathematical Physics, 40(11):6125, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1889].

Moreschi:1999:ERF

[2225] Osvaldo M. Moreschi and Sergio Dain. Erratum: “Rest frame system for asymptotically flat space–times”, [J. Math. Phys. 39, 6631 (1998)]. Journal of Mathematical Physics, 40(11):6126–6127, November 1999. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1804].

Baloitcha:1999:QDR

[2226] E. Baloïtcha and M. N. Hounkonnou. Quantum description of rigidly or adiabatically constrained four-particle systems and supersymmetry. Journal of Mathematical Physics, 40(12):6133–6144, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dzieciol:1999:CWF

[2227] Aleksander Dzieciol, Staffan Yngve, and Per Olof Fröman. Coulomb wave functions with complex values of the variable and the parameters. Journal of Mathematical Physics, 40(12):6145–6166, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Dzieciol:1999:PIF

[2228] Aleksander Dzieciol, Staffan Yngve, and Per Olof Fröman. Phase-integral formulas for Coulomb wave functions with complex values of the variable and the parameters. Journal of Mathematical Physics, 40(12):6167–6177, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garrahan:1999:CDS

[2229] J. P. Garrahan and M. Kruczenski. Collective dynamics of solitons and inequivalent quantizations. Journal of Mathematical Physics, 40(12): 6178–6188, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Geyer:1999:OCLb

[2230] B. Geyer, P. M. Lavrov, and D. Mülsch. The osp(1,2)-covariant Lagrangian quantization of reducible massive gauge theories. Journal of Mathematical Physics, 40(12):6189–6208, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hiroshima:1999:GSM

[2231] Fumio Hiroshima. Ground states of a model in nonrelativistic quantum electrodynamics. I. Journal of Mathematical Physics, 40(12):6209–6222, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioannidou:1999:SSS

[2232] T. Ioannidou, B. Piette, and W. J. Zakrzewski. Spherically symmetric solutions of the SU (N ) Skyrme models. Journal of Mathematical Physics, 40(12):6223–6233, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Jakel:1999:TAP

[2233] Christian D. Jäkel. Two algebraic properties of thermal quantum field theories. Journal of Mathematical Physics, 40(12):6234–6244, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kostelecky:1999:NQH

[2234] V. Alan Kostelecký and Charles D. Lane. Nonrelativistic quantum Hamiltonian for Lorentz violation. Journal of Mathematical Physics, 40(12):6245–6253, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lambiase:1999:CEB

[2235] G. Lambiase, V. V. Nesterenko, and M. Bordag. Casimir energy of a ball and cylinder in the zeta function technique. Journal of Mathematical Physics, 40(12):6254–6265, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mandelzweig:1999:QMV

[2236] V. B. Mandelzweig. Quasilinearization method and its verification on exactly solvable models in quantum mechanics. Journal of Mathematical Physics, 40(12):6266–6291, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ohta:1999:PFE

[2237] Yūji Ohta. Picard–Fuchs equations and Whitham hierarchy in N = 2 supersymmetric SU (r + 1 ) Yang–Mills theory. Journal of Mathematical Physics, 40(12):6292–6301, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Poirier:1999:WWC

[2238] Bill Poirier. Wigner–Weyl correspondence and semiclassical quantization in spherical coordinates. Journal of Mathematical Physics, 40(12):6302– 6318, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zhdanov:1999:SSE

[2239] Renat Zhdanov and Alexander Zhalij. On separable Schrödinger equations. Journal of Mathematical Physics, 40(12):6319–6338, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gavrilov:1999:ESE

[2240] L. Gavrilov and A. M. Perelomov. On the explicit solutions of the elliptic Calogero system. Journal of Mathematical Physics, 40(12):6339–6352, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ioannidou:1999:SSH

[2241] T. Ioannidou, B. Piette, and W. J. Zakrzewski. SU( N ) skyrmions and harmonic maps. Journal of Mathematical Physics, 40(12):6353–6365, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rauch-Wojciechowski:1999:QLS

[2242] Stefan Rauch-Wojciechowski, Krzysztof Marciniak, and Hans Lundmark. Quasi-Lagrangian systems of Newton equations. Journal of Mathematical Physics, 40(12):6366–6398, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Boillat:1999:RGM

[2243] Guy Boillat and Tommaso Ruggeri. Relativistic gas: Moment equations and maximum wave velocity. Journal of Mathematical Physics, 40(12): 6399–6406, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Sato:1999:IRT

[2244] Haru-Tada Sato. Integral representations of thermodynamic 1PI Green’s functions in the world-line formalism. Journal of Mathematical Physics, 40(12):6407–6429, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gallavotti:1999:FTA

[2245] G. Gallavotti, G. Gentile, and V. Mastropietro. A field theory approach to Lindstedt series for hyperbolic tori in three time scales problems. Journal of Mathematical Physics, 40(12):6430–6472, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gurses:1999:CRO

[2246] Metin Gürses, Atalay Karasu, and Vladimir V. Sokolov. On construction of recursion operators from Lax representation. Journal of Mathematical Physics, 40(12):6473–6490, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lou:1999:FVS

[2247] Sen yue Lou and Li-Li Chen. Formal variable separation approach for nonintegrable models. Journal of Mathematical Physics, 40(12):6491– 6500, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Willox:1999:KCR

[2248] R. Willox and I. Loris. KP constraints from reduced multi-component hierarchies. Journal of Mathematical Physics, 40(12):6501–6525, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zeng:1999:SVS

[2249] Yunbo Zeng and Wen-Xiu Ma. Separation of variables for soliton equations via their binary constrained flows. Journal of Mathematical Physics, 40(12):6526–6557, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Ivashchuk:1999:CSS

[2250] V. D. Ivashchuk and V. N. Melnikov. Cosmological and spherically symmetric solutions with intersecting p-branes. Journal of Mathematical Physics, 40(12):6558–6576, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Pengpan:1999:KCD

[2251] Teparksorn Pengpan. Kostant’s cubic Dirac operator of Lie superalgebras. Journal of Mathematical Physics, 40(12):6577–6588, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rainer:1999:CCS

[2252] M. Rainer. Cones and causal structures on topological and differentiable manifolds. Journal of Mathematical Physics, 40(12):6589–6597, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [2265].

Bikbov:1999:LPS

[2253] R. Bikbov and S. Nechaev. On the limiting power of the set of knots generated by 1 + 1- and 2 + 1-braids. Journal of Mathematical Physics, 40(12):6598–6608, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Grosse:1999:GDG

[2254] H. Grosse and G. Reiter. Graded differential geometry of graded matrix algebras. Journal of Mathematical Physics, 40(12):6609–6625, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hoffman:1999:EMP

[2255] Gary G. Hoffman. Even moments of a polymer chain distribution. Journal of Mathematical Physics, 40(12):6626–6637, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Palmer:1999:ZJM

[2256] John Palmer. Zeros of the Jimbo, Miwa, Ueno tau function. Journal of Mathematical Physics, 40(12):6638–6681, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Piccione:1999:NMI

[2257] Paolo Piccione and Daniel V. Tausk. A note on the Morse index theorem for geodesics between submanifolds in semi-Riemannian geometry. Journal of Mathematical Physics, 40(12):6682–6688, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rosengren:1999:APT

[2258] Hjalmar Rosengren. Another proof of the triple sum formula for Wigner 9j-symbols. Journal of Mathematical Physics, 40(12):6689–6691, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

VanderJeugt:1999:IGT

[2259] J. Van der Jeugt and K. Srinivasa Rao. Invariance groups of transformations of basic hypergeometric series. Journal of Mathematical Physics, 40(12):6692–6700, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Varsaie:1999:KS

[2260] S. Varsaie. Kählerian supermanifolds. Journal of Mathematical Physics, 40(12):6701–6706, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Garg:1999:EAD

[2261] Anupam Garg. Erratum: “Application of the discrete Wentzel–Kramers– Brillouin method to spin tunnelling”, [J. Math. Phys. 39, 5166 (1998)]. Journal of Mathematical Physics, 40(12):6707, December 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1695].

Byrd:2000:EDG

[2262] Mark Byrd. Erratum: “Differential geometry on SU(3) with applications to three state systems”, [J. Math. Phys. 39, 6125–6136 (1998)]. Journal of Mathematical Physics, 41(2):1026–1030, February 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1766].

Branson:2000:EVE

[2263] Thomas P. Branson, Peter B. Gilkey, and Dmitri V. Vassilevich. Erratum: “Vacuum expectation value asymptotics for second-order differential operators on manifolds with boundary”, [J. Math. Phys. 39, 1040 (1998)]. Journal of Mathematical Physics, 41(5):3301, May 2000. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1436].

Raghavan:2000:EVB

[2264] K. N. Raghavan and P. Sankaran. Erratum: “On Verma bases for representations of sl(n, C)”, [J. Math. Phys. 40, 2190 (1999)]. Journal of Mathematical Physics, 41(5):3302, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1957].

Rainer:2000:ECC

[2265] M. Rainer. Erratum: “Cones and causal structures on topological and differentiable manifolds”, [J. Math. Phys. 40, 6589 (1999)]. Journal of Mathematical Physics, 41(5):3303, May 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2252].

Saenz:2000:ELS

[2266] A. W. Sáenz. Erratum: “Lie-series approach to the evolution of resonant and nonresonant anharmonic oscillators in quantum mechanics”, [J. Math. Phys. 39, 1887 (1998)]. Journal of Mathematical Physics, 41 (7):5135, July 2000. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1486].

Hannay:2001:CMH

[2267] J. H. Hannay. Comment on “Method of handling the divergences in the radiation theory of sources that move faster than their waves”, [J. Math. Phys. 40, 4331 (1999)]. Journal of Mathematical Physics, 42(8): 3973–3974, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2115].

Krikorian:2001:CUP

[2268] R. A. Krikorian. Comment on “Uncertainty principle for proper time and mass”, [J. Math. Phys 40, 1237 (1999)]. Journal of Mathematical Physics, 42(8):3975–3977, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1897] and response [2269].

Kudaka:2001:RCU

[2269] Shoju Kudaka and Shuichi Matsumoto. Response to “Comment on ‘Uncertainty principle for proper time and mass’ ”, [J. Math. Phys. 42, 3975 (2001)]. Journal of Mathematical Physics, 42(8):3978–3979, August 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1897, 2268].

Ivashchuk:2001:EHK

[2270] V. D. Ivashchuk, S.-W. Kim, and V. N. Melnikov. Erratum: “Hyperbolic Kac–Moody algebra from intersecting p-branes”, [J. Math. Phys. 40, 4072 (1999)]. Journal of Mathematical Physics, 42(11):5493, November 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2091].

Aldaya:2003:ECC

[2271] V. Aldaya and J. Guerrero. Erratum: “Canonical coherent states for the relativistic harmonic oscillator”, [J. Math. Phys. 36, 3191 (1995)]. Journal of Mathematical Physics, 44(11):5456, November 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [232].

Chadan:2004:ESC

[2272] K. Chadan and R. Kobayashi. Erratum: “A sufficient condition for the existence of bound states in a potential”, [J. Math. Phys. 38, 4900 (1997)]. Journal of Mathematical Physics, 45(7):2987, July 2004. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v45/i7/p2987_

s1. See [1259].

Zhang:2010:CTP

[2273] Shun Zhang. Comment on “On tractability of path integration”, [J. Math. Phys. 37, 2071 (1996)]. Journal of Mathematical Physics, 51(2):024101, February 2010. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v51/i2/p024101_s1. See [633].